### v1

# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

Abstract
We develop a theory to analyze how structure in connectivity shapes the high-dimensional, internally generated activity of nonlinear recurrent neural networks. Using two complementary methods – a path-integral calculation of fluctuations around the saddle point, and a recently introduced two-site cavity approach – we derive analytic expressions that characterize important features of collective activity, including its dimensionality and temporal correlations. To model structure in the coupling matrices of real neural circuits, such as synaptic connectomes obtained through electron microscopy, we introduce the random-mode model, which parameterizes a coupling matrix using random input and output modes and a specified spectrum. This model enables systematic study of the effects of low-dimensional structure in connectivity on neural activity. These effects manifest in features of collective activity, that we calculate, and can be undetectable when analyzing only single-neuron activities. We derive a relation between the effective rank of the coupling matrix and the dimension of activity. By extending the random-mode model, we compare the effects of single-neuron heterogeneity and low-dimensional connectivity. We also investigate the impact of structured overlaps between input and output modes, a feature of biological coupling matrices. Our theory provides tools to relate neural-network architecture and collective dynamics in artificial and biological systems.
I Introduction
Nonlinear recurrent neural networks are used in neuroscience as models of neural-circuit dynamics [1, 2, 3] and in machine learning as systems for sequence processing [4, 5, 6, 7]. Such networks can generate complex, time-varying activity in the absence of external input [8, 9]. This high-dimensional chaotic activity, first analyzed by Sompolinsky et al. [10], has served as a model for asynchronous cortical dynamics observed in vivo. Moreover, initializing recurrent neural networks so that they exhibit chaotic dynamics has been shown to facilitate subsequent learning [11, 12, 13].
Theoretical studies often focus on large recurrent neural networks with independent and identically distributed (i.i.d.) couplings. However, the structure of real-world networks, including neural circuits, typically deviates substantially from this assumption [14]. Many real-world networks show approximate low-rank structure, meaning that their connectivity is well described by a number of rank-one components that is smaller than the number of units in the network [15]. Often, networks have “smooth low-rank” or “effectively low-rank” structure in which the strengths of these rank-one components do not exhibit a hard cutoff and instead decay smoothly, with a more rapid falloff than in networks with i.i.d. couplings. This can arise when connections between units depend on their distance in physical space, or in a low-dimensional feature space [16, 17, 18]. Such low-rank structure can also emerge when artificial neural networks are trained on tasks [19, 20].
To illustrate this, we examined a central-brain connectome of the fruit fly Drosophila melanogaster [21] (Fig. 1a). We estimated strengths from synaptic counts, assigned coupling signs based on predicted neurotransmitter types [22], and normalized the couplings (Fig. 1b; Appendix B). We analyzed the network’s structure using its singular value decomposition (SVD), which expresses the connectivity as a sum of rank-one components, with the strength of each component given by the corresponding singular value. The singular values decay much more quickly than those of an i.i.d. matrix of the same size, which follow a universal form that does not depend on details of the single-element distribution (Fig. 1c). We quantify this decay rate using the participation ratio of the squared singular-value spectrum (Sec. II.4.1). For this connectome, the participation ratio is $0.18$, much lower than the value of $0.5$ for an i.i.d. matrix, demonstrating that this network has a much lower-rank structure than would be expected for i.i.d. couplings.
Figure 1: Analysis of Drosophila central brain connectome. (a) Volume of fly brain for which reconstruction was performed (blue; reproduced from [21]). (b) Visualization of the full normalized coupling matrix. (c) Singular value spectra of the normalized coupling matrix (red) and an i.i.d. random matrix (gray). The fly connectome exhibits a smooth spectrum that decays quickly, corresponding to a reduced participation ratio. $N=18028$ neurons.
In neuroscience, the importance of connectivity lies in its role in controlling neural activity. To this end, dynamical mean-field theory (DMFT) has been used to analyze activity in large recurrent neural networks, focusing on single-neuron properties such as the autocovariance function. However, recent neuroscience studies have highlighted the significance of geometric or other collective properties that emerge only when examining the simultaneous activity of large neuronal populations [23, 24]. Given neural data, these properties are often analyzed using dimensionality reduction techniques like principal component analysis. Recent advances have extended DMFT techniques to analytically characterize these collective features in nonlinear network models producing high-dimensional chaotic activity [25]. Other studies have addressed simpler cases, including linear networks driven by noise [26, 27, 28, 29] and nonlinear networks producing low-dimensional activity [30, 31]. A key quantity of interest is the dimension of activity, which quantifies the number of high-variance activity modes. This metric reflects a network’s capacity for complex computations and representations, and influences its learning and generalization capabilities [32, 33]. Despite these advances, the analytical calculation of activity dimension in nonlinear networks has been limited to i.i.d. connectivity, or connectivity with simple correlations between reciprocal synapses, and thus cannot be performed for networks with more realistic structures (Fig. 1c). To address this gap, we study networks with smooth low-rank parameterizations of couplings, and develop DMFT tools to analyze collective activity in such networks. We show that, in certain cases, the effects of structure in connectivity are not evident in single-neuron activity, but instead revealed through dramatic changes in properties of collective activity that we show how to compute.
This paper has two main contributions. The first is a new path-integral calculation of a specific four-point function of network activity, originally defined by Clark et al. [25]. This function captures statistics describing collective activity, including its dimensionality. In Secs. II.1II.3, we describe the recurrent-network model and define this four-point function. We explain its connection to collective activity features, including the activity dimension and timescales of principal components, and review its calculation in i.i.d. networks. In Secs. III.1III.4, we introduce the new path-integral calculation and apply it to both the i.i.d. model and a model with correlations between reciprocal pairs of synapses.
The second contribution is a tractable model of neural-network coupling matrices that we call the random-mode model. In Sec. II.4, we define this model, in which couplings are generated as a sum of rank-one outer products of random input and output modes with component-specific strengths. In Sec. II.4.1, we demonstrate that parameterizing these strengths provides control over the singular value spectrum of the coupling matrix, enabling investigation of the interplay between global connectivity structure and network dynamics. In Sec. III.5, we calculate the two- and four-point functions for the random-mode model using the path-integral approach. This analysis reveals a bipartite coupling between neurons and latent variables. In Sec. III.6, we interpret the resulting formulas for the four-point function, which show how the effective rank of the coupling matrix controls the dimension of activity. In Sec. IV, we analyze limiting cases of the random-mode model, illustrating that a vanishing effective rank leads to vanishing activity dimension. In Sec. V, we show that the same results can be obtained using an intuitive two-site cavity method. In Sec. VI, we study the behavior of a generalization of the random-mode model featuring heterogeneity among single-neuron properties and contrast the effects with those of low-rank structure in the connectivity. Finally, in Sec. VII, we incorporate a simple form of overlaps between input and output modes into the model, which we anticipate to be crucial for describing neural computations.
II Model and summary statistics
II.1 Recurrent neural network model
We study a recurrent neural network of $N$ neurons. Each neuron is characterized by its pre-activation $x_{i}(t)$ and activation $\phi_{i}(t)$, where $i=1,\ldots,N$. The activations are related to the pre-activations through a nonlinear function $\phi(x)$, with $\phi_{i}(t)=\phi(x_{i}(t))$. While our results are agnostic to the form of the nonlinearity, when comparing theory and simulations we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$, which is similar to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals. The network dynamics are governed by
$T[x_{i}](t)=\sum_{j}J_{ij}\phi_{j}(t),$ (1)
where $J_{ij}$ is the coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional describing single-neuron dynamics. A canonical choice is
$T[x](t)=(1+\partial_{t})x(t),$ (2)
resulting in low-pass filtering of the input. We measure time in units of the time constant that would appear in front of $\partial_{t}$.
II.2 Correlation functions
II.2.1 Two-point functions
To characterize the network dynamics, we introduce several types of two-point correlation functions. Let ${a\in\{x,\phi\}}$ denote either the pre-activations or activations. We first define the neuron- and time-averaged covariance:
$C^{a}(\tau)=\frac{1}{N}\sum_{i}\left\langle a_{i}(t)a_{i}(t+\tau)\right\rangle_{t}.$ (3)
We also define the time-averaged, but not neuron-averaged, covariance; and the neuron-averaged, but not time-averaged, covariance:
$C^{a}_{ij}(\tau)$ $=\left\langle a_{i}(t)a_{j}(t+\tau)\right\rangle_{t},$ (4)
$C^{a}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i}a_{i}(t_{1})a_{i}(t_{2}).$ (5)
For a given $\tau$, the spectra of $C^{a}_{ij}(\tau)$, with indices $i$ and $j$, and of $C^{a}(t_{1},t_{2}+\tau)$, with indices $t_{1}$ and $t_{2}$, are the same (up to a scaling), so spectral properties of one parameter can be computed using the other, a form of duality. For example, the properly scaled trace operations applied to these parameters yield the same result, namely, $C^{a}(\tau)$.
We also define the response function:
$S^{a}_{ij}(t_{1},t_{2})=\left.\frac{\delta a_{i}(t_{1})}{\delta I_{j}(t_{2})}\right|_{I_{j}=0},$ (6)
where $I_{j}(t)$ is an infinitesimal source term added to the right-hand side of the network dynamics (Eq. 1). In analogy with the covariance functions, we define $S^{a}_{ij}(\tau)$, $S^{a}(t_{1},t_{2})$, and $S^{a}(\tau)$ to be the response function averaged over time, neurons, or both.
II.2.2 Four-point functions
Figure 2: Relationship between principal component timescales and two- and four-point correlation functions. (a) Example activity traces of principal components $p^{\phi}_{k}(t)$. (b) Empirical autocovariance functions for each principal component, weighted by their eigenvalues. (c) Same as (b) but weighted by the squared eigenvalue. (d) Averages over curves in (b) and (c), with comparisons to $C^{\phi}_{\star}(\tau)$ and $\Psi^{\phi}_{\star}(\tau,0)$. Dots: simulation results. Lines: theoretical predictions. All panels describe the same network of $N=4000$ neurons with i.i.d. couplings and $g=6$.
To characterize the dimension and timescales of activity, we define
$\Psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i,j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}),$ (7)
where $\bm{\tau}=(\tau_{1},\tau_{2})$. This sum includes the on-diagonal contributions, unlike the quantity computed in Clark et al. [25], which we denote by the lower-case $\psi^{a}(\bm{\tau})$:
$\psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i\neq j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}).$ (8)
For the type of connectivity we will study, the network exhibits high-dimensional activity such that pairwise covariances are $\mathcal{O}(1/\sqrt{N})$. Thus, in the definition of $\Psi^{a}(\bm{\tau})$ (Eq. 7), the on-diagonal contributions (given by $N^{-1}\sum_{i}C^{a}_{ii}(\tau_{1})C^{a}_{ii}(\tau_{2})$) and off-diagonal contributions (given by $\psi^{a}(\bm{\tau})$) are both $\mathcal{O}(1)$.
Leveraging neuron-time duality, $\Psi^{a}(\bm{\tau})$ can also be expressed in terms of the time-by-time covariance $C^{a}(t_{1},t_{2})$:
$\Psi^{a}(\bm{\tau})=N\left\langle C^{a}(t_{1},t_{2})C^{a}(t_{1}+\tau_{1},t_{2}+\tau_{2})\right\rangle_{t_{1},t_{2}}.$ (9)
The neuron-by-neuron (Eq. 7) and time-by-time (Eq. 9) definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively.
II.2.3 Relating $\Psi^{a}$ to the dimension and timescales of activity
The dimension of activity can be quantified by the participation ratio of its covariance spectrum [34, 35, 23, 36, 32, 37, 9, 38, 27]. To define the participation ratio, we diagonalize the equal-time covariance matrix $C_{ij}^{a}(0)$ to obtain eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$, as in principal components analysis. The participation ratio is
$\text{PR}^{a}=\frac{1}{N}\frac{(\sum_{k}\lambda^{a}_{k})^{2}}{\sum_{k}(\lambda^{a}_{k})^{2}},$ (10)
[…]
An advantageous property of this dimension measure is that it can be expressed in terms of the previously defined two- and four-point correlation functions as
$\text{PR}^{a}=\frac{C^{a}(0)^{2}}{\Psi^{a}(0,0)}.$ (11)
Recall that for the types of connectivity we study, cross-covariances are $\mathcal{O}(1/\sqrt{N})$, so the on- and off-diagonal contributions to $\Psi^{a}(\bm{\tau})$ are both order-one (Eq. 7). This leads to $\text{PR}^{a}$ taking a value between 0 and 1 as $N\rightarrow\infty$. Thus, the unnormalized dimension of activity $N\,\text{PR}^{a}$ is extensive, i.e., activity is high-dimensional. At the same time, activity does not uniformly occupy all available dimensions, which would require cross-covariances to be smaller than $\mathcal{O}(1/\sqrt{N})$.
Timescales of collective activity can be studied as follows. We define the principal components (PCs) of network activity as
$p^{a}_{k}(t)=\frac{(\bm{v}^{a}_{k})^{T}\bm{a}(t)}{\sqrt{\lambda^{a}_{k}}}.$ (12)
These PCs are the basis of much of modern analysis of high-dimensional neural data. The PCs all have unit variance but potentially very different characteristic timescales (Fig. 2a). To extract the timescales of just the leading PCs, we compute the PC-averaged covariance, weighted by the squares of the eigenvalues of the corresponding PCs. This gives:
$\frac{1}{N}\sum_{k}(\lambda^{a}_{k})^{2}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}$ $=\Psi^{a}(\tau,0),$ (13)
showing that $\Psi^{a}(\bm{\tau})$ captures the temporal structure of the leading PCs of network activity (Fig. 2c,d). Note that weighting by the eigenvalues themselves, rather than their squares, gives $N^{-1}\sum_{k}\lambda^{a}_{k}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}=C^{a}(\tau)$, recovering single-neuron information (Fig. 2b,d).
II.3 Computing the two- and four-point functions in the i.i.d. case
We now review how to compute these two- and four-point functions in a classic network model with i.i.d. couplings $\bm{J}$. The first- and second-order coupling statistics are:
$\left\langle J_{ij}\right\rangle_{\bm{J}}=0,\hskip 15.00002pt\left\langle J_{ij}^{2}\right\rangle_{\bm{J}}=\frac{g^{2}}{N}.$ (14)
The magnitude of a typical coupling is $g/\sqrt{N}$. We refer to $g$ as the coupling strength. For the canonical dynamics of Eq. 2, the network is quiescent for $g<1$ and chaotic for $g>1$, with a sharp phase transition as $N\rightarrow\infty$. In this paper we assume that the network is non-quiescent.
We compute $C^{a}(t_{1},t_{2})$ and $\Psi^{a}(\bm{\tau})$ in the limit $N\rightarrow\infty$ under the disorder average. We use the subscript $\star$ to denote such limiting, disorder-averaged values:
$C^{a}_{\star}(t_{1},t_{2})$ $=\lim_{N\rightarrow\infty}\left\langle C^{a}(t_{1},t_{2})\right\rangle_{\bm{J}},$ (15)
$\Psi^{a}_{\star}(\bm{\tau})$ $=\lim_{N\rightarrow\infty}\left\langle\Psi^{a}(\bm{\tau})\right\rangle_{\bm{J}}.$ (16)
These values also correspond to the saddle point of a path integral. Both of these functions are self-averaging, meaning that the same values are obtained, up to $\mathcal{O}(1/\sqrt{N})$ fluctuations, when they are computed based on a single realization of a large network. In this paper, we are interested in statistically stationary, temporally fluctuating states such that $C^{\phi}_{\star}(t_{1},t_{2})=C^{\phi}_{\star}(t_{2}-t_{1})$.
The two-point function $C^{a}_{\star}(t_{1},t_{2})$ can be computed through a single-site picture that describes the dynamics of a typical neuron embedded in the rest of the network, with pre-activation $x(t)$ and activation $\phi(t)$. The single-site dynamics are given by
$T[x](t)=\eta^{x}(t),$ (17)
where $\eta^{x}(t)$ is a Gaussian field with mean zero and covariance $g^{2}C^{\phi}_{\star}(t_{1},t_{2})$. We denote this by
$\eta^{x}\sim\mathcal{GP}(0,g^{2}C^{\phi}_{\star}).$ (18)
$C^{\phi}_{\star}(t_{1},t_{2})$ is determined self-consistently by enforcing
$C^{\phi}_{\star}(t_{1},t_{2})=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (19)
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within this single-site process, i.e., with respect to the Gaussian distribution of $\eta^{x}(t)$. Once $C^{\phi}_{\star}(t_{1},t_{2})$ has been determined, $C^{x}_{\star}(t_{1},t_{2})$ follows easily. This single-site problem can be derived through either a single-site cavity calculation (see [39]) or a saddle-point condition in a path integral (see [40, 41]). For the i.i.d. couplings considered here, there is a simpler heuristic derivation: in the neuronal input $\sum_{j}J_{ij}\phi_{j}(t)$, the correlations between the couplings $J_{ij}$ and dynamic variables $\phi_{j}(t)$ can be safely neglected to leading order in $1/N$, yielding both Gaussianity of $\eta^{x}(t)$ by the central limit theorem and the second-order statistics of Eq. 18.
Clark et al. [25] first computed $\Psi_{\star}^{a}(\bm{\tau})$ using a dynamic, two-site version of the cavity method, based on the neuron-by-neuron definition (Eq. 7). This method finds $\psi^{a}_{\star}(\bm{\tau})$, the off-diagonal contribution to $\Psi^{a}_{\star}(\bm{\tau})$, noting that the on-diagonal contribution is simply $C^{a}_{\star}(\tau_{1})C^{a}_{\star}(\tau_{2})$. A cavity is first created by removing two neurons from the network and allowing the rest of the network, the reservoir, to generate dynamic activity. The cavity neurons are then introduced, and their effect on the reservoir is treated perturbatively. This yields a pair of coupled mean-field equations for the cavity units, generalizing the single-site picture discussed above to a two-site picture. Finally, self-consistency conditions are constructed by recognizing that the cavity pair is statistically equivalent to any reservoir pair. This calculation results in expressions for the four-point function in Fourier space, given by
$\Psi_{\star}^{\phi}(\bm{\omega})=\frac{C_{\star}^{\phi}(\omega_{1})C_{\star}^{\phi}(\omega_{2})}{|1-g^{2}S_{\star}^{\phi}(\omega_{1})S_{\star}^{\phi}(\omega_{2})|^{2}}$ (20)
for the activations, and
$\Psi_{\star}^{x}(\bm{\omega})$ $=C_{\star}^{x}(\omega_{1})C_{\star}^{x}(\omega_{2})+|U(\bm{\omega})|^{2}C_{\star}^{\phi}(\omega_{1})C_{\star}^{\phi}(\omega_{2})$
[…]
$U(\bm{\omega})$ $=\frac{g^{2}S_{\star}^{x}(\omega_{1})S_{\star}^{x}(\omega_{2})}{1-g^{2}S_{\star}^{\phi}(\omega_{1})S_{\star}^{\phi}(\omega_{2})},$ (21b)
for the pre-activations, where $\bm{\omega}=(\omega_{1},\omega_{2})$ and $C^{x\phi}_{\star}(\omega)$ is a cross-covariance between the pre-activation and activation. Note that if the joint distribution of pre-activations were Gaussian, $\Psi^{x}_{\star}(\bm{\tau})$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ would differ only by a proportionality constant due to Price’s theorem [42]. The more complex relationship observed here reflects the non-Gaussian joint statistics across different neurons, which is relevant because the network is nonlinear.
II.4 Definition of the random-mode model
Figure 3: Schematic of the random-mode model. Upper: Couplings $\bm{J}$ are generated as a sum of outer products, ${\bm{L}}_{a}{\bm{R}}^{T}_{a}$, with component strengths $D_{a}$. Lower: The two-point function $C^{\phi}_{\star}(\tau)$ and four-point function $\Psi^{\phi}_{\star}(\tau)$ are calculated in terms of the statistics of $D_{a}$. The two-point function depends only on the effective gain $g_{\text{eff}}$, while the four-point function depends on both $g_{\text{eff}}$ and $\text{PR}^{D}$, the effective dimension of the connectivity determined by the $D_{a}$ distribution.
Many studies have examined the dynamics of networks with i.i.d. couplings, as described in the previous section. However, the coupling matrices of real-world networks often have approximate low-rank spectral structure, deviating dramatically from the i.i.d. assumption. To analyze networks with such properties, we introduce the random-mode model. This model generates a coupling matrix $\bm{J}$ of the form
$\bm{J}=\sum_{a=1}^{M}D_{a}\bm{L}_{a}\bm{R}_{a}^{T},$ (22)
where $\bm{R}_{a}$ and $\bm{L}_{a}$ are “input” and “output” modes in the neuron space and $D_{a}>0$ is a weighting factor called the component strength. The effect of multiplying by $\bm{J}$ is to project a neuronal activity vector onto $\bm{R}_{a}$, scale it by $D_{a}$, project the result along $\bm{L}_{a}$, and sum these contributions for all $a=1,\ldots,M$ (Fig. 3, top). Eq. 22 can be written
$\bm{J}=\bm{L}\bm{D}\bm{R}^{T},$ (23)
where $\bm{L}$ and $\bm{R}$ are $N\times M$ matrices and $\bm{D}=\text{diag}(D_{1},D_{2},\ldots,D_{M})$.
A crucial modeling choice is how the number of modes, $M$, scales relative to the network size, $N$. We consider the limit where both $M$ and $N$ approach infinity while maintaining a fixed ratio,
$\alpha=\frac{M}{N}=\mathcal{O}(1).$ (24)
This scaling results in extensive-rank connectivity, meaning that the rank of $\bm{J}$ grows linearly with $N$. Importantly, choosing $\alpha$ to be small (or choosing the component strengths to decay rapidly; see below) allows for connectivity that is low-rank compared to the number of neurons.
For now, we assume that the matrices $\bm{L}$ and $\bm{R}$ have independent and identically distributed elements with zero mean and variance $1/N$. We assume these elements are Gaussian, although our results should not depend on the exact form of the single-element density for large $N$ due to the central limit theorem.
A key feature of the random-mode model is that we have full control over the component strengths $D_{a}$, which are analogous to the singular values of $\bm{J}$ (we make this link precise in Sec. II.4.1). We assume only that these component strengths follow a distribution that remains constant as $N$ grows, ensuring a well-defined large-$N$ limit. By varying this distribution, we can explore the impact of the spectral structure of the connectivity on collective features of network activity (Fig. 3, bottom).
We denote the $n$-th moment of the $D_{a}$ distribution by
$r_{n}=\left\langle D^{n}\right\rangle_{D}.$ (25)
It will be useful to define the participation ratio,
$\text{PR}^{D}=\frac{r_{2}^{2}}{r_{4}},$ (26)
from which we define the “effective rank” of the couplings,
effective $\text{effective rank}=\alpha\text{PR}^{D}.$ (27)
This quantity is related to the participation ratio of the singular value spectrum of $\bm{J}$ via Eq. 28 below. We will show that this effective rank is a key feature of $\bm{J}$ that controls the collective activity structure.
II.4.1 Relationship to singular value decomposition
Figure 4: Distribution of singular values in the random-mode model. Histograms show singular values of $\bm{L}\bm{D}\bm{R}^{T}$ with $D_{a}=1$ for various $\alpha$. For $\alpha\ll 1$, the singular values concentrate at one. Vertical lines: prediction of Eq. 29. $N=5000$.
The random-mode model has similarities to the SVD, which decomposes a rank-$M$ matrix $\bm{J}$ as $\bm{U}\bm{S}\bm{V}^{T}$. Here, $\bm{U}$ and $\bm{V}$ are $N\times M$ matrices whose columns are the left and right singular vectors (or modes), respectively, and $\bm{S}=\text{diag}(S_{1},S_{2},\ldots,S_{M})$ is a diagonal matrix containing the singular values, $S_{a}>0$. The left singular vectors are orthonormal among themselves, as are the right singular vectors (although the left and right singular vectors are generally not orthonormal to each other). In contrast, the columns of $\bm{L}$ and $\bm{R}$ in the random-mode model have i.i.d. statistics, resulting in random, $\mathcal{O}(1/\sqrt{N})$ overlaps among both within- and across-matrix columns. This i.i.d. assumption offers analytical tractability by eliminating the need to enforce orthonormality, a complicated global constraint (but see [43], who handle random orthonormal matrices in a regression setting). Furthermore, by not explicitly modeling overlaps between $\bm{L}$ and $\bm{R}$, the random-mode model parameterizes $\bm{J}$ using only the $\mathcal{O}(N)$ component strengths $D_{a}$, although such overlaps can be introduced (see Sec. VII).
Given a matrix $\bm{J}$ drawn from the random-mode model, we can examine the relationship between the singular values, $S_{a}$, and the component strengths, ${D}_{a}$. The non-orthonormality of $\bm{L}$ and $\bm{R}$ introduces some discrepancy between the two sets of values. We first compare the participation ratio of the squared singular value spectrum, $\text{PR}^{S}$, to the effective rank, $\alpha\text{PR}^{D}$. We find:
$\text{PR}^{S}=\frac{\alpha\text{PR}^{D}}{1+2\alpha\text{PR}^{D}},$ (28)
where $\text{PR}^{S}=\langle S^{2}\rangle_{S}^{2}/\langle S^{4}\rangle_{S}$, and $S$ are the singular values of $\bm{J}$. This follows from using Wick’s theorem to evaluate the expectations of $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ and $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ for the numerator and denominator, respectively, of $\text{PR}^{S}$. For $\alpha\text{PR}^{D}\ll 1$, $\text{PR}^{S}=\alpha\text{PR}^{D}+\mathcal{O}((\alpha\text{PR}^{D})^{2}),$ demonstrating that the component strengths are closely related to the singular values in the low-dimensional regime. In this regime, the random modes approximate an orthonormal set of vectors as $N\rightarrow\infty$ since the $\mathcal{O}(1/\sqrt{N})$ overlaps between modes can be neglected when there are few of them. While we express the analytic results in this paper in terms of $\alpha\text{PR}^{D}$, Eq. 28 can always be used to translate between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$.
We further quantify the discrepancy between singular values and component strengths in the case where $D_{a}=1$ for all $a$. Using methods from free probability theory [44, 45], we compute that the $M$ singular values are distributed over a range with boundaries
$S_{\pm}=\sqrt{1+\frac{5\alpha}{2}-\frac{\alpha^{2}}{20}\pm\left(1+\frac{\alpha}{8}\right)^{\frac{3}{2}}\sqrt{8\alpha}}.$ (29)
This is in agreement with large-$N$ simulations (Fig. 4). For small $\alpha$, $S_{\pm}=1\pm\sqrt{2\alpha}+\mathcal{O}(\alpha).$ While the random-mode model spreads the nonzero singular values over a range, this spreading becomes negligible for small $\alpha$. More generally, the distributions of $D_{a}$ and $S_{a}$ coincide for small $\alpha\text{PR}^{D}$.
II.4.2 Comparison to low-rank recurrent neural networks
A related parameterization of connectivity has been explored in low-rank recurrent neural networks, where the coupling matrix is expressed as a sum of outer products of input modes $\bm{n}_{a}$ and output modes $\bm{m}_{a}$:
$\bm{J}=\frac{1}{N}\sum_{a=1}^{R}\bm{m}_{a}\bm{n}_{a}^{T},$ (30)
with $R=\mathcal{O}(1)$ being the connectivity rank [30, 46, 31, 47]. In some instances, this low-rank component is added to an i.i.d. matrix modeling unstructured “background” connectivity [30, 46].
Several features separate this approach from the random-mode model. Low-rank recurrent neural networks maintain a constant rank $R$ as $N\rightarrow\infty$, yielding intensive-rank connectivity that models low-dimensional activity, whereas the random-mode model uses extensive-rank connectivity, modeling high-dimensional activity. Furthermore, low-rank recurrent neural networks allow for the most general possible parameterization of overlaps between modes, captured by a $R\times R$ overlap matrix $\bm{m}_{a}^{T}\bm{n}_{b}$. Specifying such overlaps in the random-mode model would require $\mathcal{O}(N^{2})$ parameters, comparable to defining the full coupling matrix. To reduce parameter complexity, the random-mode model specifies only the variances of mode components via the component strengths $D_{a}$, disregarding overlaps. This substantially reduces parameter count while preserving interesting network dynamics. In Sec. VII, we introduce $\mathcal{O}(N)$ additional parameters describing a restricted overlap structure among modes, specifying only diagonal elements of the overlap matrix. Further possible overlap forms are discussed in Sec. VIII.
The mean-field descriptions of network dynamics produced by these connectivities also differ markedly. In low-rank recurrent neural networks, low-dimensional network activity is described by a finite set of order parameters representing projections of network activity onto input modes $\bm{n}_{a}$, obeying effective nonlinear dynamics. Conversely, the extensive-rank connectivity in the random-mode model generates high-dimensional chaotic activity, described by covariance functions capturing temporal fluctuations. While the random-mode model’s mean-field analysis introduces latent variables given by projections of network activity onto input modes, their extensive number precludes their use as order parameters.
A final distinction lies in the scaling of the coupling matrix $\bm{J}$ and its implications for the alignment of activity with connectivity. In low-rank recurrent neural networks, elements of $\bm{J}$ scale as $\mathcal{O}(1/N)$ and the input-output mode overlaps are $\mathcal{O}(1)$, resulting in strong $\mathcal{O}(1)$ alignment of network activity with input modes. This is reminiscent of the rich regime of neural networks [48, 49, 50, 51]. Conversely, in the random-mode model, the elements of $\bm{J}$ scale as $\mathcal{O}(1/\sqrt{N})$ and the input-output mode overlaps are $\mathcal{O}(1/\sqrt{N})$, leading to weak $\mathcal{O}(1/\sqrt{N})$ alignment of network activity with any given input mode; the pre-activations are $\mathcal{O}(1)$ since each neuron combines input from $\mathcal{O}(N)$ modes. This is reminiscent of the lazy regime of neural networks [49, 52, 53].
III Path-integral calculation of the two- and four-point functions
The path-integral approach provides a systematic framework for analyzing neural-network dynamics, allowing for controlled approximations such as $1/N$ expansions [54, 55, 41, 56]. In this section, we demonstrate how to formulate the calculation of the four-point function $\Psi^{a}_{\star}(\bm{\tau})$ within this formalism, allowing us to solve complex models more efficiently than the two-site cavity method by following a straightforward procedure whose steps we outline at the end of Sec. III.3.
III.1 Relating $\Psi^{a}$ to fluctuations of the time-by-time covariance
Figure 5: Neuron-time duality and its relation to the dimension of activity. Both plots are based on the same simulation of a network of $N=2500$ neurons with $g=2.25$ and i.i.d. couplings. The dimension can be computed either by computing the statistics of the off-diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
To cast the problem in the language of path integrals, we turn to the time-by-time definition of $\Psi^{a}(\bm{\tau})$, given by Eq. 9. If the time-by-time covariance $C^{a}(t_{1},t_{2})$ were equal to its limiting value, $C^{a}_{\star}(t_{2}-t_{1})$, we would have $\Psi_{\star}^{a}(\bm{\tau})\overset{?}{=}N\left\langle C^{a}_{\star}(t_{2}-t_{1})C^{a}_{\star}(t_{2}-t_{1}+\tau_{2}-\tau_{1})\right\rangle_{t_{1},t_{2}}=0$ since, for large time differences, $C^{a}_{\star}(t_{2}-t_{1})$, decays exponentially in $|t_{2}-t_{1}|$ [10]. However, based on the neuron-by-neuron definition, Eq. 7, $\Psi_{\star}^{a}(\bm{\tau})$ is clearly a nonzero, order-one quantity. This apparent contradiction is resolved by considering fluctuations around the saddle point. The necessity of these fluctuations becomes clear when viewed through the lens of neuron-time duality. The presence of nonzero pairwise covariances $C^{a}_{ij}(0)$ indicates low-dimensional structure in the system. Equivalently, this structure manifests as the system revisiting similar states in phase space over time, resulting in nonzero temporal covariance $C^{a}(t_{1},t_{2})$ even for large time differences $|t_{2}-t_{1}|\gg 1$ (Fig. 5). To capture this behavior, we define fluctuations around the saddle point as
$\delta C^{a}(t_{1},t_{2})=C^{a}(t_{1},t_{2})-C^{a}_{\star}(t_{2}-t_{1}).$ (31)
We express $\Psi^{a}(\bm{\tau})$ purely in terms of these fluctuations as
$\Psi^{a}(\bm{\tau})=N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{1}+\tau_{1},t_{2}+\tau_{2})\right\rangle_{t_{1},t_{2}},$ (32)
noting that terms involving $C^{a}_{\star}(t_{2}-t_{1})$ vanish at large time differences. Our task, therefore, is to compute the covariance of the fluctuations, $\delta C^{a}(t_{1},t_{2})$, captured by
$N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{3},t_{4})\right\rangle_{\star}\\ =\lim_{N\rightarrow\infty}N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{3},t_{4})\right\rangle_{\bm{J}}.$ (33)
Then, to compute $\Psi_{\star}^{a}(\bm{\tau})$, we use the temporal limits
[…]
$t_{4}$ $=t_{2}+\tau_{2},$ $\tau_{1},\tau_{2}$ $\sim\mathcal{O}(1).$
III.2 Path-integral formalism and two-point functions
Given any fixed $\bm{J}$, we express the network dynamics using a path integral in the Martin–Siggia–Rose/Janssen–De Dominicis–Peliti formalism [57]:
$Z[\bm{J}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\left(i\sum_{i}\hat{x}_{i}T[{x}_{i}]-i\sum_{i,j}J_{ij}\hat{x}_{j}\phi_{j}\right),$ (35)
[…]
For simplicity, we begin with the i.i.d. model of $\bm{J}$ with mean zero and variance $g^{2}/N$. After performing the Gaussian integration over $\bm{J}$, we introduce an auxiliary field $C^{\phi}(t_{1},t_{2})$, defined by Eq. 5, and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$. This leads to a partition function for a statistical field theory involving $C^{\phi}(t_{1},t_{2})$ and $\hat{C}^{\phi}(t_{1},t_{2})$:
$Z=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi}]\right),$ (36)
where the intensive action is
$\mathcal{S}[C^{\phi},\hat{C}^{\phi}]=-\frac{1}{2}C^{\phi}\hat{C}^{\phi}-\log W[C^{\phi},\hat{C}^{\phi}],$ (37)
and the single-site path integral is
$W[C^{\phi},\hat{C}^{\phi}]=\int\mathcal{D}x\int\mathcal{D}\hat{x}\\ \times\exp\left(ixT[\hat{x}]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi\right).$ (38)
In the limit $N\rightarrow\infty$, the saddle point dominates this integral. The derivatives of the action, which are zero at the saddle point, are
$\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+g^{2}\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W},$ (39)
$\frac{\delta\mathcal{S}}{\delta\hat{C}^{\phi}(t_{1},t_{2})}$ $=-{C}^{\phi}(t_{1},t_{2})+\left\langle{\phi}(t_{1}){\phi}(t_{2})\right\rangle_{W},$ (40)
where $\left\langle\cdots\right\rangle_{W}$ denotes an average within the dynamic process described by $W[C^{\phi},\hat{C}^{\phi}]$. In evaluating derivatives of the action, we follow the rule that derivatives with respect to $C^{\phi}(t_{1},t_{2})$ also affect $C^{\phi}(t_{2},t_{1})$, and likewise for $\hat{C}^{\phi}(t_{1},t_{2})$, due to the symmetry in the action. Using the vanishing of correlation functions involving only the conjugate field $\hat{x}(t)$, the saddle-point conditions yield
$\hat{C}_{\star}^{\phi}(t_{1},t_{2})$ $=0,$ (41)
$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (42)
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C_{\star}^{\phi},0]$. This recovers the same single-site process (Eq. 17, Eq. 18) and self-consistency condition (Eq. 19) described in Sec. II.3.
III.3 Four-point functions
We now compute $\Psi^{a}_{\star}(\bm{\tau})$ within the path-integral formalism. Fluctuations around the saddle point derived above are governed by the Hessian of the action [40, 59]. This Hessian has blocks given by
$H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ $=\frac{\delta^{2}S}{\delta C^{\phi}(t_{1},t_{2})\delta C^{\phi}(t_{3},t_{4})}=-g^{4}\left\langle\left\langle\hat{x}(t_{1})\hat{x}(t_{2}),\hat{x}(t_{3})\hat{x}(t_{4})\right\rangle\right\rangle_{\star},$ (43a)
$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ $=\frac{\delta^{2}S}{\delta C^{\phi}(t_{1},t_{2})\delta\hat{C}^{\phi}(t_{3},t_{4})}=-I(t_{1},t_{2},t_{3},t_{4})-g^{2}\left\langle\left\langle\hat{x}(t_{1})\hat{x}(t_{2}),\phi(t_{3})\phi(t_{4})\right\rangle\right\rangle_{\star},$ (43b)
$H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ $=\frac{\delta^{2}S}{\delta\hat{C}^{\phi}(t_{1},t_{2})\delta\hat{C}^{\phi}(t_{3},t_{4})}=-\left\langle\left\langle\phi(t_{1})\phi(t_{2}),\phi(t_{3})\phi(t_{4})\right\rangle\right\rangle_{\star},$ (43c)
where $\left\langle\left\langle A,B\right\rangle\right\rangle_{\star}=\left\langle AB\right\rangle_{\star}-\left\langle A\right\rangle_{\star}\left\langle B\right\rangle_{\star}$, $I(t_{1},t_{2},t_{3},t_{4})=\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})$. Expanding the action to second order around the saddle point, the path integral becomes
$Z=\exp\left(-N\mathcal{S}[C^{\phi}_{\star},0]\right)\int\mathcal{D}\delta C^{\phi}\mathcal{D}\delta\hat{C}^{\phi}$ (44)
×$\times\exp\left(-\frac{N}{2}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}^{T}\begin{pmatrix}H_{C^{\phi}C^{\phi}}&H_{C^{\phi}\hat{C}^{\phi}}\\ H_{\hat{C}^{\phi}{C}^{\phi}}^{T}&H_{\hat{C}^{\phi}\hat{C}^{\phi}}\end{pmatrix}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}\right),$
revealing the structure of the Gaussian fluctuations around the saddle point. The covariance matrix among the fluctuation variables, $\delta C^{\phi}(t_{1},t_{2})$ and $\delta\hat{C}^{\phi}(t_{1},t_{2})$, is $1/N$ times the inverse Hessian, yielding
$N\left\langle\delta C^{\phi}(t_{1},t_{2})\delta C^{\phi}(t_{3},t_{4})\right\rangle_{\star}$ (45)
$=\left[H_{C^{\phi}C^{\phi}}-H_{C^{\phi}\hat{C}^{\phi}}H_{\hat{C}^{\phi}\hat{C}^{\phi}}^{-1}H_{\hat{C}^{\phi}C^{\phi}}\right]^{-1}(t_{1},t_{2},t_{3},t_{4}).$
In principle, we have all the necessary components to evaluate this expression: the single-site dynamics are known via the saddle-point condition, and the Hessian blocks are given in terms of connected correlation functions in this single-site process. We could then take the temporal separation limits of Eq. 34 to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$. The challenge is that each Hessian block is a complicated time2-by-time2 matrix that, for general values of the time variables, has no analytic inverse. To circumvent this problem, we show in Appendix C that the temporal limits of Eq. 34 can be taken before taking the inverses. Under these temporal limits, the Hessian blocks are
$H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-C^{\phi}_{\star}(\tau_{1})C^{\phi}_{\star}(\tau_{2}),$ (46)
$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ (47)
$\hskip 40.00006pt=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(-\tau_{1})S^{\phi}_{\star}(-\tau_{2}),$
$H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0,$ (48)
where we used the fact that multiplication by $-i\hat{x}(t)$ is equivalent to a functional derivative with respect to a source at time $t$. Since the relevant quantities are now translation-invariant, i.e., depend only on $\tau_{1}$ and $\tau_{2}$, we can transform to Fourier space. In this representation, each Hessian block becomes a frequency-dependent scalar,
$H_{\hat{C}^{\phi}\hat{C}^{\phi}}(\bm{\omega})$ $=-C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2}),$ (49)
$H_{C^{\phi}\hat{C}^{\phi}}(\bm{\omega})$ $=-1+g^{2}\left(S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})\right)^{*},$ (50)
$H_{C^{\phi}C^{\phi}}(\bm{\omega})$ $=0.$ (51)
In summary, the full Hessian can be replaced by a two-by-two, frequency-dependent Hesssian given by
$\bm{H}=\begin{pmatrix}0&-1+g^{2}(S^{\phi}_{12})^{*}\\ -1+g^{2}S^{\phi}_{12}&-C^{\phi}_{12}\end{pmatrix},$ (52)
where we simplified notation by suppressing frequency arguments via the shorthand $C^{\phi}_{12}=C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2})$ and $S^{\phi}_{12}=S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})$. The Fourier-space function $\Psi_{\star}^{\phi}(\bm{\omega})$ is the upper-left element of the inverse of this matrix. Performing the two-by-two matrix inversion gives
$\Psi_{\star}^{\phi}=\frac{C^{\phi}_{12}}{|1-g^{2}S^{\phi}_{12}|^{2}},$ (53)
which agrees with the two-site cavity result (Eq. 20).
To summarize, we compute $\Psi_{\star}^{\phi}(\bm{\tau})$ through the following steps (in the next section, we show how to compute $\Psi^{x}_{\star}(\bm{\tau})$):
  1. 1.
    Write down the path integral, integrate out disorder, and introduce two-point fields and conjugate partners to factorize the action along extensive indices.
  2. 2.
    Sovle the saddle-point equations of the resulting action, yielding the covariance and response functions.
  3. 3.
    Derive the frequency-dependent Hessian by computing second derivatives of the action with respect to the two-point fields, applying the temporal limits of Eq. 34, and transforming to Fourier space. Then, determine the element of the inverse frequency-dependent Hessian corresponding to $\Psi^{\phi}_{\star}(\bm{\omega})$.
This approach allows for the efficient calculation of two- and four-point functions for a given model, requiring only a few lines of computation directly from the action of the theory. To demonstrate the flexibility of the path-integral approach in computing $\Psi^{a}_{\star}(\bm{\tau})$, we apply it to compute $\Psi^{\phi}_{\star}(\bm{\tau})$ for a generalization of the i.i.d. model in which there are correlations between reciprocal couplings, recovering the result in [25] from the two-site cavity method (Appendix E).
III.4 Adding sources
Above, computing $\Psi^{\phi}_{\star}(\bm{\tau})$ was simplified by $C^{\phi}(t_{1},t_{2})$ appearing naturally in the statistical field theory resulting from integrating out $\bm{J}$. To compute $\Psi^{x}_{\star}(\bm{\tau})$, we need to introduce a source-field term in the action.
Consider a general intensive action $\mathcal{S}[\bm{C};\mathcal{J}]$ where $\bm{C}$ is a collection of fields (e.g., $\bm{C}=\{C^{\phi},\hat{C}^{\phi}\}$), and $\mathcal{J}$ is a source. In Appendix D, we show that
$\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0}\\ =\left[-\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}+\left(\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right)^{T}\bm{H}^{-1}\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{\mathcal{J}=0},$ (54)
[…]
III.5 Path-integral analysis of the random-mode model
Having demonstrated the path-integral approach to computing two- and four-point functions in well known connectivity models, we now apply this formalism to the random-mode model. In terms of the mode matrices $\bm{L}$ and $\bm{R}$, the path integral is
$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\\ \times\exp\left(i\sum_{i}\hat{x}_{i}T[x_{i}]-i\sum_{i,j}\left[\bm{L}\bm{D}\bm{R}\right]_{ij}\hat{x}_{i}\phi_{j}\right).$ (55)
We would like to integrate out $\bm{L}$ and $\bm{R}$, but this is complicated by them appearing together in a quadratic term in the action. To simplify the integration, we introduce a set of $M$ latent variables,
$z_{a}(t)=D_{a}\sum_{i}R_{ia}\phi_{i}(t),$ (56)
and their conjugates $\hat{z}_{a}(t)$. We use indices $i,j$ for neurons and $a,b$ for latent variables. The introduction of these latent variables makes the action linear in $\bm{L}$ and $\bm{R}$:
$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\int\mathcal{D}\bm{z}\int\mathcal{D}\hat{\bm{z}}\\ \times\exp\Big{(}i\sum_{i}\hat{x}_{i}T[x_{i}]+i\sum_{a}z_{a}\hat{z}_{a}\\ -i\sum_{i,a}L_{ia}\hat{x}_{i}z_{a}-i\sum_{i,a}D_{a}R_{ia}\phi_{i}\hat{z}_{a}\Big{)}.$ (57)
We can now easily integrate out $\bm{L}$ and $\bm{R}$. To factorize the action over indices spanning extensive dimensions, we introduce the field $C^{\phi}(t_{1},t_{2})$ (Eq. 5) with conjugate $\hat{Q}(t_{1},t_{2})$ as well as
$Q(t_{1},t_{2})=\frac{1}{M}\sum_{a}z_{a}(t_{1})z_{a}(t_{2})$ (58)
with conjugate $\hat{C}^{\phi}(t_{1},t_{2})$. The disorder-averaged path integral is
$Z$ $=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}Q\int\mathcal{D}\hat{Q}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},Q,\hat{Q}]\right),$ (59)
$\mathcal{S}[C^{\phi},\hat{C}^{\phi},Q,\hat{Q}]$ $=-\frac{1}{2}C^{\phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}-\log W^{x}[Q,\hat{Q}]-\alpha\left\langle\log W^{z}_{D}[C^{\phi},\hat{C}^{\phi}]\right\rangle_{D},$ (60)
$W^{x}[Q,\hat{Q}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\phi\hat{Q}\phi-\frac{1}{2}\hat{x}Q\hat{x}\right),$ (61)
$W^{z}_{D}[C^{\phi},\hat{C}^{\phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}z\hat{C}^{\phi}z-\frac{1}{2}D^{2}\hat{z}C^{\phi}\hat{z}\right).$ (62)
Thus, analysis of the path integral is made tractable by recasting the problem as interacting neurons and latent variables. The neuronal single-site picture is described by $W^{x}[Q,\hat{Q}]$, and the latent-variable single-site picture is described by $W^{z}_{D}[C^{\phi},\hat{C}^{\phi}]$, where the subscript $D$ reflects that each latent variable has its own component strength. This formulation introduces an extensive set of latent variables $z_{a}(t)$, in contrast to the finite set of latent variables, conventionally denoted by $\kappa_{a}(t)$, in low-rank neural networks [30, 31]. The statistics of the Gaussian input to each single-site process (neurons or latent variables) are determined by the statistics of the other process, creating a bipartite, mutually referential structure that also arises in the cavity treatment of the problem.
III.5.1 Two-point functions
Following the steps outlined at the end of Sec. III.3, we first compute the DMFT by calculating the saddle point. The derivatives of the action are
$\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{Q}^{\phi}(t_{1},t_{2})+\alpha\left\langle D^{2}\left\langle\hat{z}(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D}}\right\rangle_{D},$ (63)
$\frac{\delta\mathcal{S}}{\delta\hat{C}^{\phi}(t_{1},t_{2})}$ $=-Q(t_{1},t_{2})+\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{W^{z}_{D}}\right\rangle_{D},$ (64)
$\frac{\delta\mathcal{S}}{\delta Q(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W^{x}},$ (65)
$\frac{\delta\mathcal{S}}{\delta\hat{Q}(t_{1},t_{2})}$ $=-C^{\phi}(t_{1},t_{2})+\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{W^{x}},$ (66)
where $\left\langle\cdots\right\rangle_{W^{x}}$ and $\left\langle\cdots\right\rangle_{W^{z}_{D}}$ are averages within the dynamic processes described by $W^{x}[Q,\hat{Q}]$ and $W^{z}_{D}[C^{\phi},\hat{C}^{\phi}]$, respectively. Setting these derivatives to zero yields
$\hat{C}^{\phi}_{\star}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (67)
$C^{\phi}_{\star}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (68)
$Q_{\star}(t_{1},t_{2})$ $=\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}\right\rangle_{D},$ (69)
where $\left\langle\cdots\right\rangle_{\star}$ and $\left\langle\cdots\right\rangle_{\star D}$ are averages within the dynamic processes described by $W^{x}[Q_{\star},0]$ and $W^{z}_{D}[C^{\phi}_{\star},0]$, respectively. These single-site processes are
$T[x](t)$ $=\eta^{x}(t),$ (70)
$z(t)$ $=\eta^{z}(t),$ (71)
where $\eta^{x}(t)$ and $\eta^{z}(t)$ are Gaussian fields,
$\eta^{x}$ $\sim\mathcal{GP}(0,Q_{\star}),$ (72)
$\eta^{z}$ $\sim\mathcal{GP}(0,D^{2}C^{\phi}_{\star}),$ (73)
showing that the neuronal and latent-variable single-site processes have mutually referential statistics  [^1] Note that the variance of $\eta^{z}(t)$ is scaled by $D^{2}$; equivalently, we could factor out $D^{2}$ from the variance and include $D$ as a multiplicative factor in the latent-variable single-site dynamics (Eq. 71).
Solving for $Q_{\star}(t_{1},t_{2})$ by combining Eq. 69, Eq. 71, and Eq. 73, we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq. 25. Consolidating these results, the self-consistent problem that determines $C^{\phi}_{\star}(t_{1},t_{2})$ can be expressed purely in terms of this correlation function as
$T[x](t)$ $=\eta^{x}(t),$
$\eta^{x}$ $\sim\mathcal{GP}(0,\alpha r_{2}C^{\phi}_{\star}),$
$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}.$
Comparing to Eqs. 1719, we recover the self-consistent equations of the i.i.d. model with coupling strength
$g_{\text{eff}}=\sqrt{\alpha r_{2}}.$ (74)
Thus, the low-rank coupling structure remains undetectable at the single-neuron level, where neurons experience effectively i.i.d. connections with variance $g_{\text{eff}}^{2}/N$. The structure of the random-mode model emerges only in collective neural activity, as we now demonstrate by calculating the four-point function.
III.5.2 Four-point functions
The frequency-dependent saddle-point Hessian is
$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}C_{12}^{\phi}&-1&0\\ \hline\cr 0&-1&0&(S_{12}^{\phi})^{*}\\ -1&0&S_{12}^{\phi}&-C_{12}^{\phi}\end{array}\right).$ (75)
The upper-left block, corresponding to the latent variables, is related to the Hessian of an i.i.d. network with linear dynamics and a heterogeneous distribution of single-neuron scale factors $D_{a}$; the lower-right block, corresponding to the neurons, is related to the Hessian of an i.i.d. network of nonlinear neurons (Eq. 52). They are coupled through the simple off-diagonal blocks, reflecting the interaction between neurons and latent variables. Finally, we obtain for the four-point function:
$\Psi_{\star}^{\phi}=\frac{1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}{|1-g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}C_{12}^{\phi}.$ (76)
To obtain $\Psi^{x}_{\star}$, we add $-\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral (Eq. 61) and compute
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\>\>\>\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}0\\ 0\\ S_{12}^{x}\\ -C_{12}^{x\phi}\\ \end{pmatrix}.$ (77)
Inserting these into Eq. 54 and using a formula (Eq. 149) to invert the Hessian, we obtain
$\Psi_{\star}^{x}=C^{x}_{12}+\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)|U|^{2}C^{\phi}_{12}\\ +UC^{x\phi}_{12}+\text{h.c.},$ (78)
where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ (Eq. 21b with $g=g_{\text{eff}}$).
III.6 Interpreting the formulas
Figure 6: Dimension of activity and collective timescales in the random-mode model. (a) Activity dimension $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ for various coupling strengths $g_{\text{eff}}$. Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions. Inset: Extended $g_{\text{eff}}=6$ case, showing convergence to i.i.d. coupling behavior with growing $\alpha$. (b) Normalized four-point function $\overline{\Psi^{\phi}(\tau,0)}\equiv\Psi^{\phi}(\tau,0)/\Psi^{\phi}(0,0)$ for various $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$. Inset: Theory curves for all $\alpha\text{PR}^{D}$, demonstrating relative invariance of collective timescales to the effective rank. Simulations use $N=5000$ neurons.
The analytic formulas for the four-point function of the random-mode model (Eq. 76 and Eq. 78) reveal several insights. Taking the limit $\alpha\text{PR}^{D}\rightarrow\infty$ for fixed $g_{\text{eff}}$, which leads to elements of $\bm{J}$ being i.i.d., recovers the known formulas for i.i.d. couplings (Eq. 20 and Eq. 21). Comparing to the i.i.d. formulas, any structure beyond i.i.d. connectivity (i.e., $1/\alpha\text{PR}^{D}$ being finite) strictly decreases the dimension of activity.
Our analysis reveals two key statistical features of the connectivity that characterize how connectivity shapes activity: the effective coupling strength, $g_{\text{eff}}$, and the effective rank, $\alpha\text{PR}^{D}$. The first characterizes the typical magnitude of couplings, while the second characterizes global information about the number of significant connectivity components. As in the i.i.d. model, $g_{\text{eff}}$ influences both single-neuron activity (described by $C^{a}_{\star}(\tau)$) and collective activity (described by $\Psi^{a}_{\star}(\tau)$). In contrast, $\alpha\text{PR}^{D}$ influences collective activity exclusively. The random-mode model therefore provides independent control over single-neuron and collective aspects of activity.
To validate the theoretical prediction for $\Psi^{\phi}_{\star}(\bm{\tau})$ (Eq. 76), we simulated networks and computed the participation ratio from the empirical equal-time covariance matrix of the activations. For the component strengths $D_{a}$, we used a spectrum given by
$D_{a}\propto\exp(-\beta_{D}a/M)$ (79)
for $a=1,\ldots,M$, which yields $\text{PR}^{D}=\beta^{-1}_{D}\tanh(\beta_{D})$. The corresponding density is $P(D)=1/\beta_{D}D$ with support on $[\exp(-\beta_{D}),1]$. To attain a desired effective rank $\alpha\text{PR}^{D}<1$, we set $\alpha=M/N=1$ and solved for $\beta$. For $\alpha\text{PR}^{D}>1$, we fixed $\text{PR}^{D}=1$ and increased $\alpha$. We adjusted $g_{\text{eff}}$ by rescaling $D_{a}$ (i.e., adjusting the proportionality constant in Eq. 79).
Our simulations show excellent agreement with theoretical predictions across parameter settings (Fig. 6). Fig. 6a illustrates how the dimension of activity, $\text{PR}^{\phi}$, varies with $\alpha\text{PR}^{D}$ for different values of $g_{\text{eff}}$. Consistent with i.i.d.-coupling networks, $\text{PR}^{\phi}$ increases monotonically with $g_{\text{eff}}$ [25]. It also increases monotonically with $\alpha\text{PR}^{D}$, approaching the dimension of a network with i.i.d. couplings for large $\alpha\text{PR}^{D}$ (Fig. 6a inset). As $\alpha\text{PR}^{D}$ approaches zero, the dimension of activity also approaches zero. We examine the limits $g_{\text{eff}}\rightarrow 1^{+}$ and $\alpha\text{PR}^{D}\rightarrow 0$ in detail in Sec. IV.
Beyond the dimension of activity, which depends on $\Psi^{\phi}_{\star}(\bm{\tau})$ at $\bm{\tau}=(0,0)$, we also examined the temporal profile of this function. Fig. 6b shows a normalized version of $\Psi^{\phi}(\tau,0)$, which is related to the correlation functions of leading principal components, for various values of $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$. The decay timescale of $\Psi^{\phi}_{\star}(\tau,0)$ decreases with increasing $g_{\text{eff}}$, approaching a limiting behavior at $g_{\text{eff}}\sim 10$. While $\alpha\text{PR}^{D}$ affects the overall scale of this function, and consequently the dimension of activity, it has little effect on the decay timescale, and thus little effect on the timescales of leading principal components (Fig. 6b, inset). Thus, the effective rank primarily influences the dimension of activity rather than the collective temporal structure.
IV Limiting behavior of the random-mode model
Figure 7: Analysis of dimension of activity and covariance spectrum in the low-effective-rank limit. (a) Top: Dimension of activity $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ ($\alpha=1$) for various coupling strengths $g_{\text{eff}}$. Component strengths $D_{a}$ are constant for $a/N\leq\text{PR}^{D}$ and 00 otherwise. Small dots: individual simulations; large dots: means over 10 simulations; solid lines: theoretical predictions; dotted lines: small-$\text{PR}^{D}$ expansion, $K^{\phi}(g_{\text{eff}})\mathrm{PR}^{D}$. Inset: Ratio of proportionality factor $K^{\phi}(g_{\text{eff}})$ to $\text{PR}^{\phi}_{\text{i.i.d.}}$ versus $g_{\text{eff}}$. Bottom: Same analysis for pre-activations. (b) Top: Rank-ordered eigenvalue plots for activation covariance matrix $C^{\phi}_{ij}(0)$. Dotted vertical lines indicate $N\,\text{PR}^{D}$. Bottom: Same for pre-activations. All simulations use $N=6500$ neurons.
Here we examine the behavior of the random-mode model in two limiting cases: as the system approaches the transition to chaos from above ($g_{\text{eff}}\rightarrow 1^{+}$), and in the limit of low effective rank ($\alpha\text{PR}^{D}\ll 1$).
We first consider approaching the transition to chaos from above, i.e., $g_{\text{eff}}\rightarrow 1^{+}$. The relevant expansion parameter is $g_{\text{eff}}-1\ll 1$. For networks with i.i.d. couplings, Clark et al. [25] showed that, for coupling strength $g$, $\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})\sim 1/(g-1)$ and $\text{PR}^{a}_{\text{i.i.d.}}\sim(g-1)^{3}$ in this limit. For the random-mode model, to leading order in $1/(g_{\text{eff}}-1)$, we have
$\Psi^{a}_{\star}(\bm{\tau})=\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)\underbrace{\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})}_{\sim 1/(g_{\text{eff}}-1)}.$ (80)
Consequently, to leading order in $g_{\text{eff}}-1$, the dimension of activity behaves as
$\text{PR}^{a}=\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)^{-1}\underbrace{\text{PR}^{a}_{\text{i.i.d.}}}_{\sim(g_{\text{eff}}-1)^{3}}.$ (81)
Thus, as the network approaches the transition to chaos, the effect of structure in connectivity is to reduce the dimension by a factor of $(1+1/\alpha\text{PR}^{D})^{-1}$ compared to a network with i.i.d. couplings. Taking the further limit of $\alpha\text{PR}^{D}\ll 1$ yields the simple relation
$\text{PR}^{a}=\alpha\text{PR}^{D}\text{PR}^{a}_{\text{i.i.d.}}$ (82)
The difference between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$ is a higher-order correction, so we also have
$\text{PR}^{a}=\text{PR}^{S}\text{PR}^{a}_{\text{i.i.d.}}$ (83)
That is, in the limits of both low effective rank and small effective coupling strength, the dimension of activity is equal to the dimension of the singular-value spectrum times the dimension of activity in an i.i.d. network.
We now consider the limit of low effective rank, $\alpha\text{PR}^{D}\ll 1$, for arbitrary $g_{\text{eff}}$. In this limit, the component strengths $D_{a}$ closely approximate the singular values of the coupling matrix. To leading order in $\alpha\text{PR}^{D}$,
$\text{PR}^{a}=K^{a}(g_{\text{eff}})\alpha\text{PR}^{D},$ (84)
where
[…]
We match this theory of limiting behavior to simulated networks with $\alpha=1$ and a step function for the component strengths:
$D_{a}=\begin{cases}\text{const.}&\frac{a}{N}\leq\text{PR}^{D}\\ 0&\frac{a}{N}>\text{PR}^{D}.\end{cases}$ (87)
We varied $\text{PR}^{D}$ and verified its linear relationships with $\text{PR}^{\phi}$ and $\text{PR}^{x}$ in the $\text{PR}^{D}\rightarrow 0$ limit (Fig. 7a).
Focusing on $K^{\phi}(g_{\text{eff}})$, note that its expression (Eq. 85) resembles the formula for the dimension of activity in a i.i.d. network with coupling strength $g_{\text{eff}}$, but includes an additional factor $|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}$ in the integral (compare to Eq. 20). Numerical evaluation of $K^{\phi}(g_{\text{eff}})$ shows that it approaches $\text{PR}^{\phi}_{\text{i.i.d.}}$ as $g_{\text{eff}}\rightarrow 1^{+}$ as expected from Eq. 82. As $g_{\text{eff}}$ increases, $K^{\phi}(g_{\text{eff}})$ increases monotonically to approximately $1.53\,\text{PR}^{\phi}_{\text{i.i.d.}}$ (Fig. 7a, inset). These results show that for networks with low effective rank, $\text{PR}^{\phi}\sim\alpha\text{PR}^{D}\text{PR}^{\phi}_{\text{i.i.d.}}$ up to an order-one fudge factor that becomes exactly one as $g_{\text{eff}}\rightarrow 1^{+}$.
We also investigated the relationship between the eigenvalues $\lambda^{a}_{i}$ of $C^{a}_{ij}(0)$ and the component strengths $D_{a}$. In the low-effective-rank limit, the quantities $\text{PR}^{a}$ and $\alpha\text{PR}^{D}$ computed from these spectra exhibit a linear relationship (with proportionality factor $K^{a}(g_{\text{eff}})$), implying that the ratio of their second-squared and fourth moments are tightly linked. However, we find that the full forms of the eigenvalue and component-strength spectra differ markedly. Whereas the component strengths $D_{a}$ decay according to a step function in this example, the rank-ordered eigenvalues $\lambda^{\phi}$ and $\lambda^{x}$ decay smoothly (Fig. 7b). Thus, the dynamics transform the connectivity spectrum into an activity spectrum with a different form while preserving the linear relationship between $\text{PR}^{a}$ and $\alpha\text{PR}^{D}$.
V Cavity analysis of the random-mode model
Figure 8: Schematic of the two-site cavity approach for the bipartite representation of the random-mode model. Left: Neuron cavity, considering the introduction of neurons $x_{0}(t)$ and $x_{0^{\prime}}(t)$. Right: Latent-variable cavity, considering the introduction of latent variables $z_{\hat{0}}(t)$ and $z_{\hat{0}^{\prime}}(t)$. The cavity pictures are coupled, with averages from one appearing in self-consistent equations for the other.
While the path-integral approach provides a systematic framework for analyzing neural-network dynamics, the cavity method offers an intuitive, complementary perspective. This section provides an overview of the cavity analysis for the random-mode model, with the full mathematical details available in Appendix G. While these approaches yield equivalent results, they differ in their methodology: the path-integral approach focuses on fluctuations around the saddle point in a field theory, whereas the cavity approach examines perturbations induced by a subset of held-out variables.
Leveraging Eq. 56, we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al. [25] to both groups. This leads to separate, mutually referential cavity pictures for each group, mirroring the two single-site path integrals (Sec. III.5). As in the path-integral formalism, self-consistent expressions in one picture are defined using averages from the other. Unlike the path-integral approach, which derives $\Psi^{\phi}(\bm{\tau})$ using its time-by-time definition (Eq. 9), the cavity method derives this function using its neuron-by-neuron definition (Eq. 7).
The cavity method we employ, schematized in Fig. 8, has two key features that distinguish it from simpler cavity calculations. First, it uses a two-site structure: we consider the simultaneous removal of two neurons, or two latent variables, from the network, rather than just one. This allows us to study pairwise correlations between neurons. This approach was previously used in Clark et al. [25]. Second, it incorporates a bipartite structure: we treat both neurons and latent variables as dynamic objects, creating separate but mutually referential cavity pictures for each group. Such bipartite structure, without the two-site structure, has been used in cavity calculations in the Hopfield model [61, 62]. Note that these two features are distinct: the two-site aspect refers to the number of neurons removed in each cavity, while the bipartite aspect refers to the types of variables considered (neurons and latent variables).
VI Single-neuron heterogeneity
Figure 9: Effect of single-neuron heterogeneity on dimension of activity in the random-mode model. (a) Dimension of normalized activity $\text{PR}^{\phi}$ versus participation ratio of gain distribution $\text{PR}^{G}$ for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$. Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions. (b) Fractional reduction in dimension for weighted activations relative to $\text{PR}^{\phi}$. Blue: $G_{i}\phi_{i}$; purple: $G^{\text{readout}}_{i}\phi_{i}$. $g_{\text{eff}}=10$. Thick dots: means over 10 simulations, averaging before taking the ratio; lines: theoretical predictions. All simulations use $N=5000$ neurons.
While the random-mode model introduces heterogeneity in connectivity through the component strengths $D_{a}$, individual neurons may also exhibit heterogeneous properties. For instance, cortical neurons display a broad distribution of mean firing rates [63]. To compare the effects of low-dimensional structure in connectivity with the effects of heterogeneous neuronal properties, we extend our analysis to networks with diverse single-neuron characteristics.
In particular, we generalize the neuronal nonlinearity to $\Phi_{\bm{\theta}}(x)$, which depends on a vector of parameters $\bm{\theta}$. The parameter vector for the $i$-th neuron, $\bm{\theta}_{i}$, is sampled i.i.d. from a distribution $P(\bm{\theta})$. This formulation allows us to model diverse neuronal properties such as gains or biases. When analyzing neural data, it is common to normalize neuronal activities (e.g., through $z$-scoring) to put them on the same scale. To facilitate analysis of such normalized activations, we also introduce a set of variables $\phi_{i}(t)=\phi(x_{i}(t))$ that do not depend on $\bm{\theta}_{i}$.
We incorporate the generalized nonlinearity into the path integral, as well as a source term for the normalized variables, and integrate out $\bm{L}$ and $\bm{R}$. To factorize the action over indices spanning extensive dimensions, we introduce, among others, the order parameter
$C^{\Phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i}\Phi_{\bm{\theta}_{i}}(t_{1})\Phi_{\bm{\theta}_{i}}(t_{2}),$ (88)
where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}(x_{i}(t))$. In Appendix H, we derive a single-site dynamic process $T[x](t)=\eta^{x}(t)$, where $\eta^{x}\sim\mathcal{GP}(0,\alpha r_{2}C^{\Phi}_{\star})$, with the self-consistency condition
$C^{\Phi}_{\star}(t_{1},t_{2})=\left\langle\left\langle\Phi_{\bm{\theta}}(t_{1})\Phi_{\bm{\theta}}(t_{2})\right\rangle_{\star}\right\rangle_{\bm{\theta}}.$ (89)
This condition is identical to that of an i.i.d. network, but with an average over $\bm{\theta}\sim P(\bm{\theta})$, in addition to the usual Gaussian average over $\eta^{x}(t)$, to account for single-neuron heterogeneity.
To write down the solution the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$:
$\mathfrak{C}^{\Phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\Phi_{\bm{\theta}}\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\Phi_{\bm{\theta}}\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (90a)
$\mathfrak{s}^{\Phi}_{12}$ $=\left\langle\left\langle\frac{\delta\Phi_{\bm{\theta}}}{\delta I}\right\rangle_{\star}(\omega_{1})\left\langle\frac{\delta\Phi_{\bm{\theta}}}{\delta I}\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (90b)
$\mathfrak{C}^{\Phi\phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (90c)
along with the usual shorthand $C^{\Phi}_{12}=C^{\Phi}_{\star}(\omega_{1})C^{\Phi}_{\star}(\omega_{2})$. For the unnormalized variables $\Phi_{\bm{\theta}_{i}}(t)$, we obtain
$\Psi^{\Phi}_{\star}=\frac{\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\Phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}|\alpha r_{2}\mathfrak{s}^{\Phi}_{12}|^{2}}{|1-\alpha r_{2}\mathfrak{s}^{\Phi}_{12}|^{2}}C^{\Phi}_{12}.$ (91)
For the normalized variables $\phi_{i}(t)$, we obtain
$\Psi^{\phi}_{\star}=\Bigg{[}|A|^{2}\left(\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}\frac{C^{\Phi}_{12}}{C^{\phi}_{12}}\right)\\ +\left(A\frac{\mathfrak{C}^{\Phi\phi}_{12}}{C^{\phi}_{12}}+\text{h.c.}\right)+1\Bigg{]}C^{\phi}_{12},$ (92)
where $A={\alpha r_{2}S^{\phi}_{12}}/{(1-\alpha r_{2}\mathfrak{s}^{\phi}_{12})}$.
VI.1 Diverse single-neuron firing rates
An illustrative case of single-neuron heterogeneity occurs when the $\bm{\theta}$ parameters consist of a gain parameter $G_{i}$ for each neuron, modeling diversity in mean firing rates:
$\Phi_{\bm{\theta}_{i}}(x)=G_{i}\phi(x).$ (93)
In this scenario, the dynamics of the network are determined by an effective coupling matrix $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$, where $\bm{G}=\text{diag}(G_{1},\ldots,G_{N})$. This extension of the random-mode model allows us to directly compare the effects of low-dimensional structure in the connectivity (via the component strengths $D_{a}$) with heterogeneity in single-neuron properties (via the gains $G_{i}$). We assume the gains are drawn from a distribution with moments
$q_{n}=\left\langle G^{n}\right\rangle_{G}.$ (94)
To quantify the heterogeneity in firing rates, we define the participation ratio of the gain distribution:
$\text{PR}^{G}=\frac{q_{2}^{2}}{q_{4}}.$ (95)
We analyze collective activity for both the unnormalized activations $\Phi_{\bm{\theta}_{i}}(t)$ and the normalized activations $\phi_{i}(t)$. The latter are analogous to $z$-scored firing rates in neural recordings, where $z$-scoring removes single-neuron heterogeneity while preserving nontrivial structure due to cross-neuron and temporal correlations.
For the unnormalized variables, we calculate the four-point function
$\Psi^{\Phi}_{\star}=\frac{\frac{1}{\text{PR}^{G}}+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\Phi}_{12},$ (96)
and for the normalized variables,
$\Psi^{\phi}_{\star}=\frac{1+\left(\frac{1}{\text{PR}^{G}}+\frac{1}{\alpha\text{PR}^{D}}-1\right)|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12},$ (97)
where
$g_{\text{eff}}$ $=\sqrt{\alpha r_{2}q_{2}},$ (98)
$C^{\Phi}_{12}$ $=q_{2}^{2}C^{\phi}_{12}.$ (99)
Eq. 97 demonstrates that reductions in effective rank, described by $\alpha\text{PR}^{D}$, and in the heterogeneity of gains, described by $\text{PR}^{G}$, have symmetric effects on the dimension of normalized activity, with their combined effect captured by the harmonic mean, $\left(\frac{1}{\alpha\text{PR}^{D}}+\frac{1}{\text{PR}^{G}}\right)^{-1}$. This symmetry arises because swapping the distributions of $D_{a}$ and $G_{i}$ does not change the spectrum of $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ when $\bm{L}$ and $\bm{R}$ are i.i.d., since $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ and $\bm{R}^{T}\bm{G}\bm{L}\bm{D}$ have the same nonzero eigenvalues. One might expect that normalizing the activations before computing the dimension of activity would remove the effect of heterogeneous gains on the dimension. This is not the case, however, since neuronal gains affect the recurrent dynamics, and not merely the readout.
To further explore this and related effects, we compare the activity dimensions for three sets of activations: (1) unnormalized activations, $\Phi_{\bm{\theta}_{i}}(t)$; (2) normalized activations, $\phi_{i}(t)$; and (3) normalized activations multiplied by an independent set of “readout” gains, $\Phi^{\text{readout}}_{i}(t)=G^{\text{readout}}_{i}\phi_{i}(t)$, where the readout gains have the same distribution as the actual gains. Case (3) corresponds to variables with the same distribution of firing rates as the unnormalized activations, but where this heterogeneity is unrelated to the recurrent dynamics. In this case, we have $C^{\Phi^{\text{readout}}}_{12}=C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$. The four-point function is
$\Psi^{\Phi^{\text{readout}}}=\Bigg{[}\left(\frac{1}{\text{PR}^{G}}-1\right)\left(|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}+|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\right)\\ +1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\Bigg{]}\frac{C^{\Phi^{\text{readout}}}_{12}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}.$ (100)
To validate our theoretical predictions (Eq. 96, Eq. 97, and Eq. 100), we simulated networks with component strengths and gains given by
$D_{a}$ $\propto\exp(-\beta_{D}a/M),$ (101)
$G_{i}$ $\propto\exp(-\beta_{G}i/N).$ (102)
The dimension of activity in these networks is determined by the coupling strength, $g_{\text{eff}}$; the effective rank of the coupling matrix, $\alpha\text{PR}^{D}$; and the participation ratio of the single-neuron gain distribution, $\text{PR}^{G}$. Increases in each of these parameters leads to higher activity dimension for all three sets of activations.
Normalized activations consistently exhibit the highest dimension. While scaling normalized activations by heterogeneous factors reduces dimension as expected, the magnitude of this reduction depends on whether $G_{i}$ or $G^{\text{readout}}_{i}$ is used, even when their distributions are identical. This occurs because neurons with the largest gains preferentially participate in the leading modes of the normalized activations. Further scaling these already-dominant neurons by their gains results in overrepresentation of these neurons compared to using random gains. Thus, the dimension of unnormalized activations is lowest, with the dimension of random-readout activations falling between those of the normalized and unnormalized activations. This effect demonstrates that shuffling the normalizations of recorded neural activities should increase the dimension if firing rates and input covariances are related to one another, as they are in recurrent networks. Not observing such an increase in dimension, on the other hand, would suggest that heterogeneity in firing rates is unrelated to input covariance.
VII Mode overlaps
Figure 10: Input-output mode overlaps and eigenvalue spectra of the fly-brain connectome. (a) Overlap matrix between right and left singular vectors of the connectome. First 1000 modes are shown. (b) First 3000 diagonal elements of the overlap matrix from (a). Horizontal lines indicate an approximate cutoff for $\mathcal{O}(1/\sqrt{N})$ random overlaps. (c) Eigenvalue spectra of various connectome-based matrices. Top: Actual hemibrain spectrum, exhibiting large outliers along the real axis. Middle: Eigenvalues from a realization of the random-mode model, using singular values from the connectome as component strengths $D_{a}$, resulting in a circularly symmetric spectrum. Bottom: Random-mode model realization with input-output mode overlaps incorporated, where overlaps are given by the inner products of left and right singular vectors from the fly-connectome SVD. Data from Scheffer et al. [21], $N=18028$ neurons.
Thus far, we have assumed that the matrices $\bm{L}$ and $\bm{R}$ are independent of each other. Under this assumption, the overlap $\bm{L}_{a}^{T}\bm{R}_{b}$ between input mode $b$ and output mode $a$ is a random, $\mathcal{O}(1/\sqrt{N})$ quantity. However, to implement computations, modes must interact, suggesting that the coupling matrices of real neural circuits have larger, structured overlaps between input and output modes.
To examine interactions between different modes in a real neural circuit, we returned to the connectome of Drosophila melanogaster. We computed the overlap matrix between its right and left singular vectors, which correspond to the input and output modes, respectively. Rather than the unstructured, $\mathcal{O}(1/\sqrt{N})$ overlaps expected from i.i.d. random modes, we observe structured, large overlaps (Fig. 10a). These large overlaps are particularly pronounced near the diagonal of the overlap matrix (Fig. 10b).
Motivated by these observations, we now extend the random-mode model to incorporate structured overlaps between input and output modes. Specifically, we define $\bm{L}$ and $\bm{R}$ as zero-mean Gaussian matrices with the following covariance structure:
$\left\langle L_{ia}L_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{1}{N},$ (103)
$\left\langle R_{ia}R_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{1}{N},$ (104)
$\left\langle L_{ia}R_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{O_{ab}}{N},$ (105)
where $O_{ab}$ is the overlap matrix. In this case,
$\left\langle\bm{L}^{T}\bm{R}\right\rangle=\bm{O}.$ (106)
To capture the most salient feature of the fly connectome, large diagonal overlaps, we specialize to a diagonal form for the overlap matrix:
$O_{ab}=\delta_{ab}\rho_{a},$ (107)
where $\rho_{a}$ is the correlation between the $a$-th input and output modes. This form of the overlaps implies that, for large $N$, $\bm{L}_{a}^{T}\bm{R}_{a}=\rho_{a}$, while $\bm{L}_{a}^{T}\bm{R}_{b}$ for $a\neq b$ remains a random $\mathcal{O}(1/\sqrt{N})$ quantity. Incorporating this structure allows the random-mode model to more accurately match the eigenspectrum of the connectome, which has large outliers along the real axis (Fig. 10c).
Defining an instance of this connectivity with a well-defined large-$N$ limit requires specifying a joint distribution from which $(D_{a},\rho_{a})$ are sampled for each $a$. Our analytic expressions for the two- and four-point functions depend on averages $\left\langle\cdots\right\rangle_{D,\rho}$ over this distribution. Note that for large $N$, self-connections (autapses) in $\bm{J}$ are given by
$J_{ii}=\alpha\left\langle D\rho\right\rangle_{D,\rho},$ (108)
which is, in general, $\mathcal{O}(1)$. To avoid biologically unrealistic strong self-connections, one can choose $\left\langle D\rho\right\rangle_{D,\rho}=0$, in which case $J_{ii}$ is random and $\mathcal{O}(1/\sqrt{N})$. Alternatively, if one interprets each model neuron as a cluster of biological neurons, strong self-connections can be interpreted as aggregated within-cluster connectivity [64, 65]. Furthermore, note that the covariance between reciprocal pairs of couplings is given by $N\left\langle J_{ij}J_{ji}\right\rangle_{\bm{J}}=\sigma$ ($i\neq j$) where
$\sigma=\alpha\left\langle(D\rho)^{2}\right\rangle_{D,\rho}.$ (109)
We compute the two- and four-point functions using the path integral (Appendix I). As in previous cases, we obtain single-site processes for both neurons and latent variables. The neuronal single-site process is given by
$T[x](t)=\eta^{x}(t)+[R_{\star}\ast\phi](t),$ (110)
where $\eta^{x}\sim\mathcal{GP}(0,Q_{\star})$ and $\ast$ denotes convolution. The input covariance $Q_{\star}(t_{1},t_{2})$ and convolutional kernel $R_{\star}(t_{1},t_{2})$ are defined in terms of statistics of the latent-variable single-site process. Due to the linearity of the latent-variable dynamics, we can express them as
$Q_{\star}(\omega)$ $=\alpha\left\langle\frac{D^{2}}{|1-D\rho S^{\phi}_{\star}(\omega)|^{2}}\right\rangle_{D,\rho}C^{\phi}_{\star}(\omega),$ (111)
$R_{\star}(\omega)$ $=\alpha\left\langle\frac{D\rho}{1-D\rho S^{\phi}_{\star}(\omega)}\right\rangle_{D,\rho}.$ (112)
We enforce $C^{\phi}_{\star}(t_{1},t_{2})=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}$ and $S^{\phi}_{\star}(t_{1},t_{2})=\left\langle{\delta\phi(t_{1})}/{\delta I(t_{2})}\right\rangle_{\star}$, yielding a self-consistent system that determines $C^{\phi}_{\star}(t_{1},t_{2})$ and $S^{\phi}_{\star}(t_{1},t_{2})$. Unlike previous cases, there is no effective coupling strength $g_{\text{eff}}$ describing the relationship between $Q_{\star}(t_{1},t_{2})$ and $C^{\phi}_{\star}(t_{1},t_{2})$; instead, this relationship involves a convolution with a kernel defined using the joint distribution over $(D,\rho)$. This can be viewed as a frequency-dependent effective coupling strength.
Expanding the self-coupling kernel $R_{\star}(\omega)$ (Eq. 112) in powers of $\rho$ yields the first two terms
$R_{\star}(\omega)=J_{ii}+\sigma S^{\phi}_{\star}(\omega)+\cdots$ (113)
The first term is what we get in the single-site picture for i.i.d. cross-neuron couplings and deterministic $\mathcal{O}(1)$ self couplings. The second term appears in the single-site picture if there are correlations between reciprocal pairs of couplings, but the structure is otherwise i.i.d. (Appendix E). Nonzero higher-order terms reflect that the structure of the random-mode model is not fully captured by either of these forms. Note that the first two terms being zero forces all subsequent terms to be zero. Thus, nonzero covariance among reciprocal pairs of couplings is a necessary and sufficient condition for attaining a nonzero self-coupling kernel in the single-site picture.
Figure 11: Effect of input-output mode correlations on dimension of activity in the random-mode model. Dimension of activity $\text{PR}^{\phi}$ versus correlation parameter $\rho_{1}$ in the $(D,\rho)$ joint distribution, for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$. Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions. All simulations use $N=5000$ neurons.
To compute the four-point function, we write down and invert the frequency-dependent Hessian, and obtain
$\Psi^{\phi}_{\star}$ $=\frac{V-1+\left\{\frac{(1+X_{12})(1+X_{21}^{*})}{1-W^{*}}+\text{h.c.}\right\}}{|1-U|^{2}}C^{\phi}_{12},$ (114a)
$U$ $=\alpha\left\langle D^{2}\Sigma_{1}\Sigma_{2}\right\rangle_{D,\rho},$ (114b)
$V$ $=\alpha\left\langle D^{4}|\Sigma_{1}\Sigma_{2}|^{2}\right\rangle_{D,\rho},$ (114c)
$W$ $=\alpha\left\langle D^{2}\rho^{2}\Sigma_{1}^{*}\Sigma_{2}\right\rangle_{D,\rho},$ (114d)
$X_{12}$ $=\alpha\left\langle D^{3}\rho\Sigma_{1}|\Sigma_{2}|^{2}\right\rangle_{D,\rho},$ (114e)
$\Sigma_{k}$ $=\frac{S^{\phi}_{k}}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (114f)
We validated this theory by choosing parameterizations of $D_{a}$ and $\rho_{a}$. We used $\alpha=M/N=1$ and $D_{a}\propto\exp(-\beta_{D}a/N)$ for the component strengths, as before. For the overlaps $\rho_{a}$, we use a sigmoidal shape inspired by Fig. 10b in which the leading components have larger input-output mode overlaps. In particular, we used
$\rho_{a}=\rho_{1}\left(1-\frac{2}{1+\exp(-\beta_{\rho}(a/N-u_{\text{flip}}))}\right),$ (115)
which varies from $\rho_{1}$ to $-\rho_{1}$, hitting $\rho_{a}=0$ at $a/N=u_{\text{flip}}$. We fixed $\beta_{\rho}=16$, varied $\rho_{1}$, and set $u_{\text{flip}}$ such that $\left\langle D\rho\right\rangle_{D,\rho}=0$, eliminating strong self couplings. Across values of $g_{\text{eff}}$ and $\text{PR}^{D}$, sufficiently strong overlaps of either sign reduce activity dimension (Fig. 11). The maximum dimension occurs for $\rho_{1}<0$, a regime in which activity projected along connectivity modes with the greatest strength is most strongly suppressed by the recurrent dynamics. This leftward shift is reduced for larger $g_{\text{eff}}$.
VIII Discussion
A key question in neuroscience and machine learning is how learning shapes collective dynamics through modifications of a neural network’s connectivity. While our analysis has focused on disordered networks that generate chaotic activity and are not trained to perform tasks, the smooth low-rank connectivity structure we have explored provides a starting point for understanding the impact of learned coupling matrices on collective dynamics. Recent works have shown that large neural networks trained on real-world tasks exhibit smooth spectral structures in their coupling matrices, similar to those we have studied [20]. Moreover, while networks trained on simple tasks often display low-rank updates to their initial couplings, with a finite number of outlier eigenvalues emerging from the bulk [19, 46], training networks on many tasks, as is increasingly common, may produce coupling matrices with a low but extensive number of significant components [66, 67]. In such scenarios, the random-mode model, or extensions of it, may predict the resulting activity dimension and other aspects of collective dynamics. The path-integral formalism we have developed is also amenable to analyzing cases where the coupling matrix is drawn from a Gibbs ensemble with an energy function equal to the loss function for a task. This approach represents scenarios where couplings are trained to minimize a loss function with stochastic gradient descent [52]. Finally, initializing couplings with low-rank structure could improve training procedures like FORCE learning, which are based on suppressing chaos, because the initial strange attractor is lower-dimensional [11, 12].
Historically, neural connectivity was often characterized through network motifs – recurring patterns like cycles or chains among small neuronal groups [68, 14, 69, 70]. Such local statistics can be estimated using connectivity datasets obtained from methods that sample subsets of synapses. However, recent advances in whole-brain connectome reconstruction have unveiled global synaptic connectivity structures [71]. These connectomes reveal a more complex picture, where the spectrum and modes of the connectivity matrix exhibit large-scale structure that cannot be explained by local motifs [72]. Our parameterization of coupling matrices attempts to describe this global perspective, offering control over both the spectrum and mode overlaps. This could be particularly valuable in interpreting large-scale recordings of neural activity in the context of known anatomical connectivity.
In a similar vein, Tiberi et al. [72] proposed a random matrix theory approach for noise-driven linear networks based on parameterizing the eigendecomposition of $\bm{J}$. By contrast, our use of a DMFT approach allows us to describe nonlinear dynamics. Their parameterization also does not capture relationships between eigenvalues and eigenvectors, which are likely crucial for computations. Our parameterization, which is more akin to an SVD than an eigendecomposition, allows for a simultaneous parameterization of component strengths and corresponding mode overlaps. Other parameterizations of connectivity similar to the random-mode model have been used to study the relationship between spiking and rate-based recurrent networks [73] and to train spiking networks more efficiently [74].
Our description of single-neuron heterogeneity, where each neuron is parameterized by an arbitrary vector $\bm{\theta}$, can be readily applied to explore how various forms of heterogeneity influence dimensionality and other features of collective activity. Introducing specific forms of heterogeneity constrained by observed electrophysiological properties [75] and other cell-specific characteristics across neurons could provide insight on the functional consequences of neuronal diversity [76].
On the technical side, we have demonstrated that the two-site cavity method and the analysis of fluctuations around the saddle point in the path integral approach yield equivalent results for computing the four-point function. This equivalence implies that the intricate cavity construction is equivalent to computing and inverting the Hessian of the action in the path integral. A deeper understanding of the relationship between these methods could lead to new hybrid approaches that combine the strengths of both techniques.
In conclusion, the random-mode model and associated analytical techniques provide a set of tools for exploring the relationship between connectivity and collective dynamics. Our parameterization of connectivity is a middle ground between deterministic low-rank models and completely random networks. Future work incorporating more complex mode interactions, diverse neuronal properties, and learning dynamics will further bridge neural connectivity, activity, and function.
Acknowledgements
We gratefully acknowledge L.F. Abbott for discussions and feedback on this work.
Appendix A Glossary of Terms
Number of neurons ($N$)
Total number of neurons in the network.
Number of modes ($M$)
Number of components in the random-mode model.
(Effective) coupling strength ($g$, $g_{\text{eff}}$)
Standard deviation of the couplings times $\sqrt{N}$ in the i.i.d. connectivity model (Eq. 14), random-mode model (Eq. 74), or variants thereof (Eq. 98).
Random-mode model
Network connectivity model: $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{L}_{a}\bm{R}_{a}^{T}$ (Eq. 22).
Connectivity component ($\bm{L}_{a}\bm{R}_{a}^{T}$)
Outer product of output and input modes.
Input mode ($\bm{R}_{a}$)
Neuronal pattern in the $a$-th component onto which activity patterns are projected.
Output mode ($\bm{L}_{a}$)
Neuronal pattern in the $a$-th component along which the projected activity pattern is expanded.
Component strength ($D_{a}$)
Scaling factor for the $a$-th component.
Ratio of modes to neurons ($\alpha$)
Given by $\alpha=M/N$ (Eq. 24).
Participation ratio of component strengths ($\text{PR}^{D}$)
Given by $\text{PR}^{D}=r_{2}^{2}/r_{4}$ (Eq. 26) where $r_{n}=\langle D^{n}\rangle_{D}$ (Eq. 25).
Effective rank ($\alpha\text{PR}^{D}$)
Connectivity dimensionality measure (Eq. 27).
Two-point function
Correlation function given by $C^{\phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i}\langle\phi_{i}(t_{1})\phi_{i}(t_{2})\rangle_{t}$ (Eq. 5).
Four-point function
Higher-order correlation function given by $\Psi^{\phi}(\tau_{1},\tau_{2})=\frac{1}{N}\sum_{i,j}C^{\phi}_{ij}(\tau_{1})C^{\phi}_{ij}(\tau_{2})$ (Eq. 7). See also time-by-time definition (Eq. 9).
Appendix B Preprocessing of the fly hemibrain connectome
We used the dataset originally presented in Scheffer et al. [21]. To form the connection matrix for analysis, we used synaptic counts as a measure of connection strength, determining the magnitudes of the matrix elements. Neurons were assigned neurotransmitter probabilities based on a machine-learning analysis applied to the electron-microscopy data [22]. We selected the most probable neurotransmitter assignment for each neuron. The neurotransmitter types were then mapped to signs (excitatory or inhibitory) according to Table 1.
To normalize the connectivity matrix, we used an iterative scaling procedure analogous to the Sinkhorn–Knopp algorithm that proceeds as follows. We first rescale the positive elements in each row by a row-specific factor, such that the L2-norm of the positive elements is $1/\sqrt{2}$. The same procedure is applied to the negative elements, using a separate row-specific factor. This results in each row having a L2-norm of one with balanced excitatory and inhibitory inputs. A small number of neurons with exclusively excitatory or inhibitory inputs are left with an L2-norm of $1/\sqrt{2}$ instead of one. Following the row normalization, we perform the same operations on the columns. This disrupts the previous row normalization. Thus, we iterate the row and column normalization steps in an alternating fashion until convergence was achieved. The final result is a matrix with proper scaling along both rows and columns, which we interpret as the projection of the original matrix onto the set of properly row- and column-normalized matrices. Importantly, it has the same nonzero elements with the same signs as the original matrix.
Neurotransmitter Synaptic Sign
GABA -1 (Inhibitory)
Acetylcholine +1 (Excitatory)
Glutamate -1 (Inhibitory)
Serotonin Ignored
Octopamine Ignored
Dopamine Ignored
Neither Ignored
Table 1: Mapping of neurotransmitters to synaptic signs
Appendix C Temporal limits and inverses
Here, we demonstrate, for the i.i.d. model that the temporal limits in Eq. 34 can be taken before the inverses. This allows us to work with a frequency-dependent Hessian. In this section, we use Einstein notation for integrals over time variables (i.e., repeated time indices are integrated over). We would like to evaluate
$\Psi(\tau_{1},\tau_{2})=U^{-1}(t_{1},t_{2},s_{1},s_{2})K(s_{1},s_{2},s_{3},s_{4})U^{-1}(t_{1}+\tau_{1},t_{2}+\tau_{2},s_{3},s_{4})$ (116)
in the limits Eq. 34, where $U(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ and $K(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ (note that $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0$ due to the vanishing of correlation functions involving only the conjugate field). Restating Eq. 43, we have
$U(t_{1},t_{2},t_{3},t_{4})$ $=-\left[\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})\right]-g^{2}\langle\langle\phi(t_{1})\phi(t_{2}),\hat{x}(t_{3})\hat{x}(t_{4})\rangle\rangle_{\star},$ (117)
$K(t_{1},t_{2},t_{3},t_{4})$ $=\langle\langle\phi(t_{1})\phi(t_{2}),\phi(t_{3})\phi(t_{4})\rangle\rangle_{\star}.$ (118)
We assume temporal stationarity throughout this section and consider only $a=\phi$.
We start by examining the properties of $U(t_{1},t_{2},t_{3},t_{4})$. $U$ vanishes if one of the four time points is far (measured on the timescale of $C^{\phi}_{\star}(\tau)$) from the other three. For $t_{1}$ or $t_{2}$, this is a consequence of the perturbation associated with $\hat{x}(t_{1})$ or $\hat{x}(t_{2})$ being far away; for $t_{3}$ or $t_{4}$, this is a consequence of $\langle\phi(t_{i})\rangle_{\star}=0$ separating from the expectation. $U$ also vanishes if there is more than one time point which is far from the remaining ones by the same arguments. Thus, the time points need to be close in pairs. However, if $t_{1},t_{2}$ are close and $t_{3},t_{4}$ are close, but the pairs are distant, $U$ still vanishes since the perturbations associated with $\hat{x}(t_{i})$ are both far. This leaves three possibilities for a non-vanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
Separating the three possibilities into two subdomains where all time points are close ($c$) and where the time points belong to either of the two pairwise separate ($s$) possibilities, we write $U=U_{s}+U_{c}$ where $U_{s}$ and $U_{c}$ are defined to vanish on the respective other domain. On the separated domain, $U$ simplifies to
$U_{s}(t_{1},t_{2},t_{3},t_{4})=-\left[\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})\right]\\ +g^{2}\left[S^{\phi}_{\star}(t_{1}-t_{3})S^{\phi}_{\star}(t_{2}-t_{4})+S^{\phi}_{\star}(t_{1}-t_{4})S^{\phi}_{\star}(t_{2}-t_{3})\right],$ (119)
where we used that the two distant contributions are non-overlapping to add them.
We assume that we can also decompose the inverse into a contribution for separate pairs and a contribution where all time points are close, $U^{-1}=U_{s}^{-1}+U_{c}^{-1}$, where again $U_{s}^{-1}$ and $U_{c}^{-1}$ vanish on the respective other domain. We demonstrate that $U_{s}^{-1}$ is the inverse of $U_{s}$, and that $U_{c}^{-1}$ is the inverse of $U_{c}$, by noting that $U^{-1}(t_{1},t_{2},s_{1},s_{2})U(s_{1},s_{2},t_{3},t_{4})$ reduces to the contraction of $U_{c}^{-1}$ and $U_{c}$ if $t_{3}$ and $t_{4}$, and hence $s_{1}$ and $s_{2}$, are close; or to the contraction of $U_{s}^{-1}$ and $U_{s}$ if $t_{3}$ and $t_{4}$, and hence $s_{1}$ and $s_{2}$, are distant.
For Eq. 116, we need $U^{-1}_{s}$ due to the limits Eq. 34. The fact that $U$ reduces to $U_{s}$ in the limits Eq. 34 justifies taking the limits before inverting $U$. Furthermore, these limits imply that, in Eq. 116, $s_{1}$ and $s_{2}$, as well as $s_{3}$ and $s_{4}$, are far apart. In this case, $K$ simplifies to
$K_{s}(s_{1},s_{2},s_{3},s_{4})=C^{\phi}_{\star}(s_{1}-s_{3})C^{\phi}_{\star}(s_{2}-s_{4})+C^{\phi}_{\star}(s_{1}-s_{4})C^{\phi}_{\star}(s_{2}-s_{3}).$ (120)
This justifies taking the limits Eq. 34 for $K$.
Appendix D Incorporating sources
This section outlines the general approach for incorporating source terms and then applies it to the specific case of correlations among the pre-activations. Our goal is to compute
$\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0},$ (121)
which gives the fluctuations around the mean of the quantity multiplying $\mathcal{J}$ in the action. To facilitate this calculation, we introduce a new field $U$ and set it equal to $\mathcal{J}$ using the conjugate $\hat{U}$:
$Z[\mathcal{J}]=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp\left(-N\hat{U}U+N\hat{U}\mathcal{J}-NS[\hat{C};U]\right).$ (122)
Instead of taking the second derivative with respect to $\mathcal{J}$, we set $\mathcal{J}=0$ and compute the fluctuations of $\hat{U}$ in the augmented theory whose path integral is
$Z=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp(-N\tilde{\mathcal{S}}[\bm{C},U,\hat{U}]),$ (123)
where
$\tilde{\mathcal{S}}[\bm{C},U,\hat{U}]=\hat{U}U+\mathcal{S}[\bm{C};U].$ (124)
The Hessian of the augmented action, in terms of the Hessian of the original action, is
$\left(\begin{array}[]{c|cc}\bm{H}&\frac{\partial^{2}S}{\partial\mathcal{J}\partial\bm{C}}&\bm{0}\\[2.5pt] \hline\cr\\[-10.00002pt] \left(\frac{\partial^{2}S}{\partial\mathcal{J}\partial\bm{C}}\right)^{T}&\frac{\partial^{2}S}{\partial\mathcal{J}^{2}}&1\\[5.0pt] \bm{0}^{T}&1&0\end{array}\right),$ (125)
where
$H_{ab}=\frac{\delta^{2}S}{\delta C_{a}\delta C_{b}},\quad\left[\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{a}=\frac{\delta^{2}S}{\delta\mathcal{J}\delta C_{a}}.$ (126)
The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point. Using a Schur complement to compute this element, we obtain the expression Eq. 54 stated in the main text.
We now apply Eq. 54 to the action of the i.i.d. connectivity model with a source term for correlations among pre-activations. The action is
$\mathcal{S}[C^{\phi},\hat{C}^{\phi}]=-\frac{1}{2}C^{\phi}\hat{C}^{\phi}-\log\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\bigg{(}i\hat{x}T[x]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi-\frac{1}{2}x\mathcal{J}x\bigg{)}.$ (127)
Computing the necessary quantities and taking the temporal limits of Eq. 34, we obtain, in Fourier space,
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\quad\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}g^{2}S_{12}^{x}\\ -C_{12}^{x\phi}\end{pmatrix},$ (128)
where $C^{x\phi}_{12}=C_{\star}^{x\phi}(\omega_{1})C_{\star}^{x\phi}(\omega_{2})$ and $S^{x}_{12}=S^{x}_{\star}(\omega_{1})S^{x}_{\star}(\omega_{2})$. Substituting these into Eq. 54 and using the frequency-dependent Hessian from Eq. 52, we obtain
$\Psi_{\star}^{x}=C^{x}_{12}+|U|^{2}C^{\phi}_{12}+UC^{x\phi}_{12}+\text{h.c.},\quad\text{where }U=\frac{g^{2}S^{x}_{12}}{1-g^{2}S^{\phi}_{12}},$ (129)
in agreement with the two-site cavity result of Clark et al. [25], given in Eq. 21.
Appendix E Partially symmetric disorder
To demonstrate the flexibility of the path-integral approach to computing the four-point function, we now consider a generalization of the classic i.i.d. model in which we introduce correlations between reciprocal couplings, characterized by the following statistics:
$\left\langle J_{ij}J_{k\ell}\right\rangle_{\bm{J}}=g^{2}\delta_{ik}\delta_{j\ell}+g^{2}\rho\delta_{i\ell}\delta_{jk}.$ (130)
Here, $\rho\in[-1,1]$ is a parameter that controls the degree of symmetry in the connectivity. When $\rho=1$, the connectivity is fully symmetric, while $\rho=-1$ corresponds to fully antisymmetric connectivity. The case $\rho=0$ recovers the i.i.d. model. We express this partially symmetric connectivity as a linear combination of two matrices, $\bm{X}$ and $\bm{Y}$, that are i.i.d. random with mean zero and variance $1/N$:
$J_{ij}=\sqrt{{g^{2}-|\sigma|}}X_{ij}+\sqrt{\frac{|\sigma|}{2}}(Y_{ij}+\text{sgn}(\sigma)Y_{ji}),$ (131)
where $\sigma=\rho g^{2}$. In terms of $\bm{X}$ and $\bm{Y}$, the path integral for this partially symmetric model is
$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg{(}i\sum_{i}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{ij}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{ij}Y_{ij}(\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j})\bigg{)}.$ (132)
We integrate out $\bm{X}$ and $\bm{Y}$. To factorize the action over indices spanning extensive dimensions, we introduce
[…]
We obtain, for the disorder-averaged path integral,
$Z$ $=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}R\int\mathcal{D}\hat{R}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]\right),$ (135)
$\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=-\frac{1}{2}C^{\phi}\hat{C}^{\phi}+\frac{\sigma}{2}S^{\phi}\hat{S}^{\phi}-\log W[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}],$ (136)
$W[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=\int\mathcal{D}\hat{x}\int\mathcal{D}x\exp\bigg{(}i\hat{x}T[x]-\frac{1}{2}\phi\hat{C}^{\phi}\phi-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{i\sigma}{2}\phi\hat{S}^{\phi}\hat{x}-\frac{i\sigma}{2}\hat{x}S^{\phi}\phi\bigg{)}.$ (137)
To obtain the saddle-point solution, we compute the derivatives of the action, under the rule that derivatives with respect to $S^{\phi}(t_{1},t_{2})$ also affect $\hat{S}^{\phi}(t_{2},t_{1})$, and vice versa, due to the symmetry in the action. This gives
$\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+g^{2}\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W},$ (138)
$\frac{\delta\mathcal{S}}{\delta\hat{C}^{\phi}(t_{1},t_{2})}$ $=-{C}^{\phi}(t_{1},t_{2})+\left\langle{\phi}(t_{1}){\phi}(t_{2})\right\rangle_{W},$ (139)
$\frac{\delta\mathcal{S}}{\delta S^{\phi}(t_{1},t_{2})}$ $=\sigma\hat{S}^{\phi}(t_{1},t_{2})+i\sigma\left\langle\hat{x}(t_{1})\phi(t_{2})\right\rangle_{W},$ (140)
$\frac{\delta\mathcal{S}}{\delta\hat{S}^{\phi}(t_{1},t_{2})}$ $=\sigma S^{\phi}(t_{1},t_{2})+i\sigma\left\langle\phi(t_{1})\hat{x}(t_{2})\right\rangle_{W}.$ (141)
Setting these derivatives to zero yields the saddle-point conditions:
$\hat{C}^{\phi}_{\star}(t_{1},t_{2})$ $=0,$ (142)
$C^{\phi}_{\star}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (143)
$S^{\phi}_{\star}(t_{1},t_{2})$ $=\hat{S}^{\phi}_{\star}(t_{2},t_{1})=\left\langle\frac{\delta\phi(t_{1})}{\delta I(t_{2})}\right\rangle_{\star},$ (144)
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$. The single-site process at the saddle point is described by
$T[x](t)$ $=\eta^{x}(t)+\sigma\left[S^{\phi}_{\star}\ast\phi\right](t),$ (145)
$\eta^{x}$ $\sim\mathcal{GP}(0,g^{2}C^{\phi}_{\star}),$ (146)
where $\ast$ denotes convolution. The symmetric structure provides a convolutional, nonlinear self-coupling $\sigma[S^{\phi}_{\star}\ast\phi](t)$ in the single-site dynamics. The two-point correlation and response functions, $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$, must be determined self-consistently within this single site picture.
Introducing the notation $C^{\phi}_{k}=C^{\phi}_{\star}(\omega_{k})$, $S^{\phi}_{k}=S^{\phi}_{\star}(\omega_{k})$ for $k=1,2$, the frequency-dependent Hessian at the saddle point is
$\bm{H}=\begin{pmatrix}0&-1+g^{2}(S^{\phi}_{12})^{*}&0&0\\ -1+g^{2}S^{\phi}_{12}&-C^{\phi}_{12}&\sigma S^{\phi}_{1}C^{\phi}_{2}&\sigma S^{\phi}_{2}C^{\phi}_{1}\\ 0&\sigma(S^{\phi}_{1})^{*}C^{\phi}_{2}&0&\sigma-\sigma^{2}(S^{\phi}_{1})^{*}S^{\phi}_{2}\\ 0&\sigma(S^{\phi}_{2})^{*}C^{\phi}_{1}&\sigma-\sigma^{2}S^{\phi}_{1}(S^{\phi}_{2})^{*}&0\end{pmatrix}.$ (147)
Inverting this matrix and isolating the upper-left element, we obtain, in agreement with Clark et al. [25],
$\Psi^{\phi}_{\star}=\frac{1}{|1-g^{2}S^{\phi}_{12}|^{2}}\frac{1-|\sigma(S^{\phi}_{1})^{*}S^{\phi}_{2}|^{2}}{|1-\sigma(S^{\phi}_{1})^{*}S^{\phi}_{2}|^{2}}C^{\phi}_{12}.$ (148)
Appendix F Inverse formula
In Sec. III.5.2 and Appendix H, we use the fact that
$\begin{pmatrix}0&a^{*}&0&-1\\ a&b&-1&0\\ 0&-1&0&c^{*}\\ -1&0&c&d\end{pmatrix}^{-1}=\begin{pmatrix}-\frac{b+|c|^{2}d}{|1-ac|^{2}}&-\frac{c}{1-ac}&-\frac{a^{*}d+bc}{|1-ac|^{2}}&-\frac{1}{1-ac}\\ -\frac{c^{*}}{1-(ac)^{*}}&0&-\frac{1}{1-(ac)^{*}}&0\\ -\frac{ad+bc^{*}}{|1-ac|^{2}}&-\frac{1}{1-ac}&-\frac{|a|^{2}d+b}{|1-ac|^{2}}&-\frac{a}{1-ac}\\ -\frac{1}{1-(ac)^{*}}&0&-\frac{a^{*}}{1-(ac)^{*}}&0\end{pmatrix}.$ (149)
Appendix G Two-site cavity calculation of the four-point function in the random-mode model
To apply cavity techniques, we reformulate the network as a bipartite system of neurons and latent variables,
$T[x_{i}](t)$ $=\sum_{a}L_{ia}z_{a}(t),$ (150)
$z_{a}(t)$ $=D_{a}\sum_{j}R_{ja}\phi_{j}(t),$ (151)
as done in the path-integral approach. In the two-site cavity calculation, we keep track various intermediate quantities to order $1/\sqrt{N}$, leading to an expression for $\Psi^{\phi}(\bm{\tau})$ accurate to leading order (order one), which we identify as $\Psi^{\phi}_{\star}(\bm{\tau})$.
To distinguish between different types of variables and their indices, we use the following notation. Neurons are indexed by $i,j\in\{1,\ldots,N\}$ and latent variables by $a,b\in\{1,\ldots,M\}$, as usual. For the cavity variables, we introduce special indices: neuronal cavity variables are indexed by $\mu,\nu\in\{0,0^{\prime}\}$, while latent cavity variables are indexed by $\alpha,\beta\in\{\hat{0},\hat{0}^{\prime}\}$.
G.1 Neuronal cavity
We begin by introducing two neurons, $x_{0}(t)$ and $x_{0^{\prime}}(t)$. The leading-order ($1/\sqrt{N}$) effect on the latent variables is
[…]
where $S^{z}_{ab}(t_{1},t_{2})$ is the response function of the latent variables. The dynamic equations for the cavity neurons are
[…]
where
$\eta_{\mu}(t)$ $=\sum_{a}L_{\mu a}z_{a}(t),$ (154)
$F_{\mu\nu}(t_{1},t_{2})$ $=\sqrt{N}\sum_{a,b}L_{\mu a}D_{b}R_{\nu b}S^{z}_{ab}(t_{1},t_{2}).$ (155)
Defining the cavity-field time-average cross-covariance as
$C^{\eta}_{00^{\prime}}(\omega)=\sum_{a,b}L_{0a}L_{0^{\prime}b}C^{z}_{ab}(\omega),$ (156)
we obtain the following expression for the cross-covariance of the cavity units, up to order $1/\sqrt{N}$:
$C^{\phi}_{00^{\prime}}(\omega)=|S_{\star}^{\phi}(\omega)|^{2}C^{\eta}_{00^{\prime}}(\omega)+\frac{1}{\sqrt{N}}\left[(S_{\star}^{\phi}(\omega)F_{00^{\prime}}(\omega))^{*}+S_{\star}^{\phi}(\omega)F_{0^{\prime}0}(\omega)\right]C^{\phi}_{\star}(\omega).$ (157)
Our goal is to compute the parameter
$\psi^{\phi}(\bm{\omega})=N\left\langle C^{\phi}_{00^{\prime}}(\omega_{1})C^{\phi}_{00^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}},$ (158)
where $\left\langle\cdot\right\rangle_{\bm{L},\bm{D},\bm{R}}$ denotes an average over $\bm{L}$, $\bm{D}$, and $\bm{R}$. To evaluate this, we need to square and disorder-average Eq. 157. This requires us to consider several two-frequency correlation functions, which we define as
$\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}(\bm{\omega})$ $=\left\langle F_{00^{\prime}}(\omega_{1})F_{00^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}},$ (159)
$\Gamma_{F^{*}_{00^{\prime}}F_{0^{\prime}0}}(\bm{\omega})$ $=\left\langle F^{*}_{00^{\prime}}(\omega_{1})F_{0^{\prime}0}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}},$ (160)
$\Gamma_{F^{*}_{00^{\prime}}C^{\eta}_{00^{\prime}}}(\bm{\omega})$ $=\sqrt{N}\left\langle F^{*}_{00^{\prime}}(\omega_{1})C^{\eta}_{00^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}},$ (161)
$\Gamma_{C^{\eta}_{00^{\prime}}C^{\eta}_{00^{\prime}}}(\bm{\omega})$ $=N\left\langle C^{\eta}_{00^{\prime}}(\omega_{1})C^{\eta}_{00^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}}.$ (162)
These have all been scaled to be order-one. We can evaluate these $\Gamma_{\cdots}(\bm{\omega})$ functions due to the independence of the couplings and dynamic variables in the expressions defining them, which is a consequence of the cavity construction. Of these, the only nonvanishing ones are
$\Gamma_{C^{\eta}_{00^{\prime}}C^{\eta}_{00^{\prime}}}(\bm{\omega})$ $=\alpha\left\langle C^{z}_{\hat{0}\hat{0}}(\omega_{1})C^{z}_{\hat{0}\hat{0}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}}+\alpha^{2}N\left\langle C^{z}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{z}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}},$ (163)
$\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}(\bm{\omega})$ $=\alpha r_{2}+\alpha^{2}N\left\langle D_{\hat{0}^{\prime}}^{2}S^{z}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})S^{z}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}},$ (164)
where $\hat{0},\hat{0}^{\prime}$ denote the indices in the latent-variable cavity picture.
G.2 Latent-variable cavity
We now introduce two new latent variables, $z_{\hat{0}}(t)$ and $z_{\hat{0}^{\prime}}(t)$. The leading-order ($1/\sqrt{N}$) effect on the neurons is
[…]
where $S^{\phi}_{ij}(t_{1},t_{2})$ is the response function of the neurons. The dynamic equations for the cavity latent variables are
[…]
where
$\gamma_{\alpha}(t)$ $=\sum_{i}R_{i\alpha}\phi_{i}(t),$ (167)
$G_{\alpha\beta}(t_{1},t_{2})$ $=\sqrt{N}\sum_{i,j}R_{i\alpha}L_{j\beta}S^{\phi}_{ij}(t_{1},t_{2}).$ (168)
Defining the time-average cavity-field cross-covariance as
$C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega)=\sum_{i,j}R_{i\hat{0}}R_{j\hat{0}^{\prime}}C^{\phi}_{ij}(\omega),$ (169)
we obtain the following expression for the cross-covariance of the cavity latent variables up to order $1/\sqrt{N}$:
$C^{z}_{\hat{0}\hat{0}^{\prime}}(\omega)=D_{\hat{0}}D_{\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega)+\frac{1}{\sqrt{N}}\left[D_{\hat{0}}G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega)C^{z}_{\hat{0}^{\prime}\hat{0}^{\prime}}(\omega)+D_{\hat{0}^{\prime}}G_{\hat{0}^{\prime}\hat{0}}(\omega)C^{z}_{\hat{0}\hat{0}}(\omega)\right].$ (170)
We aim to compute the parameter
$\psi^{z}(\bm{\omega})=N\left\langle C^{z}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{z}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}}$ (171)
by squaring and disorder-averaging Eq. 170. This requires us to consider several two-frequency correlation functions, which we define as
$\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\left\langle G_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (172)
$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}^{\prime}\hat{0}}}(\bm{\omega})$ $=\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}^{\prime}\hat{0}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (173)
$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\sqrt{N}\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (174)
$\Gamma_{C^{\gamma}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=N\left\langle C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}}.$ (175)
Again, we can evaluate these $\Gamma_{\cdots}(\bm{\omega})$ functions due to the independence of the couplings and dynamic variables in the expressions defining them. The nonvanishing ones are
$\Gamma_{C^{\gamma}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2})+N\left\langle C^{\phi}_{00^{\prime}}(\omega_{1})C^{\phi}_{00^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}},$ (176)
$\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})+N\left\langle S^{\phi}_{00^{\prime}}(\omega_{1})S^{\phi}_{00^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}},$ (177)
where $0,0^{\prime}$ denote the indices in the neuronal cavity picture.
G.3 Combining the two two-site cavity pictures
To combine the results from both two-site cavity pictures, we use the following relations:
$C^{z}_{\hat{0}\hat{0}}(\omega)$ $=D_{\hat{0}}^{2}C^{\phi}_{\star}(\omega),$ (178)
$S^{z}_{\hat{0}\hat{0}^{\prime}}(\omega)$ $=\frac{D_{\hat{0}}}{\sqrt{N}}G_{\hat{0}\hat{0}^{\prime}}(\omega),$ (179)
$S^{\phi}_{00^{\prime}}(\omega)$ $=\frac{1}{\sqrt{N}}(S^{\phi}_{\star}(\omega))^{2}F_{00^{\prime}}(\omega).$ (180)
Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions (Eqs. 159-162 and 172175), we obtain
$\Gamma_{C^{\eta}_{00^{\prime}}C^{\eta}_{00^{\prime}}}(\bm{\omega})$ $=\alpha r_{4}C_{\star}^{\phi}(\omega_{1})C_{\star}^{\phi}(\omega_{2})+\alpha^{2}\psi^{z}(\bm{\omega}),$ (181)
$\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}(\bm{\omega})$ $=\alpha r_{2}\left[1+\alpha r_{2}\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})\right],$ (182)
$\Gamma_{C^{\gamma}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=C_{\star}^{\phi}(\omega_{1})C_{\star}^{\phi}(\omega_{2})+\psi^{\phi}(\bm{\omega}),$ (183)
$\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=S_{\star}^{\phi}(\omega_{1})S_{\star}^{\phi}(\omega_{2})\left[1+S_{\star}^{\phi}(\omega_{1})S_{\star}^{\phi}(\omega_{2})\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}(\bm{\omega})\right].$ (184)
We solve Eq. 182 and Eq. 184 simultaneously, yielding
$\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}(\bm{\omega})$ $=\frac{\alpha r_{2}}{1-\alpha r_{2}S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})},$ (185)
$\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\frac{S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})}{1-\alpha r_{2}S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})}.$ (186)
With these solutions, we express $\psi^{z}_{\star}(\bm{\omega})$ and $\psi^{\phi}_{\star}(\bm{\omega})$ in terms of the $\Gamma_{\cdots}(\bm{\omega})$ functions. Switching to frequency-suppressed notation,
$\psi^{z}_{\star}$ $=r_{2}^{2}\Gamma_{C^{\gamma}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}+r_{2}r_{4}\left(\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}+\text{h.c.}\right)C^{\phi}_{12},$ (187)
$\psi^{\phi}_{\star}$ $=|S^{\phi}_{12}|^{2}\Gamma_{C^{\eta}_{00^{\prime}}C^{\eta}_{00^{\prime}}}+\left(\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}S^{\phi}_{12}+\text{h.c.}\right)C^{\phi}_{12}.$ (188)
Now, we substitute Eq. 187 into Eq. 188. Using $\Psi^{\phi}_{\star}=\psi^{\phi}_{\star}+C^{\phi}_{12}$, we obtain
$\Psi^{\phi}_{\star}=|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\left[\Psi^{\phi}_{\star}+\frac{1}{\alpha\text{PR}^{D}}\frac{1-|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12}\right]+\frac{1-|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12},$ (189)
where we used the previously defined quantities $g_{\text{eff}}^{2}=\alpha r_{2}$ and $\text{PR}^{D}=r_{2}^{2}/r_{4}$. Eq. 189 can be solved for $\Psi^{\phi}_{\star}$, yielding
$\Psi^{\phi}_{\star}=\frac{1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12}.$ (190)
This result is identical to Eq. 76 derived using the path-integral approach, demonstrating the consistency between the two methods.
Appendix H Single-unit heterogeneity calculations
The full set of order parameters is
[…]
The action and single-site path integrals for this system are
$\mathcal{S}[C^{\Phi},\hat{C}^{\Phi},Q,\hat{Q}]$ $=-\frac{1}{2}C^{\Phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}-\left\langle\log W_{\bm{\theta}}[Q,\hat{Q};\mathcal{J}]\right\rangle_{\bm{\theta}}-\alpha\left\langle\log W^{z}_{D}[C^{\Phi},\hat{C}^{\Phi}]\right\rangle_{D},$ (193)
$W^{x}_{\bm{\theta}}[Q,\hat{Q};\mathcal{J}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\Phi_{\bm{\theta}}\hat{Q}\Phi_{\bm{\theta}}-\frac{1}{2}\hat{x}Q\hat{x}-\frac{1}{2}\phi\mathcal{J}\phi\right),$ (194)
$W^{z}_{D}[C^{\Phi},\hat{C}^{\Phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}z\hat{C}^{\Phi}z-\frac{1}{2}D^{2}\hat{z}C^{\Phi}\hat{z}\right).$ (195)
The saddle-point calculation proceeds similarly to the non-heterogeneous case of the random-mode model. We obtain
$\hat{C}^{\Phi}_{\star}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (196)
$C^{\Phi}_{\star}(t_{1},t_{2})$ $=\left\langle\left\langle\Phi_{\bm{\theta}}(t)\Phi_{\bm{\theta}}(t^{\prime})\right\rangle_{\star}\right\rangle_{\bm{\theta}},$ (197)
$Q_{\star}(t_{1},t_{2})$ $=\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}\right\rangle_{D},$ (198)
where $\left\langle\cdots\right\rangle_{\star}$ and $\left\langle\cdots\right\rangle_{\star D}$ are averages with respect to $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ and $W^{z}_{D}[C^{\Phi}_{\star},0]$, respectively (note that $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ does not depend on $\bm{\theta}$). Using $\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}=D^{2}C^{\Phi}_{\star}(t_{1},t_{2})$ and, consequently, $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\Phi}_{\star}(t_{1},t_{2})$, the self-consistent problem that determines $C^{\Phi}_{\star}(t_{1},t_{2})$ is given by Eq. 89 and the preceding description in Sec. VI.
The frequency-dependent Hessian at the saddle point is
$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}{C_{{12}}^{\Phi}}&-1&0\\ \hline\cr 0&-1&0&(\mathfrak{s}^{\Phi}_{12})^{*}\\ -1&0&\mathfrak{s}^{\Phi}_{12}&-\mathfrak{C}^{\Phi}_{12}\end{array}\right),$ (199)
where the new $\mathfrak{F}$raktur variables are defined in Eq. 90. We invert the Hessian using Eq. 149 and identify the upper-left element as Eq. 91. Combining the inverted Hessian with the relevant quantities for computing fluctuations of normalized variables,
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C^{\phi}_{12},\>\>\>\frac{\delta^{2}S}{\delta\bm{C}\delta\mathcal{J}}=\begin{pmatrix}0\\ 0\\ (S^{\phi}_{12})^{*}\\ -\mathfrak{C}^{\Phi\phi}_{12}\end{pmatrix},$ (200)
where
$\mathfrak{C}^{\Phi\phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (201)
$S^{\phi}_{12}$ $=\left\langle\frac{\delta\phi}{\delta I}\right\rangle_{\star}(\omega_{1})\left\langle\frac{\delta\phi}{\delta I}\right\rangle_{\star}(\omega_{2}),$ (202)
we obtain Eq. 92.
Appendix I Structured mode overlaps calculations
To generate matrices $\bm{L}$ and $\bm{R}$ with the correlation structure of Eq. 106 and Eq. 107, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$:
${L}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{1}_{ia}+\sqrt{{|\rho_{a}|}}{Y}_{ia},$ (203)
${R}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{2}_{ia}+\text{sgn}(\rho_{a})\sqrt{{|\rho_{a}|}}{Y}_{ia}.$ (204)
We begin with the path integral containing latent variables, insert this parameterization of $\bm{L}$ and $\bm{R}$, then and integrate out $\bm{X}_{1}$, $\bm{X}_{2}$, and $\bm{Y}$. To factorize the action over indices spanning extensive dimensions, we introduce
[…]
The resulting path integral, action, and single-site path integrals are given by:
$\centering Z=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}S^{\phi}\@add@centering$ $\int\mathcal{D}\hat{S}^{\phi}\int\mathcal{D}Q\int\mathcal{D}\hat{Q}\int\mathcal{D}R\int\mathcal{D}\hat{R}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi},Q,\hat{Q},R,\hat{R}]\right),$ (209)
$\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi},Q,\hat{Q},R,\hat{R}]$ $=-\frac{1}{2}C^{\phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}+\frac{1}{2}S^{\phi}\hat{R}+\frac{1}{2}R\hat{S}^{\phi}$ (210)
$\hskip 100.00015pt-\log W^{x}[Q,\hat{Q},R,\hat{R}]-\alpha\left\langle\log W^{z}_{D,\rho}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]\right\rangle_{D,\rho},$ (211)
$W^{x}[Q,\hat{Q},R,\hat{R}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\hat{x}Q\hat{x}-\frac{1}{2}\phi\hat{Q}\phi-\frac{i}{2}\hat{x}R\phi-\frac{i}{2}\phi\hat{R}\hat{x}\right),$ (212)
$W^{z}_{D,\rho}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}D^{2}\hat{z}C^{\phi}\hat{z}-\frac{1}{2}z\hat{C}^{\phi}z-\frac{i}{2}D\rho\hat{z}S^{\phi}z-\frac{i}{2}D\rho z\hat{S}^{\phi}\hat{z}\right).$ (213)
We first calculate the saddle point conditions by taking the derivatives of the action:
$\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{Q}(t_{1},t_{2})+\alpha\left\langle D^{2}\left\langle\hat{z}(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (214)
$\frac{\delta\mathcal{S}}{\delta\hat{C}^{\phi}(t_{1},t_{2})}$ $=-Q(t_{1},t_{2})+\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (215)
$\frac{\delta\mathcal{S}}{\delta S^{\phi}(t_{1},t_{2})}$ $=\hat{R}(t_{1},t_{2})+i\alpha\left\langle D\rho\left\langle\hat{z}(t_{1})z(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (216)
$\frac{\delta\mathcal{S}}{\delta\hat{S}^{\phi}(t_{1},t_{2})}$ $=R(t_{1},t_{2})+i\alpha\left\langle D\rho\left\langle z(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (217)
$\frac{\delta\mathcal{S}}{\delta Q(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W^{x}},$ (218)
$\frac{\delta\mathcal{S}}{\delta\hat{Q}(t_{1},t_{2})}$ $=-C^{\phi}(t_{1},t_{2})+\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{W^{x}},$ (219)
$\frac{\delta\mathcal{S}}{\delta R(t_{1},t_{2})}$ $=\hat{S}^{\phi}(t_{1},t_{2})+i\left\langle\hat{x}(t_{1})\phi(t_{2})\right\rangle_{W^{x}},$ (220)
$\frac{\delta\mathcal{S}}{\delta\hat{R}(t_{1},t_{2})}$ $=S^{\phi}(t_{1},t_{2})+i\left\langle\phi(t_{1})\hat{x}(t_{2})\right\rangle_{W^{x}}.$ (221)
The saddle-point conditions yield:
$\hat{C}^{\phi}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (222)
$C^{\phi}_{\star}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (223)
$S^{\phi}_{\star}(t_{1},t_{2})$ $=\hat{S}^{\phi}_{\star}(t_{1},t_{2})=\left\langle\frac{\delta\phi(t_{1})}{\delta I(t_{2})}\right\rangle_{\star},$ (224)
$Q_{\star}(t_{1},t_{2})$ $=\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D,\rho}\right\rangle_{D,\rho},$ (225)
$R_{\star}(t_{1},t_{2})$ $=\hat{R}_{\star}(t_{1},t_{2})=\alpha\left\langle D\rho\left\langle\frac{\delta z(t_{1})}{\delta I(t_{2})}\right\rangle_{\star D,\rho}\right\rangle_{D,\rho},$ (226)
where $\left\langle\cdots\right\rangle_{\star D,\rho}$ and $\left\langle\cdots\right\rangle_{\star}$ are averages within the dynamic processes described by $W^{z}_{D,\rho}[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$ and $W^{x}[Q_{\star},0,R_{\star},\hat{R}_{\star}]$, respectively. The yields neuronal and latent-variable single-site processes described by
$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\ast\phi](t)$ (227)
$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\ast z](t),$ (228)
where the Gaussian fields have statistics
$\eta^{x}$ $\sim\mathcal{GP}(0,Q_{\star}),$ (229)
$\eta^{z}$ $\sim\mathcal{GP}(0,D^{2}C^{\phi}_{\star}).$ (230)
This gives rise to the single-site problem described in Sec. VII.
The frequency-dependent Hessian at the saddle point is
$\bm{H}=\left(\begin{array}[]{cccc|cccc}0&u^{*}&0&0&0&-1&0&0\\ u&-vC^{\phi}_{1}C^{\phi}_{2}&x_{12}C_{2}&x_{21}C^{\phi}_{1}&-1&0&0&0\\ 0&x_{12}^{*}C^{\phi}_{2}&0&-w&0&0&0&1\\ 0&x_{21}^{*}C^{\phi}_{1}&-w^{*}&0&0&0&1&0\\ \hline\cr 0&-1&0&0&0&(S^{\phi}_{1}S^{\phi}_{2})^{*}&0&0\\ -1&0&0&0&S^{\phi}_{1}S^{\phi}_{2}&-C^{\phi}_{1}C^{\phi}_{2}&S^{\phi}_{1}C^{\phi}_{2}&S^{\phi}_{2}C^{\phi}_{1}\\ 0&0&0&1&0&(S^{\phi}_{1})^{*}C^{\phi}_{2}&0&-(S^{\phi}_{1})^{*}S^{\phi}_{2}\\ 0&0&1&0&0&(S^{\phi}_{2})^{*}C^{\phi}_{1}&-S^{\phi}_{1}(S^{\phi}_{2})^{*}&0\end{array}\right)$ (231)
where
$u$ $=\alpha\left\langle D^{2}\sigma_{1}\sigma_{2}\right\rangle_{D,\rho},$ (232)
$v$ $=\alpha\left\langle D^{4}|\sigma_{1}|^{2}|\sigma_{2}|^{2}\right\rangle_{D,\rho},$ (233)
$w$ $=\alpha\left\langle D^{2}\rho^{2}\sigma_{1}^{*}\sigma_{2}\right\rangle_{D,\rho},$ (234)
$x_{12}$ $=\alpha\left\langle D^{3}\rho\sigma_{1}|\sigma_{2}|^{2}\right\rangle_{D,\rho},$ (235)
$\sigma_{k}$ $=\frac{1}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (236)
Isolating the upper-left element of the inverse of this matrix yields Eq. 114.
Appendix J Numerical details
J.1 Numerics for theory
To validate the theory, we used networks with dynamics defined by $T[x](t)=(1+\partial_{t})x(t)$ (Eq. 2) and with nonlinearity $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$. For all connectivity models, except those with structured $\bm{L}$-$\bm{R}$ overlaps (VII), the single-site two-point functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ depend only on $g_{\text{eff}}$ and can be numerically calculated as in i.i.d. networks (see, e.g., [10, 77]). In summary:
  1. 1.
    Since $\eta(t)$ is Gaussian, so is $x(t)$ for this linear form of $T[\cdot]$, allowing us to write $C^{\phi}_{\star}(\tau)$ in terms of Gaussian integrals over $x(t)$ and $x(t+\tau)$, which have marginal variance $C^{\phi}_{\star}(0)$ and covariance $C^{\phi}_{\star}(\tau)$. Due to our choice of $\phi(x)$, this expression simplifies to
    $C^{\phi}_{\star}(\tau)=\frac{2}{\pi}\tan^{-1}\left[\frac{C^{x}_{\star}(\tau)}{\sqrt{\left(C^{x}_{\star}(0)+\frac{2}{\pi}\right)^{2}-C^{x}_{\star}(\tau)^{2}}}\right].$ (237)
  2. 2.
    Squaring the single-site picture gives a second-order ODE $\partial^{2}_{\tau}C^{x}_{\star}(\tau)=C^{x}_{\star}(\tau)-g_{\text{eff}}^{2}C^{\phi}_{\star}(\tau)$. Since the rhs depends only on $C^{x}_{\star}(\tau)$ and $C^{x}_{\star}(0)$, we can consider it to be the negative derivative $-\partial V/\partial C^{x}_{\star}(\tau)$ of a potential $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$ with an explicit dependence on the initial condition $C^{x}_{\star}(0)$.
  3. 3.
    The rhs is integrated w.r.t. $C^{x}_{\star}(\tau)$ to provide an expression for $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$, which due to our choice of $\phi(x)$ simplifies to
    $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))=-\frac{1}{2}C^{x}_{\star}(\tau)^{2}+g_{\text{eff}}^{2}\left[\frac{2}{\pi}\sqrt{\left(C^{x}_{\star}(0)+\frac{2}{\pi}\right)^{2}-C^{x}_{\star}(\tau)^{2}}+C^{x}_{\star}(\tau)C^{\phi}_{\star}(\tau)\right].$ (238)
  4. 4.
    Restricting to solutions with $C^{x}_{\star}(\tau)\rightarrow 0$ as $\tau\rightarrow\infty$ and $\partial_{\tau}C^{x}_{\star}(\tau)|_{\tau=0}=0$, enforcing conservation of energy gives $V(C^{x}_{\star}(0),C^{x}_{\star}(0))=V(0,C^{x}_{\star}(0))$, which can be solved for $C^{x}_{\star}(0)$ by numerical root finding.
  5. 5.
    Finally, Euler integration of the $[C^{x}_{\star}(\tau),\partial_{\tau}C^{x}_{\star}(\tau)]$ dynamics with these initial conditions gives its full time course, along with $C^{\phi}_{\star}(\tau)$ via Eq. 237.
We integrated the Newtonian ODE from $\tau=0$ to $\tau_{\text{max}}=200$ with step size $d\tau=0.025$. When then use an FFT to otain the frequency-space representation. The frequency-space representation of the response function $S^{\phi}_{\star}(\omega)$ was computed directly as
$S^{\phi}_{\star}(\omega)=\frac{\left\langle\phi^{\prime}\right\rangle_{\star}}{1+i\omega},$ (239)
with $\left\langle\phi^{\prime}\right\rangle_{\star}=1/\sqrt{1+(\pi/2)C^{x}_{\star}(0)}$. Various expressions for $\Psi^{a}_{\star}(\omega_{1},\omega_{2})$ were then computed in frequency space before transforming back to the time domain by a 2D inverse FFT.
For structured $\bm{L}$-$\bm{R}$ overlaps, the $\mathcal{O}(1)$ self-coupling in the single-site process makes $x(t)$ non-Gaussian, preventing the use of Gaussian integrals in the DMFT solution. We there obtain the solution via standard numerical methods, enforcing self-consistency among the $\eta^{x}(t)$ autocovariance function, $Q_{\star}(\tau)$; the self-coupling kernel, $R_{\star}(\tau)$; the response function, $S^{\phi}_{\star}(\tau)$; and the $\phi(t)$ autocovariance function, $C^{\phi}_{\star}(\tau)$. In summary:
  1. 1.
    Seed values for kernels are set as $C^{\phi}_{0}(\tau)=\exp[-(\tau/15)^{2}]+10^{-4}\delta(\tau)$ and $S^{\phi}_{0}(\tau)=0$.
  2. 2.
    $Q_{0}(\tau)$ and $R_{0}(\tau)$ are computed in Fourier space using the current values of $C^{\phi}_{0}(\tau)$ and $S^{\phi}_{0}(\tau)$ via Eq. 112, with the average taken over $(D,\rho)$.
  3. 3.
    A set of $N_{\text{sample}}=10^{4}$ instances of $\eta^{x}$ are sampled according to the Gaussian process kernel $Q_{0}(\tau)$ by standard spectral methods.
  4. 4.
    Each of $N_{\text{sample}}$ single-site processes $x_{s}(t)$ is Euler integrated according to Eq. 110 with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi(x_{s}(t))$.
  5. 5.
    These time series $\phi(x_{s}(t))$ and $\eta^{x}_{s}(t)$ are used to compute empirical values $(C^{\phi}_{1}(\tau),S^{\phi}_{1}(\tau))$ for the autocovariance and response. The latter is computed by the Furutsu-Novikov theorem: $S^{\phi}(\omega)=C^{\phi\eta}(\omega)/C^{\eta}(\omega)$. The numerator is estimated using these time series, while $C^{\eta}(\omega)$ is simply $Q_{0}(\omega)$.
  6. 6.
    After $N_{\text{iter}}=300$ repetitions of steps (1)-(5) with smoothed updates
    $C^{\phi}_{m}(\tau)\leftarrow 0.2C^{\phi}_{m-1}(\tau)+0.8C^{\phi}_{m}(\tau)$ (240)
    $S^{\phi}_{m}(\tau)\leftarrow 0.2S^{\phi}_{m-1}(\tau)+0.8S^{\phi}_{m}(\tau)$ (241)
    as new seed values for the upcoming $m$th iteration, the final numerical estimates for $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ are computed as the averages over the last $50$ iterations.
We used a temporal discretization $\tau_{\text{max}}=120$, $d\tau=0.04$. With these quantities, the various components of Eq. 114 and the expression for $\Psi^{\phi}_{\star}$ are straightforward to compute. We took $\left\langle\cdot\right\rangle_{D,\rho}$ averages in this case by averaging over every other discrete value of $D_{a}$ and $\rho_{a}$.
J.2 Simulation details
Each network was integrated via Runga-Kutta (RK4) with $dt=0.05$. The empirical lagged covariance matrix was computed by averaging over $N_{\text{loop}}$ individual covariance estimates, each of which is based on $T=2000$ discrete, evenly spaced time samples $t^{i_{\text{loop}}}_{k}$, with $t^{i_{\text{loop}}}_{k}-t^{i_{\text{loop}}}_{k-1}=1$:
$\widehat{C^{\phi}_{ij}(\tau)}=\left\langle\frac{1}{T-1}\sum_{k=1}^{T}\phi_{i}(t^{i_{\text{loop}}}_{k})\phi_{j}(t^{i_{\text{loop}}}_{k}-\tau)\right\rangle_{i_{\text{loop}}=1}^{N_{\text{loop}}}.$ (242)
The value of $N_{\text{loop}}$ varied but was generally $\geq 50$ and always chosen to ensure saturated dimension values, i.e. $TN_{\text{loop}}\gg N$. Specific quantities of interest, such as $\text{PR}^{\phi}$ or $\Psi^{\phi}(\tau,0)$, were computed using this estimate.

### v2

@@
 Abstract
-We develop a theory to analyze how structure in connectivity shapes the high-dimensional, internally generated activity of nonlinear recurrent neural networks.
-Using two complementary methods – a path-integral calculation of fluctuations around the saddle point, and a recently introduced two-site cavity approach – we derive analytic expressions that characterize important features of collective activity, including its dimensionality and temporal correlations.
-To model structure in the coupling matrices of real neural circuits, such as synaptic connectomes obtained through electron microscopy, we introduce the random-mode model, which parameterizes a coupling matrix using random input and output modes and a specified spectrum.
-This model enables systematic study of the effects of low-dimensional structure in connectivity on neural activity.
-These effects manifest in features of collective activity, that we calculate, and can be undetectable when analyzing only single-neuron activities.
-We derive a relation between the effective rank of the coupling matrix and the dimension of activity.
-By extending the random-mode model, we compare the effects of single-neuron heterogeneity and low-dimensional connectivity.
-We also investigate the impact of structured overlaps between input and output modes, a feature of biological coupling matrices.
-Our theory provides tools to relate neural-network architecture and collective dynamics in artificial and biological systems.
+Studies of the dynamics of nonlinear recurrent neural networks often assume independent and identically distributed couplings, but large-scale connectomics data indicate that biological neural circuits exhibit markedly different connectivity properties.
+These include rapidly decaying singular-value spectra and structured singular-vector overlaps.
+Here, we develop a theory to analyze how these forms of structure shape high-dimensional collective activity in nonlinear recurrent neural networks.
+We first introduce the random-mode model, a random-matrix ensemble related to the singular-value decomposition that enables control over the spectrum and right-left mode overlaps.
+Then, using a novel path-integral calculation, we derive analytic expressions that reveal how connectivity structure affects features of collective dynamics: the dimension of activity, which quantifies the number of high-variance collective-activity fluctuations, and the temporal correlations that characterize the timescales of these fluctuations.
+We show that connectivity structure can be invisible in single-neuron activities while dramatically shaping collective activity.
+Furthermore, despite the nonlinear, high-dimensional nature of these networks, the dimension of activity depends on just two connectivity parameters—the variance of the couplings and the effective rank of the coupling matrix, which quantifies the number of dominant rank-one connectivity components.
+We contrast the effects of single-neuron heterogeneity and low-dimensional connectivity, making predictions about how $z$-scoring data affects the dimension of activity.
+Finally, we demonstrate the presence of structured overlaps between left and right modes in the Drosophila connectome, incorporate them into the theory, and show how they further shape collective dynamics.
 I Introduction
-Nonlinear recurrent neural networks are used in neuroscience as models of neural-circuit dynamics [1, 2, 3] and in machine learning as systems for sequence processing [4, 5, 6, 7].
-Such networks can generate complex, time-varying activity in the absence of external input [8, 9].
-This high-dimensional chaotic activity, first analyzed by Sompolinsky et al.
-[10], has served as a model for asynchronous cortical dynamics observed in vivo.
-Moreover, initializing recurrent neural networks so that they exhibit chaotic dynamics has been shown to facilitate subsequent learning [11, 12, 13].
-Theoretical studies often focus on large recurrent neural networks with independent and identically distributed (i.i.d.) couplings.
-However, the structure of real-world networks, including neural circuits, typically deviates substantially from this assumption [14].
-Many real-world networks show approximate low-rank structure, meaning that their connectivity is well described by a number of rank-one components that is smaller than the number of units in the network [15].
-Often, networks have “smooth low-rank” or “effectively low-rank” structure in which the strengths of these rank-one components do not exhibit a hard cutoff and instead decay smoothly, with a more rapid falloff than in networks with i.i.d.
+The collective activity of a high-dimensional nonlinear system is determined by the structure of the interactions between its elements, but the mapping from structure to activity is generally only analytically tractable for limited forms of structure.
+In systems neuroscience, structure in synaptic connectivity controls the function of a neural system by determining what patterns of activity are produced by populations of neurons.
+This paper studies this mapping in the context of nonlinear recurrent neural networks, used widely in neuroscience as models of neural-circuit dynamics [1, 2, 3] and in machine learning as systems for sequence processing [4, 5, 6, 7].
+While these models omit many features of actual neural circuits, they capture several of their fundamental features, including their large scale, nonlinear units, and recurrent interactions.
+Functionally, a key property of such systems is that, like real neural circuits, they can generate rich time-varying activity in the absence of external input [8, 9].
+Theoretical studies of nonlinear recurrent neural networks, including the seminal work of Sompolinsky et al.
+[10], often focus on the case of independent and identically distributed (i.i.d.) couplings, or simple variants thereof [11].
+Such disordered networks produce high-dimensional chaotic activity, reminiscent of asynchronous cortical activity observed in vivo [10].
+Initializing recurrent neural networks so that they exhibit chaotic dynamics has also been shown to facilitate subsequent learning [12, 13, 14, 15, 16].
+Experimental access to large-scale connectivity structure creates an opportunity and imperative to develop theories that explain models beyond those with i.i.d.
 couplings.
-This can arise when connections between units depend on their distance in physical space, or in a low-dimensional feature space [16, 17, 18].
-Such low-rank structure can also emerge when artificial neural networks are trained on tasks [19, 20].
-To illustrate this, we examined a central-brain connectome of the fruit fly Drosophila melanogaster [21] (Fig.
-1a).
-We estimated strengths from synaptic counts, assigned coupling signs based on predicted neurotransmitter types [22], and normalized the couplings (Fig.
-1b; Appendix B).
-We analyzed the network’s structure using its singular value decomposition (SVD), which expresses the connectivity as a sum of rank-one components, with the strength of each component given by the corresponding singular value.
-The singular values decay much more quickly than those of an i.i.d.
-matrix of the same size, which follow a universal form that does not depend on details of the single-element distribution (Fig.
-1c).
-We quantify this decay rate using the participation ratio of the squared singular-value spectrum (Sec.
-II.4.1).
-For this connectome, the participation ratio is $0.18$, much lower than the value of $0.5$ for an i.i.d.
-matrix, demonstrating that this network has a much lower-rank structure than would be expected for i.i.d.
-couplings.
+Recently, large-scale synaptic reconstructions of neural circuits, or “connectomes,” have become available in multiple species, including the full brain of Drosophila melanogaster [17] and increasingly large portions of mammalian circuits.
+These large-scale connectomes reveal that neural circuits deviate substantially from the i.i.d.
+assumption, as previously demonstrated via the presence of small-scale motifs [18].
+In particular, in accordance with real-world networks more generally [19], neural circuits appear to exhibit low-rank structure in which the connectivity is well-described by fewer rank-one components than the network size (we demonstrate this explicitly for the Drosophila connectome in Sec.
+II.1).
+In neural-network models, recurrent or otherwise, such structure emerges naturally in several contexts, including when connections depend on distance in physical or feature space [20, 21, 22], or when networks are trained on tasks with low-dimensional structure [23, 24, 25].
+In parallel with the development of large-scale connectivity maps, recent advances in recording technologies including silicon probes [26] and large-scale calcium imaging [27] now enable the activities of many neurons to be monitored simultaneously, with the latest datasets including hundreds of thousands of neurons.
+This technological progress has enabled researchers to characterize collective properties of neural activity rather than focusing solely on single-neuron responses.
+These collective properties capture how neurons interact and reveal distributed computational processes that are only visible with high-yield recording technologies [28].
+Such population-level features are typically analyzed using dimensionality reduction techniques like principal component analysis [29].
+Relating connectivity and collective-activity structure analytically.
+We currently lack analytical tools to predict relationships between collective activity described by large-scale recordings, and connectivity structure described by connectomics or other datasets.
+Dynamical mean-field theory (DMFT) is widely used to analytically characterize activity in large, nonlinear recurrent neural networks.
+Although recent advances in DMFT have extended these techniques to describe collective features in networks producing high-dimensional chaotic activity [30], these calculations have been limited to i.i.d.
+connectivity or connectivity with simple correlations between reciprocal couplings.
+Extending such calculations to more complicated connectivity structures rapidly becomes unwieldy.
+Consequently, such analyses have not been performed for networks whose connections are constrained by the statistics of connectomic or other datasets, limiting our ability to leverage these datasets for theoretical insight.
+One tractable approach has been to study networks with very low-dimensional connectivity and thus activity (specifically, in the limit where the rank of the coupling matrix remains finite while the network becomes large; see Sec.
+II.3).
+However, connectomes, like that of Drosophila considered in Sec.
+II.1, do not support this finite-rank assumption.
+Furthermore, while neural activity recorded in experiments is often low-dimensional, this may be inherited from the low dimensionality of experimental tasks; in more complicated tasks or spontaneous states, activity is typically higher-dimensional [31, 27].
+To address these limitations and challenges, this paper makes two contributions:
+  1. 1.
+    A tractable model of neural-network coupling matrices that we call the random-mode model.
+Motivated by the Drosophila connectome considered in Sec.
+II.1, we introduce the random-mode model in Sec.
+II.2, in which couplings are generated as a sum of rank-one outer products of random left and right modes with component-specific strengths.
+Unlike well studied low-rank recurrent neural network models, we scale the rank with the network size (Sec.
+II.3).
+We show in Sec.
+II.4 that parameterizing these strengths provides control over the singular value spectrum of the coupling matrix.
+  2. 2.
+    A new path-integral calculation of a specific four-point function of network activity, originally defined by Clark et al.
+[30], that captures features of collective activity including its dimensionality.
+We review these and other summary statistics in Sec.
+III.
+In Sec.
+IV A–C, we describe the calculation, based on fluctuations around the saddle point in a path integral, and apply it to the random-mode model.
+In Sec.
+IV.4, we use the solution to study how features of the coupling matrix control collective activity.
+Additionally, in Sec.
+V, we analyze notable limiting cases of the random-mode model.
+In Sec.
+VI, we study a generalization of the model featuring heterogeneity among single-neuron properties and contrast these effects with those of low-rank structure in connectivity.
+Finally, in Sec.
+VII, we incorporate structured overlaps between left and right modes into the model, demonstrating their presence in the Drosophila connectome and showing how they further shape collective dynamics.
+II Random-mode model
+II.1 Spectral structure of neural-circuit connectivity
 Figure 1: Analysis of Drosophila central brain connectome.
@@
 $N=18028$ neurons.
-In neuroscience, the importance of connectivity lies in its role in controlling neural activity.
-To this end, dynamical mean-field theory (DMFT) has been used to analyze activity in large recurrent neural networks, focusing on single-neuron properties such as the autocovariance function.
-However, recent neuroscience studies have highlighted the significance of geometric or other collective properties that emerge only when examining the simultaneous activity of large neuronal populations [23, 24].
-Given neural data, these properties are often analyzed using dimensionality reduction techniques like principal component analysis.
-Recent advances have extended DMFT techniques to analytically characterize these collective features in nonlinear network models producing high-dimensional chaotic activity [25].
-Other studies have addressed simpler cases, including linear networks driven by noise [26, 27, 28, 29] and nonlinear networks producing low-dimensional activity [30, 31].
-A key quantity of interest is the dimension of activity, which quantifies the number of high-variance activity modes.
-This metric reflects a network’s capacity for complex computations and representations, and influences its learning and generalization capabilities [32, 33].
-Despite these advances, the analytical calculation of activity dimension in nonlinear networks has been limited to i.i.d.
-connectivity, or connectivity with simple correlations between reciprocal synapses, and thus cannot be performed for networks with more realistic structures (Fig.
-1c).
-To address this gap, we study networks with smooth low-rank parameterizations of couplings, and develop DMFT tools to analyze collective activity in such networks.
-We show that, in certain cases, the effects of structure in connectivity are not evident in single-neuron activity, but instead revealed through dramatic changes in properties of collective activity that we show how to compute.
-This paper has two main contributions.
-The first is a new path-integral calculation of a specific four-point function of network activity, originally defined by Clark et al.
-[25].
-This function captures statistics describing collective activity, including its dimensionality.
-In Secs.
-II.1II.3, we describe the recurrent-network model and define this four-point function.
-We explain its connection to collective activity features, including the activity dimension and timescales of principal components, and review its calculation in i.i.d.
-networks.
-In Secs.
-III.1III.4, we introduce the new path-integral calculation and apply it to both the i.i.d.
-model and a model with correlations between reciprocal pairs of synapses.
-The second contribution is a tractable model of neural-network coupling matrices that we call the random-mode model.
-In Sec.
-II.4, we define this model, in which couplings are generated as a sum of rank-one outer products of random input and output modes with component-specific strengths.
-In Sec.
-II.4.1, we demonstrate that parameterizing these strengths provides control over the singular value spectrum of the coupling matrix, enabling investigation of the interplay between global connectivity structure and network dynamics.
-In Sec.
-III.5, we calculate the two- and four-point functions for the random-mode model using the path-integral approach.
-This analysis reveals a bipartite coupling between neurons and latent variables.
-In Sec.
-III.6, we interpret the resulting formulas for the four-point function, which show how the effective rank of the coupling matrix controls the dimension of activity.
-In Sec.
-IV, we analyze limiting cases of the random-mode model, illustrating that a vanishing effective rank leads to vanishing activity dimension.
-In Sec.
-V, we show that the same results can be obtained using an intuitive two-site cavity method.
-In Sec.
-VI, we study the behavior of a generalization of the random-mode model featuring heterogeneity among single-neuron properties and contrast the effects with those of low-rank structure in the connectivity.
-Finally, in Sec.
-VII, we incorporate a simple form of overlaps between input and output modes into the model, which we anticipate to be crucial for describing neural computations.
-II Model and summary statistics
-II.1 Recurrent neural network model
+Let $\bm{J}$ denote an $N\times N$ weighted, directed coupling matrix among $N$ neurons.
+Many theoretical studies have examined single-neuron [10] and collective [30] properties of activity in nonlinear recurrent neural networks with i.i.d.
+$J_{ij}$.
+However, the coupling matrices of real-world networks, including neural circuits, often exhibit approximate low-rank structure that deviates dramatically from the i.i.d.
+assumption.
+To illustrate this, we analyzed a central-brain connectome of the fruit fly Drosophila melanogaster using the singular value decomposition (SVD), which decomposes any matrix as a sum of rank-one components, with each component’s strength given by the corresponding singular value (Fig.
+1a; details in Appendix B).
+Specifically, the SVD decomposes $\bm{J}$ in the form
+$\bm{J}=\sum_{a=1}^{M}S_{a}\bm{u}_{a}\bm{v}_{a}^{T},$ (1)
+where, for $a=1,\ldots,M$, $S_{a}>0$ are the singular values; $\bm{u}_{a}$ and $\bm{v}_{a}$ are $N$-dimensional left and right singular vectors, respectively; and $M$ is the rank of $\bm{J}$.
+The ranked singular values of the Drosophila connectome decay much more rapidly than those of an i.i.d.
+matrix of the same size, which follow a universal distribution independent of single-element statistics (Fig.
+1c; similar analysis appears in Thibeault et al.
+[19]).
+We quantify this rapid decay using the participation ratio of the squared singular-value spectrum, $\text{PR}^{S}=(\sum_{a=1}^{M}S_{a}^{2})^{2}/\sum_{a=1}^{M}S_{a}^{4}$.
+The connectome’s participation ratio is 0.18, substantially lower than the 0.5 value characteristic of i.i.d.
+matrices, confirming that this biological network exhibits much more concentrated low-rank structure than would arise from random and independent couplings.
+II.2 Definition of the random-mode model
+Both the singular-value spectrum and, as we will show later, the overlaps of the left and right singular vectors of the Drosophila connectome deviate from the predictions of an i.i.d.
+coupling matrix.
+We therefore introduce the random-mode model, a generative model for coupling matrices that allows us to capture these properties within a random-matrix ensemble.
+The random-mode model has the same mathematical form as an SVD, but is a statistical generative process for $\bm{J}$ rather than a linear algebraic decomposition of it.
+The model generates a coupling matrix of the form
+$\bm{J}=\sum_{a=1}^{M}D_{a}\bm{\ell}_{a}\bm{r}_{a}^{T},$ (2)
+where, for each $a=1,\ldots,M$, $D_{a}>0$ are the component strengths; $\bm{r}_{a}$ and $\bm{\ell}_{a}$ are $N$-dimensional left and right modes, respectively; and $M$ is the number of components, where $\text{rank}=\text{min}(N,M)$.
+We now describe the properties of these quantities.
+Component strengths.
+The component strengths $D_{a}$ are analogous to the singular values $S_{a}$ of $\bm{J}$.
+We model these quantities deterministically with the requirement that they are defined for arbitrarily large $M$ and that empirical averages $M^{-1}\sum_{a=1}^{M}f(D_{a})$ converge to limiting values, denoted $\left\langle f(D)\right\rangle_{D}$, as $M\to\infty$.
+Mode vectors.
+The left and right modes $\bm{\ell}_{a}$ and $\bm{r}_{a}$ are analogous to the left and right singular vectors $\bm{u}_{a}$ and $\bm{v}_{a}$ of $\bm{J}$.
+We assume they are sampled i.i.d.
+across the neuron index $i$:
+$P(\{\bm{\ell}_{a},\bm{r}_{a}\}_{a=1}^{M})=\prod_{i=1}^{N}P(\{\ell_{ai},r_{ai}\}_{a=1}^{M}),$ (3)
+where $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ is the joint distribution over the $2M$ mode components, which must be specified. We assume that this distribution has the following moments fixed:
+$\left\langle\ell_{a}\right\rangle=\left\langle r_{a}\right\rangle=0,$ (4)
+$\left\langle\ell_{a}\ell_{b}\right\rangle=\left\langle r_{a}r_{b}\right\rangle=\frac{1}{N}\delta_{ab}.$ (5)
+This yields orthonormality of mode vectors in expectation:
+$\begin{split}\left\langle\bm{\ell}_{a}^{T}\bm{\ell}_{b}\right\rangle&=\delta_{ab},\\ \left\langle\bm{r}_{a}^{T}\bm{r}_{b}\right\rangle&=\delta_{ab},\end{split}$ (6)
+in analogy to singular vectors for which orthonormality is deterministic  [^1] This random structure, rather than exact, deterministic orthonormality, offers analytical tractability by eliminating the need to enforce orthonormality—a complicated global constraint.
+But see [82], who handle random orthonormal matrices in a regression setting..
+Individual mode components are $\mathcal{O}(1/\sqrt{N})$.
+Note that we have not constrained second moments that mix left and right mode components.
+Figure 2: Schematic of the random-mode model.
+Upper: Couplings $\bm{J}$ are generated as a sum of outer products, ${\bm{\ell}}_{a}{\bm{r}}^{T}_{a}$, with component strengths $D_{a}$.
+Lower: The two-point function $C^{\phi}_{\star}(\tau)$ and four-point function $\Psi^{\phi}_{\star}(\tau)$ are calculated in terms of the statistics of $D_{a}$.
+The two-point function depends only on the effective gain $g_{\text{eff}}$, while the four-point function depends on both $g_{\text{eff}}$ and $\text{PR}^{D}$, the effective dimension of the connectivity determined by the $D_{a}$ distribution.
+II.3 Rank scaling and comparison to low-rank recurrent neural networks
+A crucial modeling choice is how the number of modes, $M$, scales relative to the network size, $N$.
+The two main possibilities are $M=\mathcal{O}(N)$ or $M=\mathcal{O}(1)$, corresponding to extensive and intensive rank scaling, respectively.
+The random-mode model is characterized by extensive rank scaling, where both $M$ and $N$ approach infinity while maintaining a fixed ratio:
+$\alpha=\frac{M}{N}=\mathcal{O}(1).$ (7)
+By choosing $\alpha$ to be small, or by choosing the spectrum of ranked component strengths to decay rapidly, we can construct connectivity that is low-rank, or well-described by a small number of rank-one components, compared to the number of neurons.
+The alternative approach of intensive rank scaling, where $M$ remains finite while $N\to\infty$, leads to what has been termed “low-rank recurrent neural networks” in the literature [33, 34, 35, 36].
+In some instances, the intensive-rank component is added to an i.i.d.
+matrix modeling unstructured “background” connectivity [33, 34].
+Note that in this paper, when we use “low-rank” or “low-dimensional,” we refer to quantities that are small as a fraction of $N$, even if they are extensive; this differs from some prior works (such as those on low-rank recurrent neural networks) where these terms refer to intensive quantities.
+The distinction between extensive and intensive rank scaling has important implications for specifying the distribution over the $2M$ mode components $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$.
+For finite $M$, one can directly specify this distribution.
+At large $N$, such a distribution encodes the relative geometry of the left and right modes, since the inner products between modes converge to ($N$ times) the second moments of $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$, with negligible fluctuations.
+A common choice is a $2M$-dimensional multivariate Gaussian distribution.
+An important insight from prior work on the intensive-rank case is that these overlaps are crucial for shaping the dynamics and computations implemented by the network, as overlaps between connectivity modes dictate how activity modes interact [33, 34, 35, 36].
+Our goal is to go beyond the intensive-rank setup, which produces correspondingly low-dimensional activity, to model more sophisticated, higher-dimensional dynamics.
+However, specifying the full distribution $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ in the random-mode model, where $M\to\infty$, is more complicated.
+First, it must be well-defined for all $M$ to allow taking the limit.
+More fundamentally, we would like to use the random-mode model to construct generative models that capture certain essential spectral features of coupling matrices, over which we have systematic control.
+Using a joint Gaussian distribution would require specifying $M^{2}$ parameters which, since $M=\mathcal{O}(N)$, is comparable to defining the full $N\times N$ coupling matrix.
+For this reason, this paper specializes to a case where $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ factorizes across the mode index $a$:
+$P(\{\ell_{a},r_{a}\}_{a=1}^{M})=\prod_{a=1}^{M}P(\ell_{a},r_{a}),$ (8)
+where $P(\ell_{a},r_{a})$ must be specified.
+Moreover, until Sec.
+VII, we further specialize to the case where the left and right modes are fully independent: $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$, with each marginal factor $P(\ell_{a})$ or $P(r_{a})$ being a univariate Gaussian with mean zero and variance $1/N$  [^2] The Gaussian assumption for individual components is not crucial—our results depend only on first and second moments for large $N$ due to the central limit theorem; higher-order cumulants (scaling appropriately with $1/N$) are subdominant in the path integral saddle-point approximation..
+In this case, all $2MN$ components specifying the mode vectors are i.i.d.
+with mean zero and variance $1/N$; we refer to this case as “i.i.d.
+modes.” In Sec.
+VII we use a non-factorized form of $P(\ell_{a},r_{a})$ to specify correlations between $\ell_{a}$ and $r_{a}$, introducing $M$ deterministic parameters (in addition to the $M$ component strengths) to specify the left–right correlation for each $a$.
+II.4 Relationship to singular value decomposition
+In the SVD, the left and right singular vectors are orthonormal: $\bm{u}_{a}^{T}\bm{u}_{b}=\bm{v}_{a}^{T}\bm{v}_{b}=\delta_{ab}$.
+In contrast, the mode vectors are orthonormal only in expectation (Eq.
+6); once sampled, there are random $\mathcal{O}(1/\sqrt{N})$ overlaps:
+$\begin{split}\bm{\ell}_{a}^{T}\bm{\ell}_{b}&=\delta_{ab}+\mathcal{O}(1/\sqrt{N}),\\ \bm{r}_{a}^{T}\bm{r}_{b}&=\delta_{ab}+\mathcal{O}(1/\sqrt{N}).\end{split}$ (9)
+When the effective rank (defined below) is small, the random-mode model “looks like” the SVD at large $N$ since the effects of the $\mathcal{O}(1/\sqrt{N})$ deviations from orthonormality in the sampled modes are negligible.
+Indeed, a well-known result in high-dimensional geometry is that finite numbers of random vectors with i.i.d.
+components approximate orthonormal bases at large $N$ [38].
+When the effective rank is not small, overlaps between different modes produce a discrepancy between the component strengths $D_{a}$ and singular values $S_{a}$.
+We now characterize this discrepancy analytically.
+We denote the $n$-th moments and participation ratio of the component-strength distribution by
+$r_{n}$ $=\left\langle D^{n}\right\rangle_{D},$ (10)
+$\text{PR}^{D}$ $=\frac{r_{2}^{2}}{r_{4}},$ (11)
+from which we define the “effective rank” of the couplings:
+$\text{effective rank}=\alpha\text{PR}^{D}.$ (12)
+The effective rank is related to the participation ratio of the squared singular value spectrum  [^3] This follows from using Wick’s theorem to evaluate the expectations of $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ and $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ for the numerator and denominator, respectively, of $\text{PR}^{S}$.
+through
+$\text{PR}^{S}=\frac{\alpha\text{PR}^{D}}{1+2\alpha\text{PR}^{D}}.$ (13)
+Expanding in small $\alpha\text{PR}^{D}$, $\text{PR}^{S}=\alpha\text{PR}^{D}+\mathcal{O}((\alpha\text{PR}^{D})^{2})$, demonstrating that the component strengths are closely related to the singular values in the low-dimensional regime.
+While we express the analytic results in this paper in terms of $\alpha\text{PR}^{D}$, the above equation can always be used to translate between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$.
+We further quantify this discrepancy in the case where $D_{a}=1$ for all $a$.
+Using methods from free probability theory [40, 41] (Appendix C), the $M$ singular values are distributed over a range with boundaries
+$S_{\pm}=\sqrt{1+\frac{5\alpha}{2}-\frac{\alpha^{2}}{8}\pm\left(1+\frac{\alpha}{8}\right)^{3/2}\sqrt{8\alpha}}.$ (14)
+For small $\alpha$, $S_{\pm}=1\pm\sqrt{2\alpha}+\mathcal{O}(\alpha)$.
+While the random-mode model spreads the nonzero singular values over a range, this spread becomes negligible for small $\alpha$.
+More generally, the distributions of $D_{a}$ and $S_{a}$ coincide for small $\alpha\text{PR}^{D}$.
+Having presented a generative model of connectivity, we now present the network model that transforms this connectivity into neural activity.
+We then define summary statistics that characterize the structure of this activity, allowing us to analyze its dependence on the connectivity.
+III Network model and summary statistics
+III.1 Recurrent neural network model
 We study a recurrent neural network of $N$ neurons.
-Each neuron is characterized by its pre-activation $x_{i}(t)$ and activation $\phi_{i}(t)$, where $i=1,\ldots,N$.
-The activations are related to the pre-activations through a nonlinear function $\phi(x)$, with $\phi_{i}(t)=\phi(x_{i}(t))$.
-While our results are agnostic to the form of the nonlinearity, when comparing theory and simulations we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$, which is similar to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals.
+Each neuron $i\in\{1,\ldots,N\}$ is characterized by its pre-activation $x_{i}(t)$ and activation $\phi_{i}(t)=\phi(x_{i}(t))$, where $\phi(\cdot)$ is a scalar nonlinearity; we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$  [^4] This behaves similarly to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals..
 The network dynamics are governed by
-$T[x_{i}](t)=\sum_{j}J_{ij}\phi_{j}(t),$ (1)
-where $J_{ij}$ is the coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional describing single-neuron dynamics. A canonical choice is
-$T[x](t)=(1+\partial_{t})x(t),$ (2)
-resulting in low-pass filtering of the input. We measure time in units of the time constant that would appear in front of $\partial_{t}$.
-II.2 Correlation functions
-II.2.1 Two-point functions
-To characterize the network dynamics, we introduce several types of two-point correlation functions.
-Let ${a\in\{x,\phi\}}$ denote either the pre-activations or activations.
-We first define the neuron- and time-averaged covariance:
-$C^{a}(\tau)=\frac{1}{N}\sum_{i}\left\langle a_{i}(t)a_{i}(t+\tau)\right\rangle_{t}.$ (3)
-We also define the time-averaged, but not neuron-averaged, covariance; and the neuron-averaged, but not time-averaged, covariance:
+$T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t),\>\>T[x](t)=(1+\partial_{t})x(t),$ (15)
+where $J_{ij}$ denotes the synaptic coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional that specifies the single-neuron dynamics.
+We expect our results to be agnostic to the specific choice of $T[\cdot]$.
+We next introduce correlation functions that capture single-neuron and collective activity properties.
+III.2 Two-point functions and duality of neuronal and temporal covariances
+Figure 3: Duality of neuron-by-neuron and time-by-time covariances and its relation to the dimension of activity.
+Both plots are based on the same simulation of a network of $N=2500$ neurons with $g=2.25$ and i.i.d.
+couplings.
+The dimension can be computed either by computing the statistics of the off-diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
+Let $a\in\{x,\phi\}$ denote either pre-activations or activations. We define two complementary covariance functions, averaging over either time or neurons:
 $C^{a}_{ij}(\tau)$ $=\left\langle a_{i}(t)a_{j}(t+\tau)\right\rangle_{t},$ (16)
-$C^{a}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i}a_{i}(t_{1})a_{i}(t_{2}).$ (5)
-For a given $\tau$, the spectra of $C^{a}_{ij}(\tau)$, with indices $i$ and $j$, and of $C^{a}(t_{1},t_{2}+\tau)$, with indices $t_{1}$ and $t_{2}$, are the same (up to a scaling), so spectral properties of one parameter can be computed using the other, a form of duality.
-For example, the properly scaled trace operations applied to these parameters yield the same result, namely, $C^{a}(\tau)$.
-We also define the response function:
-$S^{a}_{ij}(t_{1},t_{2})=\left.\frac{\delta a_{i}(t_{1})}{\delta I_{j}(t_{2})}\right|_{I_{j}=0},$ (6)
-where $I_{j}(t)$ is an infinitesimal source term added to the right-hand side of the network dynamics (Eq.
-1).
-In analogy with the covariance functions, we define $S^{a}_{ij}(\tau)$, $S^{a}(t_{1},t_{2})$, and $S^{a}(\tau)$ to be the response function averaged over time, neurons, or both.
-II.2.2 Four-point functions
-Figure 2: Relationship between principal component timescales and two- and four-point correlation functions.
-(a) Example activity traces of principal components $p^{\phi}_{k}(t)$.
-(b) Empirical autocovariance functions for each principal component, weighted by their eigenvalues.
-(c) Same as (b) but weighted by the squared eigenvalue.
-(d) Averages over curves in (b) and (c), with comparisons to $C^{\phi}_{\star}(\tau)$ and $\Psi^{\phi}_{\star}(\tau,0)$.
-Dots: simulation results.
-Lines: theoretical predictions.
-All panels describe the same network of $N=4000$ neurons with i.i.d.
-couplings and $g=6$.
-To characterize the dimension and timescales of activity, we define
-$\Psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i,j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}),$ (7)
-where $\bm{\tau}=(\tau_{1},\tau_{2})$. This sum includes the on-diagonal contributions, unlike the quantity computed in Clark et al. [25], which we denote by the lower-case $\psi^{a}(\bm{\tau})$:
-$\psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i\neq j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}).$ (8)
-For the type of connectivity we will study, the network exhibits high-dimensional activity such that pairwise covariances are $\mathcal{O}(1/\sqrt{N})$.
-Thus, in the definition of $\Psi^{a}(\bm{\tau})$ (Eq.
-7), the on-diagonal contributions (given by $N^{-1}\sum_{i}C^{a}_{ii}(\tau_{1})C^{a}_{ii}(\tau_{2})$) and off-diagonal contributions (given by $\psi^{a}(\bm{\tau})$) are both $\mathcal{O}(1)$.
-Leveraging neuron-time duality, $\Psi^{a}(\bm{\tau})$ can also be expressed in terms of the time-by-time covariance $C^{a}(t_{1},t_{2})$:
+$C^{a}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}a_{i}(t_{1})a_{i}(t_{2}).$ (17)
+These functions exhibit a form of duality: for any lag $\tau$, the matrix $\bm{C}^{a}(\tau)$ with elements $C^{a}_{ij}(\tau)$ indexed by $(i,j)$ and the matrix with elements $C^{a}(t_{1},t_{2}+\tau)$ indexed by $(t_{1},t_{2})$ have the same eigenvalue spectrum up to scaling (to see this, note that both matrices can be computed from a time-by-neuron activity matrix $\bm{A}$ as $\bm{A}^{T}\bm{A}$ and $\bm{A}\bm{A}^{T}$, respectively, which have identical spectra)  [^5] This duality is reminiscent of ergodicity theorems, which relate temporal and spatial averages, but differs in a variety of ways, including that it holds even for finite time intervals, whereas ergodicity concerns infinite-time limits..
+We also define the neuron-averaged response function, describing propagation of infinitesimal perturbations:
+$S^{a}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\left.\frac{\delta a_{i}(t_{1})}{\delta I_{j}(t_{2})}\right|_{I_{j}=0},$ (18)
+where $I_{j}(t)$ is a source term added to the right-hand side of the network dynamics: $T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t)+I_{i}(t)$. Finally, we defined “stationarized” counterparts:
+$C^{a}(\tau)=\left\langle C^{a}(t,t\pm\tau)\right\rangle_{t},\>\>\>\>S^{a}(\tau)=\left\langle S^{a}(t,t-\tau)\right\rangle_{t}.$ (19)
+III.3 Four-point functions, dimension of activity, and principal-component timescales
+In contrast to the above two-point functions, which characterize single-neuron activity, four-point functions describe features of collective activity.
+The same four-point functions were used in [30].
+Here, we motivate these functions in more detail and express them in two ways via the duality between neuronal and temporal covariances.
+To motivate their definition, consider the dimension of activity, which can be quantified as the participation ratio of the spectrum of the equal-time covariance matrix $\bm{C}^{a}(0)$ [44, 45, 29, 46, 47, 48, 9, 49, 50].
+Given the eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$ for $k=1,\ldots,N$ (as one would compute as part of principal components analysis), the participation ratio is
+$\text{PR}^{a}=\frac{1}{N}\frac{(\sum_{k=1}^{N}\lambda^{a}_{k})^{2}}{\sum_{k=1}^{N}(\lambda^{a}_{k})^{2}}.$ (20)
+To see why $\text{PR}^{a}$ provides a meaningful measure of dimension, consider the case where $D$ eigenvalues equal a positive constant and the remaining eigenvalues are zero.
+Then, $\text{PR}^{a}=D/N$.
+More generally, when the spectrum exhibits a smooth decay rather than a hard cutoff, the participation ratio identifies the characteristic decay scale (divided by $N$) [46].
+We compare the participation ratio to alternative measures of effective dimensionality in Appendix D.
+Without the normalization factor $1/N$, this quantity would vary between $1$ and $N$; with normalization, it varies between $1/N$ and $1$.
+In the limit $N\to\infty$, this quantity could have three qualitatively different behaviors:
+  • Subextensive dimensionality: $\text{PR}^{a}=0$. The number of dimensions filled by activity grows sublinearly with $N$.
+  • Nontrivial extensive dimensionality: $0<\text{PR}^{a}<1$. The number of dimensions filled by activity grows linearly with $N$ and not all dimensions are filled equally.
+  • Trivial extensive dimensionality: $\text{PR}^{a}=1$. All dimensions are filled equally by activity.
+An advantageous property of the participation ratio is that it can be expressed as the ratio of the squared-trace and Frobenius norm of $\bm{C}^{a}(0)$, both of which can be further expressed in terms of matrix elements $C_{ij}^{a}(0)$:
+$\text{PR}^{a}$ $=\frac{\left(\text{tr}\bm{C}^{a}(0)\right)^{2}}{\lVert\bm{C}(0)\rVert_{F}^{2}}$
+$=\frac{C^{a}(0)^{2}}{\frac{1}{N}\sum_{i=1}^{N}C_{ii}(0)^{2}+\frac{1}{N}\sum_{i\neq j}C_{ij}(0)^{2}},$ (21)
+where $C^{a}(0)$ is given by Eq.
+19.
+Assuming that diagonal elements are uniform with negligible fluctuations, i.e., $C^{a}_{ii}(0)=C^{a}(0)$ for all $i$ (as occurs, for example, under i.i.d.
+connectivity [10]), this simplifies to
+$\text{PR}^{a}=\frac{C^{a}(0)^{2}}{C^{a}(0)^{2}+\psi^{a}(0,0)},$ (22)
+where, following [30], we define the four-point function
+$\psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i\neq j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}),$ (23)
+where $\bm{\tau}=(\tau_{1},\tau_{2})$.
+The three scaling behaviors of dimensionality outlined above correspond in the limit $N\to\infty$ to $\psi^{a}(0,0)\to\infty$}, $\psi^{a}(0,0)\to\text{const.}$}, or $\psi^{a}(0,0)\to 0$}, respectively.
+Since $\psi^{a}(0,0)$ is given by a sum over $\mathcal{O}(N^{2})$ squared cross-covariances times $1/N$, its magnitude is $N$ times that of an individual squared cross-covariance.
+Both i.i.d.
+matrices and the extensive-rank matrices generated by the random-mode model lead to $\mathcal{O}(1/\sqrt{N})$ cross-covariances.
+Thus, $\psi^{a}(0,0)\to\text{const.}$ as $N\to\infty$, leading to nontrivial extensive dimensionality, $0<\text{PR}^{a}<1$.
+In this paper, we focus on a related function $\Psi^{a}(\bm{\tau})$ that includes the diagonal terms:
+$\Psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i,j=1}^{N}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}).$ (24)
+As per the duality between neuronal and temporal covariances, $\Psi^{a}(\bm{\tau})$ can also be expressed in terms of the time-by-time covariance $C^{a}(t_{1},t_{2})$ as
 $\Psi^{a}(\bm{\tau})=N\left\langle C^{a}(t_{1},t_{2})C^{a}(t_{1}+\tau_{1},t_{2}+\tau_{2})\right\rangle_{t_{1},t_{2}}.$ (25)
-The neuron-by-neuron (Eq. 7) and time-by-time (Eq. 9) definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively.
-II.2.3 Relating $\Psi^{a}$ to the dimension and timescales of activity
-The dimension of activity can be quantified by the participation ratio of its covariance spectrum [34, 35, 23, 36, 32, 37, 9, 38, 27].
-To define the participation ratio, we diagonalize the equal-time covariance matrix $C_{ij}^{a}(0)$ to obtain eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$, as in principal components analysis.
-The participation ratio is
-$\text{PR}^{a}=\frac{1}{N}\frac{(\sum_{k}\lambda^{a}_{k})^{2}}{\sum_{k}(\lambda^{a}_{k})^{2}},$ (10)
-[…]
-An advantageous property of this dimension measure is that it can be expressed in terms of the previously defined two- and four-point correlation functions as
-$\text{PR}^{a}=\frac{C^{a}(0)^{2}}{\Psi^{a}(0,0)}.$ (11)
-Recall that for the types of connectivity we study, cross-covariances are $\mathcal{O}(1/\sqrt{N})$, so the on- and off-diagonal contributions to $\Psi^{a}(\bm{\tau})$ are both order-one (Eq.
-7).
-This leads to $\text{PR}^{a}$ taking a value between 0 and 1 as $N\rightarrow\infty$.
-Thus, the unnormalized dimension of activity $N\,\text{PR}^{a}$ is extensive, i.e., activity is high-dimensional.
-At the same time, activity does not uniformly occupy all available dimensions, which would require cross-covariances to be smaller than $\mathcal{O}(1/\sqrt{N})$.
-Timescales of collective activity can be studied as follows. We define the principal components (PCs) of network activity as
-$p^{a}_{k}(t)=\frac{(\bm{v}^{a}_{k})^{T}\bm{a}(t)}{\sqrt{\lambda^{a}_{k}}}.$ (12)
-These PCs are the basis of much of modern analysis of high-dimensional neural data.
-The PCs all have unit variance but potentially very different characteristic timescales (Fig.
-2a).
-To extract the timescales of just the leading PCs, we compute the PC-averaged covariance, weighted by the squares of the eigenvalues of the corresponding PCs.
+The neuron-by-neuron (Eq.
+24) and time-by-time (Eq.
+25) definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively.
+Finally, the dimension of activity is given by
+$\text{PR}^{a}=\frac{C^{a}(0)^{2}}{\Psi^{a}(0,0)},$ (26)
+which, unlike Eq. 22, holds even when single-neuron variances $C^{a}_{ii}(0)$ are nonuniform across neurons.
+To study timescales of collective activity, we consider the principal components of activity:
+$p^{a}_{k}(t)=\frac{1}{\sqrt{\lambda^{a}_{k}}}(\bm{v}^{a}_{k})^{T}\bm{a}(t).$ (27)
+These principal components are the basis of much of modern analysis of high-dimensional neural data.
+They all have unit variance due to the $1/\sqrt{\lambda^{a}_{k}}$ normalization, but potentially very different characteristic timescales (Fig.
+9a).
+To extract the timescales of just the leading components, we weight them by the squares of their corresponding eigenvalues.
 This gives:
 $\frac{1}{N}\sum_{k}(\lambda^{a}_{k})^{2}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}$ $=\Psi^{a}(\tau,0),$ (28)
-showing that $\Psi^{a}(\bm{\tau})$ captures the temporal structure of the leading PCs of network activity (Fig.
-2c,d).
+showing that $\Psi^{a}(\bm{\tau})$ captures the temporal structure of the leading principal components of activity (Fig.
+9c,d).
 Note that weighting by the eigenvalues themselves, rather than their squares, gives $N^{-1}\sum_{k}\lambda^{a}_{k}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}=C^{a}(\tau)$, recovering single-neuron information (Fig.
-2b,d).
-II.3 Computing the two- and four-point functions in the i.i.d. case
+9b,d).
+IV Path-integral analysis of the random-mode model
+We aim to compute the summary statistics $C^{a}(t_{1},t_{2})$, $S^{a}(t_{1},t_{2})$, and $\Psi^{a}(\bm{\tau})$ in the limit $N\to\infty$ under the disorder average.
+For a general function $\mathcal{F}$, we use the subscript $\star$ to denote such limiting values:
+$\mathcal{F}_{\star}=\lim_{N\to\infty}\left\langle\mathcal{F}\right\rangle_{\bm{J}}.$ (29)
+These values also correspond to the saddle point of a path integral.
+The summary statistics we have defined are self-averaging, meaning that the same values are obtained, up to $\mathcal{O}(1/\sqrt{N})$ fluctuations, when they are computed based on a single realization of a large network.
+In this paper, we are interested in statistically stationary states such that, for $N\to\infty$, the two-point functions depend on only time differences and thus can be replaced by their stationarized counterparts (Eq.
+19), i.e., for all $t$,
+$C^{a}_{\star}(t,t\pm\tau)=C^{a}_{\star}(\tau),\>\>\>\>S_{\star}^{a}(t,t-\tau)=S^{a}_{\star}(\tau).$ (30)
+Calculating two-point functions is well established “standard DMFT.” As we review in Appendix E, four-point functions were previously calculated under i.i.d.
+connectivity using a two-site version of the cavity method [30].
+In this section, we introduce a new path-integral approach to calculating four-point functions that allows for rapid solution of the i.i.d.
+model and, importantly, more complicated models including the random-mode model.
+IV.1 Relating $\Psi^{a}$ to temporal covariance fluctuations
+To cast the problem in the language of path integrals, we turn to the time-by-time definition of $\Psi^{a}(\bm{\tau})$ (Eq.
+25).
+If the time-by-time covariance $C^{a}(t_{1},t_{2})$ were equal to its limiting value, $C^{a}_{\star}(t_{2}-t_{1})$, we would have $\Psi_{\star}^{a}(\bm{\tau})\overset{?}{=}N\left\langle C^{a}_{\star}(t_{2}-t_{1})C^{a}_{\star}(t_{2}-t_{1}+\tau_{2}-\tau_{1})\right\rangle_{t_{1},t_{2}}=0$} since, for large time differences, $C^{a}_{\star}(t_{2}-t_{1})$, decays exponentially in $|t_{2}-t_{1}|$ [10].
+However, based on the neuron-by-neuron definition (Eq.
+24), $\Psi_{\star}^{a}(\bm{\tau})$ is clearly a nonzero, order-one quantity.
+This apparent contradiction is resolved by considering the $\mathcal{O}(1/\sqrt{N})$ fluctuations around the limiting value.
+The necessity of these fluctuations becomes clear when considering the duality between neuronal and temporal covariances: the presence of nonzero cross-covariances $C^{a}_{ij}(0)$ indicates low-dimensional structure in the system; equivalently, this structure manifests as the system revisiting similar states over time, resulting in nonzero temporal covariance $C^{a}(t_{1},t_{2})$ even for large time differences $|t_{2}-t_{1}|\gg 1$ (Fig.
+3).
+To capture this behavior, we define fluctuations around the saddle point as
+$\delta C^{a}(t_{1},t_{2})=C^{a}(t_{1},t_{2})-C^{a}_{\star}(t_{2}-t_{1}).$ (31)
+We express $\Psi^{a}(\bm{\tau})$ purely in terms of these fluctuations as
+$\Psi^{a}(\bm{\tau})=N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{1}+\tau_{1},t_{2}+\tau_{2})\right\rangle_{t_{1},t_{2}},$ (32)
+noting that terms involving $C^{a}_{\star}(t_{2}-t_{1})$ vanish at large time differences.
+Our task, therefore, is to compute the covariance of the fluctuations, $\delta C^{a}(t_{1},t_{2})$, captured asymtotically by
+$\left\langle N\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{3},t_{4})\right\rangle_{\star}.$ (33)
+Then, to compute $\Psi_{\star}^{a}(\bm{\tau})$, we take in Eq. 33 the limits
+$t_{3}$ $=t_{1}+\tau_{1},$ $|t_{2}-t_{1}|$ $\to\infty,$ (34)
+$t_{4}$ $=t_{2}+\tau_{2},$ $\tau_{1},\tau_{2}$ $\sim\mathcal{O}(1).$
+IV.2 Calculation of two- and four-point functions
+Details of our calculation are given in Appendix F.
+Briefly, we compute the covariance of fluctuations (Eq.
+33) using the Martin-Siggia-Rose-De Dominicis-Janssen (MSRDJ) path-integral formalism [51, 52, 53, 54, 55, 56].
+For any fixed coupling matrix $\bm{J}$, the path integral is
+$Z[\bm{J}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\\ \exp\Bigg{\{}i\sum_{i=1}^{N}\int dt\hat{x}_{i}(t)\Bigg{(}T[{x}_{i}](t)-\sum_{j=1}^{N}J_{ij}\phi_{j}(t)\Bigg{)}\Bigg{\}}.$ (35)
+The auxiliary field $\hat{x}_{i}(t)$ is conjugate to $x_{i}(t)$ and enforces the network’s equations of motion (Eq.
+15) through the delta-function representation $(2\pi)^{-1}\int d\hat{x}\,e^{i\hat{x}x}=\delta(x)$.
+Upon introduction of source terms that are not needed for our purposes, this path integral serves as a generating functional for correlation and response functions.
+In principle, to compute $\Psi^{a}_{\star}(\bm{\tau})$ in this framework, after integrating out $\bm{J}$ from $Z[\bm{J}]$ to obtain a statistical field theory governing two-point functions, we (1) find the saddle-point solution for two-point functions by extremizing the action (i.e., standard DMFT); (2) compute the time2-by-time2 Hessian matrix of the action describing fluctuations around this saddle point; (3) invert this Hessian and extract the sub-block corresponding to the covariance of $\delta C^{\phi}(t_{1},t_{2})$ (Eq.
+33); and (4) apply the temporal limits (Eq.
+34) to this sub-block to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$.
+Direct inversion of the time2-by-time2 Hessian is analytically intractable.
+However, we show that the temporal limits commute with Hessian inversion, allowing us to invert a low-dimensional frequency-dependent Hessian instead (Appendix G).
+This readily yields the Fourier-space function $\Psi^{\phi}_{\star}(\bm{\omega})$, where $\bm{\omega}=(\omega_{1},\omega_{2})$.
+A few additional steps yield $\Psi^{x}_{\star}(\bm{\omega})$.
+This approach allows for the efficient calculation of two and four point functions for a given model, requiring only a few lines of computation directly from the action of the theory.
+As a demonstration of this approach, we apply it to a generalization of the i.i.d. model with correlated reciprocal couplings, recovering results of [30] (Appendix H).
+IV.3 Solution of the random-mode model
+Applied to the random-mode model, the self-consistent equations that determine $C^{\phi}_{\star}(t_{1},t_{2})$ are (Appendix I.1)
+$T[x](t)$ $=\eta^{x}(t),$ (36a)
+$\eta^{x}$ $\sim\mathcal{GP}(0,\alpha r_{2}C^{\phi}_{\star}),$ (36b)
+$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}.$ (36c)
+These equations are equivalent those for an i.i.d. model (Eqs. 6971) with coupling strength
+$g_{\text{eff}}=\sqrt{\alpha r_{2}}.$ (37)
+The four-point function is (Appendix I.2)
+$\Psi_{\star}^{\phi}=\frac{1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}{|1-g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}C_{12}^{\phi},$ (38)
+where we have simplified notation by suppressing frequency variables via the shorthand $C^{\phi}_{12}=C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2})$ and $S^{\phi}_{12}=S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})$.
+The same solution is obtained through a more involved, but intuitive, two-site cavity calculation (Appendix L).
+This solution for the random-mode model is one of our main results that we now spend some time interpreting.
+IV.4 Interpreting the solution
+Figure 4: Dimension of activity and collective timescales in the random-mode model.
+(a) Activity dimension $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ for various coupling strengths $g_{\text{eff}}$.
+Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
+Inset: Extended $g_{\text{eff}}=6$ case, showing convergence to i.i.d.
+coupling behavior with growing $\alpha$.
+(b) Normalized four-point function $\overline{\Psi^{\phi}(\tau,0)}\equiv\Psi^{\phi}(\tau,0)/\Psi^{\phi}(0,0)$ for various $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$.
+Inset: Theory curves for all $\alpha\text{PR}^{D}$, demonstrating relative invariance of collective timescales to the effective rank.
+Simulations use $N=5000$ neurons.
+Our analysis reveals two features of connectivity that characterize how it shapes activity, namely, the effective coupling strength $g_{\text{eff}}$ and the effective rank $\alpha\text{PR}^{D}$.
+The former characterizes local structure (the size of individual couplings) while the latter characterizes global structure (the number of large connectivity components).
+In the equations determining $C_{\star}^{\phi}(\tau)$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ above, we see immediately that global structure in the connectivity is invisible at the single-neuron level but shapes collective activity.
+Specifically, the correlation function $C^{\phi}_{\star}(\tau)$ is fully determined by $g_{\text{eff}}$ and is insensitive to the effective rank of connectivity.
+This has important implications for neural data analysis.
+In particular, analyses of single-neuron, non-collective activity properties may miss signatures of structured connectivity that only become apparent when analyzing collective activity properties among simultaneously recorded neurons.
+What is the nature of the dependence of collective-activity properties, such as the dimension of activity, on the effective rank?
+To plot this relationship and validate the theory, we simulated networks and computed the participation ratio from the empirical equal-time covariance matrix of the activations.
+We used a component-strength spectrum $D_{a}\propto\exp(-\beta_{D}a/M)$ which yields $\text{PR}^{D}=\beta^{-1}_{D}\tanh(\beta_{D})$.
+To attain a desired effective rank $\alpha\text{PR}^{D}<1$, we set $\alpha=M/N=1$ and solved for $\beta$.
+For $\alpha\text{PR}^{D}>1$, we fixed $\text{PR}^{D}=1$ and increased $\alpha$.
+We adjusted $g_{\text{eff}}$ by rescaling $D_{a}$ uniformly.
+We find excellent theory-simulation agreement (Fig.
+4).
+Fig.
+4a illustrates how the dimension of activity, $\text{PR}^{\phi}$, varies with $\alpha\text{PR}^{D}$ for different values of $g_{\text{eff}}$.
+Consistent with i.i.d.-coupling networks, $\text{PR}^{\phi}$ increases monotonically with $g_{\text{eff}}$ [30].
+It also increases monotonically with $\alpha\text{PR}^{D}$.
+Taking $\alpha\text{PR}^{D}\to\infty$ for fixed $g_{\text{eff}}$ in Eq.
+38 recovers the formula for i.i.d.
+couplings (Eq.
+72) of [30].
+Indeed, in this limit, $\bm{J}$ approaches an i.i.d.
+matrix.
+Correspondingly, $\text{PR}^{\phi}$ approaches the dimension of a network with i.i.d.
+couplings (Fig.
+4a inset).
+Comparing Eq.
+72 to the corresponding i.i.d.
+formula ($1/\alpha\text{PR}^{D}=0$) and noting that single-neuron properties are unaffected by $\alpha\text{PR}^{D}$, one sees that any structure beyond i.i.d.
+connectivity ($1/\alpha\text{PR}^{D}$ finite) strictly decreases the dimension of activity.
+The dimension of activity depends on $\Psi^{\phi}_{\star}(\bm{\tau})$ at $\bm{\tau}=(0,0)$; we now examine the temporal profile of this function.
+Fig.
+4b shows a normalized version of $\Psi^{\phi}(\tau,0)$, which is related to the correlation functions of leading principal components, for various values of $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$.
+The decay timescale of $\Psi^{\phi}_{\star}(\tau,0)$ decreases with increasing $g_{\text{eff}}$, approaching a limiting behavior at $g_{\text{eff}}\approx 10$.
+While $\alpha\text{PR}^{D}$ affects the overall scale of this function, and consequently the dimension of activity, it has little effect on the decay timescale, and thus little effect on the timescales of leading principal components (Fig.
+4b, inset).
+Thus, the effective rank primarily influences the dimension of activity rather than the collective temporal structure.
+A remarkable feature of these results is that only two connectivity features, $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$, are sufficient to determine the two- and four-point functions.
+Due to the nonlinear, high-dimensional nature of the network, this reduction is unexpected.
+Indeed, this reduction breaks down when mode overlaps are introduced (Sec.
+VII), in which case the formula for the four-point function requires knowledge of the full joint distribution of component strengths and overlaps.
+V Limiting behavior of the random-mode model
+Here we examine the behavior of the random-mode model in two limiting cases: as the system approaches the transition to chaos from above ($g_{\text{eff}}\to 1^{+}$), and in the limit of low effective rank ($\alpha\text{PR}^{D}\to 0^{+}$).
+We first consider the limit $g_{\text{eff}}\to 1^{+}$ for arbitrary $\alpha\text{PR}^{D}$.
+The relevant expansion parameter is $g_{\text{eff}}-1\ll 1$.
+For networks with i.i.d.
+couplings with coupling strenght $g$, it is known that, as $g\to 1^{+}$, $\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})\sim 1/(g-1)$ and $\text{PR}^{a}_{\text{i.i.d.}}\sim(g-1)^{3}$ in this limit [30].
+For the random-mode model, to leading order in $1/(g_{\text{eff}}-1)$,
+$\Psi^{a}_{\star}(\bm{\tau})=\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)\underbrace{\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})}_{\sim 1/(g_{\text{eff}}-1)}.$ (39)
+Consequently, to leading order in $g_{\text{eff}}-1$, the dimension of activity behaves as
+$\text{PR}^{a}=\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)^{-1}\underbrace{\text{PR}^{a}_{\text{i.i.d.}}}_{\sim(g_{\text{eff}}-1)^{3}}.$ (40)
+Thus, as the network approaches the transition to chaos from above, structured connectivity reduces the dimension by a factor of $(1+1/\alpha\text{PR}^{D})^{-1}$ compared to a network with i.i.d.
+couplings.
+Taking the further limit of $\alpha\text{PR}^{D}\to 0^{+}$ yields the simple relation
+$\text{PR}^{a}=\alpha\text{PR}^{D}\text{PR}^{a}_{\text{i.i.d.}}$. (41)
+The difference between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$ is a higher-order correction, so we also have $\text{PR}^{a}=\text{PR}^{S}\text{PR}^{a}_{\text{i.i.d.}}$.
+That is, in the limits of both low effective rank and small effective coupling strength, the dimension of activity is equal to the dimension of the singular-value spectrum times the dimension of activity in an i.i.d.
+network.
+We next consider the limit $\alpha\text{PR}^{D}\to 0^{+}$, for arbitrary $g_{\text{eff}}$.
+In this limit, the component strengths closely approximate the singular values.
+To leading order in $\alpha\text{PR}^{D}$,
+$\text{PR}^{a}$ $=K^{a}(g_{\text{eff}})\alpha\text{PR}^{D},$ (42)
+$K^{a}(g_{\text{eff}})$ $={C^{a}_{\star}(0)^{2}}\Bigg{[}\int d\bm{\omega}\left|\frac{g_{\text{eff}}^{2}S^{a}_{12}}{1-g_{\text{eff}}^{2}S^{\phi}_{12}}\right|^{2}C^{\phi}_{12}\Bigg{]}^{-1}.$ (43)
+We verify this limiting behavior in simulations using $\alpha=1$ and a step function for the component strengths: $D_{a}=\text{const.}$ for $a/N\leq\text{PR}^{D}$ and $D_{a}=0$ otherwise.
+We varied $\text{PR}^{D}$, observing linear relationships with $\text{PR}^{\phi}$ and $\text{PR}^{x}$ in the $\text{PR}^{D}\to 0^{+}$ limit (Fig.
+10a).
+Focusing on $K^{\phi}(g_{\text{eff}})$, note that its expression (Eq.
+43) resembles the formula for the dimension of activity in an i.i.d.
+network with coupling strength $g_{\text{eff}}$ (Eq.
+72), but includes an additional factor $|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}$ in the integral.
+Numerical evaluation confirms that, as expected from Eq.
+41, $K^{\phi}(g_{\text{eff}})\to\text{PR}^{\phi}_{\text{i.i.d.}}$ as $g_{\text{eff}}\to 1^{+}$.
+As $g_{\text{eff}}$ increases, $K^{\phi}(g_{\text{eff}})$ increases monotonically to $\sim$$1.53\,\text{PR}^{\phi}_{\text{i.i.d.}}$ (Fig.
+10a, inset).
+Thus, for networks with low effective rank, $\text{PR}^{\phi}\sim\alpha\text{PR}^{D}\text{PR}^{\phi}_{\text{i.i.d.}}$ up to an order-one fudge factor that becomes exactly one as $g_{\text{eff}}\to 1^{+}$.
+Finally, we consider the dependence of the full spectrum eigenvalues $\lambda^{a}_{i}$ of $\bm{C}^{a}(0)$ on the spectrum of component strengths $D_{a}$.
+The latter leads to the former via a highly nonlinear relationship.
+Indeed, sticking to the step-function component-strength spectrum, the ranked eigenvalues $\lambda^{\phi}$ and $\lambda^{x}$ of the covariance matrix decay smoothly rather than exhibiting a cutoff (Fig.
+10b).
+Nevertheless, certain properties of these connectivity and activity spectra are linked.
+In particular, in the limit $\alpha\text{PR}^{D}\to 0^{+}$, $\alpha\text{PR}^{D}$ and $\text{PR}^{a}$ exhibit a linear relationship (with proportionality factor $K^{a}(g_{\text{eff}})$), implying that the ratios of the second-squared and fourth moments of these spectra are linked (despite that, even in this limit, the full spectra do not coincide).
+VI Single-neuron heterogeneity
+Figure 5: Effect of single-neuron heterogeneity on dimension of activity in the random-mode model.
+(a) Dimension of normalized activity $\text{PR}^{\phi}$ versus participation ratio of gain distribution $\text{PR}^{G}$ for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$.
+Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
+(b) Fractional reduction in dimension for weighted activations relative to $\text{PR}^{\phi}$.
+Blue: $G_{i}\phi_{i}$; purple: $G^{\text{readout}}_{i}\phi_{i}$.
+$g_{\text{eff}}=10$.
+Thick dots: means over 10 simulations, averaging before taking the ratio; lines: theoretical predictions.
+All simulations use $N=5000$ neurons.
+Cortical neurons display broad distributions of firing rates [57].
+This heterogeneity among neurons is analogous to the heterogeneity of component strengths in the random-mode model, but operates in the neuron basis rather than the component basis.
+We now compare how low-dimensional connectivity structure and heterogeneous neuronal properties affect collective activity by extending our framework to include single-neuron heterogeneity.
+In a general formulation described in Appendix J, we extend the nonlinearity of neuron $i$ to $\Phi_{\bm{\theta}_{i}}(x)$, which depends on a vector of neuron-specific parameters $\bm{\theta}_{i}$, and solve for the resulting two- and four-point functions.
+Here, to model firing-rate heterogeneity specifically, we take $\bm{\theta}_{i}$ to consist of a single gain parameter $G_{i}$, with $\Phi_{\bm{\theta}_{i}}(x)=G_{i}\phi(x)$.
+In analogy with the component-strength distribution, we define
+$q_{n}$ $=\left\langle G^{n}\right\rangle_{G},$ (44)
+$\text{PR}^{G}$ $=\frac{q_{2}^{2}}{q_{4}}.$ (45)
+The single-site picture that determines the two-point functions is largely similar to Eq. 36c.
+We analyze collective activity (four-point functions) for (1) unnormalized activations $\Phi_{\bm{\theta}_{i}}(t)=G_{i}\phi(x_{i}(t))$ and (2) normalized activations $\phi_{i}(t)=\phi(x_{i}(t))$, the latter being analogous to $z$-scored firing rates in neural recordings, where $z$-scoring removes single-neuron heterogeneity while preserving structure from cross-neuron and temporal correlations.
+The four-point functions for the unnormalized and normalized variables are found to be
+$\Psi^{\Phi}_{\star}$ $=\frac{\frac{1}{\text{PR}^{G}}+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\Phi}_{12},$ (46a)
+$\Psi^{\phi}_{\star}$ $=\frac{1+\left(\frac{1}{\text{PR}^{G}}+\frac{1}{\alpha\text{PR}^{D}}-1\right)|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12},$ (46b)
+respectively, where $C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$ and we have updated the definition of the effective coupling strength to $g_{\text{eff}}=\sqrt{\alpha r_{2}q_{2}}$.
+Eq.
+46b demonstrates that reductions in effective rank, described by $\alpha\text{PR}^{D}$, and reductions in the heterogeneity of gains, described by $\text{PR}^{G}$, have symmetric effects on the dimension of normalized activity, with their combined effect captured by the harmonic mean $\left(\frac{1}{\alpha\text{PR}^{D}}+\frac{1}{\text{PR}^{G}}\right)^{-1}$  [^6] This symmetry arises because the dynamics of the network are determined by an effective coupling matrix $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$, where $\bm{D}=\text{diag}(D_{1},\ldots,D_{M})$ and $\bm{G}=\text{diag}(G_{1},\ldots,G_{N})$.
+Swapping the distributions of $D_{a}$ and $G_{i}$ does not change the spectrum when $\bm{L}$ and $\bm{R}$ are i.i.d., since $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ and $\bm{R}^{T}\bm{G}\bm{L}\bm{D}$ have the same nonzero eigenvalues..
+One might expect that normalizing the activations before computing the dimension of activity would remove the effect of heterogeneous gains on the dimension.
+This is not the case, however, since neuronal gains affect the recurrent dynamics, not merely the observed firing rates.
+To further explore these effects, we compare the activity dimensions for three sets of activations: unnormalized activations and normalized activations (1 and 2 above); and (3) normalized activations multiplied by an independent set of “readout” gains, $\Phi^{\text{readout}}_{i}(t)=G^{\text{readout}}_{i}\phi_{i}(t)$, where the readout gains have the same distribution as the actual gains.
+Case (3), whose four-point function is given by Eq.
+164, considers variables with the same distribution of firing rates as the unnormalized activations, but where this heterogeneity is unrelated to the recurrent dynamics.
+To validate our theoretical predictions (Eq.
+46a, Eq.
+46b, and Eq.
+164), we simulated networks with component strengths and gains given by $D_{a}\propto\exp(-\beta_{D}a/M)$ and $G_{i}\propto\exp(-\beta_{G}i/N)$ (Fig.
+5).
+The dimension of activity is determined by the coupling strength $g_{\text{eff}}$, the effective rank $\alpha\text{PR}^{D}$, and the participation ratio of the gain distribution $\text{PR}^{G}$.
+Increases in each of these parameters lead to higher activity dimension for all three sets of activations.
+Normalized activations exhibit the highest dimension.
+While scaling normalized activations by heterogeneous factors reduces dimension as expected, the magnitude of this reduction depends on whether $G_{i}$ or $G^{\text{readout}}_{i}$ is used, even when their distributions are identical.
+This occurs because neurons with the largest gains preferentially participate in the leading modes of the normalized activations.
+Further scaling these already-dominant neurons by their gains results in overrepresentation of these neurons compared to using random gains.
+Thus, the dimension of unnormalized activations is lowest, with the dimension of random-readout activations falling between those of the normalized and unnormalized activations.
+In experimental data analysis, it is common to normalize single-neuron activities to control for differences in firing rates across neurons.
+Our analysis implies that such normalization may not eliminate the effects of single-neuron heterogeneity on collective-activity properties.
+The persistence of these effects through normalization provides an experimental signature for distinguishing between heterogeneity that affects recurrent dynamics and heterogeneity that affects only the observed firing rates (e.g., in a circuit that is not recurrent at all).
+VII Mode overlaps
+Thus far, we have assumed i.i.d.
+modes, with the factorization $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$ implying that left–right overlaps $\bm{\ell}_{a}^{T}\bm{r}_{a}$ are random and $\mathcal{O}(1/\sqrt{N})$.
+The same holds for $\bm{\ell}_{a}^{T}\bm{r}_{b}$ with $a\neq b$ due to the more fundamental assumption that $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ factorizes across $a$.
+However, computations in biological circuits require structured interactions between modes, suggesting stronger, non-random overlaps between left and right modes.
+To test this, we return to the Drosophila connectome and compute the overlap matrix between left and right singular vectors, $O_{ab}=\bm{u}_{a}^{T}\bm{v}_{b}$ (Fig.
+6a).
+Contrary to the small, unstructured overlaps expected under the i.i.d.
+assumption, the connectome exhibits large, structured overlaps, particularly along the diagonal (Fig.
+6b).
+Figure 6: Right-left mode overlaps and eigenvalue spectra of the fly-brain connectome.
+(a) Overlap matrix between right and left singular vectors of the connectome.
+First 1000 modes are shown.
+(b) First 3000 diagonal elements of the overlap matrix from (a).
+Horizontal lines indicate an approximate cutoff for $\mathcal{O}(1/\sqrt{N})$ random overlaps.
+(c) Eigenvalue spectra of various connectome-based matrices.
+Top: Actual hemibrain spectrum, exhibiting large outliers along the real axis.
+Middle: Eigenvalues from a realization of the random-mode model, using singular values from the connectome as component strengths $D_{a}$, resulting in a circularly symmetric spectrum.
+Bottom: Random-mode model realization with right-left mode overlaps incorporated, where overlaps are given by the inner products of left and right singular vectors from the fly connectome SVD.
+Data from Scheffer et al.
+[17], $N=18028$ neurons.
+Motivated by the presence of these large diagonal overlaps, we extend the random-mode model to include correlations between corresponding left and right modes.
+Specifically, we take $P(\ell_{a},r_{a})$ to be a zero-mean bivariate Gaussian with marginal variance $1/N$ and covariance $\rho_{a}/N$, where $\rho_{a}$ is the correlation between the $a$-th mode pair.
+Both $D_{a}$ and $\rho_{a}$ are treated deterministically, and we assume that $M^{-1}\sum_{a}f(D_{a},\rho_{a})$ converges to a limiting value $\left\langle f(D,\rho)\right\rangle_{D,\rho}$ as $N\to\infty$.
+In this setting, $\bm{\ell}_{a}^{T}\bm{r}_{a}=\rho_{a}$ for large $N$, while $\bm{\ell}_{a}^{T}\bm{r}_{b}$ for $a\neq b$ remains random and $\mathcal{O}(1/\sqrt{N})$.
+This additional structure produces eigenvalue spectra resembling that of the connectome, including large real outliers (Fig.
+6c) [59].
+This purely diagonal overlap structure, however, can only generate real eigenvalue outliers.
+The presence of complex outliers in the connectome, together with the full overlap matrix itself (Fig.
+6a), implies the presence of richer correlations.
+Capturing such features would require relaxing the full factorization of $P(\{\ell_{a},r_{a}\})$ across $a$ while retaining $\mathcal{O}(M)$ parameter scaling—for example, via blockwise factorization or Markovian correlations across $a$.
+We leave such generalizations for future work.
+We also note that by setting $\rho_{a}=1$ for all $a$, the model becomes a Hopfield network storing an extensive set of Gaussian patterns.
+In such networks, each mode can be “condensed” ($\mathcal{O}(1)$ overlap with activity) or “uncondensed” ($\mathcal{O}(1/\sqrt{N})$ overlap with activity).
+Our analysis assumes no condensed modes, but sufficiently strong positive overlaps break this assumption.
+The Gaussian case, however, is unable to generate multiple multistable attractors as in a memory system [35].
+Note that the form of the single-element density is relevant for the condensed patterns, but not for the uncondensed patterns.
+Allowing for condensed patterns within our analysis would require specifying this density.
+See [60] for a DMFT analysis of generalized Hopfield dynamics in which these ideas are implemented.
+In Appendix K, we compute the two- and four-point functions. As in previous cases, the dynamics reduce to single-site processes:
+$T[x](t)$ $=\eta^{x}(t)+\int^{t}dt^{\prime}R_{\star}(t,t^{\prime})\phi(t^{\prime}),$ (47a)
+$\eta^{x}$ $\sim\mathcal{GP}(0,Q_{\star}),$ (47b)
+$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (47c)
+$S^{\phi}_{\star}(t_{1},t_{2})$ $=\left\langle\frac{\delta\phi(t_{1})}{\delta I(t_{2})}\right\rangle_{\star}.$ (47d)
+Assuming temporal stationarity, the kernels are given in Fourier space by
+$Q_{\star}(\omega)$ $=\alpha\left\langle\frac{D^{2}}{|1-D\rho S^{\phi}_{\star}(\omega)|^{2}}\right\rangle_{D,\rho}C^{\phi}_{\star}(\omega),$ (48)
+$R_{\star}(\omega)$ $=\alpha\left\langle\frac{D\rho}{1-D\rho S^{\phi}_{\star}(\omega)}\right\rangle_{D,\rho}.$ (49)
+Unlike the “i.i.d. modes” case, the single-site picture is not equivalent to that of an i.i.d. network for some $g_{\text{eff}}$.
+To gain intuition for the self-coupling kernel $R_{\star}(\omega)$, we expand in powers of $\rho$. The coefficients encode connectivity motifs of increasing orders:
+$\begin{split}&R_{\star}(\omega)=\left\langle J_{ii}\right\rangle_{\star}+\left\langle NJ_{ij}J_{ji}\right\rangle_{\star}\>S^{\phi}_{\star}(\omega)+\cdots,\text{ where}\\ &\left\langle J_{ii}\right\rangle_{\star}=\alpha\left\langle D\rho\right\rangle_{D,\rho},\>\>\left\langle NJ_{ij}J_{ji}\right\rangle_{\star}=\alpha\left\langle(D\rho)^{2}\right\rangle_{D,\rho}.\end{split}$ (50)
+Here, $\left\langle J_{ii}\right\rangle_{\star}$ corresponds to deterministic self-couplings or autapses.
+To avoid such biologically implausible $\mathcal{O}(1)$ self-connections, we set $\left\langle D\rho\right\rangle_{D,\rho}=0$, yielding $J_{ii}=\mathcal{O}(1/\sqrt{N})$  [^7] If model neurons represent clusters of biological neurons, strong self-connections may instead be interpreted as within-cluster coupling [83, 84]..
+The second term in the expansion (note that $i\neq j$) reflects reciprocal correlations in an otherwise i.i.d.
+network (Appendix H).
+Nonzero higher-order terms indicate that this structure cannot be reduced to either of these simpler cases.
+Notably, if both $\left\langle NJ_{ii}\right\rangle_{\star}$ and $\left\langle NJ_{ij}J_{ji}\right\rangle_{\star}$ vanish, all higher-order terms vanish as well, making reciprocal covariance necessary and sufficient for a nonzero self-coupling kernel in the single-site description.
+The four-point function, meanwhile, is given by
+$\Psi^{\phi}_{\star}$ $=\frac{V-1+\left\{\frac{(1+X_{12})(1+X_{21}^{*})}{1-W^{*}}+\text{h.c.}\right\}}{|1-U|^{2}}C^{\phi}_{12},$ (51)
+$U$ $=\alpha\left\langle D^{2}\Sigma_{1}\Sigma_{2}\right\rangle_{D,\rho},\>V=\alpha\left\langle D^{4}|\Sigma_{1}\Sigma_{2}|^{2}\right\rangle_{D,\rho},$
+$W$ $=\alpha\left\langle D^{2}\rho^{2}\Sigma_{1}^{*}\Sigma_{2}\right\rangle_{D,\rho},\>X_{12}=\alpha\left\langle D^{3}\rho\Sigma_{1}|\Sigma_{2}|^{2}\right\rangle_{D,\rho},$
+$\Sigma_{k}$ $=\frac{S^{\phi}_{k}}{1-D\rho S^{\phi}_{k}}\quad(k=1,2).$
+We validate this theory using $\alpha=1$ and $D_{a}\propto\exp(-\beta_{D}a/N)$, as before.
+For overlaps, we use a sigmoidal spectrum for $\rho_{a}$ qualitatively inspired by Fig.
+6b, with multiplicative prefactor $\gamma_{\rho}$ and other parameters chosen such that $\left\langle D\rho\right\rangle_{D,\rho}=0$ (Appendix K.3).
+Across values of $g_{\text{eff}}$ and $\text{PR}^{D}$, sufficiently strong overlaps, of either sign, reduce the dimension of activity (Fig.
+7).
+The maximum dimension occurs for $\gamma_{\rho}<0$, where dynamics most strongly suppress activity aligned with dominant connectivity modes.
+This effect weakens as $g_{\text{eff}}$ increases.
+Figure 7: Effect of right-left mode correlations on dimension of activity in the random-mode model.
+Dimension of activity $\text{PR}^{\phi}$ versus correlation parameter $\gamma_{\rho}$ in the $(D,\rho)$ joint distribution, for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$.
+Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
+All simulations use $N=5000$ neurons.
+VIII Discussion
+Motifs vs.
+global structure.
+Neuronal connectivity is often described using statistics of motifs: local patterns like chains or cycles among small groups of neurons [62, 18, 63, 64].
+These can be estimated from partial observations of a weight matrix which, until recently, set the limits of experimental measurement.
+Advances in whole-brain connectome reconstruction now allow access to global connectivity, including spectral features.
+Our parameterization of coupling matrices represents this global structure, with control over both the spectrum and mode overlaps.
+These spectral features determine how activity modes interact dynamically and thus may more directly dictate neural computations compared to motifs.
+That said, in some studies the primary goal is to capture motifs or graph-theoretic features (e.g., degree distributions) over which the random-mode model does not provide control (although some of these features have spectral signatures that could be incorporated).
+When such features are the focus, other connectivity models are better suited, for example maximum-entropy random graph models constrained by fixed structural features (e.g., the configuration model for matching degree distributions).
+Linking connectivity and activity datasets.
+The random-mode model, together with its dynamical theory, could aid in the interpretation of large-scale neural activity recordings in the context of anatomical connectivity, such as in Drosophila, where both the connectome and whole-brain activity data are available [65].
+In the current work, we used this connectome to characterize the spectral structure of an actual large neural circuit, stopping short of considering the activity that it generates.
+Future work could progressively incorporate richer connectome constraints into the model to identify which features most strongly shape collective activity.
+One possible sequence would be to add the spectrum of component strengths, then diagonal mode overlaps, and finally off-diagonal overlaps, assessing at each step how much the correspondence between modeled and observed activity improves.
+This is similar to approaches in statistical physics that aim to recover higher-order correlations by constraining lower-order statistics [66, 67, 68].
+Learning.
+Learning alters collective dynamics by modifying connectivity—the core principle of machine learning.
+Our analysis focuses on structured but untrained networks that produce chaotic activity.
+Nevertheless, the extensive low-rank connectivity structures we study provide a starting point for understanding trained networks.
+Large networks trained on simple tasks often show intensive low-rank modifications to their initial couplings, with a finite number of outlier eigenvalues emerging from the bulk [23, 34].
+Existing theoretical tools can already describe such systems by taking the trained weights, performing a low-rank decomposition, modeling mode overlaps statistically, and studying the resulting low-rank DMFT [36].
+By contrast, empirical studies of large networks trained on real-world tasks suggest the emergence of extensive low-rank structure, with smooth spectral profiles with many significant components [24].
+This structure resembles that modeled by the random-mode model.
+Whether mode overlaps extracted from such networks can be incorporated into the random-mode model framework, and thereby yield a working theory for the extensive-rank case, is an open question.
+A more ambitious direction is to develop a DMFT for networks whose weights are learned from data, rather than first training networks and then retrofitting a statistical model.
+Such a theory would describe not only how data shapes weights, as addressed by classical works [69], but also how data interacts with recurrent network dynamics.
+The path-integral formalism is well-suited to this, as it allows derivation of DMFT equations for coupling matrices drawn from a Gibbs ensemble with an energy function given by a task loss.
+While this approach has been applied to low-dimensional tasks, leading to intensive-rank weights [70], applying it to high-dimensional tasks could yield a theory of how such tasks generate extensive-rank weights that in turn produce extensive-dimensional activity whose single-neuron and collective properties the theory would describe.
+Another route to extensive-rank structure is via networks that perform many tasks, which could be modeled, for example, by representing mode overlaps through blockwise factorization of $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ across $a$.
+This “multi-task” scenario is increasingly relevant in both neuroscience and machine learning [71, 72].
+Even in standard RNN training, initializing couplings with low-rank structure could improve procedures such as FORCE learning, which rely on suppressing chaos [12, 13].
+Lower-dimensional attractors may be easier to control, potentially leading to faster convergence.
+Modeling functional consequences of single-neuron heterogeneity.
+A central question in neuroscience is how neuronal diversity shapes collective dynamics and computation [73].
+Appendix J presents a general framework that assigns each neuron a parameter vector $\bm{\theta}$ and enables study of how different forms of heterogeneity affect properties such as dimensionality.
+Sec.
+VI focuses on gain heterogeneity to model firing-rate diversity, but this framework could also incorporate more intricate cell-specific features, such as electrophysiological characteristics [74], linking single-neuron heterogeneity to functional consequences for population activity.
+Transient dynamics and excitation–inhibition.
+Our analysis has focused on stationary states with translation-invariant temporal correlations.
+Many neural computations, however, rely on transient dynamics that break stationarity and depend on absolute time.
+Extending the DMFT accordingly is numerically tractable for two-point functions but would become unwieldily for four-point functions, $\Psi^{a}_{\star}(t_{1},t_{2},t_{3},t_{4})$.
+Representing and manipulating such a large object may require approximations or low-rank decomposition.
+One notable transient phenomenon is transient amplification, linked to non-normal coupling matrices that naturally arise from excitatory–inhibitory segregation (Dale’s law) [75, 76].
+Although the random-mode model does not enforce Dale’s law, the resulting matrices are nevertheless non-normal and may exhibit transient amplification; such effects may be amplified through structured off-diagonal overlaps between left and right modes.
+Non-normality also decouples the highest-variance dimensions from the slowest modes.
+Relatedly, extending the random-mode model to respect Dale’s law and incorporate a non-saturating $\phi(x)$—achieving stability via excitation–inhibition balance rather than saturation [77, 78]—could enable more direct comparisons to both connectivity and activity data.
+Random basis property.
+The random-mode model assumes that the $2M$ mode components associated with each neuron are sampled i.i.d.
+for each neuron (Eq.
+3).
+This leads to neuronal permutation symmetry of the distribution over $\bm{J}$.
+Furthermore, when the marginal distributions over left and right mode components are isotropic Gaussians, as we have assumed, the embedding into neuronal space is, by definition, random and Gaussian.
+This “random basis property” may underlie heterogeneous tuning observed in neural circuits [25], but in our case it imposes modeling limitations: while the random-mode model has configurable spectral properties, it precludes configuration of spatial or anatomical properties.
+This limitation could be addressed by conditioning the distribution over the $2M$ mode components on neuron-local properties such as brain region or spatial location, provided the number of effectively distinct possibilities is intensive.
+This would yield a regionally or spatially embedded random-mode model, apt for connectome studies.
+Because we assume zero-mean mode components (Eq.
+4), $\left\langle J_{ij}\right\rangle=0$ for $i\neq j$.
+Thus, the low-rank structure is not inherited from the expected weight matrix.
+This differs from the mechanism described by Thibeault et al.
+[19], who analyze cases where low-rank structure arises because $\left\langle\bm{J}\right\rangle$ is low-rank, with the random part treated as noise.
+By Weyl’s inequality, the singular values of $\bm{J}$ differ from those of $\left\langle\bm{J}\right\rangle$ by at most the spectral norm of the noise.
+This mechanism is fundamentally different from the random-mode model’s zero-mean, randomly embedded structure.
+Which of these mechanisms is more prevalent in real-world networks is an open question.
+Alternative generative models.
+Tiberi et al.
+[79] proposed a generative model for coupling matrices that parameterizes the eigendecomposition of $\bm{J}$ rather than its SVD.
+Their formulation does not capture relationships between eigenvalues and eigenvectors.
+Our SVD-like approach allows simultaneous control over component strengths and their corresponding mode overlaps.
+Whereas Tiberi et al.
+[79] studied noise-driven linear networks within a random matrix theory framework, our DMFT formulation applies to nonlinear dynamics.
+Connectivity parameterizations resembling the random-mode model have been used to study the correspondence between spiking and rate-based recurrent neural networks [80] and to expedite training spiking networks [81].
+Conclusion.
+The random-mode model and associated analytical techniques provide tools for exploring the relationship between connectivity and collective dynamics.
+Our parameterization of connectivity is a middle ground between intensive-rank models and i.i.d.
+random networks.
+Building on this framework stands to further bridge neuronal connectivity, activity, and function.
+Acknowledgements
+We thank L.
+F.
+Abbott for feedback.
+A.v.M.
+was supported by the Swartz Foundation; D.G.C, O.M., and A.L.-K.
+by the Kavli Foundation and Gatsby Charitable Foundation (grant no.
+GAT3708); and D.G.C.
+by the National Institutes of Health (grant no.
+T32NS064929).
+Appendix
+Appendix A Glossary of terms
+Number of neurons ($N$)
+Total number of neurons in the network.
+Number of modes ($M$)
+Number of components in the random-mode model.
+(Effective) coupling strength ($g$, $g_{\text{eff}}$)
+Standard deviation of the couplings times $\sqrt{N}$ in the i.i.d. connectivity model, random-mode model (Eq. 37), or variants thereof (e.g., see just below Eq. 46a and Eq. 46b).
+Random-mode model
+Network connectivity model: $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{\ell}_{a}\bm{r}_{a}^{T}$ (Eq. 2).
+Connectivity component ($\bm{\ell}_{a}\bm{r}_{a}^{T}$)
+Outer product of output and right modes.
+Right mode ($\bm{r}_{a}$)
+Neuronal pattern in the $a$-th component onto which activity patterns are projected.
+Left mode ($\bm{\ell}_{a}$)
+Neuronal pattern in the $a$-th component along which the projected activity pattern is expanded.
+Component strength ($D_{a}$)
+Scaling factor for the $a$-th component.
+Ratio of modes to neurons ($\alpha$)
+Given by $\alpha=M/N$ (Eq. 7).
+Participation ratio of component strengths ($\text{PR}^{D}$)
+Given by $\text{PR}^{D}=r_{2}^{2}/r_{4}$ (Eq. 11) where $r_{n}=\langle D^{n}\rangle_{D}$ (Eq. 10).
+Effective rank ($\alpha\text{PR}^{D}$)
+Connectivity dimensionality measure (Eq. 12).
+Two-point function
+Correlation function given by $C^{\phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\langle\phi_{i}(t_{1})\phi_{i}(t_{2})\rangle_{t}$ (Eq. 17).
+Four-point function
+Higher-order correlation function given by $\Psi^{\phi}(\tau_{1},\tau_{2})=\frac{1}{N}\sum_{i,j=1}^{N}C^{\phi}_{ij}(\tau_{1})C^{\phi}_{ij}(\tau_{2})$ (Eq.
+24).
+See also time-by-time definition (Eq.
+25).
+Appendix B Preprocessing of fly hemibrain connectome
+We use the dataset from Scheffer et al.
+[17], in which synaptic counts are reported for each connection.
+These counts determine the magnitudes of the coupling matrix elements.
+Neurotransmitter probabilities for each neuron are obtained from a machine-learning analysis of the electron microscopy data [85].
+Each neuron is assigned its most probable neurotransmitter, which is then mapped to a synaptic sign (excitatory or inhibitory) according to Table 1 [86, 87].
+Neurotransmitter Synaptic Sign
+GABA -1 (Inhibitory)
+Acetylcholine +1 (Excitatory)
+Glutamate -1 (Inhibitory)
+Serotonin Ignored
+Octopamine Ignored
+Dopamine Ignored
+Neither Ignored
+Table 1: Mapping of neurotransmitters to synaptic signs
+Synaptic counts do not directly correspond to effective synaptic strengths, so it is reasonable to choose “fudge factors” to relate the two.
+Our primary goal in choosing these factors is to standardize the spectrum of the coupling matrix so that it is roughly confined to the unit disk.
+At the same time, we want to preserve key structural features of the original connectivity.
+Specifically, the preprocessing is designed to preserve:
+  • the sparsity pattern (which connections are zero vs. nonzero);
+  • the sign of each connection (excitatory vs. inhibitory);
+  • approximately, the relative magnitudes of same-sign inputs to each neuron; and
+  • approximately, the relative magnitudes of same-sign outputs from each neuron.
+The last two properties are preserved only approximately because they are enforced simultaneously by the iterative normalization procedure described below.
+We normalize the coupling matrix using an iterative scaling method analogous to the Sinkhorn–Knopp algorithm.
+First, for each row, we rescale the positive elements by a row-specific factor so that their L2-norm equals $1/\sqrt{2}$.
+We then rescale the negative elements, using a separate factor, so that their L2-norm also equals $1/\sqrt{2}$.
+This yields rows with a total L2-norm of one and balanced excitatory and inhibitory contributions.
+A small number of neurons with only excitatory or only inhibitory inputs retain an L2-norm of $1/\sqrt{2}$.
+Next, we apply the same rescaling procedure to each column, which disrupts the previous row normalization.
+We alternate the row and column scaling steps until convergence.
+The result can be interpreted as the projection of the original matrix onto the set of matrices that are properly normalized along both rows and columns, while keeping the same nonzero elements and signs.
+This iterative normalization provides a principled way to choose the required scale factors while achieving our goal of a standardized spectrum.
+Appendix C Free probability calculation
+We wish to determine the limiting density of the singular values of
+$\bm{J}=\bm{L}\bm{R}^{T},$ (52)
+where $\bm{L}$ and $\bm{R}$ are independent $N\times M$ random matrices with $\alpha=M/N$ and entries of variance $1/N$. Equivalently, we can compute the eigenvalues of
+$\bm{J}^{T}\bm{J}=\bm{R}\bm{L}^{T}\bm{L}\bm{R}^{T},$ (53)
+and then take the square root to obtain the singular values. These eigenvalues are the same as those of
+$\bm{L}^{T}\bm{L}\bm{R}^{T}\bm{R},$ (54)
+which is the product of two independent Wishart matrices:
+$\bm{L}^{T}\bm{L}\bm{R}^{T}\bm{R}=\bm{W}_{L}\bm{W}_{R}.$ (55)
+The Stieltjes transform $G(z)$ of a spectral density $\rho(\lambda)$ is defined as
+$G(z)$ $=\int_{-\infty}^{\infty}\frac{d\lambda\rho(\lambda)}{z-\lambda},\quad z\in\mathbb{C}\setminus\mathbb{R},$ (56)
+$\rho(\lambda)$ $=-\frac{1}{\pi}\lim_{\epsilon\to 0^{+}}\mathrm{Im}G(\lambda+i\epsilon).$ (57)
+In a 2D electrostatics analogy, the real and imaginary parts of $G(z)$ correspond to the Cartesian components of the electric field generated by a line charge density $\rho(\lambda)$ on the real axis.
+The inversion formula is then analogous to Gauss’s law, relating the discontinuity of the normal field component to the local charge density.
+The moment generating function and S-transform are defined by
+$M(z)$ $=zG(z)-1,$ (58)
+$S(z)$ $=\frac{1+z}{zM^{-1}(z)},$ (59)
+where $M^{-1}(z)$ is the functional inverse of $M(z)$. For a Marchenko–Pastur distributed Wishart matrix $\bm{W}$ with aspect ratio $\alpha$,
+$S(z)=\frac{1}{1+\alpha z}.$ (60)
+A key fact in free probability is that the S-transforms of free variables multiply. Thus, for $\bm{W}_{1}\bm{W}_{2}$, the S-transform is
+$S(z)=\frac{1}{(1+\alpha z)^{2}}.$ (61)
+From the definition of the S-transform, the inverse moment generating function is
+$M^{-1}(z)=\frac{1+z}{zS(z)}=\frac{(1+z)(1+\alpha z)^{2}}{z}.$ (62)
+Therefore, $M(z)$ satisfies
+$z=\frac{(1+M(z))(1+\alpha M(z))^{2}}{M(z)}.$ (63)
+Using M(z)=zG(z)1M(z)=zG(z)-1, we obtain a cubic equation for $G(z)$:
+$\alpha^{2}z^{3}G(z)^{3}+2z^{2}(\alpha^{2}-\alpha)G(z)^{2}+z\big{(}z+2\alpha-\alpha^{2}-1\big{)}G(z)-z=0.$ (64)
+In principle, this can be solved using the cubic formula, but the explicit form is cumbersome.
+For the support boundaries of $\rho(\lambda)$, we only need the points where $\mathrm{Im}G(z)$ changes from zero to nonzero.
+This occurs when the discriminant of the cubic vanishes.
+The discriminant is
+$\Delta=\alpha^{2}\lambda^{7}\left[4\lambda^{2}+(\alpha^{2}-20\alpha-8)\lambda+(-4\alpha^{3}+12\alpha^{2}-12\alpha+4)\right],$ (65)
+where we replaced $z$ with $\lambda$ to indicate restriction to the real axis. Setting $\Delta=0$ gives $\lambda=0$ or the nontrivial quadratic equation
+$4\lambda^{2}+(\alpha^{2}-20\alpha-8)\lambda+(-4\alpha^{3}+12\alpha^{2}-12\alpha+4)=0.$ (66)
+The roots of this quadratic are
+$\lambda_{\pm}=1+\frac{5\alpha}{2}-\frac{\alpha^{2}}{8}\pm\left(1+\frac{\alpha}{8}\right)^{3/2}\sqrt{8\alpha}.$ (67)
+The singular values correspond to the square roots of the eigenvalues, given by Eq. 14.
+Figure 8: Distribution of singular values in the random-mode model.
+Histograms show singular values of $\bm{L}\bm{D}\bm{R}^{T}$ with $D_{a}=1$ for various $\alpha$.
+For $\alpha\ll 1$, the singular values concentrate at one.
+Vertical lines: prediction of Eq.
+14.
+$N=5000$.
+Appendix D Alternative measures of effective dimensionality
+Alternative measures of effective dimensionality, which could be applied to both connectivity and activity in place of the participation ratio, have been proposed in the literature.
+Roy and Vetterli [88] define effective rank as the exponential of the entropy of a categorical probability distribution obtained by normalizing the eigenvalues values to sum to one, i.e., $p_{i}=\lambda_{i}/\sum_{j=1}^{N}\lambda_{j}$.
+Written in terms of this distribution the participation ratio is $\text{PR}=\left(\sum_{i=1}^{N}p_{i}^{2}\right)^{-1}$, while the exponentiated entropy is $\exp(H)=\exp\left(-\sum_{i=1}^{N}p_{i}\log p_{i}\right)$.
+By Jensen’s inequality, $\text{PR}\leq\exp(H)$.
+Exact equality holds in certain cases, including when exactly $K\leq N$ eigenvalues are equal and the rest are zero, yielding $\text{PR}=\exp(H)=K$.
+Furthermore, for smoothly decaying spectra as discussed above (for which $p_{i}=f(i/Nw)/\sum_{j=1}^{N}f(j/Nw)$, with $w\ll 1$), both measures have $\text{PR}=c_{1}\>Nw$ and $\exp(H)=c_{2}\>Nw$, where $c_{1}$ and $c_{2}$ are $N$ and $w$-independent constants depending on the specific form of $f$.
+Thus, while the measures differ in detail, they provide similar characterizations of effective dimensionality.
+This entropy-based definition has been used in neuroscience applications, for example by [89] to analyze transient amplification in neural networks.
+Both measures capture the number of dominant modes in the system.
+Thibeault et al.
+[19] call (a metric closely related to) the the participation ratio-based metric the “srank” (the difference being whether the maximum eigenvalue or sum over all eigenvalues is used as a normalizing factor) and call the entropy-like metric the “erank.”
+Appendix E Review: Calculating two- and four-point functions for i.i.d. couplings
 We now review how to compute these two- and four-point functions in a classic network model with i.i.d. couplings $\bm{J}$. The first- and second-order coupling statistics are:
-$\left\langle J_{ij}\right\rangle_{\bm{J}}=0,\hskip 15.00002pt\left\langle J_{ij}^{2}\right\rangle_{\bm{J}}=\frac{g^{2}}{N}.$ (14)
+$\left\langle J_{ij}\right\rangle=0,\hskip 15.00002pt\left\langle J_{ij}^{2}\right\rangle=\frac{g^{2}}{N}.$ (68)
 The magnitude of a typical coupling is $g/\sqrt{N}$.
 We refer to $g$ as the coupling strength.
-For the canonical dynamics of Eq.
-2, the network is quiescent for $g<1$ and chaotic for $g>1$, with a sharp phase transition as $N\rightarrow\infty$.
+For the canonical dynamics $T[x](t)=(1+\partial_{t})x(t)$, the network is quiescent for $g<1$ and chaotic for $g>1$, with a sharp phase transition as $N\to\infty$.
 In this paper we assume that the network is non-quiescent.
-We compute $C^{a}(t_{1},t_{2})$ and $\Psi^{a}(\bm{\tau})$ in the limit $N\rightarrow\infty$ under the disorder average. We use the subscript $\star$ to denote such limiting, disorder-averaged values:
-$C^{a}_{\star}(t_{1},t_{2})$ $=\lim_{N\rightarrow\infty}\left\langle C^{a}(t_{1},t_{2})\right\rangle_{\bm{J}},$ (15)
-$\Psi^{a}_{\star}(\bm{\tau})$ $=\lim_{N\rightarrow\infty}\left\langle\Psi^{a}(\bm{\tau})\right\rangle_{\bm{J}}.$ (16)
-These values also correspond to the saddle point of a path integral.
-Both of these functions are self-averaging, meaning that the same values are obtained, up to $\mathcal{O}(1/\sqrt{N})$ fluctuations, when they are computed based on a single realization of a large network.
-In this paper, we are interested in statistically stationary, temporally fluctuating states such that $C^{\phi}_{\star}(t_{1},t_{2})=C^{\phi}_{\star}(t_{2}-t_{1})$.
 The two-point function $C^{a}_{\star}(t_{1},t_{2})$ can be computed through a single-site picture that describes the dynamics of a typical neuron embedded in the rest of the network, with pre-activation $x(t)$ and activation $\phi(t)$.
@@
 For the i.i.d.
-couplings considered here, there is a simpler heuristic derivation: in the neuronal input $\sum_{j}J_{ij}\phi_{j}(t)$, the correlations between the couplings $J_{ij}$ and dynamic variables $\phi_{j}(t)$ can be safely neglected to leading order in $1/N$, yielding both Gaussianity of $\eta^{x}(t)$ by the central limit theorem and the second-order statistics of Eq.
-18.
+couplings considered here, there is a simpler heuristic derivation: in the neuronal input $\sum_{j=1}^{N}J_{ij}\phi_{j}(t)$, the correlations between the couplings $J_{ij}$ and dynamic variables $\phi_{j}(t)$ can be safely neglected to leading order in $1/N$, yielding both Gaussianity of $\eta^{x}(t)$ by the central limit theorem and the second-order statistics of Eq.
+70.
 Clark et al.
 [30] first computed $\Psi_{\star}^{a}(\bm{\tau})$ using a dynamic, two-site version of the cavity method, based on the neuron-by-neuron definition (Eq.
-7).
+24).
 This method finds $\psi^{a}_{\star}(\bm{\tau})$, the off-diagonal contribution to $\Psi^{a}_{\star}(\bm{\tau})$, noting that the on-diagonal contribution is simply $C^{a}_{\star}(\tau_{1})C^{a}_{\star}(\tau_{2})$.
@@
 $\Psi_{\star}^{x}(\bm{\omega})$ $=C_{\star}^{x}(\omega_{1})C_{\star}^{x}(\omega_{2})+|U(\bm{\omega})|^{2}C_{\star}^{\phi}(\omega_{1})C_{\star}^{\phi}(\omega_{2})$
-[…]
+$+U(\bm{\omega})C_{\star}^{x\phi}(\omega_{1})C_{\star}^{x\phi}(\omega_{2})+\text{h.c.},$ (73a)
 $U(\bm{\omega})$ $=\frac{g^{2}S_{\star}^{x}(\omega_{1})S_{\star}^{x}(\omega_{2})}{1-g^{2}S_{\star}^{\phi}(\omega_{1})S_{\star}^{\phi}(\omega_{2})},$ (73b)
@@
 The more complex relationship observed here reflects the non-Gaussian joint statistics across different neurons, which is relevant because the network is nonlinear.
-II.4 Definition of the random-mode model
-Figure 3: Schematic of the random-mode model.
-Upper: Couplings $\bm{J}$ are generated as a sum of outer products, ${\bm{L}}_{a}{\bm{R}}^{T}_{a}$, with component strengths $D_{a}$.
-Lower: The two-point function $C^{\phi}_{\star}(\tau)$ and four-point function $\Psi^{\phi}_{\star}(\tau)$ are calculated in terms of the statistics of $D_{a}$.
-The two-point function depends only on the effective gain $g_{\text{eff}}$, while the four-point function depends on both $g_{\text{eff}}$ and $\text{PR}^{D}$, the effective dimension of the connectivity determined by the $D_{a}$ distribution.
-Many studies have examined the dynamics of networks with i.i.d.
-couplings, as described in the previous section.
-However, the coupling matrices of real-world networks often have approximate low-rank spectral structure, deviating dramatically from the i.i.d.
-assumption.
-To analyze networks with such properties, we introduce the random-mode model.
-This model generates a coupling matrix $\bm{J}$ of the form
-$\bm{J}=\sum_{a=1}^{M}D_{a}\bm{L}_{a}\bm{R}_{a}^{T},$ (22)
-where $\bm{R}_{a}$ and $\bm{L}_{a}$ are “input” and “output” modes in the neuron space and $D_{a}>0$ is a weighting factor called the component strength.
-The effect of multiplying by $\bm{J}$ is to project a neuronal activity vector onto $\bm{R}_{a}$, scale it by $D_{a}$, project the result along $\bm{L}_{a}$, and sum these contributions for all $a=1,\ldots,M$ (Fig.
-3, top).
-Eq.
-22 can be written
-$\bm{J}=\bm{L}\bm{D}\bm{R}^{T},$ (23)
-where $\bm{L}$ and $\bm{R}$ are $N\times M$ matrices and $\bm{D}=\text{diag}(D_{1},D_{2},\ldots,D_{M})$.
-A crucial modeling choice is how the number of modes, $M$, scales relative to the network size, $N$. We consider the limit where both $M$ and $N$ approach infinity while maintaining a fixed ratio,
-$\alpha=\frac{M}{N}=\mathcal{O}(1).$ (24)
-This scaling results in extensive-rank connectivity, meaning that the rank of $\bm{J}$ grows linearly with $N$.
-Importantly, choosing $\alpha$ to be small (or choosing the component strengths to decay rapidly; see below) allows for connectivity that is low-rank compared to the number of neurons.
-For now, we assume that the matrices $\bm{L}$ and $\bm{R}$ have independent and identically distributed elements with zero mean and variance $1/N$.
-We assume these elements are Gaussian, although our results should not depend on the exact form of the single-element density for large $N$ due to the central limit theorem.
-A key feature of the random-mode model is that we have full control over the component strengths $D_{a}$, which are analogous to the singular values of $\bm{J}$ (we make this link precise in Sec.
-II.4.1).
-We assume only that these component strengths follow a distribution that remains constant as $N$ grows, ensuring a well-defined large-$N$ limit.
-By varying this distribution, we can explore the impact of the spectral structure of the connectivity on collective features of network activity (Fig.
-3, bottom).
-We denote the $n$-th moment of the $D_{a}$ distribution by
-$r_{n}=\left\langle D^{n}\right\rangle_{D}.$ (25)
-It will be useful to define the participation ratio,
-$\text{PR}^{D}=\frac{r_{2}^{2}}{r_{4}},$ (26)
-from which we define the “effective rank” of the couplings,
-effective $\text{effective rank}=\alpha\text{PR}^{D}.$ (27)
-This quantity is related to the participation ratio of the singular value spectrum of $\bm{J}$ via Eq.
-28 below.
-We will show that this effective rank is a key feature of $\bm{J}$ that controls the collective activity structure.
-II.4.1 Relationship to singular value decomposition
-Figure 4: Distribution of singular values in the random-mode model.
-Histograms show singular values of $\bm{L}\bm{D}\bm{R}^{T}$ with $D_{a}=1$ for various $\alpha$.
-For $\alpha\ll 1$, the singular values concentrate at one.
-Vertical lines: prediction of Eq.
-29.
-$N=5000$.
-The random-mode model has similarities to the SVD, which decomposes a rank-$M$ matrix $\bm{J}$ as $\bm{U}\bm{S}\bm{V}^{T}$.
-Here, $\bm{U}$ and $\bm{V}$ are $N\times M$ matrices whose columns are the left and right singular vectors (or modes), respectively, and $\bm{S}=\text{diag}(S_{1},S_{2},\ldots,S_{M})$ is a diagonal matrix containing the singular values, $S_{a}>0$.
-The left singular vectors are orthonormal among themselves, as are the right singular vectors (although the left and right singular vectors are generally not orthonormal to each other).
-In contrast, the columns of $\bm{L}$ and $\bm{R}$ in the random-mode model have i.i.d.
-statistics, resulting in random, $\mathcal{O}(1/\sqrt{N})$ overlaps among both within- and across-matrix columns.
-This i.i.d.
-assumption offers analytical tractability by eliminating the need to enforce orthonormality, a complicated global constraint (but see [43], who handle random orthonormal matrices in a regression setting).
-Furthermore, by not explicitly modeling overlaps between $\bm{L}$ and $\bm{R}$, the random-mode model parameterizes $\bm{J}$ using only the $\mathcal{O}(N)$ component strengths $D_{a}$, although such overlaps can be introduced (see Sec.
-VII).
-Given a matrix $\bm{J}$ drawn from the random-mode model, we can examine the relationship between the singular values, $S_{a}$, and the component strengths, ${D}_{a}$.
-The non-orthonormality of $\bm{L}$ and $\bm{R}$ introduces some discrepancy between the two sets of values.
-We first compare the participation ratio of the squared singular value spectrum, $\text{PR}^{S}$, to the effective rank, $\alpha\text{PR}^{D}$.
-We find:
-$\text{PR}^{S}=\frac{\alpha\text{PR}^{D}}{1+2\alpha\text{PR}^{D}},$ (28)
-where $\text{PR}^{S}=\langle S^{2}\rangle_{S}^{2}/\langle S^{4}\rangle_{S}$, and $S$ are the singular values of $\bm{J}$.
-This follows from using Wick’s theorem to evaluate the expectations of $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ and $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ for the numerator and denominator, respectively, of $\text{PR}^{S}$.
-For $\alpha\text{PR}^{D}\ll 1$, $\text{PR}^{S}=\alpha\text{PR}^{D}+\mathcal{O}((\alpha\text{PR}^{D})^{2}),$ demonstrating that the component strengths are closely related to the singular values in the low-dimensional regime.
-In this regime, the random modes approximate an orthonormal set of vectors as $N\rightarrow\infty$ since the $\mathcal{O}(1/\sqrt{N})$ overlaps between modes can be neglected when there are few of them.
-While we express the analytic results in this paper in terms of $\alpha\text{PR}^{D}$, Eq.
-28 can always be used to translate between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$.
-We further quantify the discrepancy between singular values and component strengths in the case where $D_{a}=1$ for all $a$.
-Using methods from free probability theory [44, 45], we compute that the $M$ singular values are distributed over a range with boundaries
-$S_{\pm}=\sqrt{1+\frac{5\alpha}{2}-\frac{\alpha^{2}}{20}\pm\left(1+\frac{\alpha}{8}\right)^{\frac{3}{2}}\sqrt{8\alpha}}.$ (29)
-This is in agreement with large-$N$ simulations (Fig.
-4).
-For small $\alpha$, $S_{\pm}=1\pm\sqrt{2\alpha}+\mathcal{O}(\alpha).$ While the random-mode model spreads the nonzero singular values over a range, this spreading becomes negligible for small $\alpha$.
-More generally, the distributions of $D_{a}$ and $S_{a}$ coincide for small $\alpha\text{PR}^{D}$.
-II.4.2 Comparison to low-rank recurrent neural networks
-A related parameterization of connectivity has been explored in low-rank recurrent neural networks, where the coupling matrix is expressed as a sum of outer products of input modes $\bm{n}_{a}$ and output modes $\bm{m}_{a}$:
-$\bm{J}=\frac{1}{N}\sum_{a=1}^{R}\bm{m}_{a}\bm{n}_{a}^{T},$ (30)
-with $R=\mathcal{O}(1)$ being the connectivity rank [30, 46, 31, 47]. In some instances, this low-rank component is added to an i.i.d. matrix modeling unstructured “background” connectivity [30, 46].
-Several features separate this approach from the random-mode model.
-Low-rank recurrent neural networks maintain a constant rank $R$ as $N\rightarrow\infty$, yielding intensive-rank connectivity that models low-dimensional activity, whereas the random-mode model uses extensive-rank connectivity, modeling high-dimensional activity.
-Furthermore, low-rank recurrent neural networks allow for the most general possible parameterization of overlaps between modes, captured by a $R\times R$ overlap matrix $\bm{m}_{a}^{T}\bm{n}_{b}$.
-Specifying such overlaps in the random-mode model would require $\mathcal{O}(N^{2})$ parameters, comparable to defining the full coupling matrix.
-To reduce parameter complexity, the random-mode model specifies only the variances of mode components via the component strengths $D_{a}$, disregarding overlaps.
-This substantially reduces parameter count while preserving interesting network dynamics.
-In Sec.
-VII, we introduce $\mathcal{O}(N)$ additional parameters describing a restricted overlap structure among modes, specifying only diagonal elements of the overlap matrix.
-Further possible overlap forms are discussed in Sec.
-VIII.
-The mean-field descriptions of network dynamics produced by these connectivities also differ markedly.
-In low-rank recurrent neural networks, low-dimensional network activity is described by a finite set of order parameters representing projections of network activity onto input modes $\bm{n}_{a}$, obeying effective nonlinear dynamics.
-Conversely, the extensive-rank connectivity in the random-mode model generates high-dimensional chaotic activity, described by covariance functions capturing temporal fluctuations.
-While the random-mode model’s mean-field analysis introduces latent variables given by projections of network activity onto input modes, their extensive number precludes their use as order parameters.
-A final distinction lies in the scaling of the coupling matrix $\bm{J}$ and its implications for the alignment of activity with connectivity.
-In low-rank recurrent neural networks, elements of $\bm{J}$ scale as $\mathcal{O}(1/N)$ and the input-output mode overlaps are $\mathcal{O}(1)$, resulting in strong $\mathcal{O}(1)$ alignment of network activity with input modes.
-This is reminiscent of the rich regime of neural networks [48, 49, 50, 51].
-Conversely, in the random-mode model, the elements of $\bm{J}$ scale as $\mathcal{O}(1/\sqrt{N})$ and the input-output mode overlaps are $\mathcal{O}(1/\sqrt{N})$, leading to weak $\mathcal{O}(1/\sqrt{N})$ alignment of network activity with any given input mode; the pre-activations are $\mathcal{O}(1)$ since each neuron combines input from $\mathcal{O}(N)$ modes.
-This is reminiscent of the lazy regime of neural networks [49, 52, 53].
-III Path-integral calculation of the two- and four-point functions
-The path-integral approach provides a systematic framework for analyzing neural-network dynamics, allowing for controlled approximations such as $1/N$ expansions [54, 55, 41, 56].
-In this section, we demonstrate how to formulate the calculation of the four-point function $\Psi^{a}_{\star}(\bm{\tau})$ within this formalism, allowing us to solve complex models more efficiently than the two-site cavity method by following a straightforward procedure whose steps we outline at the end of Sec.
-III.3.
-III.1 Relating $\Psi^{a}$ to fluctuations of the time-by-time covariance
-Figure 5: Neuron-time duality and its relation to the dimension of activity.
-Both plots are based on the same simulation of a network of $N=2500$ neurons with $g=2.25$ and i.i.d.
-couplings.
-The dimension can be computed either by computing the statistics of the off-diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
-To cast the problem in the language of path integrals, we turn to the time-by-time definition of $\Psi^{a}(\bm{\tau})$, given by Eq.
-9.
-If the time-by-time covariance $C^{a}(t_{1},t_{2})$ were equal to its limiting value, $C^{a}_{\star}(t_{2}-t_{1})$, we would have $\Psi_{\star}^{a}(\bm{\tau})\overset{?}{=}N\left\langle C^{a}_{\star}(t_{2}-t_{1})C^{a}_{\star}(t_{2}-t_{1}+\tau_{2}-\tau_{1})\right\rangle_{t_{1},t_{2}}=0$ since, for large time differences, $C^{a}_{\star}(t_{2}-t_{1})$, decays exponentially in $|t_{2}-t_{1}|$ [10].
-However, based on the neuron-by-neuron definition, Eq.
-7, $\Psi_{\star}^{a}(\bm{\tau})$ is clearly a nonzero, order-one quantity.
-This apparent contradiction is resolved by considering fluctuations around the saddle point.
-The necessity of these fluctuations becomes clear when viewed through the lens of neuron-time duality.
-The presence of nonzero pairwise covariances $C^{a}_{ij}(0)$ indicates low-dimensional structure in the system.
-Equivalently, this structure manifests as the system revisiting similar states in phase space over time, resulting in nonzero temporal covariance $C^{a}(t_{1},t_{2})$ even for large time differences $|t_{2}-t_{1}|\gg 1$ (Fig.
-5).
-To capture this behavior, we define fluctuations around the saddle point as
-$\delta C^{a}(t_{1},t_{2})=C^{a}(t_{1},t_{2})-C^{a}_{\star}(t_{2}-t_{1}).$ (31)
-We express $\Psi^{a}(\bm{\tau})$ purely in terms of these fluctuations as
-$\Psi^{a}(\bm{\tau})=N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{1}+\tau_{1},t_{2}+\tau_{2})\right\rangle_{t_{1},t_{2}},$ (32)
-noting that terms involving $C^{a}_{\star}(t_{2}-t_{1})$ vanish at large time differences.
-Our task, therefore, is to compute the covariance of the fluctuations, $\delta C^{a}(t_{1},t_{2})$, captured by
-$N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{3},t_{4})\right\rangle_{\star}\\ =\lim_{N\rightarrow\infty}N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{3},t_{4})\right\rangle_{\bm{J}}.$ (33)
-Then, to compute $\Psi_{\star}^{a}(\bm{\tau})$, we use the temporal limits
-[…]
-$t_{4}$ $=t_{2}+\tau_{2},$ $\tau_{1},\tau_{2}$ $\sim\mathcal{O}(1).$
-III.2 Path-integral formalism and two-point functions
-Given any fixed $\bm{J}$, we express the network dynamics using a path integral in the Martin–Siggia–Rose/Janssen–De Dominicis–Peliti formalism [57]:
-$Z[\bm{J}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\left(i\sum_{i}\hat{x}_{i}T[{x}_{i}]-i\sum_{i,j}J_{ij}\hat{x}_{j}\phi_{j}\right),$ (35)
-[…]
-For simplicity, we begin with the i.i.d.
-model of $\bm{J}$ with mean zero and variance $g^{2}/N$.
+Appendix F Path-integral calculation of two- and four-point functions for i.i.d. couplings
+Figure 9: Relationship between principal component timescales and two- and four-point correlation functions.
+(a) Example activity traces of principal components $p^{\phi}_{k}(t)$.
+(b) Empirical autocovariance functions for each principal component, weighted by their eigenvalues.
+(c) Same as (b) but weighted by the squared eigenvalue.
+(d) Averages over curves in (b) and (c), with comparisons to $C^{\phi}_{\star}(\tau)$ and $\Psi^{\phi}_{\star}(\tau,0)$.
+Dots: simulation results.
+Lines: theoretical predictions.
+All panels describe the same network of $N=4000$ neurons with i.i.d.
+couplings with variance $g^{2}/N$ with coupling strength $g=6$.
+The general program, for i.i.d. or structured couplings, is as follows:
+  1. 1.
+    Formulate the field theory.
+Begin with the path integral $Z[\bm{J}]$ and average over the connectivity disorder.
+Introduce auxiliary fields representing two-point functions (such as $C^{\phi}(t_{1},t_{2})$) along with their conjugate partners.
+Use integral representations of delta functions to enforce field definitions, factorizing the exponential over extensive dimensions.
+This yields a statistical field theory with action $N$ times an order-one function of the auxiliary fields.
+  2. 2.
+    Solve the saddle-point equations.
+Find the saddle point by setting action derivatives to zero.
+This produces self-consistent equations for the two-point functions, completing the standard dynamical mean-field theory (DMFT) analysis.
+  3. 3.
+    Compute the Hessian. Calculate the time2-by-time2 Hessian matrix, which depends on four time variables $(t_{1},t_{2},t_{3},t_{4})$ and characterizes fluctuations around the saddle point.
+  4. 4.
+    Apply temporal separation limits.
+Impose the temporal limits (Eq.
+34), making the Hessian blocks translation-invariant functions of time differences $(\tau_{1},\tau_{2})$.
+Transform to Fourier space to obtain a low-dimensional, frequency-dependent Hessian whose elements are functions of $(\omega_{1},\omega_{2})$.
+  5. 5.
+    Invert the frequency-space Hessian.
+Perform matrix inversion to directly obtain $\Psi^{a}_{\star}(\bm{\omega})$.
+If needed, apply inverse Fourier transform to recover $\Psi^{a}_{\star}(\bm{\tau})$.
+Here we consider i.i.d.
+$\bm{J}$ with mean zero and variance $g^{2}/N$.
 After performing the Gaussian integration over $\bm{J}$, we introduce an auxiliary field $C^{\phi}(t_{1},t_{2})$, defined by Eq.
-5, and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$.
+17, and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$.
 This leads to a partition function for a statistical field theory involving $C^{\phi}(t_{1},t_{2})$ and $\hat{C}^{\phi}(t_{1},t_{2})$:
@@
 and the single-site path integral is
-$W[C^{\phi},\hat{C}^{\phi}]=\int\mathcal{D}x\int\mathcal{D}\hat{x}\\ \times\exp\left(ixT[\hat{x}]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi\right).$ (38)
-In the limit $N\rightarrow\infty$, the saddle point dominates this integral. The derivatives of the action, which are zero at the saddle point, are
+$W[C^{\phi},\hat{C}^{\phi}]=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi\right).$ (76)
+F.1 Two-point functions
+In the limit $N\to\infty$, the saddle point dominates this integral. The derivatives of the action, which are zero at the saddle point, are
 $\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+g^{2}\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W},$ (77)
@@
 This recovers the same single-site process (Eq.
-17, Eq.
-18) and self-consistency condition (Eq.
-19) described in Sec.
-II.3.
-III.3 Four-point functions
+69, Eq.
+70) and self-consistency condition (Eq.
+71) described in Sec.
+E.
+F.2 Four-point functions
 We now compute $\Psi^{a}_{\star}(\bm{\tau})$ within the path-integral formalism.
@@
 Expanding the action to second order around the saddle point, the path integral becomes
-$Z=\exp\left(-N\mathcal{S}[C^{\phi}_{\star},0]\right)\int\mathcal{D}\delta C^{\phi}\mathcal{D}\delta\hat{C}^{\phi}$ (44)
-×$\times\exp\left(-\frac{N}{2}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}^{T}\begin{pmatrix}H_{C^{\phi}C^{\phi}}&H_{C^{\phi}\hat{C}^{\phi}}\\ H_{\hat{C}^{\phi}{C}^{\phi}}^{T}&H_{\hat{C}^{\phi}\hat{C}^{\phi}}\end{pmatrix}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}\right),$
+$Z=\exp\left(-N\mathcal{S}[C^{\phi}_{\star},0]\right)\int\mathcal{D}\delta C^{\phi}\mathcal{D}\delta\hat{C}^{\phi}\exp\left(-\frac{N}{2}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}^{T}\begin{pmatrix}H_{C^{\phi}C^{\phi}}&H_{C^{\phi}\hat{C}^{\phi}}\\ H_{\hat{C}^{\phi}{C}^{\phi}}^{T}&H_{\hat{C}^{\phi}\hat{C}^{\phi}}\end{pmatrix}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}\right)$, (82)
 revealing the structure of the Gaussian fluctuations around the saddle point.
 The covariance matrix among the fluctuation variables, $\delta C^{\phi}(t_{1},t_{2})$ and $\delta\hat{C}^{\phi}(t_{1},t_{2})$, is $1/N$ times the inverse Hessian, yielding
-$N\left\langle\delta C^{\phi}(t_{1},t_{2})\delta C^{\phi}(t_{3},t_{4})\right\rangle_{\star}$ (45)
-$=\left[H_{C^{\phi}C^{\phi}}-H_{C^{\phi}\hat{C}^{\phi}}H_{\hat{C}^{\phi}\hat{C}^{\phi}}^{-1}H_{\hat{C}^{\phi}C^{\phi}}\right]^{-1}(t_{1},t_{2},t_{3},t_{4}).$
+$N\left\langle\delta C^{\phi}(t_{1},t_{2})\delta C^{\phi}(t_{3},t_{4})\right\rangle_{\star}=\left[H_{C^{\phi}C^{\phi}}-H_{C^{\phi}\hat{C}^{\phi}}H_{\hat{C}^{\phi}\hat{C}^{\phi}}^{-1}H_{\hat{C}^{\phi}C^{\phi}}\right]^{-1}(t_{1},t_{2},t_{3},t_{4}).$
 In principle, we have all the necessary components to evaluate this expression: the single-site dynamics are known via the saddle-point condition, and the Hessian blocks are given in terms of connected correlation functions in this single-site process.
@@
 The challenge is that each Hessian block is a complicated time2-by-time2 matrix that, for general values of the time variables, has no analytic inverse.
-To circumvent this problem, we show in Appendix C that the temporal limits of Eq.
+To circumvent this problem, we show in Appendix G that the temporal limits of Eq.
 34 can be taken before taking the inverses.
@@
 $H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-C^{\phi}_{\star}(\tau_{1})C^{\phi}_{\star}(\tau_{2}),$ (83)
-$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ (47)
-$\hskip 40.00006pt=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(-\tau_{1})S^{\phi}_{\star}(-\tau_{2}),$
+$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(\tau_{1})S^{\phi}_{\star}(\tau_{2}),$
 $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0,$ (84)
-where we used the fact that multiplication by $-i\hat{x}(t)$ is equivalent to a functional derivative with respect to a source at time $t$.
+where we used the fact that multiplication by $i\hat{x}(t)$ is equivalent to a functional derivative with respect to a source at time $t$.
 Since the relevant quantities are now translation-invariant, i.e., depend only on $\tau_{1}$ and $\tau_{2}$, we can transform to Fourier space.
@@
 $H_{\hat{C}^{\phi}\hat{C}^{\phi}}(\bm{\omega})$ $=-C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2}),$ (85)
-$H_{C^{\phi}\hat{C}^{\phi}}(\bm{\omega})$ $=-1+g^{2}\left(S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})\right)^{*},$ (50)
+$H_{C^{\phi}\hat{C}^{\phi}}(\bm{\omega})$ $=-1+g^{2}S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2}),$ (86)
 $H_{C^{\phi}C^{\phi}}(\bm{\omega})$ $=0.$ (87)
@@
 which agrees with the two-site cavity result (Eq. 72).
-To summarize, we compute $\Psi_{\star}^{\phi}(\bm{\tau})$ through the following steps (in the next section, we show how to compute $\Psi^{x}_{\star}(\bm{\tau})$):
-  1. 1.
-    Write down the path integral, integrate out disorder, and introduce two-point fields and conjugate partners to factorize the action along extensive indices.
-  2. 2.
-    Sovle the saddle-point equations of the resulting action, yielding the covariance and response functions.
-  3. 3.
-    Derive the frequency-dependent Hessian by computing second derivatives of the action with respect to the two-point fields, applying the temporal limits of Eq.
-34, and transforming to Fourier space.
-Then, determine the element of the inverse frequency-dependent Hessian corresponding to $\Psi^{\phi}_{\star}(\bm{\omega})$.
-This approach allows for the efficient calculation of two- and four-point functions for a given model, requiring only a few lines of computation directly from the action of the theory.
-To demonstrate the flexibility of the path-integral approach in computing $\Psi^{a}_{\star}(\bm{\tau})$, we apply it to compute $\Psi^{\phi}_{\star}(\bm{\tau})$ for a generalization of the i.i.d.
-model in which there are correlations between reciprocal couplings, recovering the result in [25] from the two-site cavity method (Appendix E).
-III.4 Adding sources
+F.3 Adding sources
 Above, computing $\Psi^{\phi}_{\star}(\bm{\tau})$ was simplified by $C^{\phi}(t_{1},t_{2})$ appearing naturally in the statistical field theory resulting from integrating out $\bm{J}$.
 To compute $\Psi^{x}_{\star}(\bm{\tau})$, we need to introduce a source-field term in the action.
+Our goal is to compute
+$\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0},$ (90)
+which gives the fluctuations around the mean of the quantity multiplying $\mathcal{J}$ in the action.
+To facilitate this calculation, we introduce a new field $U$ and set it equal to $\mathcal{J}$ using the conjugate $\hat{U}$:
+$Z[\mathcal{J}]=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp\left(-N\hat{U}U+N\hat{U}\mathcal{J}-NS[\hat{C};U]\right).$ (91)
+Instead of taking the second derivative with respect to $\mathcal{J}$, we set $\mathcal{J}=0$ and compute the fluctuations of $\hat{U}$ in the augmented theory whose path integral is
+[…]
+where
+$\tilde{\mathcal{S}}[\bm{C},U,\hat{U}]=\hat{U}U+\mathcal{S}[\bm{C};U].$ (93)
+The Hessian of the augmented action, in terms of the Hessian of the original action, is
+$\left(\begin{array}[]{c|cc}\bm{H}&\frac{\partial^{2}S}{\partial\mathcal{J}\partial\bm{C}}&\bm{0}\\[2.5pt] \hline\cr\\[-10.00002pt] \left(\frac{\partial^{2}S}{\partial\mathcal{J}\partial\bm{C}}\right)^{T}&\frac{\partial^{2}S}{\partial\mathcal{J}^{2}}&1\\[5.0pt] \bm{0}^{T}&1&0\end{array}\right),$ (94)
+where
+$H_{ab}=\frac{\delta^{2}S}{\delta C_{a}\delta C_{b}},\quad\left[\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{a}=\frac{\delta^{2}S}{\delta\mathcal{J}\delta C_{a}}.$ (95)
+The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point.
 Consider a general intensive action $\mathcal{S}[\bm{C};\mathcal{J}]$ where $\bm{C}$ is a collection of fields (e.g., $\bm{C}=\{C^{\phi},\hat{C}^{\phi}\}$), and $\mathcal{J}$ is a source.
-In Appendix D, we show that
-$\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0}\\ =\left[-\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}+\left(\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right)^{T}\bm{H}^{-1}\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{\mathcal{J}=0},$ (54)
-[…]
-III.5 Path-integral analysis of the random-mode model
-Having demonstrated the path-integral approach to computing two- and four-point functions in well known connectivity models, we now apply this formalism to the random-mode model.
+Using a Schur complement to compute this element, we obtain
+$\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0}=\left[-\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}+\left(\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right)^{T}\bm{H}^{-1}\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{\mathcal{J}=0},$ (96)
+with all quantities evaluated at the saddle point.
+By applying this formula to the action of the i.i.d.
+model with the source term $\frac{1}{2}x\mathcal{J}x$ added to the single-site path integral (Eq.
+76), we recover the expression for $\Psi^{x}_{\star}(\bm{\tau})$ from the two-site cavity method (Clark et al.
+[30]; Eq.
+73).
+For actions of the form $\mathcal{S}[\bm{C};\mathcal{J}]=-\mathcal{J}C_{a}+\text{$\mathcal{J}$-independent terms}$, this formula reduces to $({\delta^{2}}/{\delta\mathcal{J}^{2}})\log Z|_{\mathcal{J}=0}=[\bm{H}^{-1}]_{aa},$ as expected.
+We now apply Eq. 96 to the action of the i.i.d. connectivity model with a source term for correlations among pre-activations. The action is
+$\mathcal{S}[C^{\phi},\hat{C}^{\phi}]=-\frac{1}{2}C^{\phi}\hat{C}^{\phi}-\log\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\bigg{(}i\hat{x}T[x]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi-\frac{1}{2}x\mathcal{J}x\bigg{)}.$ (97)
+Computing the necessary quantities and taking the temporal limits of Eq. 34, we obtain, in Fourier space,
+$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\quad\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}g^{2}\left(S_{12}^{x}\right)^{*}\\ -C_{12}^{x\phi}\end{pmatrix}$, (98)
+where $C^{x\phi}_{12}=C_{\star}^{x\phi}(\omega_{1})C_{\star}^{x\phi}(\omega_{2})$ and $S^{x}_{12}=S^{x}_{\star}(\omega_{1})S^{x}_{\star}(\omega_{2})$.
+Substituting these into Eq.
+96 and using the frequency-dependent Hessian from Eq.
+88, we obtain
+$\Psi_{\star}^{x}=C^{x}_{12}+|U|^{2}C^{\phi}_{12}+UC^{x\phi}_{12}+\text{h.c.},\quad\text{where }U=\frac{g^{2}S^{x}_{12}}{1-g^{2}S^{\phi}_{12}},$ (99)
+in agreement with the two-site cavity result of Clark et al. [30], given in Eq. 73.
+F.4 Other kinds of path integrals
+Different path-integral formalisms have been applied to neural network models depending on the underlying dynamics.
+The Doi-Peliti formalism [93] derives path integrals from master equations for systems with non-negative integer degrees of freedom (such as spiking neural networks) that evolve according to master equations, using an operator formalism as an intermediate step.
+In contrast, we use the MSRDJ formalism for continuous-time stochastic differential equations, which naturally describes our rate-based network dynamics.
+For a comprehensive discussion of different path-integral approaches in neural network theory, see [60].
+Appendix G Temporal limits and inverses
+Here, we demonstrate, for the i.i.d.
+model that the temporal limits in Eq.
+34 can be taken before the inverses.
+This allows us to work with a frequency-dependent Hessian.
+In this section, we use Einstein notation for integrals over time variables (i.e., repeated time indices are integrated over).
+We would like to evaluate
+$\Psi(\tau_{1},\tau_{2})=U^{-1}(t_{1},t_{2},s_{1},s_{2})K(s_{1},s_{2},s_{3},s_{4})U^{-1}(t_{1}+\tau_{1},t_{2}+\tau_{2},s_{3},s_{4})$ (100)
+in the limits Eq.
+34, where $U(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ and $K(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ (note that $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0$ due to the vanishing of correlation functions involving only the conjugate field).
+Restating Eq.
+81, we have
+$U(t_{1},t_{2},t_{3},t_{4})$ $=-\left[\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})\right]-g^{2}\langle\langle\phi(t_{1})\phi(t_{2}),\hat{x}(t_{3})\hat{x}(t_{4})\rangle\rangle_{\star},$ (101)
+$K(t_{1},t_{2},t_{3},t_{4})$ $=\langle\langle\phi(t_{1})\phi(t_{2}),\phi(t_{3})\phi(t_{4})\rangle\rangle_{\star}.$ (102)
+We assume temporal stationarity throughout this section and consider only $a=\phi$.
+We start by examining the properties of $(t_{1},t_{2},t_{3},t_{4}).
+UU$ vanishes if one of the four time points is far (measured on the timescale of $C^{\phi}_{\star}(\tau)$) from the other three.
+For $t_{1}$ or $t_{2}$, this is a consequence of the perturbation associated with $\hat{x}(t_{1})$ or $\hat{x}(t_{2})$ being far away; for $t_{3}$ or $t_{4}$, this is a consequence of $\langle\phi(t_{i})\rangle_{\star}=0$ separating from the expectation.
+$U$ also vanishes if there is more than one time point which is far from the remaining ones by the same arguments.
+Thus, the time points need to be close in pairs.
+However, if $t_{1},t_{2}$ are close and $t_{3},t_{4}$ are close, but the pairs are distant, $U$ still vanishes since the perturbations associated with $\hat{x}(t_{i})$ are both far.
+This leaves three possibilities for a non-vanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
+Separating the three possibilities into two subdomains where all time points are close ($c$) and where the time points belong to either of the two pairwise separate ($s$) possibilities, we write $U=U_{s}+U_{c}$ where $U_{s}$ and $U_{c}$ are defined to vanish on the respective other domain.
+On the separated domain, $U$ simplifies to
+$U_{s}(t_{1},t_{2},t_{3},t_{4})=-\left[\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})\right]\\ +g^{2}\left[S^{\phi}_{\star}(t_{1}-t_{3})S^{\phi}_{\star}(t_{2}-t_{4})+S^{\phi}_{\star}(t_{1}-t_{4})S^{\phi}_{\star}(t_{2}-t_{3})\right],$ (103)
+where we used that the two distant contributions are non-overlapping to add them.
+We assume that we can also decompose the inverse into a contribution for separate pairs and a contribution where all time points are close, $U^{-1}=U_{s}^{-1}+U_{c}^{-1}$, where again $U_{s}^{-1}$ and $U_{c}^{-1}$ vanish on the respective other domain.
+We demonstrate that $U_{s}^{-1}$ is the inverse of $U_{s}$, and that $U_{c}^{-1}$ is the inverse of $U_{c}$, by noting that $U^{-1}(t_{1},t_{2},s_{1},s_{2})U(s_{1},s_{2},t_{3},t_{4})$ reduces to the contraction of $U_{c}^{-1}$ and $U_{c}$ if $t_{3}$ and $t_{4}$, and hence $s_{1}$ and $s_{2}$, are close; or to the contraction of $U_{s}^{-1}$ and $U_{s}$ if $t_{3}$ and $t_{4}$, and hence $s_{1}$ and $s_{2}$, are distant.
+For Eq.
+100, we need $U^{-1}_{s}$ due to the limits Eq.
+34.
+The fact that $U$ reduces to $U_{s}$ in the limits Eq.
+34 justifies taking the limits before inverting $U$.
+Furthermore, these limits imply that, in Eq.
+100, $s_{1}$ and $s_{2}$, as well as $s_{3}$ and $s_{4}$, are far apart.
+In this case, $K$ simplifies to
+$K_{s}(s_{1},s_{2},s_{3},s_{4})=C^{\phi}_{\star}(s_{1}-s_{3})C^{\phi}_{\star}(s_{2}-s_{4})+C^{\phi}_{\star}(s_{1}-s_{4})C^{\phi}_{\star}(s_{2}-s_{3}).$ (104)
+This justifies taking the limits Eq. 34 for $K$.
+Appendix H Partially symmetric disorder
+To demonstrate the flexibility of the path-integral approach to computing the four-point function, we now consider a generalization of the classic i.i.d.
+model in which we introduce correlations between reciprocal couplings, characterized by the following statistics:
+$\left\langle J_{ij}J_{k\ell}\right\rangle=g^{2}\delta_{ik}\delta_{j\ell}+g^{2}\rho\delta_{i\ell}\delta_{jk}.$ (105)
+Here, $\rho\in[-1,1]$ is a parameter that controls the degree of symmetry in the connectivity.
+When $\rho=1$, the connectivity is fully symmetric, while $\rho=-1$ corresponds to fully antisymmetric connectivity.
+The case $\rho=0$ recovers the i.i.d.
+model.
+We express this partially symmetric connectivity as a linear combination of two matrices, $\bm{X}$ and $\bm{Y}$, that are i.i.d.
+random with mean zero and variance $1/N$:
+$J_{ij}=\sqrt{{g^{2}-|\sigma|}}X_{ij}+\sqrt{\frac{|\sigma|}{2}}(Y_{ij}+\text{sgn}(\sigma)Y_{ji}),$ (106)
+where $\sigma=\rho g^{2}$. In terms of $\bm{X}$ and $\bm{Y}$, the path integral for this partially symmetric model is
+$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg{(}i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{ij}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{ij}Y_{ij}(\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j})\bigg{)}.$ (107)
+We integrate out $\bm{X}$ and $\bm{Y}$. To factorize the action over indices spanning extensive dimensions, we introduce
+$C^{\phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\phi_{i}(t_{1})\phi_{i}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{C}^{\phi}(t_{1},t_{2});$ (108)
+$S^{\phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\delta\phi_{i}(t_{1})}{\delta I_{i}(t_{2})},$ $\hskip-72.26999pt\text{conjugate: }\hat{S}^{\phi}(t_{1},t_{2}).$ (109)
+We obtain, for the disorder-averaged path integral,
+$Z$ $=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}R\int\mathcal{D}\hat{R}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]\right),$ (110)
+$\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=-\frac{1}{2}C^{\phi}\hat{C}^{\phi}+\frac{\sigma}{2}S^{\phi}\hat{S}^{\phi}-\log W[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}],$ (111)
+$W[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=\int\mathcal{D}\hat{x}\int\mathcal{D}x\exp\bigg{(}i\hat{x}T[x]-\frac{1}{2}\phi\hat{C}^{\phi}\phi-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{i\sigma}{2}\phi\hat{S}^{\phi}\hat{x}-\frac{i\sigma}{2}\hat{x}S^{\phi}\phi\bigg{)}.$ (112)
+H.1 Two-point functions
+To obtain the saddle-point solution, we compute the derivatives of the action, under the rule that derivatives with respect to $S^{\phi}(t_{1},t_{2})$ also affect $\hat{S}^{\phi}(t_{2},t_{1})$, and vice versa, due to the symmetry in the action.
+This gives
+$\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+g^{2}\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W},$ (113)
+$\frac{\delta\mathcal{S}}{\delta\hat{C}^{\phi}(t_{1},t_{2})}$ $=-{C}^{\phi}(t_{1},t_{2})+\left\langle{\phi}(t_{1}){\phi}(t_{2})\right\rangle_{W},$ (114)
+$\frac{\delta\mathcal{S}}{\delta S^{\phi}(t_{1},t_{2})}$ $=\sigma\hat{S}^{\phi}(t_{1},t_{2})+i\sigma\left\langle\hat{x}(t_{1})\phi(t_{2})\right\rangle_{W},$ (115)
+$\frac{\delta\mathcal{S}}{\delta\hat{S}^{\phi}(t_{1},t_{2})}$ $=\sigma S^{\phi}(t_{1},t_{2})+i\sigma\left\langle\phi(t_{1})\hat{x}(t_{2})\right\rangle_{W}.$ (116)
+Setting these derivatives to zero yields the saddle-point conditions:
+$\hat{C}^{\phi}_{\star}(t_{1},t_{2})$ $=0,$ (117)
+$C^{\phi}_{\star}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (118)
+$S^{\phi}_{\star}(t_{1},t_{2})$ $=\hat{S}^{\phi}_{\star}(t_{2},t_{1})=\left\langle\frac{\delta\phi(t_{1})}{\delta I(t_{2})}\right\rangle_{\star},$ (119)
+where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$.
+The single-site process at the saddle point is described by
+$T[x](t)$ $=\eta^{x}(t)+\sigma\left[S^{\phi}_{\star}\ast\phi\right](t),$ (120)
+$\eta^{x}$ $\sim\mathcal{GP}(0,g^{2}C^{\phi}_{\star}),$ (121)
+where $\ast$ denotes convolution.
+The symmetric structure provides a convolutional, nonlinear self-coupling $\sigma[S^{\phi}_{\star}\ast\phi](t)$ in the single-site dynamics.
+The two-point correlation and response functions, $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$, must be determined self-consistently within this single site picture.
+H.2 Four-point functions
+Introducing the notation $C^{\phi}_{k}=C^{\phi}_{\star}(\omega_{k})$, $S^{\phi}_{k}=S^{\phi}_{\star}(\omega_{k})$ for $k=1,2$, the frequency-dependent Hessian at the saddle point is
+$\bm{H}=\begin{pmatrix}0&-1+g^{2}S^{\phi}_{12}&0&0\\ -1+g^{2}(S^{\phi}_{12})^{*}&-C^{\phi}_{12}&\sigma(S^{\phi}_{1})^{*}C^{\phi}_{2}&\sigma(S^{\phi}_{2})^{*}C^{\phi}_{1}\\ 0&\sigma S^{\phi}_{1}C^{\phi}_{2}&0&\sigma-\sigma^{2}S^{\phi}_{1}(S^{\phi}_{2})^{*}\\ 0&\sigma S^{\phi}_{2}C^{\phi}_{1}&\sigma-\sigma^{2}(S^{\phi}_{1})^{*}S^{\phi}_{2}&0\end{pmatrix}$.
+(122)
+Inverting this matrix and isolating the upper-left element, we obtain, in agreement with Clark et al. [30],
+$\Psi^{\phi}_{\star}=\frac{1}{|1-g^{2}S^{\phi}_{12}|^{2}}\frac{1-|\sigma(S^{\phi}_{1})^{*}S^{\phi}_{2}|^{2}}{|1-\sigma(S^{\phi}_{1})^{*}S^{\phi}_{2}|^{2}}C^{\phi}_{12}.$ (123)
+Appendix I Path-integral calculation of two- and four-point functions for the random-mode model
+Having demonstrated the path-integral approach to computing two- and four-point functions for i.i.d.
+couplings, we now apply this formalism to the random-mode model.
 In terms of the mode matrices $\bm{L}$ and $\bm{R}$, the path integral is
-$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\\ \times\exp\left(i\sum_{i}\hat{x}_{i}T[x_{i}]-i\sum_{i,j}\left[\bm{L}\bm{D}\bm{R}\right]_{ij}\hat{x}_{i}\phi_{j}\right).$ (55)
+$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\left(i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]-i\sum_{i,j=1}^{N}\left[\bm{L}\bm{D}\bm{R}\right]_{ij}\hat{x}_{i}\phi_{j}\right).$ (124)
 We would like to integrate out $\bm{L}$ and $\bm{R}$, but this is complicated by them appearing together in a quadratic term in the action.
 To simplify the integration, we introduce a set of $M$ latent variables,
-$z_{a}(t)=D_{a}\sum_{i}R_{ia}\phi_{i}(t),$ (56)
+$z_{a}(t)=D_{a}\sum_{i=1}^{N}R_{ia}\phi_{i}(t),$ (125)
 and their conjugates $\hat{z}_{a}(t)$. We use indices $i,j$ for neurons and $a,b$ for latent variables. The introduction of these latent variables makes the action linear in $\bm{L}$ and $\bm{R}$:
-$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\int\mathcal{D}\bm{z}\int\mathcal{D}\hat{\bm{z}}\\ \times\exp\Big{(}i\sum_{i}\hat{x}_{i}T[x_{i}]+i\sum_{a}z_{a}\hat{z}_{a}\\ -i\sum_{i,a}L_{ia}\hat{x}_{i}z_{a}-i\sum_{i,a}D_{a}R_{ia}\phi_{i}\hat{z}_{a}\Big{)}.$ (57)
+$Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\int\mathcal{D}\bm{z}\int\mathcal{D}\hat{\bm{z}}\exp\Big{(}i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]+i\sum_{a=1}^{M}z_{a}\hat{z}_{a}-i\sum_{i,a}L_{ia}\hat{x}_{i}z_{a}-i\sum_{i,a}D_{a}R_{ia}\phi_{i}\hat{z}_{a}\Big{)}.$ (126)
 We can now easily integrate out $\bm{L}$ and $\bm{R}$.
 To factorize the action over indices spanning extensive dimensions, we introduce the field $C^{\phi}(t_{1},t_{2})$ (Eq.
-5) with conjugate $\hat{Q}(t_{1},t_{2})$ as well as
-$Q(t_{1},t_{2})=\frac{1}{M}\sum_{a}z_{a}(t_{1})z_{a}(t_{2})$ (58)
+17) with conjugate $\hat{Q}(t_{1},t_{2})$ as well as
+$Q(t_{1},t_{2})=\frac{1}{N}\sum_{a=1}^{M}z_{a}(t_{1})z_{a}(t_{2})$ (127)
 with conjugate $\hat{C}^{\phi}(t_{1},t_{2})$. The disorder-averaged path integral is
-$Z$ $=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}Q\int\mathcal{D}\hat{Q}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},Q,\hat{Q}]\right),$ (59)
+[…]
 $\mathcal{S}[C^{\phi},\hat{C}^{\phi},Q,\hat{Q}$] $=-\frac{1}{2}C^{\phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}-\log W^{x}[Q,\hat{Q}]-\alpha\left\langle\log W^{z}_{D}[C^{\phi},\hat{C}^{\phi}]\right\rangle_{D},$ (129)
-$W^{x}[Q,\hat{Q}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\phi\hat{Q}\phi-\frac{1}{2}\hat{x}Q\hat{x}\right),$ (61)
-$W^{z}_{D}[C^{\phi},\hat{C}^{\phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}z\hat{C}^{\phi}z-\frac{1}{2}D^{2}\hat{z}C^{\phi}\hat{z}\right).$ (62)
+[…]
+[…]
 Thus, analysis of the path integral is made tractable by recasting the problem as interacting neurons and latent variables.
@@
 The statistics of the Gaussian input to each single-site process (neurons or latent variables) are determined by the statistics of the other process, creating a bipartite, mutually referential structure that also arises in the cavity treatment of the problem.
-III.5.1 Two-point functions
-Following the steps outlined at the end of Sec. III.3, we first compute the DMFT by calculating the saddle point. The derivatives of the action are
+I.1 Two-point functions
+Following the steps outlined at the end of Sec. F.2, we first compute the DMFT by calculating the saddle point. The derivatives of the action are
 $\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{Q}^{\phi}(t_{1},t_{2})+\alpha\left\langle D^{2}\left\langle\hat{z}(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D}}\right\rangle_{D},$ (132)
@@
 showing that the neuronal and latent-variable single-site processes have mutually referential statistics  [^8] Note that the variance of $\eta^{z}(t)$ is scaled by $D^{2}$; equivalently, we could factor out $D^{2}$ from the variance and include $D$ as a multiplicative factor in the latent-variable single-site dynamics (Eq.
-71).
+140).
 Solving for $Q_{\star}(t_{1},t_{2})$ by combining Eq.
-69, Eq.
-71, and Eq.
-73, we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq.
-25.
-Consolidating these results, the self-consistent problem that determines $C^{\phi}_{\star}(t_{1},t_{2})$ can be expressed purely in terms of this correlation function as
-$T[x](t)$ $=\eta^{x}(t),$
-$\eta^{x}$ $\sim\mathcal{GP}(0,\alpha r_{2}C^{\phi}_{\star}),$
-$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}.$
-Comparing to Eqs. 1719, we recover the self-consistent equations of the i.i.d. model with coupling strength
-$g_{\text{eff}}=\sqrt{\alpha r_{2}}.$ (74)
-Thus, the low-rank coupling structure remains undetectable at the single-neuron level, where neurons experience effectively i.i.d.
-connections with variance $g_{\text{eff}}^{2}/N$.
-The structure of the random-mode model emerges only in collective neural activity, as we now demonstrate by calculating the four-point function.
-III.5.2 Four-point functions
-The frequency-dependent saddle-point Hessian is
-$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}C_{12}^{\phi}&-1&0\\ \hline\cr 0&-1&0&(S_{12}^{\phi})^{*}\\ -1&0&S_{12}^{\phi}&-C_{12}^{\phi}\end{array}\right).$ (75)
-The upper-left block, corresponding to the latent variables, is related to the Hessian of an i.i.d.
-network with linear dynamics and a heterogeneous distribution of single-neuron scale factors $D_{a}$; the lower-right block, corresponding to the neurons, is related to the Hessian of an i.i.d.
-network of nonlinear neurons (Eq.
-52).
-They are coupled through the simple off-diagonal blocks, reflecting the interaction between neurons and latent variables.
-Finally, we obtain for the four-point function:
-$\Psi_{\star}^{\phi}=\frac{1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}{|1-g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}C_{12}^{\phi}.$ (76)
-To obtain $\Psi^{x}_{\star}$, we add $-\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral (Eq. 61) and compute
-$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\>\>\>\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}0\\ 0\\ S_{12}^{x}\\ -C_{12}^{x\phi}\\ \end{pmatrix}.$ (77)
-Inserting these into Eq. 54 and using a formula (Eq. 149) to invert the Hessian, we obtain
-$\Psi_{\star}^{x}=C^{x}_{12}+\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)|U|^{2}C^{\phi}_{12}\\ +UC^{x\phi}_{12}+\text{h.c.},$ (78)
-where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ (Eq. 21b with $g=g_{\text{eff}}$).
-III.6 Interpreting the formulas
-Figure 6: Dimension of activity and collective timescales in the random-mode model.
-(a) Activity dimension $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ for various coupling strengths $g_{\text{eff}}$.
-Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
-Inset: Extended $g_{\text{eff}}=6$ case, showing convergence to i.i.d.
-coupling behavior with growing $\alpha$.
-(b) Normalized four-point function $\overline{\Psi^{\phi}(\tau,0)}\equiv\Psi^{\phi}(\tau,0)/\Psi^{\phi}(0,0)$ for various $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$.
-Inset: Theory curves for all $\alpha\text{PR}^{D}$, demonstrating relative invariance of collective timescales to the effective rank.
-Simulations use $N=5000$ neurons.
-The analytic formulas for the four-point function of the random-mode model (Eq.
-76 and Eq.
-78) reveal several insights.
-Taking the limit $\alpha\text{PR}^{D}\rightarrow\infty$ for fixed $g_{\text{eff}}$, which leads to elements of $\bm{J}$ being i.i.d., recovers the known formulas for i.i.d.
-couplings (Eq.
-20 and Eq.
-21).
-Comparing to the i.i.d.
-formulas, any structure beyond i.i.d.
-connectivity (i.e., $1/\alpha\text{PR}^{D}$ being finite) strictly decreases the dimension of activity.
-Our analysis reveals two key statistical features of the connectivity that characterize how connectivity shapes activity: the effective coupling strength, $g_{\text{eff}}$, and the effective rank, $\alpha\text{PR}^{D}$.
-The first characterizes the typical magnitude of couplings, while the second characterizes global information about the number of significant connectivity components.
-As in the i.i.d.
-model, $g_{\text{eff}}$ influences both single-neuron activity (described by $C^{a}_{\star}(\tau)$) and collective activity (described by $\Psi^{a}_{\star}(\tau)$).
-In contrast, $\alpha\text{PR}^{D}$ influences collective activity exclusively.
-The random-mode model therefore provides independent control over single-neuron and collective aspects of activity.
-To validate the theoretical prediction for $\Psi^{\phi}_{\star}(\bm{\tau})$ (Eq.
-76), we simulated networks and computed the participation ratio from the empirical equal-time covariance matrix of the activations.
-For the component strengths $D_{a}$, we used a spectrum given by
-$D_{a}\propto\exp(-\beta_{D}a/M)$ (79)
-for $a=1,\ldots,M$, which yields $\text{PR}^{D}=\beta^{-1}_{D}\tanh(\beta_{D})$.
-The corresponding density is $P(D)=1/\beta_{D}D$ with support on $[\exp(-\beta_{D}),1]$.
-To attain a desired effective rank $\alpha\text{PR}^{D}<1$, we set $\alpha=M/N=1$ and solved for $\beta$.
-For $\alpha\text{PR}^{D}>1$, we fixed $\text{PR}^{D}=1$ and increased $\alpha$.
-We adjusted $g_{\text{eff}}$ by rescaling $D_{a}$ (i.e., adjusting the proportionality constant in Eq.
-79).
-Our simulations show excellent agreement with theoretical predictions across parameter settings (Fig.
-6).
-Fig.
-6a illustrates how the dimension of activity, $\text{PR}^{\phi}$, varies with $\alpha\text{PR}^{D}$ for different values of $g_{\text{eff}}$.
-Consistent with i.i.d.-coupling networks, $\text{PR}^{\phi}$ increases monotonically with $g_{\text{eff}}$ [25].
-It also increases monotonically with $\alpha\text{PR}^{D}$, approaching the dimension of a network with i.i.d.
-couplings for large $\alpha\text{PR}^{D}$ (Fig.
-6a inset).
-As $\alpha\text{PR}^{D}$ approaches zero, the dimension of activity also approaches zero.
-We examine the limits $g_{\text{eff}}\rightarrow 1^{+}$ and $\alpha\text{PR}^{D}\rightarrow 0$ in detail in Sec.
-IV.
-Beyond the dimension of activity, which depends on $\Psi^{\phi}_{\star}(\bm{\tau})$ at $\bm{\tau}=(0,0)$, we also examined the temporal profile of this function.
-Fig.
-6b shows a normalized version of $\Psi^{\phi}(\tau,0)$, which is related to the correlation functions of leading principal components, for various values of $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$.
-The decay timescale of $\Psi^{\phi}_{\star}(\tau,0)$ decreases with increasing $g_{\text{eff}}$, approaching a limiting behavior at $g_{\text{eff}}\sim 10$.
-While $\alpha\text{PR}^{D}$ affects the overall scale of this function, and consequently the dimension of activity, it has little effect on the decay timescale, and thus little effect on the timescales of leading principal components (Fig.
-6b, inset).
-Thus, the effective rank primarily influences the dimension of activity rather than the collective temporal structure.
-IV Limiting behavior of the random-mode model
+138, Eq.
+140, and Eq.
+142, we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq.
+10.
+Consolidating these results yields Eq.
+36c.
+I.2 Four-point functions
 Figure 10: Analysis of dimension of activity and covariance spectrum in the low-effective-rank limit.
 (a) Top: Dimension of activity $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ ($\alpha=1$) for various coupling strengths $g_{\text{eff}}$.
-Component strengths $D_{a}$ are constant for $a/N\leq\text{PR}^{D}$ and 00 otherwise.
+Component strengths $D_{a}$ are constant for $a/N\leq\text{PR}^{D}$ and 0 otherwise.
 Small dots: individual simulations; large dots: means over 10 simulations; solid lines: theoretical predictions; dotted lines: small-$\text{PR}^{D}$ expansion, $K^{\phi}(g_{\text{eff}})\mathrm{PR}^{D}$.
@@
 All simulations use $N=6500$ neurons.
-Here we examine the behavior of the random-mode model in two limiting cases: as the system approaches the transition to chaos from above ($g_{\text{eff}}\rightarrow 1^{+}$), and in the limit of low effective rank ($\alpha\text{PR}^{D}\ll 1$).
-We first consider approaching the transition to chaos from above, i.e., $g_{\text{eff}}\rightarrow 1^{+}$.
-The relevant expansion parameter is $g_{\text{eff}}-1\ll 1$.
-For networks with i.i.d.
-couplings, Clark et al.
-[25] showed that, for coupling strength $g$, $\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})\sim 1/(g-1)$ and $\text{PR}^{a}_{\text{i.i.d.}}\sim(g-1)^{3}$ in this limit.
-For the random-mode model, to leading order in $1/(g_{\text{eff}}-1)$, we have
-$\Psi^{a}_{\star}(\bm{\tau})=\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)\underbrace{\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})}_{\sim 1/(g_{\text{eff}}-1)}.$ (80)
-Consequently, to leading order in $g_{\text{eff}}-1$, the dimension of activity behaves as
-$\text{PR}^{a}=\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)^{-1}\underbrace{\text{PR}^{a}_{\text{i.i.d.}}}_{\sim(g_{\text{eff}}-1)^{3}}.$ (81)
-Thus, as the network approaches the transition to chaos, the effect of structure in connectivity is to reduce the dimension by a factor of $(1+1/\alpha\text{PR}^{D})^{-1}$ compared to a network with i.i.d.
-couplings.
-Taking the further limit of $\alpha\text{PR}^{D}\ll 1$ yields the simple relation
-$\text{PR}^{a}=\alpha\text{PR}^{D}\text{PR}^{a}_{\text{i.i.d.}}$ (82)
-The difference between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$ is a higher-order correction, so we also have
-$\text{PR}^{a}=\text{PR}^{S}\text{PR}^{a}_{\text{i.i.d.}}$ (83)
-That is, in the limits of both low effective rank and small effective coupling strength, the dimension of activity is equal to the dimension of the singular-value spectrum times the dimension of activity in an i.i.d.
-network.
-We now consider the limit of low effective rank, $\alpha\text{PR}^{D}\ll 1$, for arbitrary $g_{\text{eff}}$.
-In this limit, the component strengths $D_{a}$ closely approximate the singular values of the coupling matrix.
-To leading order in $\alpha\text{PR}^{D}$,
-$\text{PR}^{a}=K^{a}(g_{\text{eff}})\alpha\text{PR}^{D},$ (84)
+The frequency-dependent saddle-point Hessian is
+$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}C_{12}^{\phi}&-1&0\\ \hline\cr 0&-1&0&S_{12}^{\phi}\\ -1&0&(S_{12}^{\phi})^{*}&-C_{12}^{\phi}\end{array}\right).$ (143)
+The upper-left block, corresponding to the latent variables, is related to the Hessian of an i.i.d.
+network with linear dynamics and a heterogeneous distribution of single-neuron scale factors $D_{a}$; the lower-right block, corresponding to the neurons, is related to the Hessian of an i.i.d.
+network of nonlinear neurons (Eq.
+88).
+They are coupled through the simple off-diagonal blocks, reflecting the interaction between neurons and latent variables.
+This yields the expression for $\Psi^{\phi}_{\star}$ (Eq. 38). To obtain $\Psi^{x}_{\star}$, we add $\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral (Eq. 130) and compute
+$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\>\>\>\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}0\\ 0\\ (S_{12}^{x})^{*}\\ -C_{12}^{x\phi}\\ \end{pmatrix}$. (144)
+Inserting these into Eq. 96 we obtain
+$\Psi_{\star}^{x}=C^{x}_{12}+\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)|U|^{2}C^{\phi}_{12}+UC^{x\phi}_{12}+\text{h.c.},$ (145)
+where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ (Eq. 73b with $g=g_{\text{eff}}$). Here, we have used the following formula to invert the Hessian:
+$\begin{pmatrix}0&a&0&-1\\ a^{*}&b&-1&0\\ 0&-1&0&c\\ -1&0&c^{*}&d\end{pmatrix}^{-1}=\begin{pmatrix}-\frac{b+|c|^{2}d}{|1-ac|^{2}}&-\frac{c^{*}}{1-(ac)^{*}}&-\frac{ad+bc^{*}}{|1-ac|^{2}}&-\frac{1}{1-(ac)^{*}}\\ -\frac{c}{1-ac}&0&-\frac{1}{1-ac}&0\\ -\frac{a^{*}d+bc}{|1-ac|^{2}}&-\frac{1}{1-(ac)^{*}}&-\frac{|a|^{2}d+b}{|1-ac|^{2}}&-\frac{a^{*}}{1-(ac)^{*}}\\ -\frac{1}{1-ac}&0&-\frac{a}{1-ac}&0\end{pmatrix}$.
+(146)
+Appendix J Path-integral calculation of two- and four-point functions for the random-mode model with single-unit heterogeneity
+We incorporate the generalized nonlinearity $\Phi_{\bm{\theta}_{i}}(x)$ into the path integral, as well as a source term for the normalized variables, and integrate out $\bm{L}$ and $\bm{R}$.
+To factorize the action over indices spanning extensive dimensions, we introduce the order parameters
+$C^{\Phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\Phi_{\bm{\theta}_{i}}(t_{1})\Phi_{\bm{\theta}_{i}}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{Q}(t_{1},t_{2});$ (147)
+$Q(t_{1},t_{2})$ $=\frac{1}{N}\sum_{a=1}^{M}z_{a}(t_{1})z_{a}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{C}^{\phi}(t_{1},t_{2}),$ (148)
+where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}(x_{i}(t))$. The resulting action and single-site path integrals for this system are
+$\mathcal{S}[C^{\Phi},\hat{C}^{\Phi},Q,\hat{Q}$] $=-\frac{1}{2}C^{\Phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}-\left\langle\log W_{\bm{\theta}}[Q,\hat{Q};\mathcal{J}]\right\rangle_{\bm{\theta}}-\alpha\left\langle\log W^{z}_{D}[C^{\Phi},\hat{C}^{\Phi}]\right\rangle_{D},$ (149)
+$W^{x}_{\bm{\theta}}[Q,\hat{Q};\mathcal{J}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\Phi_{\bm{\theta}}\hat{Q}\Phi_{\bm{\theta}}-\frac{1}{2}\hat{x}Q\hat{x}-\frac{1}{2}\phi\mathcal{J}\phi\right),$ (150)
+[…]
+J.1 Two-point functions
+The saddle-point calculation proceeds similarly to the non-heterogeneous case of the random-mode model. We obtain
+$\hat{C}^{\Phi}_{\star}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (152)
+$C^{\Phi}_{\star}(t_{1},t_{2})$ $=\left\langle\left\langle\Phi_{\bm{\theta}}(t_{1})\Phi_{\bm{\theta}}(t_{2})\right\rangle_{\star}\right\rangle_{\bm{\theta}},$ (153)
+$Q_{\star}(t_{1},t_{2})$ $=\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}\right\rangle_{D},$ (154)
+where $\left\langle\cdots\right\rangle_{\star}$ and $\left\langle\cdots\right\rangle_{\star D}$ are averages with respect to $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ and $W^{z}_{D}[C^{\Phi}_{\star},0]$, respectively (note that $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ does not depend on $\bm{\theta}$).
+Using $\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}=D^{2}C^{\Phi}_{\star}(t_{1},t_{2})$ and, consequently, $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\Phi}_{\star}(t_{1},t_{2})$, we obtain the self-consistent single-site dynamic process
+$T[x](t)$ $=\eta^{x}(t),$ (155a)
+$\eta^{x}$ $\sim\mathcal{GP}(0,\alpha r_{2}C^{\Phi}_{\star}),$ (155b)
+$C^{\Phi}_{\star}(t_{1},t_{2})$ $=\left\langle\left\langle\Phi_{\bm{\theta}}(t_{1})\Phi_{\bm{\theta}}(t_{2})\right\rangle_{\star}\right\rangle_{\bm{\theta}}.$ (155c)
+The self-consistency condition Eq.
+155c is identical to that of an i.i.d.
+network, but with an average over $\bm{\theta}\sim P(\bm{\theta})$, in addition to the usual Gaussian average over $\eta^{x}(t)$, to account for single-neuron heterogeneity.
+J.2 Four-point functions
+To write down the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$:
+$\mathfrak{C}^{\Phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\Phi_{\bm{\theta}}\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\Phi_{\bm{\theta}}\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (156a)
+$\mathfrak{s}^{\Phi}_{12}$ $=\left\langle\left\langle\frac{\delta\Phi_{\bm{\theta}}}{\delta I}\right\rangle_{\star}(\omega_{1})\left\langle\frac{\delta\Phi_{\bm{\theta}}}{\delta I}\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (156b)
+$\mathfrak{C}^{\Phi\phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (156c)
+along with the usual shorthand $C^{\Phi}_{12}=C^{\Phi}_{\star}(\omega_{1})C^{\Phi}_{\star}(\omega_{2})$.
+The frequency-dependent Hessian at the saddle point is
+$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}{C_{{12}}^{\Phi}}&-1&0\\ \hline\cr 0&-1&0&\mathfrak{s}^{\Phi}_{12}\\ -1&0&(\mathfrak{s}^{\Phi}_{12})^{*}&-\mathfrak{C}^{\Phi}_{12}\end{array}\right),$ (157)
+where the new $\mathfrak{F}$raktur variables are defined in Eq. 156. We invert the Hessian using Eq. 146 and identify the upper-left element as $\Phi_{\bm{\theta}_{i}}(\bm{\omega})$:
+$\Psi^{\Phi}_{\star}=\frac{\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\Phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}|\alpha r_{2}\mathfrak{s}^{\Phi}_{12}|^{2}}{|1-\alpha r_{2}\mathfrak{s}^{\Phi}_{12}|^{2}}C^{\Phi}_{12}.$ (158)
+Combining the inverted Hessian with the relevant quantities for computing fluctuations of normalized variables,
+$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C^{\phi}_{12},\>\>\>\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}0\\ 0\\ (S^{\phi}_{12})^{*}\\ -\mathfrak{C}^{\Phi\phi}_{12}\end{pmatrix}$, (159)
 where
+$\mathfrak{C}^{\Phi\phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (160)
+$S^{\phi}_{12}$ $=\left\langle\frac{\delta\phi}{\delta I}\right\rangle_{\star}(\omega_{1})\left\langle\frac{\delta\phi}{\delta I}\right\rangle_{\star}(\omega_{2}),$ (161)
+we obtain for the normalized variables $\phi_{i}(t)$:
+$\Psi^{\phi}_{\star}=\Bigg{[}|A|^{2}\left(\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}\frac{C^{\Phi}_{12}}{C^{\phi}_{12}}\right)+\left(A\frac{\mathfrak{C}^{\Phi\phi}_{12}}{C^{\phi}_{12}}+\text{h.c.}\right)+1\Bigg{]}C^{\phi}_{12},$ (162)
+where
+$A=\frac{\alpha r_{2}S^{\phi}_{12}}{1-\alpha r_{2}\mathfrak{s}^{\phi}_{12}}.$ (163)
+J.3 Four-point functions for firing-rate heterogeneity
+Upon specializing the above general results to the case where $\bm{\theta}_{i}$ contains a single gain parameter, $G_{i}$, we obtain the four-point functions for unnormalized (Eq.
+46a) and normalized (Eq.
+46b) variables given in the main text.
+Note that, like the distribution over component strengths $D_{a}$, two-point functions of activity depend only on the second moment of $G_{i}$, and four-point functions depend only on the second and fourth moments of $G_{i}$.
+For “random readout” gains with the same distribution as the recurrent gains (see main text), the four-point function is
+$\Psi^{\Phi^{\text{readout}}}=\Bigg{[}\left(\frac{1}{\text{PR}^{G}}-1\right)\left(|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}+|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\right)+1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\Bigg{]}\frac{C^{\Phi^{\text{readout}}}_{12}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}},$ (164)
+where $C^{\Phi^{\text{readout}}}_{12}=C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$
+We conclude with a note on gain modulation in recurrent neural networks.
+It has been shown that modulating single-neuron gains while keeping synaptic weights fixed has great expressive power over the dynamics of recurrent neural networks [95].
+Here, we consider random gains rather than judiciously chosen or learned gains.
+In particular, they are sampled independently of the connectivity.
+However, our framework could be extended to model gains that are related to the connectivity and, more ambitiously, could be extended to model learned gain parameters (or, as explored in the Discussion, learned connectivity).
+Appendix K Path-integral calculation of two- and four-point functions for the random-mode model with diagonally structured overlaps
+To generate matrices $\bm{L}$ and $\bm{R}$ with appropriate correlation structure, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$:
+${L}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{1}_{ia}+\sqrt{{|\rho_{a}|}}{Y}_{ia},$ (165)
+${R}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{2}_{ia}+\text{sgn}(\rho_{a})\sqrt{{|\rho_{a}|}}{Y}_{ia}.$ (166)
+We begin with the path integral containing latent variables, insert this parameterization of $\bm{L}$ and $\bm{R}$, then and integrate out $\bm{X}_{1}$, $\bm{X}_{2}$, and $\bm{Y}$.
+To factorize the action over indices spanning extensive dimensions, we introduce
+$C^{\phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\phi_{i}(t_{1})\phi_{i}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{Q}(t_{1},t_{2});$ (167)
+$S^{\phi}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\delta\phi_{i}(t_{1})}{\delta I_{i}(t_{2})},$ $\hskip-72.26999pt\text{conjugate: }\hat{R}(t_{1},t_{2});$ (168)
+$Q(t_{1},t_{2})$ $=\frac{1}{N}\sum_{a=1}^{M}z_{a}(t_{1})z_{a}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{C}^{\phi}(t_{1},t_{2});$ (169)
+$R(t_{1},t_{2})$ $=\frac{1}{N}\sum_{a=1}^{M}D_{a}\rho_{a}\frac{\delta z_{a}(t_{1})}{\delta I_{a}(t_{2})},$ $\hskip-72.26999pt\text{conjugate: }\hat{S}^{\phi}(t_{1},t_{2}).$ (170)
+The resulting path integral, action, and single-site path integrals are given by:
+[…]
+$\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi},Q,\hat{Q},R,\hat{R}$] $=-\frac{1}{2}C^{\phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}+\frac{1}{2}S^{\phi}\hat{R}+\frac{1}{2}R\hat{S}^{\phi}$ (172)
+$\hskip 100.00015pt-\log W^{x}[Q,\hat{Q},R,\hat{R}]-\alpha\left\langle\log W^{z}_{D,\rho}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]\right\rangle_{D,\rho},$ (173)
 […]
 […]
-We match this theory of limiting behavior to simulated networks with $\alpha=1$ and a step function for the component strengths:
-$D_{a}=\begin{cases}\text{const.}&\frac{a}{N}\leq\text{PR}^{D}\\ 0&\frac{a}{N}>\text{PR}^{D}.\end{cases}$ (87)
-We varied $\text{PR}^{D}$ and verified its linear relationships with $\text{PR}^{\phi}$ and $\text{PR}^{x}$ in the $\text{PR}^{D}\rightarrow 0$ limit (Fig. 7a).
-Focusing on $K^{\phi}(g_{\text{eff}})$, note that its expression (Eq.
-85) resembles the formula for the dimension of activity in a i.i.d.
-network with coupling strength $g_{\text{eff}}$, but includes an additional factor $|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}$ in the integral (compare to Eq.
-20).
-Numerical evaluation of $K^{\phi}(g_{\text{eff}})$ shows that it approaches $\text{PR}^{\phi}_{\text{i.i.d.}}$ as $g_{\text{eff}}\rightarrow 1^{+}$ as expected from Eq.
-82.
-As $g_{\text{eff}}$ increases, $K^{\phi}(g_{\text{eff}})$ increases monotonically to approximately $1.53\,\text{PR}^{\phi}_{\text{i.i.d.}}$ (Fig.
-7a, inset).
-These results show that for networks with low effective rank, $\text{PR}^{\phi}\sim\alpha\text{PR}^{D}\text{PR}^{\phi}_{\text{i.i.d.}}$ up to an order-one fudge factor that becomes exactly one as $g_{\text{eff}}\rightarrow 1^{+}$.
-We also investigated the relationship between the eigenvalues $\lambda^{a}_{i}$ of $C^{a}_{ij}(0)$ and the component strengths $D_{a}$.
-In the low-effective-rank limit, the quantities $\text{PR}^{a}$ and $\alpha\text{PR}^{D}$ computed from these spectra exhibit a linear relationship (with proportionality factor $K^{a}(g_{\text{eff}})$), implying that the ratio of their second-squared and fourth moments are tightly linked.
-However, we find that the full forms of the eigenvalue and component-strength spectra differ markedly.
-Whereas the component strengths $D_{a}$ decay according to a step function in this example, the rank-ordered eigenvalues $\lambda^{\phi}$ and $\lambda^{x}$ decay smoothly (Fig.
-7b).
-Thus, the dynamics transform the connectivity spectrum into an activity spectrum with a different form while preserving the linear relationship between $\text{PR}^{a}$ and $\alpha\text{PR}^{D}$.
-V Cavity analysis of the random-mode model
+K.1 Two-point functions
+We first calculate the saddle point conditions by taking the derivatives of the action:
+$\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{Q}(t_{1},t_{2})+\alpha\left\langle D^{2}\left\langle\hat{z}(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (176)
+$\frac{\delta\mathcal{S}}{\delta\hat{C}^{\phi}(t_{1},t_{2})}$ $=-Q(t_{1},t_{2})+\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (177)
+$\frac{\delta\mathcal{S}}{\delta S^{\phi}(t_{1},t_{2})}$ $=\hat{R}(t_{1},t_{2})+i\alpha\left\langle D\rho\left\langle\hat{z}(t_{1})z(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (178)
+$\frac{\delta\mathcal{S}}{\delta\hat{S}^{\phi}(t_{1},t_{2})}$ $=R(t_{1},t_{2})+i\alpha\left\langle D\rho\left\langle z(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (179)
+$\frac{\delta\mathcal{S}}{\delta Q(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W^{x}},$ (180)
+$\frac{\delta\mathcal{S}}{\delta\hat{Q}(t_{1},t_{2})}$ $=-C^{\phi}(t_{1},t_{2})+\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{W^{x}},$ (181)
+$\frac{\delta\mathcal{S}}{\delta R(t_{1},t_{2})}$ $=\hat{S}^{\phi}(t_{1},t_{2})+i\left\langle\hat{x}(t_{1})\phi(t_{2})\right\rangle_{W^{x}},$ (182)
+$\frac{\delta\mathcal{S}}{\delta\hat{R}(t_{1},t_{2})}$ $=S^{\phi}(t_{1},t_{2})+i\left\langle\phi(t_{1})\hat{x}(t_{2})\right\rangle_{W^{x}}.$ (183)
+The saddle-point conditions yield:
+$\hat{C}^{\phi}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (184)
+$C^{\phi}_{\star}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (185)
+$S^{\phi}_{\star}(t_{1},t_{2})$ $=\hat{S}^{\phi}_{\star}(t_{2},t_{1})=\left\langle\frac{\delta\phi(t_{1})}{\delta I(t_{2})}\right\rangle_{\star},$ (186)
+$Q_{\star}(t_{1},t_{2})$ $=\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D,\rho}\right\rangle_{D,\rho},$ (187)
+$R_{\star}(t_{1},t_{2})$ $=\hat{R}_{\star}(t_{2},t_{1})=\alpha\left\langle D\rho\left\langle\frac{\delta z(t_{1})}{\delta I(t_{2})}\right\rangle_{\star D,\rho}\right\rangle_{D,\rho},$ (188)
+where $\left\langle\cdots\right\rangle_{\star D,\rho}$ and $\left\langle\cdots\right\rangle_{\star}$ are averages within the dynamic processes described by $W^{z}_{D,\rho}[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$ and $W^{x}[Q_{\star},0,R_{\star},\hat{R}_{\star}$], respectively.
+The yields neuronal and latent-variable single-site processes described by
+$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\ast\phi](t)$ (189)
+$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\ast z](t),$ (190)
+where the Gaussian fields have statistics
+$\eta^{x}$ $\sim\mathcal{GP}(0,Q_{\star}),$ (191)
+$\eta^{z}$ $\sim\mathcal{GP}(0,D^{2}C^{\phi}_{\star}).$ (192)
+This gives rise to the single-site problem stated in Sec. VII.
+K.2 Four-point functions
+The frequency-dependent Hessian at the saddle point is
+$\bm{H}=\left(\begin{array}[]{cccc|cccc}0&u&0&0&0&-1&0&0\\ u^{*}&-vC^{\phi}_{1}C^{\phi}_{2}&x^{*}_{12}C^{\phi}_{2}&x^{*}_{21}C^{\phi}_{1}&-1&0&0&0\\ 0&x_{12}C^{\phi}_{2}&0&-w^{*}&0&0&0&1\\ 0&x_{21}C^{\phi}_{1}&-w&0&0&0&1&0\\ \hline\cr 0&-1&0&0&0&S^{\phi}_{1}S^{\phi}_{2}&0&0\\ -1&0&0&0&(S^{\phi}_{1}S^{\phi}_{2})^{*}&-C^{\phi}_{1}C^{\phi}_{2}&(S^{\phi}_{1})^{*}C^{\phi}_{2}&(S^{\phi}_{2})^{*}C^{\phi}_{1}\\ 0&0&0&1&0&S^{\phi}_{1}C^{\phi}_{2}&0&-S^{\phi}_{1}(S^{\phi}_{2})^{*}\\ 0&0&1&0&0&S^{\phi}_{2}C^{\phi}_{1}&-(S^{\phi}_{1})^{*}S^{\phi}_{2}&0\end{array}\right)$ (193)
+where
+$u$ $=\alpha\left\langle D^{2}\sigma_{1}\sigma_{2}\right\rangle_{D,\rho},$ (194)
+$v$ $=\alpha\left\langle D^{4}|\sigma_{1}|^{2}|\sigma_{2}|^{2}\right\rangle_{D,\rho},$ (195)
+$w$ $=\alpha\left\langle D^{2}\rho^{2}\sigma_{1}^{*}\sigma_{2}\right\rangle_{D,\rho},$ (196)
+$x_{12}$ $=\alpha\left\langle D^{3}\rho\sigma_{1}|\sigma_{2}|^{2}\right\rangle_{D,\rho},$ (197)
+$\sigma_{k}$ $=\frac{1}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (198)
+Isolating the upper-left element of the inverse of this matrix yields Eq. 51.
+K.3 Sigmoidal parameterization of overlaps
+To model structured left–right mode overlaps, we used the following sigmoidal form:
+$\rho_{a}=\gamma_{\rho}\left(1-\frac{2}{1+\exp(-\beta_{\rho}(a/N-u_{\text{flip}}))}\right),$ (199)
+where $\rho_{a}=0$ at $a/N=u_{\text{flip}}$.
+The parameter $\gamma_{\rho}$ sets the overall magnitude and sign of the overlaps, while $\beta_{\rho}$ controls the steepness of the transition.
+We fixed $\beta_{\rho}=16$, varied $\gamma_{\rho}$, and set $u_{\text{flip}}$ such that $\left\langle D\rho\right\rangle_{D,\rho}=0$, thereby eliminating strong self-couplings.
+Appendix L Two-site cavity calculation of four-point functions for the random-mode model
 Figure 11: Schematic of the two-site cavity approach for the bipartite representation of the random-mode model.
@@
 While the path-integral approach provides a systematic framework for analyzing neural-network dynamics, the cavity method offers an intuitive, complementary perspective.
-This section provides an overview of the cavity analysis for the random-mode model, with the full mathematical details available in Appendix G.
+This section provides an overview of the cavity analysis for the random-mode model, with the full mathematical details available in Appendix L.
 While these approaches yield equivalent results, they differ in their methodology: the path-integral approach focuses on fluctuations around the saddle point in a field theory, whereas the cavity approach examines perturbations induced by a subset of held-out variables.
+Here we compute the four-point function using a two-site cavity approach, noting that the two-point function (which, if desired, could be computed using a single-site cavity method) is the same as that of an i.i.d.
+network with appropriate $g_{\text{eff}}$.
 Leveraging Eq.
-56, we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al.
+125, we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al.
 [30] to both groups.
 This leads to separate, mutually referential cavity pictures for each group, mirroring the two single-site path integrals (Sec.
-III.5).
+IV).
 As in the path-integral formalism, self-consistent expressions in one picture are defined using averages from the other.
 Unlike the path-integral approach, which derives $\Psi^{\phi}(\bm{\tau})$ using its time-by-time definition (Eq.
-9), the cavity method derives this function using its neuron-by-neuron definition (Eq.
-7).
+25), the cavity method derives this function using its neuron-by-neuron definition (Eq.
+24).
 The cavity method we employ, schematized in Fig.
-8, has two key features that distinguish it from simpler cavity calculations.
+11, has two key features that distinguish it from simpler cavity calculations.
 First, it uses a two-site structure: we consider the simultaneous removal of two neurons, or two latent variables, from the network, rather than just one.
-This allows us to study pairwise correlations between neurons.
+This allows us to study cross-correlations between neurons.
 This approach was previously used in Clark et al.
@@
 Note that these two features are distinct: the two-site aspect refers to the number of neurons removed in each cavity, while the bipartite aspect refers to the types of variables considered (neurons and latent variables).
-VI Single-neuron heterogeneity
-Figure 9: Effect of single-neuron heterogeneity on dimension of activity in the random-mode model.
-(a) Dimension of normalized activity $\text{PR}^{\phi}$ versus participation ratio of gain distribution $\text{PR}^{G}$ for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$.
-Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
-(b) Fractional reduction in dimension for weighted activations relative to $\text{PR}^{\phi}$.
-Blue: $G_{i}\phi_{i}$; purple: $G^{\text{readout}}_{i}\phi_{i}$.
-$g_{\text{eff}}=10$.
-Thick dots: means over 10 simulations, averaging before taking the ratio; lines: theoretical predictions.
-All simulations use $N=5000$ neurons.
-While the random-mode model introduces heterogeneity in connectivity through the component strengths $D_{a}$, individual neurons may also exhibit heterogeneous properties.
-For instance, cortical neurons display a broad distribution of mean firing rates [63].
-To compare the effects of low-dimensional structure in connectivity with the effects of heterogeneous neuronal properties, we extend our analysis to networks with diverse single-neuron characteristics.
-In particular, we generalize the neuronal nonlinearity to $\Phi_{\bm{\theta}}(x)$, which depends on a vector of parameters $\bm{\theta}$.
-The parameter vector for the $i$-th neuron, $\bm{\theta}_{i}$, is sampled i.i.d.
-from a distribution $P(\bm{\theta})$.
-This formulation allows us to model diverse neuronal properties such as gains or biases.
-When analyzing neural data, it is common to normalize neuronal activities (e.g., through $z$-scoring) to put them on the same scale.
-To facilitate analysis of such normalized activations, we also introduce a set of variables $\phi_{i}(t)=\phi(x_{i}(t))$ that do not depend on $\bm{\theta}_{i}$.
-We incorporate the generalized nonlinearity into the path integral, as well as a source term for the normalized variables, and integrate out $\bm{L}$ and $\bm{R}$.
-To factorize the action over indices spanning extensive dimensions, we introduce, among others, the order parameter
-$C^{\Phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i}\Phi_{\bm{\theta}_{i}}(t_{1})\Phi_{\bm{\theta}_{i}}(t_{2}),$ (88)
-where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}(x_{i}(t))$.
-In Appendix H, we derive a single-site dynamic process $T[x](t)=\eta^{x}(t)$, where $\eta^{x}\sim\mathcal{GP}(0,\alpha r_{2}C^{\Phi}_{\star})$, with the self-consistency condition
-$C^{\Phi}_{\star}(t_{1},t_{2})=\left\langle\left\langle\Phi_{\bm{\theta}}(t_{1})\Phi_{\bm{\theta}}(t_{2})\right\rangle_{\star}\right\rangle_{\bm{\theta}}.$ (89)
-This condition is identical to that of an i.i.d.
-network, but with an average over $\bm{\theta}\sim P(\bm{\theta})$, in addition to the usual Gaussian average over $\eta^{x}(t)$, to account for single-neuron heterogeneity.
-To write down the solution the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$:
-$\mathfrak{C}^{\Phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\Phi_{\bm{\theta}}\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\Phi_{\bm{\theta}}\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (90a)
-$\mathfrak{s}^{\Phi}_{12}$ $=\left\langle\left\langle\frac{\delta\Phi_{\bm{\theta}}}{\delta I}\right\rangle_{\star}(\omega_{1})\left\langle\frac{\delta\Phi_{\bm{\theta}}}{\delta I}\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (90b)
-$\mathfrak{C}^{\Phi\phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (90c)
-along with the usual shorthand $C^{\Phi}_{12}=C^{\Phi}_{\star}(\omega_{1})C^{\Phi}_{\star}(\omega_{2})$. For the unnormalized variables $\Phi_{\bm{\theta}_{i}}(t)$, we obtain
-$\Psi^{\Phi}_{\star}=\frac{\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\Phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}|\alpha r_{2}\mathfrak{s}^{\Phi}_{12}|^{2}}{|1-\alpha r_{2}\mathfrak{s}^{\Phi}_{12}|^{2}}C^{\Phi}_{12}.$ (91)
-For the normalized variables $\phi_{i}(t)$, we obtain
-$\Psi^{\phi}_{\star}=\Bigg{[}|A|^{2}\left(\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}\frac{C^{\Phi}_{12}}{C^{\phi}_{12}}\right)\\ +\left(A\frac{\mathfrak{C}^{\Phi\phi}_{12}}{C^{\phi}_{12}}+\text{h.c.}\right)+1\Bigg{]}C^{\phi}_{12},$ (92)
-where $A={\alpha r_{2}S^{\phi}_{12}}/{(1-\alpha r_{2}\mathfrak{s}^{\phi}_{12})}$.
-VI.1 Diverse single-neuron firing rates
-An illustrative case of single-neuron heterogeneity occurs when the $\bm{\theta}$ parameters consist of a gain parameter $G_{i}$ for each neuron, modeling diversity in mean firing rates:
-$\Phi_{\bm{\theta}_{i}}(x)=G_{i}\phi(x).$ (93)
-In this scenario, the dynamics of the network are determined by an effective coupling matrix $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$, where $\bm{G}=\text{diag}(G_{1},\ldots,G_{N})$.
-This extension of the random-mode model allows us to directly compare the effects of low-dimensional structure in the connectivity (via the component strengths $D_{a}$) with heterogeneity in single-neuron properties (via the gains $G_{i}$).
-We assume the gains are drawn from a distribution with moments
-$q_{n}=\left\langle G^{n}\right\rangle_{G}.$ (94)
-To quantify the heterogeneity in firing rates, we define the participation ratio of the gain distribution:
-$\text{PR}^{G}=\frac{q_{2}^{2}}{q_{4}}.$ (95)
-We analyze collective activity for both the unnormalized activations $\Phi_{\bm{\theta}_{i}}(t)$ and the normalized activations $\phi_{i}(t)$.
-The latter are analogous to $z$-scored firing rates in neural recordings, where $z$-scoring removes single-neuron heterogeneity while preserving nontrivial structure due to cross-neuron and temporal correlations.
-For the unnormalized variables, we calculate the four-point function
-$\Psi^{\Phi}_{\star}=\frac{\frac{1}{\text{PR}^{G}}+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\Phi}_{12},$ (96)
-and for the normalized variables,
-$\Psi^{\phi}_{\star}=\frac{1+\left(\frac{1}{\text{PR}^{G}}+\frac{1}{\alpha\text{PR}^{D}}-1\right)|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12},$ (97)
-where
-$g_{\text{eff}}$ $=\sqrt{\alpha r_{2}q_{2}},$ (98)
-$C^{\Phi}_{12}$ $=q_{2}^{2}C^{\phi}_{12}.$ (99)
-Eq.
-97 demonstrates that reductions in effective rank, described by $\alpha\text{PR}^{D}$, and in the heterogeneity of gains, described by $\text{PR}^{G}$, have symmetric effects on the dimension of normalized activity, with their combined effect captured by the harmonic mean, $\left(\frac{1}{\alpha\text{PR}^{D}}+\frac{1}{\text{PR}^{G}}\right)^{-1}$.
-This symmetry arises because swapping the distributions of $D_{a}$ and $G_{i}$ does not change the spectrum of $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ when $\bm{L}$ and $\bm{R}$ are i.i.d., since $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ and $\bm{R}^{T}\bm{G}\bm{L}\bm{D}$ have the same nonzero eigenvalues.
-One might expect that normalizing the activations before computing the dimension of activity would remove the effect of heterogeneous gains on the dimension.
-This is not the case, however, since neuronal gains affect the recurrent dynamics, and not merely the readout.
-To further explore this and related effects, we compare the activity dimensions for three sets of activations: (1) unnormalized activations, $\Phi_{\bm{\theta}_{i}}(t)$; (2) normalized activations, $\phi_{i}(t)$; and (3) normalized activations multiplied by an independent set of “readout” gains, $\Phi^{\text{readout}}_{i}(t)=G^{\text{readout}}_{i}\phi_{i}(t)$, where the readout gains have the same distribution as the actual gains.
-Case (3) corresponds to variables with the same distribution of firing rates as the unnormalized activations, but where this heterogeneity is unrelated to the recurrent dynamics.
-In this case, we have $C^{\Phi^{\text{readout}}}_{12}=C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$.
-The four-point function is
-$\Psi^{\Phi^{\text{readout}}}=\Bigg{[}\left(\frac{1}{\text{PR}^{G}}-1\right)\left(|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}+|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\right)\\ +1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\Bigg{]}\frac{C^{\Phi^{\text{readout}}}_{12}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}.$ (100)
-To validate our theoretical predictions (Eq. 96, Eq. 97, and Eq. 100), we simulated networks with component strengths and gains given by
-$D_{a}$ $\propto\exp(-\beta_{D}a/M),$ (101)
-$G_{i}$ $\propto\exp(-\beta_{G}i/N).$ (102)
-The dimension of activity in these networks is determined by the coupling strength, $g_{\text{eff}}$; the effective rank of the coupling matrix, $\alpha\text{PR}^{D}$; and the participation ratio of the single-neuron gain distribution, $\text{PR}^{G}$.
-Increases in each of these parameters leads to higher activity dimension for all three sets of activations.
-Normalized activations consistently exhibit the highest dimension.
-While scaling normalized activations by heterogeneous factors reduces dimension as expected, the magnitude of this reduction depends on whether $G_{i}$ or $G^{\text{readout}}_{i}$ is used, even when their distributions are identical.
-This occurs because neurons with the largest gains preferentially participate in the leading modes of the normalized activations.
-Further scaling these already-dominant neurons by their gains results in overrepresentation of these neurons compared to using random gains.
-Thus, the dimension of unnormalized activations is lowest, with the dimension of random-readout activations falling between those of the normalized and unnormalized activations.
-This effect demonstrates that shuffling the normalizations of recorded neural activities should increase the dimension if firing rates and input covariances are related to one another, as they are in recurrent networks.
-Not observing such an increase in dimension, on the other hand, would suggest that heterogeneity in firing rates is unrelated to input covariance.
-VII Mode overlaps
-Figure 10: Input-output mode overlaps and eigenvalue spectra of the fly-brain connectome.
-(a) Overlap matrix between right and left singular vectors of the connectome.
-First 1000 modes are shown.
-(b) First 3000 diagonal elements of the overlap matrix from (a).
-Horizontal lines indicate an approximate cutoff for $\mathcal{O}(1/\sqrt{N})$ random overlaps.
-(c) Eigenvalue spectra of various connectome-based matrices.
-Top: Actual hemibrain spectrum, exhibiting large outliers along the real axis.
-Middle: Eigenvalues from a realization of the random-mode model, using singular values from the connectome as component strengths $D_{a}$, resulting in a circularly symmetric spectrum.
-Bottom: Random-mode model realization with input-output mode overlaps incorporated, where overlaps are given by the inner products of left and right singular vectors from the fly-connectome SVD.
-Data from Scheffer et al.
-[21], $N=18028$ neurons.
-Thus far, we have assumed that the matrices $\bm{L}$ and $\bm{R}$ are independent of each other.
-Under this assumption, the overlap $\bm{L}_{a}^{T}\bm{R}_{b}$ between input mode $b$ and output mode $a$ is a random, $\mathcal{O}(1/\sqrt{N})$ quantity.
-However, to implement computations, modes must interact, suggesting that the coupling matrices of real neural circuits have larger, structured overlaps between input and output modes.
-To examine interactions between different modes in a real neural circuit, we returned to the connectome of Drosophila melanogaster.
-We computed the overlap matrix between its right and left singular vectors, which correspond to the input and output modes, respectively.
-Rather than the unstructured, $\mathcal{O}(1/\sqrt{N})$ overlaps expected from i.i.d.
-random modes, we observe structured, large overlaps (Fig.
-10a).
-These large overlaps are particularly pronounced near the diagonal of the overlap matrix (Fig.
-10b).
-Motivated by these observations, we now extend the random-mode model to incorporate structured overlaps between input and output modes.
-Specifically, we define $\bm{L}$ and $\bm{R}$ as zero-mean Gaussian matrices with the following covariance structure:
-$\left\langle L_{ia}L_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{1}{N},$ (103)
-$\left\langle R_{ia}R_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{1}{N},$ (104)
-$\left\langle L_{ia}R_{jb}\right\rangle$ $=\delta_{ij}\delta_{ab}\frac{O_{ab}}{N},$ (105)
-where $O_{ab}$ is the overlap matrix. In this case,
-$\left\langle\bm{L}^{T}\bm{R}\right\rangle=\bm{O}.$ (106)
-To capture the most salient feature of the fly connectome, large diagonal overlaps, we specialize to a diagonal form for the overlap matrix:
-$O_{ab}=\delta_{ab}\rho_{a},$ (107)
-where $\rho_{a}$ is the correlation between the $a$-th input and output modes.
-This form of the overlaps implies that, for large $N$, $\bm{L}_{a}^{T}\bm{R}_{a}=\rho_{a}$, while $\bm{L}_{a}^{T}\bm{R}_{b}$ for $a\neq b$ remains a random $\mathcal{O}(1/\sqrt{N})$ quantity.
-Incorporating this structure allows the random-mode model to more accurately match the eigenspectrum of the connectome, which has large outliers along the real axis (Fig.
-10c).
-Defining an instance of this connectivity with a well-defined large-$N$ limit requires specifying a joint distribution from which $(D_{a},\rho_{a})$ are sampled for each $a$.
-Our analytic expressions for the two- and four-point functions depend on averages $\left\langle\cdots\right\rangle_{D,\rho}$ over this distribution.
-Note that for large $N$, self-connections (autapses) in $\bm{J}$ are given by
-$J_{ii}=\alpha\left\langle D\rho\right\rangle_{D,\rho},$ (108)
-which is, in general, $\mathcal{O}(1)$.
-To avoid biologically unrealistic strong self-connections, one can choose $\left\langle D\rho\right\rangle_{D,\rho}=0$, in which case $J_{ii}$ is random and $\mathcal{O}(1/\sqrt{N})$.
-Alternatively, if one interprets each model neuron as a cluster of biological neurons, strong self-connections can be interpreted as aggregated within-cluster connectivity [64, 65].
-Furthermore, note that the covariance between reciprocal pairs of couplings is given by $N\left\langle J_{ij}J_{ji}\right\rangle_{\bm{J}}=\sigma$ ($i\neq j$) where
-$\sigma=\alpha\left\langle(D\rho)^{2}\right\rangle_{D,\rho}.$ (109)
-We compute the two- and four-point functions using the path integral (Appendix I).
-As in previous cases, we obtain single-site processes for both neurons and latent variables.
-The neuronal single-site process is given by
-$T[x](t)=\eta^{x}(t)+[R_{\star}\ast\phi](t),$ (110)
-where $\eta^{x}\sim\mathcal{GP}(0,Q_{\star})$ and $\ast$ denotes convolution.
-The input covariance $Q_{\star}(t_{1},t_{2})$ and convolutional kernel $R_{\star}(t_{1},t_{2})$ are defined in terms of statistics of the latent-variable single-site process.
-Due to the linearity of the latent-variable dynamics, we can express them as
-$Q_{\star}(\omega)$ $=\alpha\left\langle\frac{D^{2}}{|1-D\rho S^{\phi}_{\star}(\omega)|^{2}}\right\rangle_{D,\rho}C^{\phi}_{\star}(\omega),$ (111)
-$R_{\star}(\omega)$ $=\alpha\left\langle\frac{D\rho}{1-D\rho S^{\phi}_{\star}(\omega)}\right\rangle_{D,\rho}.$ (112)
-We enforce $C^{\phi}_{\star}(t_{1},t_{2})=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}$ and $S^{\phi}_{\star}(t_{1},t_{2})=\left\langle{\delta\phi(t_{1})}/{\delta I(t_{2})}\right\rangle_{\star}$, yielding a self-consistent system that determines $C^{\phi}_{\star}(t_{1},t_{2})$ and $S^{\phi}_{\star}(t_{1},t_{2})$.
-Unlike previous cases, there is no effective coupling strength $g_{\text{eff}}$ describing the relationship between $Q_{\star}(t_{1},t_{2})$ and $C^{\phi}_{\star}(t_{1},t_{2})$; instead, this relationship involves a convolution with a kernel defined using the joint distribution over $(D,\rho)$.
-This can be viewed as a frequency-dependent effective coupling strength.
-Expanding the self-coupling kernel $R_{\star}(\omega)$ (Eq. 112) in powers of $\rho$ yields the first two terms
-$R_{\star}(\omega)=J_{ii}+\sigma S^{\phi}_{\star}(\omega)+\cdots$ (113)
-The first term is what we get in the single-site picture for i.i.d.
-cross-neuron couplings and deterministic $\mathcal{O}(1)$ self couplings.
-The second term appears in the single-site picture if there are correlations between reciprocal pairs of couplings, but the structure is otherwise i.i.d.
-(Appendix E).
-Nonzero higher-order terms reflect that the structure of the random-mode model is not fully captured by either of these forms.
-Note that the first two terms being zero forces all subsequent terms to be zero.
-Thus, nonzero covariance among reciprocal pairs of couplings is a necessary and sufficient condition for attaining a nonzero self-coupling kernel in the single-site picture.
-Figure 11: Effect of input-output mode correlations on dimension of activity in the random-mode model.
-Dimension of activity $\text{PR}^{\phi}$ versus correlation parameter $\rho_{1}$ in the $(D,\rho)$ joint distribution, for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$.
-Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
-All simulations use $N=5000$ neurons.
-To compute the four-point function, we write down and invert the frequency-dependent Hessian, and obtain
-$\Psi^{\phi}_{\star}$ $=\frac{V-1+\left\{\frac{(1+X_{12})(1+X_{21}^{*})}{1-W^{*}}+\text{h.c.}\right\}}{|1-U|^{2}}C^{\phi}_{12},$ (114a)
-$U$ $=\alpha\left\langle D^{2}\Sigma_{1}\Sigma_{2}\right\rangle_{D,\rho},$ (114b)
-$V$ $=\alpha\left\langle D^{4}|\Sigma_{1}\Sigma_{2}|^{2}\right\rangle_{D,\rho},$ (114c)
-$W$ $=\alpha\left\langle D^{2}\rho^{2}\Sigma_{1}^{*}\Sigma_{2}\right\rangle_{D,\rho},$ (114d)
-$X_{12}$ $=\alpha\left\langle D^{3}\rho\Sigma_{1}|\Sigma_{2}|^{2}\right\rangle_{D,\rho},$ (114e)
-$\Sigma_{k}$ $=\frac{S^{\phi}_{k}}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (114f)
-We validated this theory by choosing parameterizations of $D_{a}$ and $\rho_{a}$.
-We used $\alpha=M/N=1$ and $D_{a}\propto\exp(-\beta_{D}a/N)$ for the component strengths, as before.
-For the overlaps $\rho_{a}$, we use a sigmoidal shape inspired by Fig.
-10b in which the leading components have larger input-output mode overlaps.
-In particular, we used
-$\rho_{a}=\rho_{1}\left(1-\frac{2}{1+\exp(-\beta_{\rho}(a/N-u_{\text{flip}}))}\right),$ (115)
-which varies from $\rho_{1}$ to $-\rho_{1}$, hitting $\rho_{a}=0$ at $a/N=u_{\text{flip}}$.
-We fixed $\beta_{\rho}=16$, varied $\rho_{1}$, and set $u_{\text{flip}}$ such that $\left\langle D\rho\right\rangle_{D,\rho}=0$, eliminating strong self couplings.
-Across values of $g_{\text{eff}}$ and $\text{PR}^{D}$, sufficiently strong overlaps of either sign reduce activity dimension (Fig.
-11).
-The maximum dimension occurs for $\rho_{1}<0$, a regime in which activity projected along connectivity modes with the greatest strength is most strongly suppressed by the recurrent dynamics.
-This leftward shift is reduced for larger $g_{\text{eff}}$.
-VIII Discussion
-A key question in neuroscience and machine learning is how learning shapes collective dynamics through modifications of a neural network’s connectivity.
-While our analysis has focused on disordered networks that generate chaotic activity and are not trained to perform tasks, the smooth low-rank connectivity structure we have explored provides a starting point for understanding the impact of learned coupling matrices on collective dynamics.
-Recent works have shown that large neural networks trained on real-world tasks exhibit smooth spectral structures in their coupling matrices, similar to those we have studied [20].
-Moreover, while networks trained on simple tasks often display low-rank updates to their initial couplings, with a finite number of outlier eigenvalues emerging from the bulk [19, 46], training networks on many tasks, as is increasingly common, may produce coupling matrices with a low but extensive number of significant components [66, 67].
-In such scenarios, the random-mode model, or extensions of it, may predict the resulting activity dimension and other aspects of collective dynamics.
-The path-integral formalism we have developed is also amenable to analyzing cases where the coupling matrix is drawn from a Gibbs ensemble with an energy function equal to the loss function for a task.
-This approach represents scenarios where couplings are trained to minimize a loss function with stochastic gradient descent [52].
-Finally, initializing couplings with low-rank structure could improve training procedures like FORCE learning, which are based on suppressing chaos, because the initial strange attractor is lower-dimensional [11, 12].
-Historically, neural connectivity was often characterized through network motifs – recurring patterns like cycles or chains among small neuronal groups [68, 14, 69, 70].
-Such local statistics can be estimated using connectivity datasets obtained from methods that sample subsets of synapses.
-However, recent advances in whole-brain connectome reconstruction have unveiled global synaptic connectivity structures [71].
-These connectomes reveal a more complex picture, where the spectrum and modes of the connectivity matrix exhibit large-scale structure that cannot be explained by local motifs [72].
-Our parameterization of coupling matrices attempts to describe this global perspective, offering control over both the spectrum and mode overlaps.
-This could be particularly valuable in interpreting large-scale recordings of neural activity in the context of known anatomical connectivity.
-In a similar vein, Tiberi et al.
-[72] proposed a random matrix theory approach for noise-driven linear networks based on parameterizing the eigendecomposition of $\bm{J}$.
-By contrast, our use of a DMFT approach allows us to describe nonlinear dynamics.
-Their parameterization also does not capture relationships between eigenvalues and eigenvectors, which are likely crucial for computations.
-Our parameterization, which is more akin to an SVD than an eigendecomposition, allows for a simultaneous parameterization of component strengths and corresponding mode overlaps.
-Other parameterizations of connectivity similar to the random-mode model have been used to study the relationship between spiking and rate-based recurrent networks [73] and to train spiking networks more efficiently [74].
-Our description of single-neuron heterogeneity, where each neuron is parameterized by an arbitrary vector $\bm{\theta}$, can be readily applied to explore how various forms of heterogeneity influence dimensionality and other features of collective activity.
-Introducing specific forms of heterogeneity constrained by observed electrophysiological properties [75] and other cell-specific characteristics across neurons could provide insight on the functional consequences of neuronal diversity [76].
-On the technical side, we have demonstrated that the two-site cavity method and the analysis of fluctuations around the saddle point in the path integral approach yield equivalent results for computing the four-point function.
-This equivalence implies that the intricate cavity construction is equivalent to computing and inverting the Hessian of the action in the path integral.
-A deeper understanding of the relationship between these methods could lead to new hybrid approaches that combine the strengths of both techniques.
-In conclusion, the random-mode model and associated analytical techniques provide a set of tools for exploring the relationship between connectivity and collective dynamics.
-Our parameterization of connectivity is a middle ground between deterministic low-rank models and completely random networks.
-Future work incorporating more complex mode interactions, diverse neuronal properties, and learning dynamics will further bridge neural connectivity, activity, and function.
-Acknowledgements
-We gratefully acknowledge L.F. Abbott for discussions and feedback on this work.
-Appendix A Glossary of Terms
-Number of neurons ($N$)
-Total number of neurons in the network.
-Number of modes ($M$)
-Number of components in the random-mode model.
-(Effective) coupling strength ($g$, $g_{\text{eff}}$)
-Standard deviation of the couplings times $\sqrt{N}$ in the i.i.d. connectivity model (Eq. 14), random-mode model (Eq. 74), or variants thereof (Eq. 98).
-Random-mode model
-Network connectivity model: $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{L}_{a}\bm{R}_{a}^{T}$ (Eq. 22).
-Connectivity component ($\bm{L}_{a}\bm{R}_{a}^{T}$)
-Outer product of output and input modes.
-Input mode ($\bm{R}_{a}$)
-Neuronal pattern in the $a$-th component onto which activity patterns are projected.
-Output mode ($\bm{L}_{a}$)
-Neuronal pattern in the $a$-th component along which the projected activity pattern is expanded.
-Component strength ($D_{a}$)
-Scaling factor for the $a$-th component.
-Ratio of modes to neurons ($\alpha$)
-Given by $\alpha=M/N$ (Eq. 24).
-Participation ratio of component strengths ($\text{PR}^{D}$)
-Given by $\text{PR}^{D}=r_{2}^{2}/r_{4}$ (Eq. 26) where $r_{n}=\langle D^{n}\rangle_{D}$ (Eq. 25).
-Effective rank ($\alpha\text{PR}^{D}$)
-Connectivity dimensionality measure (Eq. 27).
-Two-point function
-Correlation function given by $C^{\phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i}\langle\phi_{i}(t_{1})\phi_{i}(t_{2})\rangle_{t}$ (Eq. 5).
-Four-point function
-Higher-order correlation function given by $\Psi^{\phi}(\tau_{1},\tau_{2})=\frac{1}{N}\sum_{i,j}C^{\phi}_{ij}(\tau_{1})C^{\phi}_{ij}(\tau_{2})$ (Eq. 7). See also time-by-time definition (Eq. 9).
-Appendix B Preprocessing of the fly hemibrain connectome
-We used the dataset originally presented in Scheffer et al.
-[21].
-To form the connection matrix for analysis, we used synaptic counts as a measure of connection strength, determining the magnitudes of the matrix elements.
-Neurons were assigned neurotransmitter probabilities based on a machine-learning analysis applied to the electron-microscopy data [22].
-We selected the most probable neurotransmitter assignment for each neuron.
-The neurotransmitter types were then mapped to signs (excitatory or inhibitory) according to Table 1.
-To normalize the connectivity matrix, we used an iterative scaling procedure analogous to the Sinkhorn–Knopp algorithm that proceeds as follows.
-We first rescale the positive elements in each row by a row-specific factor, such that the L2-norm of the positive elements is $1/\sqrt{2}$.
-The same procedure is applied to the negative elements, using a separate row-specific factor.
-This results in each row having a L2-norm of one with balanced excitatory and inhibitory inputs.
-A small number of neurons with exclusively excitatory or inhibitory inputs are left with an L2-norm of $1/\sqrt{2}$ instead of one.
-Following the row normalization, we perform the same operations on the columns.
-This disrupts the previous row normalization.
-Thus, we iterate the row and column normalization steps in an alternating fashion until convergence was achieved.
-The final result is a matrix with proper scaling along both rows and columns, which we interpret as the projection of the original matrix onto the set of properly row- and column-normalized matrices.
-Importantly, it has the same nonzero elements with the same signs as the original matrix.
-Neurotransmitter Synaptic Sign
-GABA -1 (Inhibitory)
-Acetylcholine +1 (Excitatory)
-Glutamate -1 (Inhibitory)
-Serotonin Ignored
-Octopamine Ignored
-Dopamine Ignored
-Neither Ignored
-Table 1: Mapping of neurotransmitters to synaptic signs
-Appendix C Temporal limits and inverses
-Here, we demonstrate, for the i.i.d.
-model that the temporal limits in Eq.
-34 can be taken before the inverses.
-This allows us to work with a frequency-dependent Hessian.
-In this section, we use Einstein notation for integrals over time variables (i.e., repeated time indices are integrated over).
-We would like to evaluate
-$\Psi(\tau_{1},\tau_{2})=U^{-1}(t_{1},t_{2},s_{1},s_{2})K(s_{1},s_{2},s_{3},s_{4})U^{-1}(t_{1}+\tau_{1},t_{2}+\tau_{2},s_{3},s_{4})$ (116)
-in the limits Eq.
-34, where $U(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ and $K(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ (note that $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0$ due to the vanishing of correlation functions involving only the conjugate field).
-Restating Eq.
-43, we have
-$U(t_{1},t_{2},t_{3},t_{4})$ $=-\left[\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})\right]-g^{2}\langle\langle\phi(t_{1})\phi(t_{2}),\hat{x}(t_{3})\hat{x}(t_{4})\rangle\rangle_{\star},$ (117)
-$K(t_{1},t_{2},t_{3},t_{4})$ $=\langle\langle\phi(t_{1})\phi(t_{2}),\phi(t_{3})\phi(t_{4})\rangle\rangle_{\star}.$ (118)
-We assume temporal stationarity throughout this section and consider only $a=\phi$.
-We start by examining the properties of $U(t_{1},t_{2},t_{3},t_{4})$.
-$U$ vanishes if one of the four time points is far (measured on the timescale of $C^{\phi}_{\star}(\tau)$) from the other three.
-For $t_{1}$ or $t_{2}$, this is a consequence of the perturbation associated with $\hat{x}(t_{1})$ or $\hat{x}(t_{2})$ being far away; for $t_{3}$ or $t_{4}$, this is a consequence of $\langle\phi(t_{i})\rangle_{\star}=0$ separating from the expectation.
-$U$ also vanishes if there is more than one time point which is far from the remaining ones by the same arguments.
-Thus, the time points need to be close in pairs.
-However, if $t_{1},t_{2}$ are close and $t_{3},t_{4}$ are close, but the pairs are distant, $U$ still vanishes since the perturbations associated with $\hat{x}(t_{i})$ are both far.
-This leaves three possibilities for a non-vanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
-Separating the three possibilities into two subdomains where all time points are close ($c$) and where the time points belong to either of the two pairwise separate ($s$) possibilities, we write $U=U_{s}+U_{c}$ where $U_{s}$ and $U_{c}$ are defined to vanish on the respective other domain.
-On the separated domain, $U$ simplifies to
-$U_{s}(t_{1},t_{2},t_{3},t_{4})=-\left[\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})\right]\\ +g^{2}\left[S^{\phi}_{\star}(t_{1}-t_{3})S^{\phi}_{\star}(t_{2}-t_{4})+S^{\phi}_{\star}(t_{1}-t_{4})S^{\phi}_{\star}(t_{2}-t_{3})\right],$ (119)
-where we used that the two distant contributions are non-overlapping to add them.
-We assume that we can also decompose the inverse into a contribution for separate pairs and a contribution where all time points are close, $U^{-1}=U_{s}^{-1}+U_{c}^{-1}$, where again $U_{s}^{-1}$ and $U_{c}^{-1}$ vanish on the respective other domain.
-We demonstrate that $U_{s}^{-1}$ is the inverse of $U_{s}$, and that $U_{c}^{-1}$ is the inverse of $U_{c}$, by noting that $U^{-1}(t_{1},t_{2},s_{1},s_{2})U(s_{1},s_{2},t_{3},t_{4})$ reduces to the contraction of $U_{c}^{-1}$ and $U_{c}$ if $t_{3}$ and $t_{4}$, and hence $s_{1}$ and $s_{2}$, are close; or to the contraction of $U_{s}^{-1}$ and $U_{s}$ if $t_{3}$ and $t_{4}$, and hence $s_{1}$ and $s_{2}$, are distant.
-For Eq.
-116, we need $U^{-1}_{s}$ due to the limits Eq.
-34.
-The fact that $U$ reduces to $U_{s}$ in the limits Eq.
-34 justifies taking the limits before inverting $U$.
-Furthermore, these limits imply that, in Eq.
-116, $s_{1}$ and $s_{2}$, as well as $s_{3}$ and $s_{4}$, are far apart.
-In this case, $K$ simplifies to
-$K_{s}(s_{1},s_{2},s_{3},s_{4})=C^{\phi}_{\star}(s_{1}-s_{3})C^{\phi}_{\star}(s_{2}-s_{4})+C^{\phi}_{\star}(s_{1}-s_{4})C^{\phi}_{\star}(s_{2}-s_{3}).$ (120)
-This justifies taking the limits Eq. 34 for $K$.
-Appendix D Incorporating sources
-This section outlines the general approach for incorporating source terms and then applies it to the specific case of correlations among the pre-activations. Our goal is to compute
-$\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0},$ (121)
-which gives the fluctuations around the mean of the quantity multiplying $\mathcal{J}$ in the action.
-To facilitate this calculation, we introduce a new field $U$ and set it equal to $\mathcal{J}$ using the conjugate $\hat{U}$:
-$Z[\mathcal{J}]=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp\left(-N\hat{U}U+N\hat{U}\mathcal{J}-NS[\hat{C};U]\right).$ (122)
-Instead of taking the second derivative with respect to $\mathcal{J}$, we set $\mathcal{J}=0$ and compute the fluctuations of $\hat{U}$ in the augmented theory whose path integral is
-$Z=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp(-N\tilde{\mathcal{S}}[\bm{C},U,\hat{U}]),$ (123)
-where
-$\tilde{\mathcal{S}}[\bm{C},U,\hat{U}]=\hat{U}U+\mathcal{S}[\bm{C};U].$ (124)
-The Hessian of the augmented action, in terms of the Hessian of the original action, is
-$\left(\begin{array}[]{c|cc}\bm{H}&\frac{\partial^{2}S}{\partial\mathcal{J}\partial\bm{C}}&\bm{0}\\[2.5pt] \hline\cr\\[-10.00002pt] \left(\frac{\partial^{2}S}{\partial\mathcal{J}\partial\bm{C}}\right)^{T}&\frac{\partial^{2}S}{\partial\mathcal{J}^{2}}&1\\[5.0pt] \bm{0}^{T}&1&0\end{array}\right),$ (125)
-where
-$H_{ab}=\frac{\delta^{2}S}{\delta C_{a}\delta C_{b}},\quad\left[\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{a}=\frac{\delta^{2}S}{\delta\mathcal{J}\delta C_{a}}.$ (126)
-The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point.
-Using a Schur complement to compute this element, we obtain the expression Eq.
-54 stated in the main text.
-We now apply Eq. 54 to the action of the i.i.d. connectivity model with a source term for correlations among pre-activations. The action is
-$\mathcal{S}[C^{\phi},\hat{C}^{\phi}]=-\frac{1}{2}C^{\phi}\hat{C}^{\phi}-\log\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\bigg{(}i\hat{x}T[x]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi-\frac{1}{2}x\mathcal{J}x\bigg{)}.$ (127)
-Computing the necessary quantities and taking the temporal limits of Eq. 34, we obtain, in Fourier space,
-$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\quad\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}g^{2}S_{12}^{x}\\ -C_{12}^{x\phi}\end{pmatrix},$ (128)
-where $C^{x\phi}_{12}=C_{\star}^{x\phi}(\omega_{1})C_{\star}^{x\phi}(\omega_{2})$ and $S^{x}_{12}=S^{x}_{\star}(\omega_{1})S^{x}_{\star}(\omega_{2})$.
-Substituting these into Eq.
-54 and using the frequency-dependent Hessian from Eq.
-52, we obtain
-$\Psi_{\star}^{x}=C^{x}_{12}+|U|^{2}C^{\phi}_{12}+UC^{x\phi}_{12}+\text{h.c.},\quad\text{where }U=\frac{g^{2}S^{x}_{12}}{1-g^{2}S^{\phi}_{12}},$ (129)
-in agreement with the two-site cavity result of Clark et al. [25], given in Eq. 21.
-Appendix E Partially symmetric disorder
-To demonstrate the flexibility of the path-integral approach to computing the four-point function, we now consider a generalization of the classic i.i.d.
-model in which we introduce correlations between reciprocal couplings, characterized by the following statistics:
-$\left\langle J_{ij}J_{k\ell}\right\rangle_{\bm{J}}=g^{2}\delta_{ik}\delta_{j\ell}+g^{2}\rho\delta_{i\ell}\delta_{jk}.$ (130)
-Here, $\rho\in[-1,1]$ is a parameter that controls the degree of symmetry in the connectivity.
-When $\rho=1$, the connectivity is fully symmetric, while $\rho=-1$ corresponds to fully antisymmetric connectivity.
-The case $\rho=0$ recovers the i.i.d.
-model.
-We express this partially symmetric connectivity as a linear combination of two matrices, $\bm{X}$ and $\bm{Y}$, that are i.i.d.
-random with mean zero and variance $1/N$:
-$J_{ij}=\sqrt{{g^{2}-|\sigma|}}X_{ij}+\sqrt{\frac{|\sigma|}{2}}(Y_{ij}+\text{sgn}(\sigma)Y_{ji}),$ (131)
-where $\sigma=\rho g^{2}$. In terms of $\bm{X}$ and $\bm{Y}$, the path integral for this partially symmetric model is
-$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg{(}i\sum_{i}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{ij}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{ij}Y_{ij}(\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j})\bigg{)}.$ (132)
-We integrate out $\bm{X}$ and $\bm{Y}$. To factorize the action over indices spanning extensive dimensions, we introduce
-[…]
-[…]
-We obtain, for the disorder-averaged path integral,
-$Z$ $=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}R\int\mathcal{D}\hat{R}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]\right),$ (135)
-$\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=-\frac{1}{2}C^{\phi}\hat{C}^{\phi}+\frac{\sigma}{2}S^{\phi}\hat{S}^{\phi}-\log W[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}],$ (136)
-$W[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=\int\mathcal{D}\hat{x}\int\mathcal{D}x\exp\bigg{(}i\hat{x}T[x]-\frac{1}{2}\phi\hat{C}^{\phi}\phi-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{i\sigma}{2}\phi\hat{S}^{\phi}\hat{x}-\frac{i\sigma}{2}\hat{x}S^{\phi}\phi\bigg{)}.$ (137)
-To obtain the saddle-point solution, we compute the derivatives of the action, under the rule that derivatives with respect to $S^{\phi}(t_{1},t_{2})$ also affect $\hat{S}^{\phi}(t_{2},t_{1})$, and vice versa, due to the symmetry in the action.
-This gives
-$\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+g^{2}\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W},$ (138)
-$\frac{\delta\mathcal{S}}{\delta\hat{C}^{\phi}(t_{1},t_{2})}$ $=-{C}^{\phi}(t_{1},t_{2})+\left\langle{\phi}(t_{1}){\phi}(t_{2})\right\rangle_{W},$ (139)
-$\frac{\delta\mathcal{S}}{\delta S^{\phi}(t_{1},t_{2})}$ $=\sigma\hat{S}^{\phi}(t_{1},t_{2})+i\sigma\left\langle\hat{x}(t_{1})\phi(t_{2})\right\rangle_{W},$ (140)
-$\frac{\delta\mathcal{S}}{\delta\hat{S}^{\phi}(t_{1},t_{2})}$ $=\sigma S^{\phi}(t_{1},t_{2})+i\sigma\left\langle\phi(t_{1})\hat{x}(t_{2})\right\rangle_{W}.$ (141)
-Setting these derivatives to zero yields the saddle-point conditions:
-$\hat{C}^{\phi}_{\star}(t_{1},t_{2})$ $=0,$ (142)
-$C^{\phi}_{\star}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (143)
-$S^{\phi}_{\star}(t_{1},t_{2})$ $=\hat{S}^{\phi}_{\star}(t_{2},t_{1})=\left\langle\frac{\delta\phi(t_{1})}{\delta I(t_{2})}\right\rangle_{\star},$ (144)
-where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$.
-The single-site process at the saddle point is described by
-$T[x](t)$ $=\eta^{x}(t)+\sigma\left[S^{\phi}_{\star}\ast\phi\right](t),$ (145)
-$\eta^{x}$ $\sim\mathcal{GP}(0,g^{2}C^{\phi}_{\star}),$ (146)
-where $\ast$ denotes convolution.
-The symmetric structure provides a convolutional, nonlinear self-coupling $\sigma[S^{\phi}_{\star}\ast\phi](t)$ in the single-site dynamics.
-The two-point correlation and response functions, $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$, must be determined self-consistently within this single site picture.
-Introducing the notation $C^{\phi}_{k}=C^{\phi}_{\star}(\omega_{k})$, $S^{\phi}_{k}=S^{\phi}_{\star}(\omega_{k})$ for $k=1,2$, the frequency-dependent Hessian at the saddle point is
-$\bm{H}=\begin{pmatrix}0&-1+g^{2}(S^{\phi}_{12})^{*}&0&0\\ -1+g^{2}S^{\phi}_{12}&-C^{\phi}_{12}&\sigma S^{\phi}_{1}C^{\phi}_{2}&\sigma S^{\phi}_{2}C^{\phi}_{1}\\ 0&\sigma(S^{\phi}_{1})^{*}C^{\phi}_{2}&0&\sigma-\sigma^{2}(S^{\phi}_{1})^{*}S^{\phi}_{2}\\ 0&\sigma(S^{\phi}_{2})^{*}C^{\phi}_{1}&\sigma-\sigma^{2}S^{\phi}_{1}(S^{\phi}_{2})^{*}&0\end{pmatrix}.$ (147)
-Inverting this matrix and isolating the upper-left element, we obtain, in agreement with Clark et al. [25],
-$\Psi^{\phi}_{\star}=\frac{1}{|1-g^{2}S^{\phi}_{12}|^{2}}\frac{1-|\sigma(S^{\phi}_{1})^{*}S^{\phi}_{2}|^{2}}{|1-\sigma(S^{\phi}_{1})^{*}S^{\phi}_{2}|^{2}}C^{\phi}_{12}.$ (148)
-Appendix F Inverse formula
-In Sec. III.5.2 and Appendix H, we use the fact that
-$\begin{pmatrix}0&a^{*}&0&-1\\ a&b&-1&0\\ 0&-1&0&c^{*}\\ -1&0&c&d\end{pmatrix}^{-1}=\begin{pmatrix}-\frac{b+|c|^{2}d}{|1-ac|^{2}}&-\frac{c}{1-ac}&-\frac{a^{*}d+bc}{|1-ac|^{2}}&-\frac{1}{1-ac}\\ -\frac{c^{*}}{1-(ac)^{*}}&0&-\frac{1}{1-(ac)^{*}}&0\\ -\frac{ad+bc^{*}}{|1-ac|^{2}}&-\frac{1}{1-ac}&-\frac{|a|^{2}d+b}{|1-ac|^{2}}&-\frac{a}{1-ac}\\ -\frac{1}{1-(ac)^{*}}&0&-\frac{a^{*}}{1-(ac)^{*}}&0\end{pmatrix}.$ (149)
-Appendix G Two-site cavity calculation of the four-point function in the random-mode model
 To apply cavity techniques, we reformulate the network as a bipartite system of neurons and latent variables,
-$T[x_{i}](t)$ $=\sum_{a}L_{ia}z_{a}(t),$ (150)
-$z_{a}(t)$ $=D_{a}\sum_{j}R_{ja}\phi_{j}(t),$ (151)
+$T[x_{i}](t)$ $=\sum_{a=1}^{M}L_{ia}z_{a}(t),$ (200)
+$z_{a}(t)$ $=D_{a}\sum_{j=1}^{N}R_{ja}\phi_{j}(t),$ (201)
 as done in the path-integral approach.
@@
 For the cavity variables, we introduce special indices: neuronal cavity variables are indexed by $\mu,\nu\in\{0,0^{\prime}\}$, while latent cavity variables are indexed by $\alpha,\beta\in\{\hat{0},\hat{0}^{\prime}\}$.
-G.1 Neuronal cavity
+L.1 Neuronal cavity
 We begin by introducing two neurons, $x_{0}(t)$ and $x_{0^{\prime}}(t)$. The leading-order ($1/\sqrt{N}$) effect on the latent variables is
-[…]
+$\delta z_{a}(t)=\int^{t}dt^{\prime}\sum_{b}S^{z}_{ab}(t_{1},t_{2})D_{b}\sum_{\mu}R_{\mu b}\phi_{\mu}(t^{\prime}),$ (202)
 where $S^{z}_{ab}(t_{1},t_{2})$ is the response function of the latent variables. The dynamic equations for the cavity neurons are
-[…]
+$T[x_{\mu}](t)=\eta_{\mu}(t)+\sum_{\nu}\int^{t}dt^{\prime}F_{\mu\nu}(t_{1},t_{2})\phi_{\nu}(t^{\prime}),$ (203)
 where
@@
 To evaluate this, we need to square and disorder-average Eq.
-157.
+207.
 This requires us to consider several two-frequency correlation functions, which we define as
@@
 where $\hat{0},\hat{0}^{\prime}$ denote the indices in the latent-variable cavity picture.
-G.2 Latent-variable cavity
+L.2 Latent-variable cavity
 We now introduce two new latent variables, $z_{\hat{0}}(t)$ and $z_{\hat{0}^{\prime}}(t)$. The leading-order ($1/\sqrt{N}$) effect on the neurons is
-[…]
+$\delta\phi_{i}(t)=\int^{t}dt^{\prime}\sum_{j=1}^{N}S^{\phi}_{ij}(t_{1},t_{2})\sum_{\alpha}L_{j\alpha}z_{\alpha}(t^{\prime}),$ (215)
 where $S^{\phi}_{ij}(t_{1},t_{2})$ is the response function of the neurons. The dynamic equations for the cavity latent variables are
-[…]
+$z_{\alpha}(t)=D_{\alpha}\left[\gamma_{\alpha}(t)+\frac{1}{\sqrt{N}}\sum_{\beta}\int^{t}dt^{\prime}G_{\alpha\beta}(t_{1},t_{2})z_{\beta}(t^{\prime})\right],$ (216)
 where
-$\gamma_{\alpha}(t)$ $=\sum_{i}R_{i\alpha}\phi_{i}(t),$ (167)
-$G_{\alpha\beta}(t_{1},t_{2})$ $=\sqrt{N}\sum_{i,j}R_{i\alpha}L_{j\beta}S^{\phi}_{ij}(t_{1},t_{2}).$ (168)
+$\gamma_{\alpha}(t)$ $=\sum_{i=1}^{N}R_{i\alpha}\phi_{i}(t),$ (217)
+$G_{\alpha\beta}(t_{1},t_{2})$ $=\sqrt{N}\sum_{i,j=1}^{N}R_{i\alpha}L_{j\beta}S^{\phi}_{ij}(t_{1},t_{2}).$ (218)
 Defining the time-average cavity-field cross-covariance as
-$C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega)=\sum_{i,j}R_{i\hat{0}}R_{j\hat{0}^{\prime}}C^{\phi}_{ij}(\omega),$ (169)
+$C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega)=\sum_{i,j=1}^{N}R_{i\hat{0}}R_{j\hat{0}^{\prime}}C^{\phi}_{ij}(\omega),$ (219)
 we obtain the following expression for the cross-covariance of the cavity latent variables up to order $1/\sqrt{N}$:
@@
 by squaring and disorder-averaging Eq. 220. This requires us to consider several two-frequency correlation functions, which we define as
-$\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\left\langle G_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (172)
-$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}^{\prime}\hat{0}}}(\bm{\omega})$ $=\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}^{\prime}\hat{0}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (173)
-$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\sqrt{N}\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (174)
-$\Gamma_{C^{\gamma}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=N\left\langle C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}}.$ (175)
+$\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\left\langle G_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (222)
+$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}^{\prime}\hat{0}}}(\bm{\omega})$ $=\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}^{\prime}\hat{0}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (223)
+$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\sqrt{N}\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (224)
+$\Gamma_{C^{\gamma}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=N\left\langle C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}}.$ (225)
 Again, we can evaluate these $\Gamma_{\cdots}(\bm{\omega})$ functions due to the independence of the couplings and dynamic variables in the expressions defining them. The nonvanishing ones are
@@
 where $0,0^{\prime}$ denote the indices in the neuronal cavity picture.
-G.3 Combining the two two-site cavity pictures
+L.3 Combining the two two-site cavity pictures
 To combine the results from both two-site cavity pictures, we use the following relations:
@@
 $S^{\phi}_{00^{\prime}}(\omega)$ $=\frac{1}{\sqrt{N}}(S^{\phi}_{\star}(\omega))^{2}F_{00^{\prime}}(\omega).$ (230)
-Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions (Eqs. 159-162 and 172175), we obtain
+Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions (Eqs. 209-212 and 222225), we obtain
 $\Gamma_{C^{\eta}_{00^{\prime}}C^{\eta}_{00^{\prime}}}(\bm{\omega})$ $=\alpha r_{4}C_{\star}^{\phi}(\omega_{1})C_{\star}^{\phi}(\omega_{2})+\alpha^{2}\psi^{z}(\bm{\omega}),$ (231)
@@
 This result is identical to Eq. 38 derived using the path-integral approach, demonstrating the consistency between the two methods.
-Appendix H Single-unit heterogeneity calculations
-The full set of order parameters is
-[…]
-[…]
-The action and single-site path integrals for this system are
-$\mathcal{S}[C^{\Phi},\hat{C}^{\Phi},Q,\hat{Q}]$ $=-\frac{1}{2}C^{\Phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}-\left\langle\log W_{\bm{\theta}}[Q,\hat{Q};\mathcal{J}]\right\rangle_{\bm{\theta}}-\alpha\left\langle\log W^{z}_{D}[C^{\Phi},\hat{C}^{\Phi}]\right\rangle_{D},$ (193)
-$W^{x}_{\bm{\theta}}[Q,\hat{Q};\mathcal{J}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\Phi_{\bm{\theta}}\hat{Q}\Phi_{\bm{\theta}}-\frac{1}{2}\hat{x}Q\hat{x}-\frac{1}{2}\phi\mathcal{J}\phi\right),$ (194)
-$W^{z}_{D}[C^{\Phi},\hat{C}^{\Phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}z\hat{C}^{\Phi}z-\frac{1}{2}D^{2}\hat{z}C^{\Phi}\hat{z}\right).$ (195)
-The saddle-point calculation proceeds similarly to the non-heterogeneous case of the random-mode model. We obtain
-$\hat{C}^{\Phi}_{\star}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (196)
-$C^{\Phi}_{\star}(t_{1},t_{2})$ $=\left\langle\left\langle\Phi_{\bm{\theta}}(t)\Phi_{\bm{\theta}}(t^{\prime})\right\rangle_{\star}\right\rangle_{\bm{\theta}},$ (197)
-$Q_{\star}(t_{1},t_{2})$ $=\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}\right\rangle_{D},$ (198)
-where $\left\langle\cdots\right\rangle_{\star}$ and $\left\langle\cdots\right\rangle_{\star D}$ are averages with respect to $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ and $W^{z}_{D}[C^{\Phi}_{\star},0]$, respectively (note that $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ does not depend on $\bm{\theta}$).
-Using $\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}=D^{2}C^{\Phi}_{\star}(t_{1},t_{2})$ and, consequently, $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\Phi}_{\star}(t_{1},t_{2})$, the self-consistent problem that determines $C^{\Phi}_{\star}(t_{1},t_{2})$ is given by Eq.
-89 and the preceding description in Sec.
-VI.
-The frequency-dependent Hessian at the saddle point is
-$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}{C_{{12}}^{\Phi}}&-1&0\\ \hline\cr 0&-1&0&(\mathfrak{s}^{\Phi}_{12})^{*}\\ -1&0&\mathfrak{s}^{\Phi}_{12}&-\mathfrak{C}^{\Phi}_{12}\end{array}\right),$ (199)
-where the new $\mathfrak{F}$raktur variables are defined in Eq.
-90.
-We invert the Hessian using Eq.
-149 and identify the upper-left element as Eq.
-91.
-Combining the inverted Hessian with the relevant quantities for computing fluctuations of normalized variables,
-$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C^{\phi}_{12},\>\>\>\frac{\delta^{2}S}{\delta\bm{C}\delta\mathcal{J}}=\begin{pmatrix}0\\ 0\\ (S^{\phi}_{12})^{*}\\ -\mathfrak{C}^{\Phi\phi}_{12}\end{pmatrix},$ (200)
-where
-$\mathfrak{C}^{\Phi\phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\phi\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (201)
-$S^{\phi}_{12}$ $=\left\langle\frac{\delta\phi}{\delta I}\right\rangle_{\star}(\omega_{1})\left\langle\frac{\delta\phi}{\delta I}\right\rangle_{\star}(\omega_{2}),$ (202)
-we obtain Eq. 92.
-Appendix I Structured mode overlaps calculations
-To generate matrices $\bm{L}$ and $\bm{R}$ with the correlation structure of Eq.
-106 and Eq.
-107, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$:
-${L}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{1}_{ia}+\sqrt{{|\rho_{a}|}}{Y}_{ia},$ (203)
-${R}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{2}_{ia}+\text{sgn}(\rho_{a})\sqrt{{|\rho_{a}|}}{Y}_{ia}.$ (204)
-We begin with the path integral containing latent variables, insert this parameterization of $\bm{L}$ and $\bm{R}$, then and integrate out $\bm{X}_{1}$, $\bm{X}_{2}$, and $\bm{Y}$.
-To factorize the action over indices spanning extensive dimensions, we introduce
-[…]
-[…]
-[…]
-[…]
-The resulting path integral, action, and single-site path integrals are given by:
-$\centering Z=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}S^{\phi}\@add@centering$ $\int\mathcal{D}\hat{S}^{\phi}\int\mathcal{D}Q\int\mathcal{D}\hat{Q}\int\mathcal{D}R\int\mathcal{D}\hat{R}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi},Q,\hat{Q},R,\hat{R}]\right),$ (209)
-$\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi},Q,\hat{Q},R,\hat{R}]$ $=-\frac{1}{2}C^{\phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}+\frac{1}{2}S^{\phi}\hat{R}+\frac{1}{2}R\hat{S}^{\phi}$ (210)
-$\hskip 100.00015pt-\log W^{x}[Q,\hat{Q},R,\hat{R}]-\alpha\left\langle\log W^{z}_{D,\rho}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]\right\rangle_{D,\rho},$ (211)
-$W^{x}[Q,\hat{Q},R,\hat{R}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\hat{x}Q\hat{x}-\frac{1}{2}\phi\hat{Q}\phi-\frac{i}{2}\hat{x}R\phi-\frac{i}{2}\phi\hat{R}\hat{x}\right),$ (212)
-$W^{z}_{D,\rho}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}D^{2}\hat{z}C^{\phi}\hat{z}-\frac{1}{2}z\hat{C}^{\phi}z-\frac{i}{2}D\rho\hat{z}S^{\phi}z-\frac{i}{2}D\rho z\hat{S}^{\phi}\hat{z}\right).$ (213)
-We first calculate the saddle point conditions by taking the derivatives of the action:
-$\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{Q}(t_{1},t_{2})+\alpha\left\langle D^{2}\left\langle\hat{z}(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (214)
-$\frac{\delta\mathcal{S}}{\delta\hat{C}^{\phi}(t_{1},t_{2})}$ $=-Q(t_{1},t_{2})+\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (215)
-$\frac{\delta\mathcal{S}}{\delta S^{\phi}(t_{1},t_{2})}$ $=\hat{R}(t_{1},t_{2})+i\alpha\left\langle D\rho\left\langle\hat{z}(t_{1})z(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (216)
-$\frac{\delta\mathcal{S}}{\delta\hat{S}^{\phi}(t_{1},t_{2})}$ $=R(t_{1},t_{2})+i\alpha\left\langle D\rho\left\langle z(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (217)
-$\frac{\delta\mathcal{S}}{\delta Q(t_{1},t_{2})}$ $=-\hat{C}^{\phi}(t_{1},t_{2})+\left\langle\hat{x}(t_{1})\hat{x}(t_{2})\right\rangle_{W^{x}},$ (218)
-$\frac{\delta\mathcal{S}}{\delta\hat{Q}(t_{1},t_{2})}$ $=-C^{\phi}(t_{1},t_{2})+\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{W^{x}},$ (219)
-$\frac{\delta\mathcal{S}}{\delta R(t_{1},t_{2})}$ $=\hat{S}^{\phi}(t_{1},t_{2})+i\left\langle\hat{x}(t_{1})\phi(t_{2})\right\rangle_{W^{x}},$ (220)
-$\frac{\delta\mathcal{S}}{\delta\hat{R}(t_{1},t_{2})}$ $=S^{\phi}(t_{1},t_{2})+i\left\langle\phi(t_{1})\hat{x}(t_{2})\right\rangle_{W^{x}}.$ (221)
-The saddle-point conditions yield:
-$\hat{C}^{\phi}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (222)
-$C^{\phi}_{\star}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (223)
-$S^{\phi}_{\star}(t_{1},t_{2})$ $=\hat{S}^{\phi}_{\star}(t_{1},t_{2})=\left\langle\frac{\delta\phi(t_{1})}{\delta I(t_{2})}\right\rangle_{\star},$ (224)
-$Q_{\star}(t_{1},t_{2})$ $=\alpha\left\langle\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D,\rho}\right\rangle_{D,\rho},$ (225)
-$R_{\star}(t_{1},t_{2})$ $=\hat{R}_{\star}(t_{1},t_{2})=\alpha\left\langle D\rho\left\langle\frac{\delta z(t_{1})}{\delta I(t_{2})}\right\rangle_{\star D,\rho}\right\rangle_{D,\rho},$ (226)
-where $\left\langle\cdots\right\rangle_{\star D,\rho}$ and $\left\langle\cdots\right\rangle_{\star}$ are averages within the dynamic processes described by $W^{z}_{D,\rho}[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$ and $W^{x}[Q_{\star},0,R_{\star},\hat{R}_{\star}]$, respectively.
-The yields neuronal and latent-variable single-site processes described by
-$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\ast\phi](t)$ (227)
-$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\ast z](t),$ (228)
-where the Gaussian fields have statistics
-$\eta^{x}$ $\sim\mathcal{GP}(0,Q_{\star}),$ (229)
-$\eta^{z}$ $\sim\mathcal{GP}(0,D^{2}C^{\phi}_{\star}).$ (230)
-This gives rise to the single-site problem described in Sec. VII.
-The frequency-dependent Hessian at the saddle point is
-$\bm{H}=\left(\begin{array}[]{cccc|cccc}0&u^{*}&0&0&0&-1&0&0\\ u&-vC^{\phi}_{1}C^{\phi}_{2}&x_{12}C_{2}&x_{21}C^{\phi}_{1}&-1&0&0&0\\ 0&x_{12}^{*}C^{\phi}_{2}&0&-w&0&0&0&1\\ 0&x_{21}^{*}C^{\phi}_{1}&-w^{*}&0&0&0&1&0\\ \hline\cr 0&-1&0&0&0&(S^{\phi}_{1}S^{\phi}_{2})^{*}&0&0\\ -1&0&0&0&S^{\phi}_{1}S^{\phi}_{2}&-C^{\phi}_{1}C^{\phi}_{2}&S^{\phi}_{1}C^{\phi}_{2}&S^{\phi}_{2}C^{\phi}_{1}\\ 0&0&0&1&0&(S^{\phi}_{1})^{*}C^{\phi}_{2}&0&-(S^{\phi}_{1})^{*}S^{\phi}_{2}\\ 0&0&1&0&0&(S^{\phi}_{2})^{*}C^{\phi}_{1}&-S^{\phi}_{1}(S^{\phi}_{2})^{*}&0\end{array}\right)$ (231)
-where
-$u$ $=\alpha\left\langle D^{2}\sigma_{1}\sigma_{2}\right\rangle_{D,\rho},$ (232)
-$v$ $=\alpha\left\langle D^{4}|\sigma_{1}|^{2}|\sigma_{2}|^{2}\right\rangle_{D,\rho},$ (233)
-$w$ $=\alpha\left\langle D^{2}\rho^{2}\sigma_{1}^{*}\sigma_{2}\right\rangle_{D,\rho},$ (234)
-$x_{12}$ $=\alpha\left\langle D^{3}\rho\sigma_{1}|\sigma_{2}|^{2}\right\rangle_{D,\rho},$ (235)
-$\sigma_{k}$ $=\frac{1}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (236)
-Isolating the upper-left element of the inverse of this matrix yields Eq. 114.
-Appendix J Numerical details
-J.1 Numerics for theory
-To validate the theory, we used networks with dynamics defined by $T[x](t)=(1+\partial_{t})x(t)$ (Eq.
-2) and with nonlinearity $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$.
-For all connectivity models, except those with structured $\bm{L}$-$\bm{R}$ overlaps (VII), the single-site two-point functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ depend only on $g_{\text{eff}}$ and can be numerically calculated as in i.i.d.
+Appendix M Numerical details
+M.1 Numerics for theory
+To validate the theory, we used networks with dynamics defined by $T[x](t)=(1+\partial_{t})x(t)$ and with nonlinearity $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$.
+For all connectivity models, except those with structured $\bm{L}$-$\bm{R}$ overlaps (Sec.
+VII), the single-site two-point functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ depend only on $g_{\text{eff}}$ and can be numerically calculated as in i.i.d.
 networks (see, e.g., [10, 98]).
@@
     Squaring the single-site picture gives a second-order ODE $\partial^{2}_{\tau}C^{x}_{\star}(\tau)=C^{x}_{\star}(\tau)-g_{\text{eff}}^{2}C^{\phi}_{\star}(\tau)$.
-Since the rhs depends only on $C^{x}_{\star}(\tau)$ and $C^{x}_{\star}(0)$, we can consider it to be the negative derivative $-\partial V/\partial C^{x}_{\star}(\tau)$ of a potential $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$ with an explicit dependence on the initial condition $C^{x}_{\star}(0)$.
+Since the rhs depends only on $C^{x}_{\star}(\tau)$ and $C^{x}_{\star}(0)$, we can consider it to be the negative derivative $\partial V/\partial C^{x}_{\star}(\tau)$ of a potential $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$ with an explicit dependence on the initial condition $C^{x}_{\star}(0)$.
   3. 3.
@@
   4. 4.
-    Restricting to solutions with $C^{x}_{\star}(\tau)\rightarrow 0$ as $\tau\rightarrow\infty$ and $\partial_{\tau}C^{x}_{\star}(\tau)|_{\tau=0}=0$, enforcing conservation of energy gives $V(C^{x}_{\star}(0),C^{x}_{\star}(0))=V(0,C^{x}_{\star}(0))$, which can be solved for $C^{x}_{\star}(0)$ by numerical root finding.
+    Restricting to solutions with $C^{x}_{\star}(\tau)\to 0$ as $\tau\to\infty$ and $\partial_{\tau}C^{x}_{\star}(\tau)|_{\tau=0}=0$, enforcing conservation of energy gives $V(C^{x}_{\star}(0),C^{x}_{\star}(0))=V(0,C^{x}_{\star}(0))$, which can be solved for $C^{x}_{\star}(0)$ by numerical root finding.
   5. 5.
-    Finally, Euler integration of the $[C^{x}_{\star}(\tau),\partial_{\tau}C^{x}_{\star}(\tau)]$ dynamics with these initial conditions gives its full time course, along with $C^{\phi}_{\star}(\tau)$ via Eq.
-237.
+    Finally, Euler integration of the $C^{x}_{\star}(\tau),\partial_{\tau}C^{x}_{\star}(\tau)]$ dynamics with these initial conditions gives its full time course, along with $C^{\phi}_{\star}(\tau)$ via Eq.
+241.
 We integrated the Newtonian ODE from $\tau=0$ to $\tau_{\text{max}}=200$ with step size $d\tau=0.025$.
-When then use an FFT to otain the frequency-space representation.
+When then use an FFT to obtain the frequency-space representation.
 The frequency-space representation of the response function $S^{\phi}_{\star}(\omega)$ was computed directly as
@@
     Each of $N_{\text{sample}}$ single-site processes $x_{s}(t)$ is Euler integrated according to Eq.
-110 with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi(x_{s}(t))$.
+47a with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi(x_{s}(t))$.
   5. 5.
@@
 We took $\left\langle\cdot\right\rangle_{D,\rho}$ averages in this case by averaging over every other discrete value of $D_{a}$ and $\rho_{a}$.
-J.2 Simulation details
+M.2 Simulation details
 Each network was integrated via Runga-Kutta (RK4) with $dt=0.05$.

### v3

@@
 We first introduce the random-mode model, a random-matrix ensemble related to the singular-value decomposition that enables control over the spectrum and right-left mode overlaps.
-Then, using a novel path-integral calculation, we derive analytic expressions that reveal how connectivity structure affects features of collective dynamics: the dimension of activity, which quantifies the number of high-variance collective-activity fluctuations, and the temporal correlations that characterize the timescales of these fluctuations.
+Then, using a novel path-integral calculation, we derive analytical expressions that reveal how connectivity structure affects features of collective dynamics: the dimension of activity, which quantifies the number of high-variance collective-activity fluctuations, and the temporal correlations that characterize the timescales of these fluctuations.
 We show that connectivity structure can be invisible in single-neuron activities while dramatically shaping collective activity.
 Furthermore, despite the nonlinear, high-dimensional nature of these networks, the dimension of activity depends on just two connectivity parameters—the variance of the couplings and the effective rank of the coupling matrix, which quantifies the number of dominant rank-one connectivity components.
-We contrast the effects of single-neuron heterogeneity and low-dimensional connectivity, making predictions about how $z$-scoring data affects the dimension of activity.
+We contrast the effects of single-neuron heterogeneity and low dimensional connectivity, making predictions about how z-scoring data affects the dimension of activity.
 Finally, we demonstrate the presence of structured overlaps between left and right modes in the Drosophila connectome, incorporate them into the theory, and show how they further shape collective dynamics.
 I Introduction
-The collective activity of a high-dimensional nonlinear system is determined by the structure of the interactions between its elements, but the mapping from structure to activity is generally only analytically tractable for limited forms of structure.
+The collective activity of a high-dimensional nonlinear system is determined by the structure of the interactions between its elements, but the mapping from structure to activity is generally analytically tractable only for limited forms of structure.
 In systems neuroscience, structure in synaptic connectivity controls the function of a neural system by determining what patterns of activity are produced by populations of neurons.
 This paper studies this mapping in the context of nonlinear recurrent neural networks, used widely in neuroscience as models of neural-circuit dynamics [1, 2, 3] and in machine learning as systems for sequence processing [4, 5, 6, 7].
-While these models omit many features of actual neural circuits, they capture several of their fundamental features, including their large scale, nonlinear units, and recurrent interactions.
+While these models omit many aspects of actual neural circuits, they capture several of their fundamental features, including their large scale, nonlinear units, and recurrent interactions.
 Functionally, a key property of such systems is that, like real neural circuits, they can generate rich time-varying activity in the absence of external input [8, 9].
-Theoretical studies of nonlinear recurrent neural networks, including the seminal work of Sompolinsky et al.
-[10], often focus on the case of independent and identically distributed (i.i.d.) couplings, or simple variants thereof [11].
+Theoretical studies of nonlinear recurrent neural networks, including the pioneering work of Sompolinsky et al.
+(Ref.
+[10]), often focus on the case of independent and identically distributed (i.i.d.) couplings, or simple variants thereof [11].
 Such disordered networks produce high-dimensional chaotic activity, reminiscent of asynchronous cortical activity observed in vivo [10].
@@
 couplings.
-Recently, large-scale synaptic reconstructions of neural circuits, or “connectomes,” have become available in multiple species, including the full brain of Drosophila melanogaster [17] and increasingly large portions of mammalian circuits.
+Recently, large-scale synaptic reconstructions of neural circuits, or “connectomes,” have become available in multiple species, including the full brain of Drosophila melanogaster [17] and increasingly large portions of mammalian circuits [18, 19, 20, 21].
 These large-scale connectomes reveal that neural circuits deviate substantially from the i.i.d.
 assumption, as previously demonstrated via the presence of small-scale motifs [22].
-In particular, in accordance with real-world networks more generally [19], neural circuits appear to exhibit low-rank structure in which the connectivity is well-described by fewer rank-one components than the network size (we demonstrate this explicitly for the Drosophila connectome in Sec.
+In particular, in accordance with real-world networks more generally [23], neural circuits appear to exhibit low-rank structure in which the connectivity is well described by fewer rank-one components than the network size (we demonstrate this explicitly for the Drosophila connectome in Sec.
 II.1).
-In neural-network models, recurrent or otherwise, such structure emerges naturally in several contexts, including when connections depend on distance in physical or feature space [20, 21, 22], or when networks are trained on tasks with low-dimensional structure [23, 24, 25].
-In parallel with the development of large-scale connectivity maps, recent advances in recording technologies including silicon probes [26] and large-scale calcium imaging [27] now enable the activities of many neurons to be monitored simultaneously, with the latest datasets including hundreds of thousands of neurons.
+In neural-network models, recurrent or otherwise, such structure emerges naturally in several contexts, including when connections depend on distance in physical or feature spaces [24, 25, 26], or when networks are trained on tasks with low dimensional structure [27, 28, 29].
+In parallel with the development of large-scale connectivity maps, recent advances in recording technologies, including silicon probes [30] and large-scale calcium imaging [31, 32], now enable the activities of many neurons to be monitored simultaneously, with the latest datasets including hundreds of thousands of neurons or more [33].
 This technological progress has enabled researchers to characterize collective properties of neural activity rather than focusing solely on single-neuron responses.
-These collective properties capture how neurons interact and reveal distributed computational processes that are only visible with high-yield recording technologies [28].
-Such population-level features are typically analyzed using dimensionality reduction techniques like principal component analysis [29].
-Relating connectivity and collective-activity structure analytically.
-We currently lack analytical tools to predict relationships between collective activity described by large-scale recordings, and connectivity structure described by connectomics or other datasets.
-Dynamical mean-field theory (DMFT) is widely used to analytically characterize activity in large, nonlinear recurrent neural networks.
+These collective properties capture how neurons interact and reveal distributed computational processes that are visible only with high-yield recording technologies [34].
+Such population-level features are typically analyzed using dimensionality reduction techniques like principal components analysis [35, 36, 37, 38].
+I.1 Relating connectivity and collective-activity structure analytically.
+We currently lack analytical tools to predict relationships between connectivity structure, described by connectomics or other datasets, and collective activity, described by large-scale neural recordings.
+Dynamical mean-field theory (DMFT) is a theoretical tool from statistical physics widely used to analytically characterize activity in large nonlinear recurrent neural networks.
 Although recent advances in DMFT have extended these techniques to describe collective features in networks producing high-dimensional chaotic activity [39], these calculations have been limited to i.i.d.
-connectivity or connectivity with simple correlations between reciprocal couplings.
+connectivity, or connectivity with simple correlations between reciprocal couplings.
 Extending such calculations to more complicated connectivity structures rapidly becomes unwieldy.
 Consequently, such analyses have not been performed for networks whose connections are constrained by the statistics of connectomic or other datasets, limiting our ability to leverage these datasets for theoretical insight.
-One tractable approach has been to study networks with very low-dimensional connectivity and thus activity (specifically, in the limit where the rank of the coupling matrix remains finite while the network becomes large; see Sec.
+One tractable approach has been to study networks with very low dimensional connectivity and thus activity (specifically, in the limit where the rank of the coupling matrix remains finite while the network becomes large; see Sec.
 II.3).
@@
 II.1, do not support this finite-rank assumption.
-Furthermore, while neural activity recorded in experiments is often low-dimensional, this may be inherited from the low dimensionality of experimental tasks; in more complicated tasks or spontaneous states, activity is typically higher-dimensional [31, 27].
-To address these limitations and challenges, this paper makes two contributions:
+Furthermore, while neural activity recorded in experiments is often low dimensional, this may be inherited from the low dimensionality of experimental tasks; in more complicated tasks or spontaneous states, activity is typically higher dimensional [31, 32, 33].
+To address these limitations and challenges, this paper makes two contributions.
   1. 1.
@@
 II.1, we introduce the random-mode model in Sec.
-II.2, in which couplings are generated as a sum of rank-one outer products of random left and right modes with component-specific strengths.
-Unlike well studied low-rank recurrent neural network models, we scale the rank with the network size (Sec.
+II.2, in which the couplings are generated as a sum of rank-one outer products of random left and right modes with component-specific strengths.
+Unlike well-studied low-rank recurrent neural-network models, we scale the rank with the network size (Sec.
 II.3).
 We show in Sec.
-II.4 that parameterizing these strengths provides control over the singular value spectrum of the coupling matrix.
+II.4 that parametrizing these strengths provides control over the singular-value spectrum of the coupling matrix.
   2. 2.
     A new path-integral calculation of a specific four-point function of network activity, originally defined by Clark et al.
-[30], that captures features of collective activity including its dimensionality.
+(Ref.
+[39]), that captures features of collective activity including its dimensionality.
 We review these and other summary statistics in Sec.
 III.
-In Sec.
-IV A–C, we describe the calculation, based on fluctuations around the saddle point in a path integral, and apply it to the random-mode model.
+In Secs.
+IV.1IV.3, we describe the calculation, based on fluctuations around the saddle point in a path integral, and apply it to the random-mode model.
 In Sec.
@@
 Finally, in Sec.
-VII, we incorporate structured overlaps between left and right modes into the model, demonstrating their presence in the Drosophila connectome and showing how they further shape collective dynamics.
+VII, we incorporate structured overlaps between the left and right modes into the model, demonstrating their presence in the Drosophila connectome and showing how they further shape collective dynamics.
 II Random-mode model
 II.1 Spectral structure of neural-circuit connectivity
-Figure 1: Analysis of Drosophila central brain connectome.
+Figure 1: Analysis of Drosophila central-brain connectome.
 (a) Volume of fly brain for which reconstruction was performed (blue; reproduced from [17]).
-(b) Visualization of the full normalized coupling matrix.
-(c) Singular value spectra of the normalized coupling matrix (red) and an i.i.d.
+(b) Normalized coupling matrix (elements summed within $10\times 10$ blocks to aid visualization).
+(c) Singular-value spectra of the normalized coupling matrix (red) and an i.i.d.
 random matrix (gray).
@@
 assumption.
-To illustrate this, we analyzed a central-brain connectome of the fruit fly Drosophila melanogaster using the singular value decomposition (SVD), which decomposes any matrix as a sum of rank-one components, with each component’s strength given by the corresponding singular value (Fig.
-1a; details in Appendix B).
+To illustrate this, we analyzed a central-brain connectome of the fruit fly Drosophila melanogaster using the singular-value decomposition (SVD), which decomposes any matrix as a sum of rank-one components, with each component’s strength given by the corresponding singular value [Figs.
+1(a) and 1(b); details in Appendix B].
 Specifically, the SVD decomposes $\bm{J}$ in the form
@@
 The ranked singular values of the Drosophila connectome decay much more rapidly than those of an i.i.d.
-matrix of the same size, which follow a universal distribution independent of single-element statistics (Fig.
-1c; similar analysis appears in Thibeault et al.
-[19]).
-We quantify this rapid decay using the participation ratio of the squared singular-value spectrum, $\text{PR}^{S}=(\sum_{a=1}^{M}S_{a}^{2})^{2}/\sum_{a=1}^{M}S_{a}^{4}$.
-The connectome’s participation ratio is 0.18, substantially lower than the 0.5 value characteristic of i.i.d.
-matrices, confirming that this biological network exhibits much more concentrated low-rank structure than would arise from random and independent couplings.
+matrix of the same size, which follow a universal distribution independent of single-element statistics [Fig.
+1(c); similar analysis appears in Thibeault et al.
+(Ref.
+[23])].
+We quantify this rapid decay using the participation ratio (PR) of the squared singular-value spectrum, $\text{PR}^{S}=(\sum_{a=1}^{M}S_{a}^{2})^{2}/\sum_{a=1}^{M}S_{a}^{4}$.
+The connectome’s participation ratio is 0.18, substantially lower than the value 0.5 characteristic of i.i.d.
+matrices, confirming that this biological network exhibits a much more concentrated low-rank structure than would arise from i.i.d.
+couplings.
 II.2 Definition of the random-mode model
@@
 coupling matrix.
-We therefore introduce the random-mode model, a generative model for coupling matrices that allows us to capture these properties within a random-matrix ensemble.
-The random-mode model has the same mathematical form as an SVD, but is a statistical generative process for $\bm{J}$ rather than a linear algebraic decomposition of it.
+We therefore introduce the random-mode model, a generative model for coupling matrices that allows us to capture these properties within a random-matrix ensemble (Fig.
+2).
+The random-mode model has the same mathematical form as an SVD, but is a statistical generative process for $\bm{J}$ rather than a matrix factorization.
 The model generates a coupling matrix of the form
 $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{\ell}_{a}\bm{r}_{a}^{T},$ (2)
-where, for each $a=1,\ldots,M$, $D_{a}>0$ are the component strengths; $\bm{r}_{a}$ and $\bm{\ell}_{a}$ are $N$-dimensional left and right modes, respectively; and $M$ is the number of components, where $\text{rank}=\text{min}(N,M)$.
+where, for each $a=1,\ldots,M$, $D_{a}>0$ are the component strengths; $\bm{r}_{a}$ and $\bm{\ell}_{a}$ are $N$-dimensional right and left modes, respectively; and $M$ is the number of components, where $\text{rank}\leq\text{min}(N,M)$.
 We now describe the properties of these quantities.
-Component strengths.
+See Appendix A for a glossary of terms.
+II.2.1 Component strengths.
 The component strengths $D_{a}$ are analogous to the singular values $S_{a}$ of $\bm{J}$.
 We model these quantities deterministically with the requirement that they are defined for arbitrarily large $M$ and that empirical averages $M^{-1}\sum_{a=1}^{M}f(D_{a})$ converge to limiting values, denoted $\left\langle f(D)\right\rangle_{D}$, as $M\to\infty$.
-Mode vectors.
+II.2.2 Left and right modes.
 The left and right modes $\bm{\ell}_{a}$ and $\bm{r}_{a}$ are analogous to the left and right singular vectors $\bm{u}_{a}$ and $\bm{v}_{a}$ of $\bm{J}$.
@@
 $\left\langle\ell_{a}\ell_{b}\right\rangle=\left\langle r_{a}r_{b}\right\rangle=\frac{1}{N}\delta_{ab}.$ (5)
-This yields orthonormality of mode vectors in expectation:
+This yields orthonormality of modes in expectation,
 $\begin{split}\left\langle\bm{\ell}_{a}^{T}\bm{\ell}_{b}\right\rangle&=\delta_{ab},\\ \left\langle\bm{r}_{a}^{T}\bm{r}_{b}\right\rangle&=\delta_{ab},\end{split}$ (6)
-in analogy to singular vectors for which orthonormality is deterministic  [^1] This random structure, rather than exact, deterministic orthonormality, offers analytical tractability by eliminating the need to enforce orthonormality—a complicated global constraint.
-But see [82], who handle random orthonormal matrices in a regression setting..
+in analogy to singular vectors, for which orthonormality holds exactly  [^1] This random structure, rather than exact, deterministic orthonormality, offers analytical tractability by eliminating the need to enforce orthonormality—a complicated global constraint; but see also Ref.
+[41], which handles random orthonormal matrices in a regression setting..
 Individual mode components are $\mathcal{O}(1/\sqrt{N})$.
@@
 II.3 Rank scaling and comparison to low-rank recurrent neural networks
-A crucial modeling choice is how the number of modes, $M$, scales relative to the network size, $N$.
+A crucial modeling choice is how the number of modes $M$ scales relative to the network size $N$.
 The two main possibilities are $M=\mathcal{O}(N)$ or $M=\mathcal{O}(1)$, corresponding to extensive and intensive rank scaling, respectively.
@@
 $\alpha=\frac{M}{N}=\mathcal{O}(1).$ (7)
-By choosing $\alpha$ to be small, or by choosing the spectrum of ranked component strengths to decay rapidly, we can construct connectivity that is low-rank, or well-described by a small number of rank-one components, compared to the number of neurons.
+By choosing $\alpha$ to be small, or by choosing the spectrum of ranked component strengths to decay rapidly, we can construct connectivity that is low rank or well described by a small number of rank-one components, compared to the number of neurons.
 The alternative approach of intensive rank scaling, where $M$ remains finite while $N\to\infty$, leads to what has been termed “low-rank recurrent neural networks” in the literature [42, 43, 44, 45].
@@
 matrix modeling unstructured “background” connectivity [42, 43].
-Note that in this paper, when we use “low-rank” or “low-dimensional,” we refer to quantities that are small as a fraction of $N$, even if they are extensive; this differs from some prior works (such as those on low-rank recurrent neural networks) where these terms refer to intensive quantities.
-The distinction between extensive and intensive rank scaling has important implications for specifying the distribution over the $2M$ mode components $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$.
+Note that, in this paper, when we use “low rank” or “low dimensional,” we refer to quantities that are small as a fraction of $N$, even if they are extensive; this differs from some prior works (such as those on low-rank recurrent neural networks) where these terms refer to intensive quantities.
+The distinction between extensive- and intensive-rank scaling has important implications for specifying the distribution over the $2M$ mode components $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$.
 For finite $M$, one can directly specify this distribution.
@@
 An important insight from prior work on the intensive-rank case is that these overlaps are crucial for shaping the dynamics and computations implemented by the network, as overlaps between connectivity modes dictate how activity modes interact [42, 43, 44, 45].
-Our goal is to go beyond the intensive-rank setup, which produces correspondingly low-dimensional activity, to model more sophisticated, higher-dimensional dynamics.
+Our goal is to go beyond the intensive-rank setup, which produces correspondingly low dimensional activity, to model more sophisticated, higher dimensional dynamics.
 However, specifying the full distribution $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ in the random-mode model, where $M\to\infty$, is more complicated.
-First, it must be well-defined for all $M$ to allow taking the limit.
+First, it must be well defined for all $M$ to allow taking the limit.
 More fundamentally, we would like to use the random-mode model to construct generative models that capture certain essential spectral features of coupling matrices, over which we have systematic control.
@@
 Moreover, until Sec.
-VII, we further specialize to the case where the left and right modes are fully independent: $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$, with each marginal factor $P(\ell_{a})$ or $P(r_{a})$ being a univariate Gaussian with mean zero and variance $1/N$  [^2] The Gaussian assumption for individual components is not crucial—our results depend only on first and second moments for large $N$ due to the central limit theorem; higher-order cumulants (scaling appropriately with $1/N$) are subdominant in the path integral saddle-point approximation..
-In this case, all $2MN$ components specifying the mode vectors are i.i.d.
+VII, we further specialize to the case where the left and right modes are fully independent: $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$, with each marginal factor $P(\ell_{a})$ or $P(r_{a})$ being a univariate Gaussian with mean zero and variance $1/N$  [^2] The Gaussian assumption for individual components is not crucial—our results depend on only first and second moments for large $N$ due to the central limit theorem; higher-order cumulants (scaling appropriately with $1/N$) are subdominant in the path-integral saddle-point approximation..
+In this case, all $2MN$ components specifying the modes are i.i.d.
 with mean zero and variance $1/N$; we refer to this case as “i.i.d.
 modes.” In Sec.
-VII we use a non-factorized form of $P(\ell_{a},r_{a})$ to specify correlations between $\ell_{a}$ and $r_{a}$, introducing $M$ deterministic parameters (in addition to the $M$ component strengths) to specify the left–right correlation for each $a$.
-II.4 Relationship to singular value decomposition
+VII we use a nonfactorized form of $P(\ell_{a},r_{a})$ to specify correlations between $\ell_{a}$ and $r_{a}$, introducing $M$ deterministic parameters (in addition to the $M$ component strengths) to specify the left-right correlation for each $a$.
+II.4 Relationship to singular-value decomposition
 In the SVD, the left and right singular vectors are orthonormal: $\bm{u}_{a}^{T}\bm{u}_{b}=\bm{v}_{a}^{T}\bm{v}_{b}=\delta_{ab}$.
-In contrast, the mode vectors are orthonormal only in expectation (Eq.
-6); once sampled, there are random $\mathcal{O}(1/\sqrt{N})$ overlaps:
+In contrast, the left and right modes are orthonormal only in expectation [Eq.
+(6)]; once sampled, there are random $\mathcal{O}(1/\sqrt{N})$ overlaps,
 $\begin{split}\bm{\ell}_{a}^{T}\bm{\ell}_{b}&=\delta_{ab}+\mathcal{O}(1/\sqrt{N}),\\ \bm{r}_{a}^{T}\bm{r}_{b}&=\delta_{ab}+\mathcal{O}(1/\sqrt{N}).\end{split}$ (9)
-When the effective rank (defined below) is small, the random-mode model “looks like” the SVD at large $N$ since the effects of the $\mathcal{O}(1/\sqrt{N})$ deviations from orthonormality in the sampled modes are negligible.
+When the effective rank (defined below) is small and $N$ is large, the component strengths and modes approximate singular values and vectors, respectively, since the effects of the $\mathcal{O}(1/\sqrt{N})$ deviations from orthonormality in the sampled modes are negligible.
 Indeed, a well-known result in high-dimensional geometry is that finite numbers of random vectors with i.i.d.
@@
 We now characterize this discrepancy analytically.
-We denote the $n$-th moments and participation ratio of the component-strength distribution by
+We denote the $n$th moments and participation ratio of the component-strength distribution by
 $r_{n}$ $=\left\langle D^{n}\right\rangle_{D},$ (10)
@@
 $\text{effective rank}=\alpha\text{PR}^{D}.$ (12)
-The effective rank is related to the participation ratio of the squared singular value spectrum  [^3] This follows from using Wick’s theorem to evaluate the expectations of $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ and $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ for the numerator and denominator, respectively, of $\text{PR}^{S}$.
+The effective rank is related to the participation ratio of the squared singular-value spectrum at large $N$  [^3] This follows from using Wick’s theorem to evaluate the expectations of $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ and $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ for the numerator and denominator, respectively, of $\text{PR}^{S}$.
 through
 $\text{PR}^{S}=\frac{\alpha\text{PR}^{D}}{1+2\alpha\text{PR}^{D}}.$ (13)
-Expanding in small $\alpha\text{PR}^{D}$, $\text{PR}^{S}=\alpha\text{PR}^{D}+\mathcal{O}((\alpha\text{PR}^{D})^{2})$, demonstrating that the component strengths are closely related to the singular values in the low-dimensional regime.
+Expanding in small $\alpha\text{PR}^{D}$, $\text{PR}^{S}=\alpha\text{PR}^{D}+\mathcal{O}\bm{(}(\alpha\text{PR}^{D})^{2}\bm{)}$, demonstrating that the component strengths are closely related to the singular values in the low dimensional regime.
 While we express the analytic results in this paper in terms of $\alpha\text{PR}^{D}$, the above equation can always be used to translate between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$.
 We further quantify this discrepancy in the case where $D_{a}=1$ for all $a$.
-Using methods from free probability theory [40, 41] (Appendix C), the $M$ singular values are distributed over a range with boundaries
+Using methods from free probability theory [49, 50] (Appendix C; Fig.
+8), the $M$ singular values are distributed over a range with boundaries
 $S_{\pm}=\sqrt{1+\frac{5\alpha}{2}-\frac{\alpha^{2}}{8}\pm\left(1+\frac{\alpha}{8}\right)^{3/2}\sqrt{8\alpha}}.$ (14)
 For small $\alpha$, $S_{\pm}=1\pm\sqrt{2\alpha}+\mathcal{O}(\alpha)$.
-While the random-mode model spreads the nonzero singular values over a range, this spread becomes negligible for small $\alpha$.
+Thus, while the random-mode model spreads the nonzero singular values over a range, this spread becomes negligible for small $\alpha$.
 More generally, the distributions of $D_{a}$ and $S_{a}$ coincide for small $\alpha\text{PR}^{D}$.
@@
 We study a recurrent neural network of $N$ neurons.
-Each neuron $i\in\{1,\ldots,N\}$ is characterized by its pre-activation $x_{i}(t)$ and activation $\phi_{i}(t)=\phi(x_{i}(t))$, where $\phi(\cdot)$ is a scalar nonlinearity; we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$  [^4] This behaves similarly to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals..
+Each neuron $i\in\{1,\ldots,N\}$ is characterized by its preactivation $x_{i}(t)$ and activation $\phi_{i}(t)=\phi\bm{(}x_{i}(t)\bm{)}$, where $\phi(\cdot)$ is a scalar nonlinearity; we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$, which is sigmoid-shaped and has $\phi^{\prime}(0)=1$  [^4] This behaves similarly to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals..
 The network dynamics are governed by
-$T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t),\>\>T[x](t)=(1+\partial_{t})x(t),$ (15)
-where $J_{ij}$ denotes the synaptic coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional that specifies the single-neuron dynamics.
+$T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t),$ (15a)
+$\text{where}\quad T[x](t)=(1+\partial_{t})x(t).$ (15b)
+Here, $J_{ij}$ denotes the synaptic coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional that specifies the single-neuron dynamics for which Eq.
+(15b) is a canonical choice.
 We expect our results to be agnostic to the specific choice of $T[\cdot]$.
@@
 couplings.
-The dimension can be computed either by computing the statistics of the off-diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
-Let $a\in\{x,\phi\}$ denote either pre-activations or activations. We define two complementary covariance functions, averaging over either time or neurons:
+The dimension can be computed either by computing the statistics of the off diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
+Let $a\in\{x,\phi\}$ denote either preactivations or activations. We define two complementary covariance functions, averaging over either time or neurons,
 $C^{a}_{ij}(\tau)$ $=\left\langle a_{i}(t)a_{j}(t+\tau)\right\rangle_{t},$ (16)
 $C^{a}(t_{1},t_{2})$ $=\frac{1}{N}\sum_{i=1}^{N}a_{i}(t_{1})a_{i}(t_{2}).$ (17)
-These functions exhibit a form of duality: for any lag $\tau$, the matrix $\bm{C}^{a}(\tau)$ with elements $C^{a}_{ij}(\tau)$ indexed by $(i,j)$ and the matrix with elements $C^{a}(t_{1},t_{2}+\tau)$ indexed by $(t_{1},t_{2})$ have the same eigenvalue spectrum up to scaling (to see this, note that both matrices can be computed from a time-by-neuron activity matrix $\bm{A}$ as $\bm{A}^{T}\bm{A}$ and $\bm{A}\bm{A}^{T}$, respectively, which have identical spectra)  [^5] This duality is reminiscent of ergodicity theorems, which relate temporal and spatial averages, but differs in a variety of ways, including that it holds even for finite time intervals, whereas ergodicity concerns infinite-time limits..
-We also define the neuron-averaged response function, describing propagation of infinitesimal perturbations:
-$S^{a}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\left.\frac{\delta a_{i}(t_{1})}{\delta I_{j}(t_{2})}\right|_{I_{j}=0},$ (18)
-where $I_{j}(t)$ is a source term added to the right-hand side of the network dynamics: $T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t)+I_{i}(t)$. Finally, we defined “stationarized” counterparts:
+These functions exhibit a form of duality: for any lag $\tau$, the matrix $\bm{C}^{a}(\tau)$ with elements $C^{a}_{ij}(\tau)$ indexed by $(i,j)$ and the matrix with elements $C^{a}(t_{1},t_{2}+\tau)$ indexed by $(t_{1},t_{2})$ have the same eigenvalue spectrum, up to scaling (to see this, note that both matrices can be computed from a time-by-neuron activity matrix $\bm{A}$ as $\bm{A}^{T}\bm{A}$ and $\bm{A}\bm{A}^{T}$, respectively, which have identical spectra)  [^5] This duality is reminiscent of ergodicity theorems, which relate temporal and spatial averages, but differs in a variety of ways, including that it holds even for finite-time intervals, whereas ergodicity concerns infinite-time limits..
+We also define the neuron-averaged response function, describing the propagation of infinitesimal perturbations:
+$S^{a}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\left.\frac{\delta a_{i}(t_{1})}{\delta I_{i}(t_{2})}\right|_{\bm{I}=0},$ (18)
+where $I_{i}(t)$ is a source term added to the right-hand side of the network dynamics: $T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t)+I_{i}(t)$. Finally, we define “stationarized” counterparts,
 $C^{a}(\tau)=\left\langle C^{a}(t,t\pm\tau)\right\rangle_{t},\>\>\>\>S^{a}(\tau)=\left\langle S^{a}(t,t-\tau)\right\rangle_{t}.$ (19)
@@
 In contrast to the above two-point functions, which characterize single-neuron activity, four-point functions describe features of collective activity.
-The same four-point functions were used in [30].
+The same four-point functions were used in Ref.
+[39].
 Here, we motivate these functions in more detail and express them in two ways via the duality between neuronal and temporal covariances.
 To motivate their definition, consider the dimension of activity, which can be quantified as the participation ratio of the spectrum of the equal-time covariance matrix $\bm{C}^{a}(0)$ [53, 54, 36, 55, 56, 57, 9, 58, 59].
-Given the eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$ for $k=1,\ldots,N$ (as one would compute as part of principal components analysis), the participation ratio is
+Given the eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$ for $k=1,\ldots,N$, as one would compute as part of principal components analysis, the participation ratio is
 $\text{PR}^{a}=\frac{1}{N}\frac{(\sum_{k=1}^{N}\lambda^{a}_{k})^{2}}{\sum_{k=1}^{N}(\lambda^{a}_{k})^{2}}.$ (20)
@@
   • Trivial extensive dimensionality: $\text{PR}^{a}=1$. All dimensions are filled equally by activity.
-An advantageous property of the participation ratio is that it can be expressed as the ratio of the squared-trace and Frobenius norm of $\bm{C}^{a}(0)$, both of which can be further expressed in terms of matrix elements $C_{ij}^{a}(0)$:
-$\text{PR}^{a}$ $=\frac{\left(\text{tr}\bm{C}^{a}(0)\right)^{2}}{\lVert\bm{C}(0)\rVert_{F}^{2}}$
-$=\frac{C^{a}(0)^{2}}{\frac{1}{N}\sum_{i=1}^{N}C_{ii}(0)^{2}+\frac{1}{N}\sum_{i\neq j}C_{ij}(0)^{2}},$ (21)
+An advantageous property of the participation ratio is that it can be expressed as the ratio of the squared trace and Frobenius norm of $\bm{C}^{a}(0)$, both of which can be further expressed in terms of matrix elements $C_{ij}^{a}(0)$,
+$\text{PR}^{a}$ $=\frac{\left(\text{tr}\bm{C}^{a}(0)\right)^{2}}{\lVert\bm{C}^{a}(0)\rVert_{F}^{2}}$
+$=\frac{C^{a}(0)^{2}}{\frac{1}{N}\sum_{i=1}^{N}C^{a}_{ii}(0)^{2}+\frac{1}{N}\sum_{i\neq j}C^{a}_{ij}(0)^{2}},$ (21)
 where $C^{a}(0)$ is given by Eq.
-19.
+(19).
 Assuming that diagonal elements are uniform with negligible fluctuations, i.e., $C^{a}_{ii}(0)=C^{a}(0)$ for all $i$ (as occurs, for example, under i.i.d.
@@
 $\text{PR}^{a}=\frac{C^{a}(0)^{2}}{C^{a}(0)^{2}+\psi^{a}(0,0)},$ (22)
-where, following [30], we define the four-point function
+where, following Ref. [39], we define the four-point function
 $\psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i\neq j}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}),$ (23)
@@
 Thus, $\psi^{a}(0,0)\to\text{const.}$ as $N\to\infty$, leading to nontrivial extensive dimensionality, $0<\text{PR}^{a}<1$.
-In this paper, we focus on a related function $\Psi^{a}(\bm{\tau})$ that includes the diagonal terms:
+In this paper, we focus on a related function $\Psi^{a}(\bm{\tau})$ that includes the diagonal terms,
 $\Psi^{a}(\bm{\tau})=\frac{1}{N}\sum_{i,j=1}^{N}C^{a}_{ij}(\tau_{1})C^{a}_{ij}(\tau_{2}).$ (24)
@@
 $\Psi^{a}(\bm{\tau})=N\left\langle C^{a}(t_{1},t_{2})C^{a}(t_{1}+\tau_{1},t_{2}+\tau_{2})\right\rangle_{t_{1},t_{2}}.$ (25)
-The neuron-by-neuron (Eq.
-24) and time-by-time (Eq.
-25) definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively.
+The neuron-by-neuron [Eq.
+(24)] and time-by-time [Eq.
+(25)] definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively.
 Finally, the dimension of activity is given by
 $\text{PR}^{a}=\frac{C^{a}(0)^{2}}{\Psi^{a}(0,0)},$ (26)
-which, unlike Eq. 22, holds even when single-neuron variances $C^{a}_{ii}(0)$ are nonuniform across neurons.
-To study timescales of collective activity, we consider the principal components of activity:
+which, unlike Eq. (22), holds even when single-neuron variances $C^{a}_{ii}(0)$ are nonuniform across neurons.
+To study timescales of collective activity, we consider the principal components of activity,
 $p^{a}_{k}(t)=\frac{1}{\sqrt{\lambda^{a}_{k}}}(\bm{v}^{a}_{k})^{T}\bm{a}(t).$ (27)
-These principal components are the basis of much of modern analysis of high-dimensional neural data.
-They all have unit variance due to the $1/\sqrt{\lambda^{a}_{k}}$ normalization, but potentially very different characteristic timescales (Fig.
-9a).
+These principal components are the basis of much of modern analysis of high-dimensional neural data [35, 36, 37, 38].
+They all have unit variance due to the $1/\sqrt{\lambda^{a}_{k}}$ normalization, but potentially very different characteristic timescales [Fig.
+9(a)].
 To extract the timescales of just the leading components, we weight them by the squares of their corresponding eigenvalues.
-This gives:
+This gives
 $\frac{1}{N}\sum_{k}(\lambda^{a}_{k})^{2}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}$ $=\Psi^{a}(\tau,0),$ (28)
-showing that $\Psi^{a}(\bm{\tau})$ captures the temporal structure of the leading principal components of activity (Fig.
-9c,d).
-Note that weighting by the eigenvalues themselves, rather than their squares, gives $N^{-1}\sum_{k}\lambda^{a}_{k}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}=C^{a}(\tau)$, recovering single-neuron information (Fig.
-9b,d).
+showing that $\Psi^{a}(\bm{\tau})$ captures the temporal structure of the leading principal components of activity [Figs.
+9(c) and 9(d)].
+Note that weighting by the eigenvalues themselves, rather than their squares, gives $N^{-1}\sum_{k}\lambda^{a}_{k}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}=C^{a}(\tau)$, recovering single-neuron information [Figs.
+9(b) and 9(d)].
 IV Path-integral analysis of the random-mode model
@@
 The summary statistics we have defined are self-averaging, meaning that the same values are obtained, up to $\mathcal{O}(1/\sqrt{N})$ fluctuations, when they are computed based on a single realization of a large network.
-In this paper, we are interested in statistically stationary states such that, for $N\to\infty$, the two-point functions depend on only time differences and thus can be replaced by their stationarized counterparts (Eq.
-19), i.e., for all $t$,
+In this paper, we are interested in statistically stationary states such that, for $N\to\infty$, the two-point functions depend on time differences only and thus can be replaced by their stationarized counterparts [Eq.
+(19)], i.e., for all $t$,
 $C^{a}_{\star}(t,t\pm\tau)=C^{a}_{\star}(\tau),\>\>\>\>S_{\star}^{a}(t,t-\tau)=S^{a}_{\star}(\tau).$ (30)
-Calculating two-point functions is well established “standard DMFT.” As we review in Appendix E, four-point functions were previously calculated under i.i.d.
-connectivity using a two-site version of the cavity method [30].
-In this section, we introduce a new path-integral approach to calculating four-point functions that allows for rapid solution of the i.i.d.
-model and, importantly, more complicated models including the random-mode model.
+Calculating two-point functions is well-established “standard DMFT” [60, 61, 62].
+As we review in Appendix E, four-point functions were previously calculated for networks with i.i.d.
+couplings using a two-site version of the cavity method [39].
+In this section, we introduce a new path-integral approach to calculating four-point functions that allows for the rapid solution of the i.i.d.
+model and generalizes to more complicated models, including the random-mode model.
 IV.1 Relating $\Psi^{a}$ to temporal covariance fluctuations
-To cast the problem in the language of path integrals, we turn to the time-by-time definition of $\Psi^{a}(\bm{\tau})$ (Eq.
-25).
-If the time-by-time covariance $C^{a}(t_{1},t_{2})$ were equal to its limiting value, $C^{a}_{\star}(t_{2}-t_{1})$, we would have $\Psi_{\star}^{a}(\bm{\tau})\overset{?}{=}N\left\langle C^{a}_{\star}(t_{2}-t_{1})C^{a}_{\star}(t_{2}-t_{1}+\tau_{2}-\tau_{1})\right\rangle_{t_{1},t_{2}}=0$} since, for large time differences, $C^{a}_{\star}(t_{2}-t_{1})$, decays exponentially in $|t_{2}-t_{1}|$ [10].
-However, based on the neuron-by-neuron definition (Eq.
-24), $\Psi_{\star}^{a}(\bm{\tau})$ is clearly a nonzero, order-one quantity.
+To cast the problem in the language of path integrals, we turn to the time-by-time definition of $\Psi^{a}(\bm{\tau})$ [Eq.
+(25)].
+If the time-by-time covariance $C^{a}(t_{1},t_{2})$ were equal to its limiting value, $C^{a}_{\star}(t_{2}-t_{1})$, we would have
+$\Psi_{\star}^{a}(\bm{\tau})\overset{?}{=}N\left\langle C^{a}_{\star}(t_{2}-t_{1})C^{a}_{\star}(t_{2}-t_{1}+\tau_{2}-\tau_{1})\right\rangle_{t_{1},t_{2}}=0$} (31)
+since, for large time differences, $C^{a}_{\star}(t_{2}-t_{1})$ decays exponentially in $|t_{2}-t_{1}|$ [10].
+However, based on the neuron-by-neuron definition [Eq.
+(24)], $\Psi_{\star}^{a}(\bm{\tau})$ is clearly a nonzero, order-one quantity.
 This apparent contradiction is resolved by considering the $\mathcal{O}(1/\sqrt{N})$ fluctuations around the limiting value.
-The necessity of these fluctuations becomes clear when considering the duality between neuronal and temporal covariances: the presence of nonzero cross-covariances $C^{a}_{ij}(0)$ indicates low-dimensional structure in the system; equivalently, this structure manifests as the system revisiting similar states over time, resulting in nonzero temporal covariance $C^{a}(t_{1},t_{2})$ even for large time differences $|t_{2}-t_{1}|\gg 1$ (Fig.
+The necessity of these fluctuations becomes clear when considering the duality between neuronal and temporal covariances: the presence of nonzero cross-covariances $C^{a}_{ij}(0)$ indicates low dimensional structure in the system; equivalently, this structure manifests as the system revisiting similar states over time, resulting in nonzero temporal covariance $C^{a}(t_{1},t_{2})$ even for large time differences $|t_{2}-t_{1}|\gg 1$ (Fig.
 3).
@@
 noting that terms involving $C^{a}_{\star}(t_{2}-t_{1})$ vanish at large time differences.
-Our task, therefore, is to compute the covariance of the fluctuations, $\delta C^{a}(t_{1},t_{2})$, captured asymtotically by
+Our task, therefore, is to compute the covariance of the fluctuations, $\delta C^{a}(t_{1},t_{2})$, captured asymptotically by
 $\left\langle N\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{3},t_{4})\right\rangle_{\star}.$ (34)
-Then, to compute $\Psi_{\star}^{a}(\bm{\tau})$, we take in Eq. 33 the limits
+Then, to compute $\Psi_{\star}^{a}(\bm{\tau})$, we take in Eq. (34) the limits
 $t_{3}$ $=t_{1}+\tau_{1},$ $|t_{2}-t_{1}|$ $\to\infty,$ (35)
@@
 IV.2 Calculation of two- and four-point functions
-Details of our calculation are given in Appendix F.
-Briefly, we compute the covariance of fluctuations (Eq.
-33) using the Martin-Siggia-Rose-De Dominicis-Janssen (MSRDJ) path-integral formalism [51, 52, 53, 54, 55, 56].
+Details of the calculation are given in Appendix F.
+Briefly, we compute the covariance of fluctuations [Eq.
+(34)] using the Martin-Siggia-Rose-Janssen–de Dominicis (MSRJD) path-integral formalism [63, 64, 65, 60, 61, 62].
 For any fixed coupling matrix $\bm{J}$, the path integral is
 $Z[\bm{J}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\\ \exp\Bigg\{i\sum_{i=1}^{N}\int dt\hat{x}_{i}(t)\Bigg(T[{x}_{i}](t)-\sum_{j=1}^{N}J_{ij}\phi_{j}(t)\Bigg)\Bigg\}.$ (36)
-The auxiliary field $\hat{x}_{i}(t)$ is conjugate to $x_{i}(t)$ and enforces the network’s equations of motion (Eq.
-15) through the delta-function representation $(2\pi)^{-1}\int d\hat{x}\,e^{i\hat{x}x}=\delta(x)$.
-Upon introduction of source terms that are not needed for our purposes, this path integral serves as a generating functional for correlation and response functions.
-In principle, to compute $\Psi^{a}_{\star}(\bm{\tau})$ in this framework, after integrating out $\bm{J}$ from $Z[\bm{J}]$ to obtain a statistical field theory governing two-point functions, we (1) find the saddle-point solution for two-point functions by extremizing the action (i.e., standard DMFT); (2) compute the time2-by-time2 Hessian matrix of the action describing fluctuations around this saddle point; (3) invert this Hessian and extract the sub-block corresponding to the covariance of $\delta C^{\phi}(t_{1},t_{2})$ (Eq.
-33); and (4) apply the temporal limits (Eq.
-34) to this sub-block to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$.
+The auxiliary field $\hat{x}_{i}(t)$ is conjugate to $x_{i}(t)$ and enforces the network’s equations of motion [Eq.
+(15a)] through the $\delta$-function representation $(2\pi)^{-1}\int d\hat{x}\,e^{i\hat{x}x}=\delta(x)$.
+Upon introduction of source terms (that are not needed for our purposes), this path integral serves as a generating functional for correlation and response functions.
+In principle, to compute $\Psi^{a}_{\star}(\bm{\tau})$ in this framework, after integrating out $\bm{J}$ from $Z[\bm{J}]$ to obtain a statistical field theory governing two-point functions, we (1) find the saddle-point solution for two-point functions by extremizing the action (i.e., standard DMFT); (2) compute the time2-by-time2 Hessian matrix of the action describing fluctuations around this saddle point; (3) invert this Hessian and extract the sub-block corresponding to the covariance of $\delta C^{\phi}(t_{1},t_{2})$ [Eq.
+(34)]; and (4) apply the temporal limits [Eq.
+(35)] to this sub-block to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$.
 Direct inversion of the time2-by-time2 Hessian is analytically intractable.
-However, we show that the temporal limits commute with Hessian inversion, allowing us to invert a low-dimensional frequency-dependent Hessian instead (Appendix G).
+However, we show that the temporal limits commute with Hessian inversion, allowing us to invert a low dimensional frequency-dependent Hessian instead (Appendix G).
 This readily yields the Fourier-space function $\Psi^{\phi}_{\star}(\bm{\omega})$, where $\bm{\omega}=(\omega_{1},\omega_{2})$.
 A few additional steps yield $\Psi^{x}_{\star}(\bm{\omega})$.
-This approach allows for the efficient calculation of two and four point functions for a given model, requiring only a few lines of computation directly from the action of the theory.
-As a demonstration of this approach, we apply it to a generalization of the i.i.d. model with correlated reciprocal couplings, recovering results of [30] (Appendix H).
+This approach allows for the efficient calculation of two- and four-point functions for a given model, requiring only a few lines of computation directly from the action of the theory.
+As a demonstration of this approach, we apply it to a generalization of the i.i.d. model with correlated reciprocal couplings, recovering results of Ref. [39] (Appendix H).
 IV.3 Solution of the random-mode model
@@
 $\eta^{x}$ $\sim\mathcal{GP}(0,\alpha r_{2}C^{\phi}_{\star}),$ (37b)
-$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star}.$ (36c)
-These equations are equivalent those for an i.i.d. model (Eqs. 6971) with coupling strength
+$C_{\star}^{\phi}(t_{1},t_{2})$ $=\left\langle\phi(t_{1})\phi(t_{2})\right\rangle_{\star},$ (37c)
+where Eq.
+(37b) indicates that $\eta^{x}(t)$ is a Gaussian field with zero mean and correlation function $\alpha r_{2}C^{\phi}_{\star}(t,t^{\prime})$.
+These equations are equivalent to those for an i.i.d.
+model [Eqs.
+(70)–(72)] with coupling strength
 $g_{\text{eff}}=\sqrt{\alpha r_{2}}.$ (38)
-The four-point function is (Appendix I.2)
+The four-point function, our main object of interest, is (Appendix I.2)
 $\Psi_{\star}^{\phi}=\frac{1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}{|1-g_{\text{eff}}^{2}S_{12}^{\phi}|^{2}}C_{12}^{\phi},$ (39)
 where we have simplified notation by suppressing frequency variables via the shorthand $C^{\phi}_{12}=C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2})$ and $S^{\phi}_{12}=S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})$.
-The same solution is obtained through a more involved, but intuitive, two-site cavity calculation (Appendix L).
-This solution for the random-mode model is one of our main results that we now spend some time interpreting.
+The same solution is obtained through a more involved, but intuitive, two-site cavity calculation (Appendix L; Fig.
+11).
+This solution for the random-mode model is one of our main results, which we now spend some time interpreting.
 IV.4 Interpreting the solution
@@
 (a) Activity dimension $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ for various coupling strengths $g_{\text{eff}}$.
-Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
+Thin dots, individual simulations; thick dots, means over ten simulations; lines, theoretical predictions.
 Inset: extended $g_{\text{eff}}=6$ case, showing convergence to i.i.d.
@@
 Our analysis reveals two features of connectivity that characterize how it shapes activity, namely, the effective coupling strength $g_{\text{eff}}$ and the effective rank $\alpha\text{PR}^{D}$.
-The former characterizes local structure (the size of individual couplings) while the latter characterizes global structure (the number of large connectivity components).
-In the equations determining $C_{\star}^{\phi}(\tau)$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ above, we see immediately that global structure in the connectivity is invisible at the single-neuron level but shapes collective activity.
+The former characterizes local structure (the magnitudes of individual couplings), while the latter characterizes global structure (the number of large connectivity components).
+The above equations determining $C_{\star}^{\phi}(\tau)$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ immediately reveal that global structure in the connectivity is invisible at the single-neuron level but shapes collective activity.
 Specifically, the correlation function $C^{\phi}_{\star}(\tau)$ is fully determined by $g_{\text{eff}}$ and is insensitive to the effective rank of connectivity.
 This has important implications for neural data analysis.
-In particular, analyses of single-neuron, non-collective activity properties may miss signatures of structured connectivity that only become apparent when analyzing collective activity properties among simultaneously recorded neurons.
+In particular, analyses of single-neuron activity properties may miss signatures of structured connectivity that become apparent only when analyzing collective activity properties among simultaneously recorded neurons.
 What is the nature of the dependence of collective-activity properties, such as the dimension of activity, on the effective rank?
-To plot this relationship and validate the theory, we simulated networks and computed the participation ratio from the empirical equal-time covariance matrix of the activations.
-We used a component-strength spectrum $D_{a}\propto\exp(-\beta_{D}a/M)$ which yields $\text{PR}^{D}=\beta^{-1}_{D}\tanh(\beta_{D})$.
-To attain a desired effective rank $\alpha\text{PR}^{D}<1$, we set $\alpha=M/N=1$ and solved for $\beta$.
-For $\alpha\text{PR}^{D}>1$, we fixed $\text{PR}^{D}=1$ and increased $\alpha$.
+To plot this relationship and validate the theory, we simulated networks and computed the participation ratio from the empirical equal-time covariance matrix of the activations (see Appendix M for numerical details).
+We used a component-strength spectrum $D_{a}\propto\exp(-\beta_{D}a/M)$, which yields $\text{PR}^{D}=\beta^{-1}_{D}\tanh(\beta_{D})$.
+To attain a desired effective rank $\alpha\text{PR}^{D}<1$, we set $\alpha=M/N=1$ and solved for $\beta$; for $\alpha\text{PR}^{D}>1$, we fixed $\text{PR}^{D}=1$ and increased $\alpha$.
 We adjusted $g_{\text{eff}}$ by rescaling $D_{a}$ uniformly.
@@
 4).
-Fig.
-4a illustrates how the dimension of activity, $\text{PR}^{\phi}$, varies with $\alpha\text{PR}^{D}$ for different values of $g_{\text{eff}}$.
+Figure 4(a) illustrates how the dimension of activity, $\text{PR}^{\phi}$, varies with $\alpha\text{PR}^{D}$ for different values of $g_{\text{eff}}$.
 Consistent with i.i.d.-coupling networks, $\text{PR}^{\phi}$ increases monotonically with $g_{\text{eff}}$ [39].
@@
 Taking $\alpha\text{PR}^{D}\to\infty$ for fixed $g_{\text{eff}}$ in Eq.
-38 recovers the formula for i.i.d.
-couplings (Eq.
-72) of [30].
+(39) recovers the formula for i.i.d.
+couplings [Eq.
+(73)] of Ref.
+[39].
 Indeed, in this limit, $\bm{J}$ approaches an i.i.d.
 matrix.
-Correspondingly, $\text{PR}^{\phi}$ approaches the dimension of a network with i.i.d.
-couplings (Fig.
-4a inset).
+Correspondingly, $\text{PR}^{\phi}$ approaches the activity dimension of a network with i.i.d.
+couplings [Fig.
+4(a) inset].
 Comparing Eq.
-72 to the corresponding i.i.d.
+(73) to the corresponding i.i.d.
 formula ($1/\alpha\text{PR}^{D}=0$) and noting that single-neuron properties are unaffected by $\alpha\text{PR}^{D}$, one sees that any structure beyond i.i.d.
@@
 The dimension of activity depends on $\Psi^{\phi}_{\star}(\bm{\tau})$ at $\bm{\tau}=(0,0)$; we now examine the temporal profile of this function.
-Fig.
-4b shows a normalized version of $\Psi^{\phi}(\tau,0)$, which is related to the correlation functions of leading principal components, for various values of $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$.
+Figure 4(b) shows a normalized version of $\Psi^{\phi}(\tau,0)$, which is related to the correlation functions of leading principal components, for various values of $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$.
 The decay timescale of $\Psi^{\phi}_{\star}(\tau,0)$ decreases with increasing $g_{\text{eff}}$, approaching a limiting behavior at $g_{\text{eff}}\approx 10$.
-While $\alpha\text{PR}^{D}$ affects the overall scale of this function, and consequently the dimension of activity, it has little effect on the decay timescale, and thus little effect on the timescales of leading principal components (Fig.
-4b, inset).
+While $\alpha\text{PR}^{D}$ affects the overall scale of this function and, consequently, the dimension of activity, it has little effect on the decay timescale and, hence, little effect on the timescales of leading principal components (Fig.
+4(b), inset).
 Thus, the effective rank primarily influences the dimension of activity rather than the collective temporal structure.
-A remarkable feature of these results is that only two connectivity features, $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$, are sufficient to determine the two- and four-point functions.
-Due to the nonlinear, high-dimensional nature of the network, this reduction is unexpected.
-Indeed, this reduction breaks down when mode overlaps are introduced (Sec.
+A remarkable feature of these results is that only two connectivity statistics, $g_{\text{eff}}$ and $\alpha\text{PR}^{D}$, are sufficient to determine the two- and four-point functions.
+Because of the nonlinear, high-dimensional nature of the network, this reduction is unexpected.
+This reduction breaks down when structured mode overlaps are introduced (Sec.
 VII), in which case the formula for the four-point function requires knowledge of the full joint distribution of component strengths and overlaps.
 V Limiting behavior of the random-mode model
-Here we examine the behavior of the random-mode model in two limiting cases: as the system approaches the transition to chaos from above ($g_{\text{eff}}\to 1^{+}$), and in the limit of low effective rank ($\alpha\text{PR}^{D}\to 0^{+}$).
+Here we examine the behavior of the random-mode model in two limiting cases: as the system approaches the transition to chaos from above ($g_{\text{eff}}\to 1^{+}$) and in the limit of low effective rank ($\alpha\text{PR}^{D}\to 0^{+}$).
 We first consider the limit $g_{\text{eff}}\to 1^{+}$ for arbitrary $\alpha\text{PR}^{D}$.
-The relevant expansion parameter is $g_{\text{eff}}-1\ll 1$.
+The relevant expansion parameter is $0<g_{\text{eff}}-1\ll 1$.
 For networks with i.i.d.
-couplings with coupling strenght $g$, it is known that, as $g\to 1^{+}$, $\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})\sim 1/(g-1)$ and $\text{PR}^{a}_{\text{i.i.d.}}\sim(g-1)^{3}$ in this limit [30].
+couplings with coupling strength $g$, it is known that, as $g\to 1^{+}$, $\Psi^{a}_{\star,\text{i.i.d.}}(\bm{\tau})\sim 1/(g-1)$ and $\text{PR}^{a}_{\text{i.i.d.}}\sim(g-1)^{3}$ [39].
 For the random-mode model, to leading order in $1/(g_{\text{eff}}-1)$,
@@
 $\text{PR}^{a}=\alpha\text{PR}^{D}\text{PR}^{a}_{\text{i.i.d.}}$. (42)
-The difference between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$ is a higher-order correction, so we also have $\text{PR}^{a}=\text{PR}^{S}\text{PR}^{a}_{\text{i.i.d.}}$.
-That is, in the limits of both low effective rank and small effective coupling strength, the dimension of activity is equal to the dimension of the singular-value spectrum times the dimension of activity in an i.i.d.
+The difference between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$ is a higher-order correction, so we also have $\text{PR}^{a}=\text{PR}^{S}\text{PR}^{a}_{\text{i.i.d.}}$; that is, in the limits of both low effective rank and small effective coupling strength, the dimension of activity is equal to the dimension of the singular-value spectrum multiplied by the dimension of activity for an i.i.d.
 network.
@@
 We verify this limiting behavior in simulations using $\alpha=1$ and a step function for the component strengths: $D_{a}=\text{const.}$ for $a/N\leq\text{PR}^{D}$ and $D_{a}=0$ otherwise.
-We varied $\text{PR}^{D}$, observing linear relationships with $\text{PR}^{\phi}$ and $\text{PR}^{x}$ in the $\text{PR}^{D}\to 0^{+}$ limit (Fig.
-10a).
-Focusing on $K^{\phi}(g_{\text{eff}})$, note that its expression (Eq.
-43) resembles the formula for the dimension of activity in an i.i.d.
-network with coupling strength $g_{\text{eff}}$ (Eq.
-72), but includes an additional factor $|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}$ in the integral.
+We varied $\text{PR}^{D}$, observing linear relationships with $\text{PR}^{\phi}$ and $\text{PR}^{x}$ in the $\text{PR}^{D}\to 0^{+}$ limit [Fig.
+10(a)].
+Focusing on $K^{\phi}(g_{\text{eff}})$, note that its expression [Eq.
+(44)] resembles the formula for the dimension of activity in an i.i.d.
+network with coupling strength $g_{\text{eff}}$ [Eq.
+(73)], but includes an additional factor $|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}$ in the integral.
 Numerical evaluation confirms that, as expected from Eq.
-41, $K^{\phi}(g_{\text{eff}})\to\text{PR}^{\phi}_{\text{i.i.d.}}$ as $g_{\text{eff}}\to 1^{+}$.
-As $g_{\text{eff}}$ increases, $K^{\phi}(g_{\text{eff}})$ increases monotonically to $\sim$$1.53\,\text{PR}^{\phi}_{\text{i.i.d.}}$ (Fig.
-10a, inset).
+(42), $K^{\phi}(g_{\text{eff}})\to\text{PR}^{\phi}_{\text{i.i.d.}}$ as $g_{\text{eff}}\to 1^{+}$.
+As $g_{\text{eff}}$ increases, $K^{\phi}(g_{\text{eff}})$ increases monotonically to approximately $1.53\,\text{PR}^{\phi}_{\text{i.i.d.}}$ [Fig.
+10(a), inset].
 Thus, for networks with low effective rank, $\text{PR}^{\phi}\sim\alpha\text{PR}^{D}\text{PR}^{\phi}_{\text{i.i.d.}}$ up to an order-one fudge factor that becomes exactly one as $g_{\text{eff}}\to 1^{+}$.
-Finally, we consider the dependence of the full spectrum eigenvalues $\lambda^{a}_{i}$ of $\bm{C}^{a}(0)$ on the spectrum of component strengths $D_{a}$.
+Finally, we consider the dependence of the full eigenvalue spectrum $\lambda^{a}_{i}$ of $\bm{C}^{a}(0)$ on the spectrum of component strengths $D_{a}$.
 The latter leads to the former via a highly nonlinear relationship.
-Indeed, sticking to the step-function component-strength spectrum, the ranked eigenvalues $\lambda^{\phi}$ and $\lambda^{x}$ of the covariance matrix decay smoothly rather than exhibiting a cutoff (Fig.
-10b).
+Indeed, using the same step-function component-strength spectrum, the ranked eigenvalues $\lambda^{\phi}$ and $\lambda^{x}$ of the covariance matrix decay smoothly rather than exhibiting a cutoff [Fig.
+10(b)].
 Nevertheless, certain properties of these connectivity and activity spectra are linked.
-In particular, in the limit $\alpha\text{PR}^{D}\to 0^{+}$, $\alpha\text{PR}^{D}$ and $\text{PR}^{a}$ exhibit a linear relationship (with proportionality factor $K^{a}(g_{\text{eff}})$), implying that the ratios of the second-squared and fourth moments of these spectra are linked (despite that, even in this limit, the full spectra do not coincide).
+In particular, in the limit $\alpha\text{PR}^{D}\to 0^{+}$, $\alpha\text{PR}^{D}$ and $\text{PR}^{a}$ exhibit a linear relationship [with proportionality factor $K^{a}(g_{\text{eff}})$], implying that the ratios of the squared second moments and fourth moments of these spectra are linked (despite, even in this limit, the full spectra not coinciding).
 VI Single-neuron heterogeneity
@@
 (a) Dimension of normalized activity $\text{PR}^{\phi}$ versus participation ratio of gain distribution $\text{PR}^{G}$ for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$.
-Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
+Thin dots, individual simulations; thick dots, means over ten simulations; lines, theoretical predictions.
 (b) Fractional reduction in dimension for weighted activations relative to $\text{PR}^{\phi}$.
-Blue: $G_{i}\phi_{i}$; purple: $G^{\text{readout}}_{i}\phi_{i}$.
+Blue, $G_{i}\phi_{i}$; purple, $G^{\text{readout}}_{i}\phi_{i}$.
 $g_{\text{eff}}=10$.
-Thick dots: means over 10 simulations, averaging before taking the ratio; lines: theoretical predictions.
+Thick dots, means over ten simulations, averaging before taking the ratio; lines, theoretical predictions.
 All simulations use $N=5000$ neurons.
@@
 This heterogeneity among neurons is analogous to the heterogeneity of component strengths in the random-mode model, but operates in the neuron basis rather than the component basis.
-We now compare how low-dimensional connectivity structure and heterogeneous neuronal properties affect collective activity by extending our framework to include single-neuron heterogeneity.
+We now compare how low dimensional connectivity structure and heterogeneous neuronal properties affect collective activity by extending our framework to include single-neuron heterogeneity.
 In a general formulation described in Appendix J, we extend the nonlinearity of neuron $i$ to $\Phi_{\bm{\theta}_{i}}(x)$, which depends on a vector of neuron-specific parameters $\bm{\theta}_{i}$, and solve for the resulting two- and four-point functions.
@@
 $\text{PR}^{G}$ $=\frac{q_{2}^{2}}{q_{4}}.$ (46)
-The single-site picture that determines the two-point functions is largely similar to Eq. 36c.
-We analyze collective activity (four-point functions) for (1) unnormalized activations $\Phi_{\bm{\theta}_{i}}(t)=G_{i}\phi(x_{i}(t))$ and (2) normalized activations $\phi_{i}(t)=\phi(x_{i}(t))$, the latter being analogous to $z$-scored firing rates in neural recordings, where $z$-scoring removes single-neuron heterogeneity while preserving structure from cross-neuron and temporal correlations.
+The single-site picture that determines the two-point functions is largely similar to Eq. (37c).
+We analyze collective activity (four-point functions) for (1) unnormalized activations $\Phi_{\bm{\theta}_{i}}(t)=G_{i}\phi\bm{(}x_{i}(t)\bm{)}$ and (2) normalized activations $\phi_{i}(t)=\phi\bm{(}x_{i}(t)\bm{)}$, the latter being analogous to $z$-scored firing rates in neural recordings, where $z$ scoring removes single-neuron heterogeneity while preserving structure from cross-neuron and temporal correlations.
 The four-point functions for the unnormalized and normalized variables are found to be
@@
 respectively, where $C^{\Phi}_{12}=q_{2}^{2}C^{\phi}_{12}$ and we have updated the definition of the effective coupling strength to $g_{\text{eff}}=\sqrt{\alpha r_{2}q_{2}}$.
-Eq.
-46b demonstrates that reductions in effective rank, described by $\alpha\text{PR}^{D}$, and reductions in the heterogeneity of gains, described by $\text{PR}^{G}$, have symmetric effects on the dimension of normalized activity, with their combined effect captured by the harmonic mean $\left(\frac{1}{\alpha\text{PR}^{D}}+\frac{1}{\text{PR}^{G}}\right)^{-1}$  [^6] This symmetry arises because the dynamics of the network are determined by an effective coupling matrix $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$, where $\bm{D}=\text{diag}(D_{1},\ldots,D_{M})$ and $\bm{G}=\text{diag}(G_{1},\ldots,G_{N})$.
+Equation (47b) demonstrates that reductions in effective rank, described by $\alpha\text{PR}^{D}$, and reductions in the heterogeneity of gains, described by $\text{PR}^{G}$, have symmetric effects on the dimension of normalized activity, with their combined effect captured by the harmonic mean $\left(\frac{1}{\alpha\text{PR}^{D}}+\frac{1}{\text{PR}^{G}}\right)^{-1}$  [^6] This symmetry arises because the dynamics of the network are determined by an effective coupling matrix $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$, where $\bm{D}=\text{diag}(D_{1},\ldots,D_{M})$ and $\bm{G}=\text{diag}(G_{1},\ldots,G_{N})$.
 Swapping the distributions of $D_{a}$ and $G_{i}$ does not change the spectrum when $\bm{L}$ and $\bm{R}$ are i.i.d., since $\bm{L}\bm{D}\bm{R}^{T}\bm{G}$ and $\bm{R}^{T}\bm{G}\bm{L}\bm{D}$ have the same nonzero eigenvalues..
@@
 Case (3), whose four-point function is given by Eq.
-164, considers variables with the same distribution of firing rates as the unnormalized activations, but where this heterogeneity is unrelated to the recurrent dynamics.
-To validate our theoretical predictions (Eq.
-46a, Eq.
-46b, and Eq.
-164), we simulated networks with component strengths and gains given by $D_{a}\propto\exp(-\beta_{D}a/M)$ and $G_{i}\propto\exp(-\beta_{G}i/N)$ (Fig.
+(166), considers variables with the same distribution of firing rates as the unnormalized activations, but where this heterogeneity is unrelated to the recurrent dynamics.
+To validate our theoretical predictions [Eqs.
+(47a), (47b), and (166)], we simulated networks with component strengths and gains given by $D_{a}\propto\exp(-\beta_{D}a/M)$ and $G_{i}\propto\exp(-\beta_{G}i/N)$ (Fig.
 5).
 The dimension of activity is determined by the coupling strength $g_{\text{eff}}$, the effective rank $\alpha\text{PR}^{D}$, and the participation ratio of the gain distribution $\text{PR}^{G}$.
-Increases in each of these parameters lead to higher activity dimension for all three sets of activations.
+Increases in each of these parameters lead to a higher activity dimension for all three sets of activations.
 Normalized activations exhibit the highest dimension.
@@
 Our analysis implies that such normalization may not eliminate the effects of single-neuron heterogeneity on collective-activity properties.
-The persistence of these effects through normalization provides an experimental signature for distinguishing between heterogeneity that affects recurrent dynamics and heterogeneity that affects only the observed firing rates (e.g., in a circuit that is not recurrent at all).
+The persistence of these effects through normalization provides an experimental signature for distinguishing between heterogeneity that affects recurrent dynamics and heterogeneity that affects only the observed firing rates, for example, in a circuit that is not recurrent.
 VII Mode overlaps
 Thus far, we have assumed i.i.d.
-modes, with the factorization $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$ implying that left–right overlaps $\bm{\ell}_{a}^{T}\bm{r}_{a}$ are random and $\mathcal{O}(1/\sqrt{N})$.
+modes, with the factorization $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$ implying that left-right overlaps $\bm{\ell}_{a}^{T}\bm{r}_{a}$ are random and $\mathcal{O}(1/\sqrt{N})$.
 The same holds for $\bm{\ell}_{a}^{T}\bm{r}_{b}$ with $a\neq b$ due to the more fundamental assumption that $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ factorizes across $a$.
-However, computations in biological circuits require structured interactions between modes, suggesting stronger, non-random overlaps between left and right modes.
-To test this, we return to the Drosophila connectome and compute the overlap matrix between left and right singular vectors, $O_{ab}=\bm{u}_{a}^{T}\bm{v}_{b}$ (Fig.
-6a).
+However, computations in biological circuits require structured interactions between modes, suggesting stronger, nonrandom overlaps between left and right modes.
+To test this, we return to the Drosophila connectome and compute the overlap matrix between left and right singular vectors, $O_{ab}=\bm{u}_{a}^{T}\bm{v}_{b}$ [Fig.
+6(a)].
 Contrary to the small, unstructured overlaps expected under the i.i.d.
-assumption, the connectome exhibits large, structured overlaps, particularly along the diagonal (Fig.
-6b).
+assumption, the connectome exhibits large, structured overlaps, particularly along the diagonal [Fig.
+6(b)].
 Figure 6: Right-left mode overlaps and eigenvalue spectra of the fly-brain connectome.
 (a) Overlap matrix between right and left singular vectors of the connectome.
-First 1000 modes are shown.
+First 500 modes are shown.
 (b) First 3000 diagonal elements of the overlap matrix from (a).
@@
 Top: actual hemibrain spectrum, exhibiting large outliers along the real axis.
-Middle: Eigenvalues from a realization of the random-mode model, using singular values from the connectome as component strengths $D_{a}$, resulting in a circularly symmetric spectrum.
+Middle: eigenvalues from a realization of the random-mode model (RMM), using singular values from the connectome as component strengths $D_{a}$, resulting in a circularly symmetric spectrum.
 Bottom: random-mode model realization with right-left mode overlaps incorporated, where overlaps are given by the inner products of left and right singular vectors from the fly connectome SVD.
@@
 Motivated by the presence of these large diagonal overlaps, we extend the random-mode model to include correlations between corresponding left and right modes.
-Specifically, we take $P(\ell_{a},r_{a})$ to be a zero-mean bivariate Gaussian with marginal variance $1/N$ and covariance $\rho_{a}/N$, where $\rho_{a}$ is the correlation between the $a$-th mode pair.
+Specifically, we take $P(\ell_{a},r_{a})$ to be a zero-mean bivariate Gaussian with marginal variance $1/N$ and covariance $\rho_{a}/N$, where $\rho_{a}$ is the correlation between the $a$th mode pair.
 Both $D_{a}$ and $\rho_{a}$ are treated deterministically, and we assume that $M^{-1}\sum_{a}f(D_{a},\rho_{a})$ converges to a limiting value $\left\langle f(D,\rho)\right\rangle_{D,\rho}$ as $N\to\infty$.
-In this setting, $\bm{\ell}_{a}^{T}\bm{r}_{a}=\rho_{a}$ for large $N$, while $\bm{\ell}_{a}^{T}\bm{r}_{b}$ for $a\neq b$ remains random and $\mathcal{O}(1/\sqrt{N})$.
-This additional structure produces eigenvalue spectra resembling that of the connectome, including large real outliers (Fig.
-6c) [59].
-This purely diagonal overlap structure, however, can only generate real eigenvalue outliers.
-The presence of complex outliers in the connectome, together with the full overlap matrix itself (Fig.
-6a), implies the presence of richer correlations.
+In this setting, $\bm{\ell}_{a}^{T}\bm{r}_{a}=\rho_{a}$, with negligible $\mathcal{O}(1/\sqrt{N})$ fluctuations around this structured overlap, while $\bm{\ell}_{a}^{T}\bm{r}_{b}$ for $a\neq b$ remains unstructured and $\mathcal{O}(1/\sqrt{N})$.
+This additional structure produces eigenvalue spectra resembling that of the connectome, including large real outliers [Fig.
+6(c)] [68].
+This purely diagonal overlap structure, however, can generate only real eigenvalue outliers.
+The presence of complex outliers in the connectome, together with the full overlap matrix itself [Fig.
+6(a)], implies the presence of richer correlations.
 Capturing such features would require relaxing the full factorization of $P(\{\ell_{a},r_{a}\})$ across $a$ while retaining $\mathcal{O}(M)$ parameter scaling—for example, via blockwise factorization or Markovian correlations across $a$.
@@
 We also note that by setting $\rho_{a}=1$ for all $a$, the model becomes a Hopfield network storing an extensive set of Gaussian patterns.
-In such networks, each mode can be “condensed” ($\mathcal{O}(1)$ overlap with activity) or “uncondensed” ($\mathcal{O}(1/\sqrt{N})$ overlap with activity).
+In such networks, each mode can be “condensed” $\mathcal{O}(1)$ overlap with activity] or “uncondensed” $\mathcal{O}(1/\sqrt{N})$ overlap with activity].
 Our analysis assumes no condensed modes, but sufficiently strong positive overlaps break this assumption.
@@
 Allowing for condensed patterns within our analysis would require specifying this density.
-See [60] for a DMFT analysis of generalized Hopfield dynamics in which these ideas are implemented.
-In Appendix K, we compute the two- and four-point functions. As in previous cases, the dynamics reduce to single-site processes:
+See Ref.
+[69] for a DMFT analysis of generalized Hopfield dynamics in which these ideas are implemented.
+In Appendix K, we compute the two- and four-point functions. As in previous cases, the dynamics reduce to single-site processes,
 $T[x](t)$ $=\eta^{x}(t)+\int^{t}dt^{\prime}R_{\star}(t,t^{\prime})\phi(t^{\prime}),$ (48a)
@@
 Unlike the “i.i.d. modes” case, the single-site picture is not equivalent to that of an i.i.d. network for some $g_{\text{eff}}$.
-To gain intuition for the self-coupling kernel $R_{\star}(\omega)$, we expand in powers of $\rho$. The coefficients encode connectivity motifs of increasing orders:
+To better understand the self-coupling kernel $R_{\star}(\omega)$, we expand in powers of $\rho$. The coefficients encode connectivity motifs of increasing orders,
 $\begin{split}&R_{\star}(\omega)=\left\langle J_{ii}\right\rangle_{\star}+\left\langle NJ_{ij}J_{ji}\right\rangle_{\star}\>S^{\phi}_{\star}(\omega)+\cdots,\text{ where}\\ &\left\langle J_{ii}\right\rangle_{\star}=\alpha\left\langle D\rho\right\rangle_{D,\rho},\>\>\left\langle NJ_{ij}J_{ji}\right\rangle_{\star}=\alpha\left\langle(D\rho)^{2}\right\rangle_{D,\rho}.\end{split}$ (51)
@@
 For overlaps, we use a sigmoidal spectrum for $\rho_{a}$ qualitatively inspired by Fig.
-6b, with multiplicative prefactor $\gamma_{\rho}$ and other parameters chosen such that $\left\langle D\rho\right\rangle_{D,\rho}=0$ (Appendix K.3).
+6(b), with multiplicative prefactor $\gamma_{\rho}$ and other parameters chosen such that $\left\langle D\rho\right\rangle_{D,\rho}=0$ (Appendix K.3).
 Across values of $g_{\text{eff}}$ and $\text{PR}^{D}$, sufficiently strong overlaps, of either sign, reduce the dimension of activity (Fig.
@@
 This effect weakens as $g_{\text{eff}}$ increases.
+In summary, our analysis reveals how connectivity structure shapes collective activity and its dimensionality.
+When mode overlaps are unstructured, connectivity structure reduces to two parameters: the effective rank $\alpha\text{PR}^{D}$ and effective coupling strength $g_{\text{eff}}$.
+Structured left-right mode overlaps break this simplification: Collective dynamics then depend on the full joint distribution of component strengths and overlaps, providing an additional mechanism for controlling activity dimensionality.
 Figure 7: Effect of right-left mode correlations on dimension of activity in the random-mode model.
 Dimension of activity $\text{PR}^{\phi}$ versus correlation parameter $\gamma_{\rho}$ in the $(D,\rho)$ joint distribution, for various coupling strengths $g_{\text{eff}}$ and effective ranks $\alpha\text{PR}^{D}$.
-Thin dots: individual simulations; thick dots: means over 10 simulations; lines: theoretical predictions.
+Thin dots, individual simulations; thick dots, means over ten simulations; lines, theoretical predictions.
 All simulations use $N=5000$ neurons.
 VIII Discussion
-Motifs vs.
-global structure.
+VIII.1 Motifs versus global structure
 Neuronal connectivity is often described using statistics of motifs: local patterns like chains or cycles among small groups of neurons [73, 22, 74, 75].
@@
 Advances in whole-brain connectome reconstruction now allow access to global connectivity, including spectral features.
-Our parameterization of coupling matrices represents this global structure, with control over both the spectrum and mode overlaps.
+Our parametrization of coupling matrices represents this global structure, with control over both the spectrum and mode overlaps.
 These spectral features determine how activity modes interact dynamically and thus may more directly dictate neural computations compared to motifs.
 That said, in some studies the primary goal is to capture motifs or graph-theoretic features (e.g., degree distributions) over which the random-mode model does not provide control (although some of these features have spectral signatures that could be incorporated).
-When such features are the focus, other connectivity models are better suited, for example maximum-entropy random graph models constrained by fixed structural features (e.g., the configuration model for matching degree distributions).
-Linking connectivity and activity datasets.
+When such features are the focus, other connectivity models are better suited—for instance, maximum-entropy random graph models constrained by fixed structural features (e.g., the configuration model for matching degree distributions).
+VIII.1.1 Linking connectivity and activity datasets.
 The random-mode model, together with its dynamical theory, could aid in the interpretation of large-scale neural activity recordings in the context of anatomical connectivity, such as in Drosophila, where both the connectome and whole-brain activity data are available [76].
@@
 Future work could progressively incorporate richer connectome constraints into the model to identify which features most strongly shape collective activity.
-One possible sequence would be to add the spectrum of component strengths, then diagonal mode overlaps, and finally off-diagonal overlaps, assessing at each step how much the correspondence between modeled and observed activity improves.
+One possible sequence would be to add the spectrum of component strengths, then diagonal mode overlaps, and finally off diagonal overlaps, assessing at each step how much the correspondence between modeled and observed activity improves.
 This is similar to approaches in statistical physics that aim to recover higher-order correlations by constraining lower-order statistics [77, 78, 38].
-Learning.
-Learning alters collective dynamics by modifying connectivity—the core principle of machine learning.
+VIII.2 Learning
+Learning alters collective dynamics by modifying connectivity, the core principle of machine learning.
 Our analysis focuses on structured but untrained networks that produce chaotic activity.
 Nevertheless, the extensive low-rank connectivity structures we study provide a starting point for understanding trained networks.
-Large networks trained on simple tasks often show intensive low-rank modifications to their initial couplings, with a finite number of outlier eigenvalues emerging from the bulk [23, 34].
+Large networks trained on simple tasks often show intensive low-rank modifications to their initial couplings, resulting in a finite number of task-relevant eigenvalues [27, 43].
 Existing theoretical tools can already describe such systems by taking the trained weights, performing a low-rank decomposition, modeling mode overlaps statistically, and studying the resulting low-rank DMFT [45].
@@
 A more ambitious direction is to develop a DMFT for networks whose weights are learned from data, rather than first training networks and then retrofitting a statistical model.
-Such a theory would describe not only how data shapes weights, as addressed by classical works [69], but also how data interacts with recurrent network dynamics.
-The path-integral formalism is well-suited to this, as it allows derivation of DMFT equations for coupling matrices drawn from a Gibbs ensemble with an energy function given by a task loss.
-While this approach has been applied to low-dimensional tasks, leading to intensive-rank weights [70], applying it to high-dimensional tasks could yield a theory of how such tasks generate extensive-rank weights that in turn produce extensive-dimensional activity whose single-neuron and collective properties the theory would describe.
+Such a theory would describe not only how data shape the weights, as addressed by classical works [79], but also how data interact with recurrent network dynamics.
+The path-integral formalism is well suited to this, as it allows derivation of DMFT equations for coupling matrices drawn from a Gibbs ensemble with an energy function given by a task loss.
+While this approach has been applied to low dimensional tasks [80], applying it to high-dimensional tasks could yield a theory of how such tasks generate extensive-rank weights that, in turn, produce extensive-dimensional activity whose single-neuron and collective properties the theory would describe.
 Another route to extensive-rank structure is via networks that perform many tasks, which could be modeled, for example, by representing mode overlaps through blockwise factorization of $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ across $a$.
-This “multi-task” scenario is increasingly relevant in both neuroscience and machine learning [71, 72].
-Even in standard RNN training, initializing couplings with low-rank structure could improve procedures such as FORCE learning, which rely on suppressing chaos [12, 13].
+This multitask scenario is increasingly relevant in both neuroscience and machine learning [81, 82].
+Even in standard recurrent neural network training, initializing couplings with low-rank structure could improve procedures such as FORCE learning, which rely on suppressing chaos [12, 13].
 Lower-dimensional attractors may be easier to control, potentially leading to faster convergence.
-Modeling functional consequences of single-neuron heterogeneity.
+VIII.3 Modeling functional consequences of single-neuron heterogeneity
 A central question in neuroscience is how neuronal diversity shapes collective dynamics and computation [83].
-Appendix J presents a general framework that assigns each neuron a parameter vector $\bm{\theta}$ and enables study of how different forms of heterogeneity affect properties such as dimensionality.
-Sec.
-VI focuses on gain heterogeneity to model firing-rate diversity, but this framework could also incorporate more intricate cell-specific features, such as electrophysiological characteristics [74], linking single-neuron heterogeneity to functional consequences for population activity.
-Transient dynamics and excitation–inhibition.
+Appendix J presents a general framework that assigns each neuron a parameter vector $\bm{\theta}$ and enables the study of how different forms of heterogeneity affect properties such as dimensionality.
+Section VI focuses on gain heterogeneity to model firing-rate diversity, but this framework could also incorporate more intricate cell-specific features, such as electrophysiological characteristics [84], linking single-neuron heterogeneity to functional consequences for population activity.
+VIII.4 Transient dynamics and excitation-inhibition
 Our analysis has focused on stationary states with translation-invariant temporal correlations.
 Many neural computations, however, rely on transient dynamics that break stationarity and depend on absolute time.
-Extending the DMFT accordingly is numerically tractable for two-point functions but would become unwieldily for four-point functions, $\Psi^{a}_{\star}(t_{1},t_{2},t_{3},t_{4})$.
+Extending the DMFT accordingly is numerically tractable for two-point functions but would become unwieldy for four-point functions, $\Psi^{a}_{\star}(t_{1},t_{2},t_{3},t_{4})$.
 Representing and manipulating such a large object may require approximations or low-rank decomposition.
-One notable transient phenomenon is transient amplification, linked to non-normal coupling matrices that naturally arise from excitatory–inhibitory segregation (Dale’s law) [75, 76].
-Although the random-mode model does not enforce Dale’s law, the resulting matrices are nevertheless non-normal and may exhibit transient amplification; such effects may be amplified through structured off-diagonal overlaps between left and right modes.
+One notable transient phenomenon is transient amplification, linked to non-normal coupling matrices that naturally arise from excitatory-inhibitory segregation (Dale’s law) [85, 86].
+Although the random-mode model does not enforce Dale’s law, the resulting matrices are nevertheless non-normal and may exhibit transient amplification; such effects may be amplified through structured off diagonal overlaps between left and right modes.
 Non-normality also decouples the highest-variance dimensions from the slowest modes.
-Relatedly, extending the random-mode model to respect Dale’s law and incorporate a non-saturating $\phi(x)$—achieving stability via excitation–inhibition balance rather than saturation [77, 78]—could enable more direct comparisons to both connectivity and activity data.
-Random basis property.
+Relatedly, extending the random-mode model to respect Dale’s law and incorporating a nonsaturating $\phi(x)$—achieving stability via excitation-inhibition balance rather than saturation [87, 88]—could enable more direct comparisons to both connectivity and activity data.
+VIII.5 Random basis property
 The random-mode model assumes that the $2M$ mode components associated with each neuron are sampled i.i.d.
-for each neuron (Eq.
-3).
+for each neuron [Eq.
+(3)].
 This leads to neuronal permutation symmetry of the distribution over $\bm{J}$.
@@
 This would yield a regionally or spatially embedded random-mode model, apt for connectome studies.
-Because we assume zero-mean mode components (Eq.
-4), $\left\langle J_{ij}\right\rangle=0$ for $i\neq j$.
+Because we assume zero-mean mode components [Eq.
+(4)], $\left\langle J_{ij}\right\rangle=0$ for $i\neq j$.
 Thus, the low-rank structure is not inherited from the expected weight matrix.
 This differs from the mechanism described by Thibeault et al.
-[19], who analyze cases where low-rank structure arises because $\left\langle\bm{J}\right\rangle$ is low-rank, with the random part treated as noise.
-By Weyl’s inequality, the singular values of $\bm{J}$ differ from those of $\left\langle\bm{J}\right\rangle$ by at most the spectral norm of the noise.
+[23], who analyze cases where low-rank structure arises because $\left\langle\bm{J}\right\rangle$ is low rank, with the random part treated as noise.
+By Weyl’s inequality, the singular values of $\bm{J}$ differ from those of $\left\langle\bm{J}\right\rangle$ by, at most, the spectral norm of the noise.
 This mechanism is fundamentally different from the random-mode model’s zero-mean, randomly embedded structure.
 Which of these mechanisms is more prevalent in real-world networks is an open question.
-Alternative generative models.
+VIII.6 Alternative generative models
 Tiberi et al.
-[79] proposed a generative model for coupling matrices that parameterizes the eigendecomposition of $\bm{J}$ rather than its SVD.
+[89] proposed a generative model for coupling matrices that parametrizes the eigendecomposition of $\bm{J}$ rather than its SVD.
 Their formulation does not capture relationships between eigenvalues and eigenvectors.
@@
 Whereas Tiberi et al.
-[79] studied noise-driven linear networks within a random matrix theory framework, our DMFT formulation applies to nonlinear dynamics.
-Connectivity parameterizations resembling the random-mode model have been used to study the correspondence between spiking and rate-based recurrent neural networks [80] and to expedite training spiking networks [81].
-Conclusion.
+[89] studied noise-driven linear networks within a random-matrix-theory framework, our DMFT formulation applies to nonlinear dynamics.
+Connectivity parametrizations resembling the random-mode model have been used to study the correspondence between spiking and rate-based recurrent neural networks [90] and to expedite training spiking networks [91].
+IX Conclusion
 The random-mode model and associated analytical techniques provide tools for exploring the relationship between connectivity and collective dynamics.
-Our parameterization of connectivity is a middle ground between intensive-rank models and i.i.d.
+Our parametrization of connectivity is a middle ground between intensive-rank models and i.i.d.
 random networks.
 Building on this framework stands to further bridge neuronal connectivity, activity, and function.
-Acknowledgements
+Acknowledgments
 We thank L.
@@
 Abbott for feedback.
-A.v.M.
-was supported by the Swartz Foundation; D.G.C, O.M., and A.L.-K.
-by the Kavli Foundation and Gatsby Charitable Foundation (grant no.
-GAT3708); and D.G.C.
-by the National Institutes of Health (grant no.
+A.
+v.
+M.
+was supported by the Swartz Foundation and the Kempner Institute for the Study of Natural and Artificial Intelligence.
+D.
+G.
+C, O.
+M., and A.
+L.-K.
+were supported by the Kavli Foundation and Gatsby Charitable Foundation (Grant No.
+GAT3708).
+D.
+G.
+C.
+was also supported by the National Institutes of Health (Grant No.
 T32NS064929).
+A.
+L.-K.
+was also supported by the National Science Foundation (Grant No.
+2443158) and the National Institutes of Health (Grant No.
+RF1DA060772).
 Appendix
 Appendix A Glossary of terms
-Number of neurons ($N$)
-Total number of neurons in the network.
-Number of modes ($M$)
-Number of components in the random-mode model.
-(Effective) coupling strength ($g$, $g_{\text{eff}}$)
-Standard deviation of the couplings times $\sqrt{N}$ in the i.i.d. connectivity model, random-mode model (Eq. 37), or variants thereof (e.g., see just below Eq. 46a and Eq. 46b).
-Random-mode model
-Network connectivity model: $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{\ell}_{a}\bm{r}_{a}^{T}$ (Eq. 2).
-Connectivity component ($\bm{\ell}_{a}\bm{r}_{a}^{T}$)
-Outer product of output and right modes.
-Right mode ($\bm{r}_{a}$)
-Neuronal pattern in the $a$-th component onto which activity patterns are projected.
-Left mode ($\bm{\ell}_{a}$)
-Neuronal pattern in the $a$-th component along which the projected activity pattern is expanded.
-Component strength ($D_{a}$)
-Scaling factor for the $a$-th component.
-Ratio of modes to neurons ($\alpha$)
-Given by $\alpha=M/N$ (Eq. 7).
-Participation ratio of component strengths ($\text{PR}^{D}$)
-Given by $\text{PR}^{D}=r_{2}^{2}/r_{4}$ (Eq. 11) where $r_{n}=\langle D^{n}\rangle_{D}$ (Eq. 10).
-Effective rank ($\alpha\text{PR}^{D}$)
-Connectivity dimensionality measure (Eq. 12).
-Two-point function
-Correlation function given by $C^{\phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\langle\phi_{i}(t_{1})\phi_{i}(t_{2})\rangle_{t}$ (Eq. 17).
-Four-point function
-Higher-order correlation function given by $\Psi^{\phi}(\tau_{1},\tau_{2})=\frac{1}{N}\sum_{i,j=1}^{N}C^{\phi}_{ij}(\tau_{1})C^{\phi}_{ij}(\tau_{2})$ (Eq.
-24).
-See also time-by-time definition (Eq.
-25).
+Number of neurons ($N$): Total number of neurons in the network.
+Number of modes ($M$): Number of components in the random-mode model.
+(Effective) coupling strength ($g$, $g_{\text{eff}}$): Standard deviation of the couplings times $\sqrt{N}$ in the i.i.d.
+connectivity model, random-mode model [Eq.
+(38)], or variants thereof [e.g., see just below Eqs.
+(47a) and (47b)].
+Random-mode model: Network connectivity model: $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{\ell}_{a}\bm{r}_{a}^{T}$ [Eq. (2)].
+Connectivity component ($\bm{\ell}_{a}\bm{r}_{a}^{T}$): Outer product of output and right modes.
+Right mode ($\bm{r}_{a}$): Neuronal pattern in the $a$th component onto which activity patterns are projected.
+Left mode ($\bm{\ell}_{a}$): Neuronal pattern in the $a$th component along which the projected activity pattern is expanded.
+Component strength ($D_{a}$): Scaling factor for the $a$th component.
+Ratio of modes to neurons ($\alpha$): Given by $\alpha=M/N$ [Eq. (7)].
+Participation ratio of component strengths ($\text{PR}^{D}$): Given by $\text{PR}^{D}=r_{2}^{2}/r_{4}$ [Eq. (11)] where $r_{n}=\langle D^{n}\rangle_{D}$ [Eq. (10)].
+Effective rank ($\alpha\text{PR}^{D}$): Connectivity dimensionality measure [Eq. (12)].
+Two-point function: Correlation function given by $C^{\phi}(t_{1},t_{2})=(1/N)\sum_{i=1}^{N}\phi_{i}(t_{1})\phi_{i}(t_{2})$ [Eq. (17)].
+Four-point function: Higher-order correlation function given by $\Psi^{\phi}(\tau_{1},\tau_{2})=(1/N)\sum_{i,j=1}^{N}C^{\phi}_{ij}(\tau_{1})C^{\phi}_{ij}(\tau_{2})$ [Eq.
+(24)].
+See also time-by-time definition [Eq.
+(25)].
 Appendix B Preprocessing of fly hemibrain connectome
@@
 Neurotransmitter Synaptic sign
-GABA -1 (Inhibitory)
-Acetylcholine +1 (Excitatory)
-Glutamate -1 (Inhibitory)
+GABA 1-1 (Inhibitory)
+Acetylcholine +1+1 (Excitatory)
+Glutamate 1-1 (Inhibitory)
 Serotonin Ignored
@@
 Neither Ignored
-Table 1: Mapping of neurotransmitters to synaptic signs
+Table 1: Mapping of neurotransmitters to synaptic signs.
 Synaptic counts do not directly correspond to effective synaptic strengths, so it is reasonable to choose “fudge factors” to relate the two.
@@
 Specifically, the preprocessing is designed to preserve:
-  • the sparsity pattern (which connections are zero vs. nonzero);
-  • the sign of each connection (excitatory vs. inhibitory);
+  1. (i)
+    the sparsity pattern (for which the connections are zero versus nonzero);
+  2. (ii)
+    the sign of each connection (excitatory versus inhibitory);
+  3. (iii)
     approximately, the relative magnitudes of same-sign inputs to each neuron; and
+  4. (iv)
     approximately, the relative magnitudes of same-sign outputs from each neuron.
 The last two properties are preserved only approximately because they are enforced simultaneously by the iterative normalization procedure described below.
-We normalize the coupling matrix using an iterative scaling method analogous to the Sinkhorn–Knopp algorithm.
-First, for each row, we rescale the positive elements by a row-specific factor so that their L2-norm equals $1/\sqrt{2}$.
-We then rescale the negative elements, using a separate factor, so that their L2-norm also equals $1/\sqrt{2}$.
-This yields rows with a total L2-norm of one and balanced excitatory and inhibitory contributions.
-A small number of neurons with only excitatory or only inhibitory inputs retain an L2-norm of $1/\sqrt{2}$.
+We normalize the coupling matrix using an iterative scaling method analogous to the Sinkhorn-Knopp algorithm.
+First, for each row, we rescale the positive elements by a row-specific factor so that their $L2$ norm equals $1/\sqrt{2}$.
+We then rescale the negative elements, using a separate factor, so that their $L2$ norm also equals $1/\sqrt{2}$.
+This yields rows with a total $L2$ norm of one and balanced excitatory and inhibitory contributions.
+A small number of neurons with only excitatory or only inhibitory inputs retain an $L2$ norm of $1/\sqrt{2}$.
 Next, we apply the same rescaling procedure to each column, which disrupts the previous row normalization.
@@
 where $\bm{L}$ and $\bm{R}$ are independent $N\times M$ random matrices with $\alpha=M/N$ and entries of variance $1/N$. Equivalently, we can compute the eigenvalues of
-$\bm{J}^{T}\bm{J}=\bm{R}\bm{L}^{T}\bm{L}\bm{R}^{T},$ (53)
+$\bm{J}^{T}\bm{J}=\bm{R}\bm{L}^{T}\bm{L}\bm{R}^{T}$ (54)
 and then take the square root to obtain the singular values. These eigenvalues are the same as those of
 $\bm{L}^{T}\bm{L}\bm{R}^{T}\bm{R},$ (55)
-which is the product of two independent Wishart matrices:
+which is the product of two independent Wishart matrices,
 $\bm{L}^{T}\bm{L}\bm{R}^{T}\bm{R}=\bm{W}_{L}\bm{W}_{R}.$ (56)
@@
 $\rho(\lambda)$ $=-\frac{1}{\pi}\lim_{\epsilon\to 0^{+}}\mathrm{Im}G(\lambda+i\epsilon).$ (58)
-In a 2D electrostatics analogy, the real and imaginary parts of $G(z)$ correspond to the Cartesian components of the electric field generated by a line charge density $\rho(\lambda)$ on the real axis.
+In a two-dimensional electrostatics analogy, the real and imaginary parts of $G(z)$ correspond to the Cartesian components of the electric field generated by a line charge density $\rho(\lambda)$ on the real axis.
 The inversion formula is then analogous to Gauss’s law, relating the discontinuity of the normal field component to the local charge density.
-The moment generating function and S-transform are defined by
+The moment-generating function and $S$ transform are defined by
 $M(z)$ $=zG(z)-1,$ (59)
 $S(z)$ $=\frac{1+z}{zM^{-1}(z)},$ (60)
-where $M^{-1}(z)$ is the functional inverse of $M(z)$. For a Marchenko–Pastur distributed Wishart matrix $\bm{W}$ with aspect ratio $\alpha$,
+where $M^{-1}(z)$ is the functional inverse of $M(z)$. For a Marchenko-Pastur distributed Wishart matrix $\bm{W}$ with aspect ratio $\alpha$,
 $S(z)=\frac{1}{1+\alpha z}.$ (61)
-A key fact in free probability is that the S-transforms of free variables multiply. Thus, for $\bm{W}_{1}\bm{W}_{2}$, the S-transform is
+A key fact in free probability is that the $S$ transforms of free variables multiply. Thus, for $\bm{W}_{1}\bm{W}_{2}$, the $S$ transform is
 $S(z)=\frac{1}{(1+\alpha z)^{2}}.$ (62)
-From the definition of the S-transform, the inverse moment generating function is
+From the definition of the $S$ transform, the inverse moment-generating function is
 $M^{-1}(z)=\frac{1+z}{zS(z)}=\frac{(1+z)(1+\alpha z)^{2}}{z}.$ (63)
 Therefore, $M(z)$ satisfies
-$z=\frac{(1+M(z))(1+\alpha M(z))^{2}}{M(z)}.$ (63)
-Using M(z)=zG(z)1M(z)=zG(z)-1, we obtain a cubic equation for $G(z)$:
+$z=\frac{[1+M(z)][1+\alpha M(z)]^{2}}{M(z)}.$ (64)
+Using M(z)=zG(z)1M(z)=zG(z)-1, we obtain a cubic equation for $G(z)$,
 $\alpha^{2}z^{3}G(z)^{3}+2z^{2}(\alpha^{2}-\alpha)G(z)^{2}+z\big(z+2\alpha-\alpha^{2}-1\big)G(z)-z=0.$ (65)
 In principle, this can be solved using the cubic formula, but the explicit form is cumbersome.
-For the support boundaries of $\rho(\lambda)$, we only need the points where $\mathrm{Im}G(z)$ changes from zero to nonzero.
+For the support boundaries of $\rho(\lambda)$, we need only the points where $\mathrm{Im}G(z)$ changes from zero to nonzero.
 This occurs when the discriminant of the cubic vanishes.
@@
 $\lambda_{\pm}=1+\frac{5\alpha}{2}-\frac{\alpha^{2}}{8}\pm\left(1+\frac{\alpha}{8}\right)^{3/2}\sqrt{8\alpha}.$ (68)
-The singular values correspond to the square roots of the eigenvalues, given by Eq. 14.
+The singular values correspond to the square roots of the eigenvalues, given by Eq. (14). We confirm this formula numerically in Fig. 8.
 Figure 8: Distribution of singular values in the random-mode model.
 Histograms show singular values of $\bm{L}\bm{D}\bm{R}^{T}$ with $D_{a}=1$ for various $\alpha$.
-For $\alpha\ll 1$, the singular values concentrate at one.
-Vertical lines: prediction of Eq.
-14.
+For $\alpha\ll 1$, the singular values concentrate at 1.
+Vertical lines, prediction of Eq.
+(14).
 $N=5000$.
@@
 Alternative measures of effective dimensionality, which could be applied to both connectivity and activity in place of the participation ratio, have been proposed in the literature.
-Roy and Vetterli [88] define effective rank as the exponential of the entropy of a categorical probability distribution obtained by normalizing the eigenvalues values to sum to one, i.e., $p_{i}=\lambda_{i}/\sum_{j=1}^{N}\lambda_{j}$.
-Written in terms of this distribution the participation ratio is $\text{PR}=\left(\sum_{i=1}^{N}p_{i}^{2}\right)^{-1}$, while the exponentiated entropy is $\exp(H)=\exp\left(-\sum_{i=1}^{N}p_{i}\log p_{i}\right)$.
+Roy and Vetterli [95] define effective rank as the exponential of the entropy of a categorical probability distribution obtained by normalizing the eigenvalues to sum to 1, i.e., $p_{i}=\lambda_{i}/\sum_{j=1}^{N}\lambda_{j}$.
+Written in terms of this distribution, the participation ratio is $\text{PR}=\left(\sum_{i=1}^{N}p_{i}^{2}\right)^{-1}$, while the exponentiated entropy is $\exp(H)=\exp\left(-\sum_{i=1}^{N}p_{i}\log p_{i}\right)$.
 By Jensen’s inequality, $\text{PR}\leq\exp(H)$.
 Exact equality holds in certain cases, including when exactly $K\leq N$ eigenvalues are equal and the rest are zero, yielding $\text{PR}=\exp(H)=K$.
-Furthermore, for smoothly decaying spectra as discussed above (for which $p_{i}=f(i/Nw)/\sum_{j=1}^{N}f(j/Nw)$, with $w\ll 1$), both measures have $\text{PR}=c_{1}\>Nw$ and $\exp(H)=c_{2}\>Nw$, where $c_{1}$ and $c_{2}$ are $N$ and $w$-independent constants depending on the specific form of $f$.
+Furthermore, for smoothly decaying spectra as discussed above [for which $p_{i}=f(i/Nw)/\sum_{j=1}^{N}f(j/Nw)$, with $w\ll 1$], both measures have $\text{PR}=c_{1}\>Nw$ and $\exp(H)=c_{2}\>Nw$, where $c_{1}$ and $c_{2}$ are $N$- and $w$-independent constants depending on the specific form of $f$.
 Thus, while the measures differ in detail, they provide similar characterizations of effective dimensionality.
-This entropy-based definition has been used in neuroscience applications, for example by [89] to analyze transient amplification in neural networks.
+This entropy-based definition has been used in neuroscience applications, for example, in Ref.
+[96] to analyze transient amplification in neural networks.
 Both measures capture the number of dominant modes in the system.
 Thibeault et al.
-[19] call (a metric closely related to) the the participation ratio-based metric the “srank” (the difference being whether the maximum eigenvalue or sum over all eigenvalues is used as a normalizing factor) and call the entropy-like metric the “erank.”
+(Ref.
+[23]) call (a metric closely related to) the the participation ratio-based metric the “srank” (the difference being whether the maximum eigenvalue or sum over all eigenvalues is used as a normalizing factor) and call the entropylike metric the “erank.”
 Appendix E Review: Calculating two- and four-point functions for i.i.d. couplings
-We now review how to compute these two- and four-point functions in a classic network model with i.i.d. couplings $\bm{J}$. The first- and second-order coupling statistics are:
+We now review how to compute these two- and four-point functions in a classic network model with i.i.d. couplings $\bm{J}$. The first- and second-order coupling statistics are
 $\left\langle J_{ij}\right\rangle=0,\hskip 15.00002pt\left\langle J_{ij}^{2}\right\rangle=\frac{g^{2}}{N}.$ (69)
@@
 For the canonical dynamics $T[x](t)=(1+\partial_{t})x(t)$, the network is quiescent for $g<1$ and chaotic for $g>1$, with a sharp phase transition as $N\to\infty$.
-In this paper we assume that the network is non-quiescent.
-The two-point function $C^{a}_{\star}(t_{1},t_{2})$ can be computed through a single-site picture that describes the dynamics of a typical neuron embedded in the rest of the network, with pre-activation $x(t)$ and activation $\phi(t)$.
+In this paper we assume that the network is nonquiescent.
+The two-point function $C^{a}_{\star}(t_{1},t_{2})$ can be computed through a single-site picture that describes the dynamics of a typical neuron embedded in the rest of the network, with preactivation $x(t)$ and activation $\phi(t)$.
 The single-site dynamics are given by
@@
 Once $C^{\phi}_{\star}(t_{1},t_{2})$ has been determined, $C^{x}_{\star}(t_{1},t_{2})$ follows easily.
-This single-site problem can be derived through either a single-site cavity calculation (see [90]) or a saddle-point condition in a path integral (see [54, 55]).
+This single-site problem can be derived through either a single-site cavity calculation (see Ref.
+[97]) or a saddle-point condition in a path integral (see Refs.
+[60, 61]).
 For the i.i.d.
 couplings considered here, there is a simpler heuristic derivation: In the neuronal input $\sum_{j=1}^{N}J_{ij}\phi_{j}(t)$, the correlations between the couplings $J_{ij}$ and dynamic variables $\phi_{j}(t)$ can be safely neglected to leading order in $1/N$, yielding both Gaussianity of $\eta^{x}(t)$ by the central limit theorem and the second-order statistics of Eq.
-70.
+(71).
 Clark et al.
-[30] first computed $\Psi_{\star}^{a}(\bm{\tau})$ using a dynamic, two-site version of the cavity method, based on the neuron-by-neuron definition (Eq.
-24).
-This method finds $\psi^{a}_{\star}(\bm{\tau})$, the off-diagonal contribution to $\Psi^{a}_{\star}(\bm{\tau})$, noting that the on-diagonal contribution is simply $C^{a}_{\star}(\tau_{1})C^{a}_{\star}(\tau_{2})$.
+[39] first computed $\Psi_{\star}^{a}(\bm{\tau})$ using a dynamic, two-site version of the cavity method, based on the neuron-by-neuron definition [Eq.
+(24)].
+This method finds $\psi^{a}_{\star}(\bm{\tau})$, the off diagonal contribution to $\Psi^{a}_{\star}(\bm{\tau})$, noting that the on-diagonal contribution is simply $C^{a}_{\star}(\tau_{1})C^{a}_{\star}(\tau_{2})$.
 A cavity is first created by removing two neurons from the network and allowing the rest of the network, the reservoir, to generate dynamic activity.
@@
 $U(\bm{\omega})$ $=\frac{g^{2}S_{\star}^{x}(\omega_{1})S_{\star}^{x}(\omega_{2})}{1-g^{2}S_{\star}^{\phi}(\omega_{1})S_{\star}^{\phi}(\omega_{2})},$ (74b)
-for the pre-activations, where $\bm{\omega}=(\omega_{1},\omega_{2})$ and $C^{x\phi}_{\star}(\omega)$ is a cross-covariance between the pre-activation and activation.
-Note that if the joint distribution of pre-activations were Gaussian, $\Psi^{x}_{\star}(\bm{\tau})$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ would differ only by a proportionality constant due to Price’s theorem [91].
+for the preactivations, where $\bm{\omega}=(\omega_{1},\omega_{2})$ and $C^{x\phi}_{\star}(\omega)$ is a cross-covariance between the preactivation and activation.
+Note that, if the joint distribution of preactivations were Gaussian, $\Psi^{x}_{\star}(\bm{\tau})$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ would differ only by a proportionality constant due to Price’s theorem [98].
 The more complex relationship observed here reflects the non-Gaussian joint statistics across different neurons, which is relevant because the network is nonlinear.
@@
 (a) Example activity traces of principal components $p^{\phi}_{k}(t)$.
-(b) Empirical autocovariance functions for each principal component, weighted by their eigenvalues.
+(b) Empirical correlation functions for each principal component (PC), weighted by their eigenvalues.
 (c) Same as (b) but weighted by the squared eigenvalue.
 (d) Averages over curves in (b) and (c), with comparisons to $C^{\phi}_{\star}(\tau)$ and $\Psi^{\phi}_{\star}(\tau,0)$.
-Dots: simulation results.
-Lines: theoretical predictions.
+Dots, simulation results.
+Lines, theoretical predictions.
 All panels describe the same network of $N=4000$ neurons with i.i.d.
@@
 Begin with the path integral $Z[\bm{J}]$ and average over the connectivity disorder.
-Introduce auxiliary fields representing two-point functions (such as $C^{\phi}(t_{1},t_{2})$) along with their conjugate partners.
-Use integral representations of delta functions to enforce field definitions, factorizing the exponential over extensive dimensions.
+Introduce auxiliary fields representing two-point functions [such as $C^{\phi}(t_{1},t_{2})$] along with their conjugate partners.
+Use integral representations of $\delta$ functions to enforce field definitions, factorizing the exponential over extensive dimensions.
 This yields a statistical field theory with action $N$ times an order-one function of the auxiliary fields.
@@
 Find the saddle point by setting action derivatives to zero.
-This produces self-consistent equations for the two-point functions, completing the standard dynamical mean-field theory (DMFT) analysis.
+This produces self-consistent equations for the two-point functions, completing the standard DMFT analysis.
   3. 3.
@@
     Apply temporal separation limits.
-Impose the temporal limits (Eq.
-34), making the Hessian blocks translation-invariant functions of time differences $(\tau_{1},\tau_{2})$.
-Transform to Fourier space to obtain a low-dimensional, frequency-dependent Hessian whose elements are functions of $(\omega_{1},\omega_{2})$.
+Impose the temporal limits [Eq.
+(35)], making the Hessian blocks translation-invariant functions of time differences $(\tau_{1},\tau_{2})$.
+Transform to Fourier space to obtain a low dimensional, frequency-dependent Hessian whose elements are functions of $(\omega_{1},\omega_{2})$.
   5. 5.
@@
 After performing the Gaussian integration over $\bm{J}$, we introduce an auxiliary field $C^{\phi}(t_{1},t_{2})$, defined by Eq.
-17, and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$.
+(17), and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$.
 This leads to a partition function for a statistical field theory involving $C^{\phi}(t_{1},t_{2})$ and $\hat{C}^{\phi}(t_{1},t_{2})$:
@@
 where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C_{\star}^{\phi},0]$.
-This recovers the same single-site process (Eq.
-69, Eq.
-70) and self-consistency condition (Eq.
-71) described in Sec.
-E.
+This recovers the same single-site process [Eqs.
+(70) and (71)] and self-consistency condition [Eq.
+(72)] described in Appendix E.
 F.2 Four-point functions
@@
 $H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ $=\frac{\delta^{2}S}{\delta\hat{C}^{\phi}(t_{1},t_{2})\delta\hat{C}^{\phi}(t_{3},t_{4})}=-\left\langle\left\langle\phi(t_{1})\phi(t_{2}),\phi(t_{3})\phi(t_{4})\right\rangle\right\rangle_{\star},$ (82c)
-where $\left\langle\left\langle A,B\right\rangle\right\rangle_{\star}=\left\langle AB\right\rangle_{\star}-\left\langle A\right\rangle_{\star}\left\langle B\right\rangle_{\star}$, $I(t_{1},t_{2},t_{3},t_{4})=\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})$.
+where $\left\langle\left\langle A,B\right\rangle\right\rangle_{\star}=\left\langle AB\right\rangle_{\star}-\left\langle A\right\rangle_{\star}\left\langle B\right\rangle_{\star}$ and $I(t_{1},t_{2},t_{3},t_{4})=\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})$.
 Expanding the action to second order around the saddle point, the path integral becomes
@@
 We could then take the temporal separation limits of Eq.
-34 to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$.
+(35) to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$.
 The challenge is that each Hessian block is a complicated time2-by-time2 matrix that, for general values of the time variables, has no analytic inverse.
 To circumvent this problem, we show in Appendix G that the temporal limits of Eq.
-34 can be taken before taking the inverses.
+(35) can be taken before taking the inverses.
 Under these temporal limits, the Hessian blocks are
 $H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-C^{\phi}_{\star}(\tau_{1})C^{\phi}_{\star}(\tau_{2}),$ (84)
-$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(\tau_{1})S^{\phi}_{\star}(\tau_{2}),$
+$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(\tau_{1})S^{\phi}_{\star}(\tau_{2}),$ (85)
 $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0,$ (86)
 where we used the fact that multiplication by $i\hat{x}(t)$ is equivalent to a functional derivative with respect to a source at time $t$.
-Since the relevant quantities are now translation-invariant, i.e., depend only on $\tau_{1}$ and $\tau_{2}$, we can transform to Fourier space.
+Since the relevant quantities are now translation invariant, i.e., depend only on $\tau_{1}$ and $\tau_{2}$, we can transform to Fourier space.
 In this representation, each Hessian block becomes a frequency-dependent scalar,
@@
 $H_{C^{\phi}C^{\phi}}(\bm{\omega})$ $=0.$ (89)
-In summary, the full Hessian can be replaced by a two-by-two, frequency-dependent Hesssian given by
+In summary, the full Hessian can be replaced by a $2\times 2$, frequency-dependent Hessian given by
 $\bm{H}=\begin{pmatrix}0&-1+g^{2}(S^{\phi}_{12})^{*}\\ -1+g^{2}S^{\phi}_{12}&-C^{\phi}_{12}\end{pmatrix}$, (90)
@@
 The Fourier-space function $\Psi_{\star}^{\phi}(\bm{\omega})$ is the upper-left element of the inverse of this matrix.
-Performing the two-by-two matrix inversion gives
+Performing the $2\times 2$ matrix inversion gives
 $\Psi_{\star}^{\phi}=\frac{C^{\phi}_{12}}{|1-g^{2}S^{\phi}_{12}|^{2}},$ (91)
-which agrees with the two-site cavity result (Eq. 72).
+which agrees with the two-site cavity result [Eq. (73)].
 F.3 Adding sources
@@
 To compute $\Psi^{x}_{\star}(\bm{\tau})$, we need to introduce a source-field term in the action.
+Consider a general intensive action $\mathcal{S}[\bm{C};\mathcal{J}]$, where $\bm{C}$ is a collection of fields (e.g., $\bm{C}=\{C^{\phi},\hat{C}^{\phi}\}$), and $\mathcal{J}$ is a source.
 Our goal is to compute
@@
 To facilitate this calculation, we introduce a new field $U$ and set it equal to $\mathcal{J}$ using the conjugate $\hat{U}$:
-$Z[\mathcal{J}]=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp\left(-N\hat{U}U+N\hat{U}\mathcal{J}-NS[\hat{C};U]\right).$ (91)
-Instead of taking the second derivative with respect to $\mathcal{J}$, we set $\mathcal{J}=0$ and compute the fluctuations of $\hat{U}$ in the augmented theory whose path integral is
+$Z[\mathcal{J}]=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp\left(-N\hat{U}U+N\hat{U}\mathcal{J}-NS[\bm{C};U]\right).$ (93)
+Instead of taking the second derivative with respect to $\mathcal{J}$, we set $\mathcal{J}=0$ and compute the fluctuations of $\hat{U}$, measured by Eq.
+(92), in the augmented theory whose path integral is
 […]
@@
 $H_{ab}=\frac{\delta^{2}S}{\delta C_{a}\delta C_{b}},\quad\left[\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{a}=\frac{\delta^{2}S}{\delta\mathcal{J}\delta C_{a}}.$ (97)
-The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point.
-Consider a general intensive action $\mathcal{S}[\bm{C};\mathcal{J}]$ where $\bm{C}$ is a collection of fields (e.g., $\bm{C}=\{C^{\phi},\hat{C}^{\phi}\}$), and $\mathcal{J}$ is a source.
-Using a Schur complement to compute this element, we obtain
+The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point. Using a Schur complement to compute this element, we obtain
 $\left.\frac{\delta^{2}}{\delta\mathcal{J}^{2}}\log Z\right|_{\mathcal{J}=0}=\left[-\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}+\left(\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right)^{T}\bm{H}^{-1}\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}\right]_{\mathcal{J}=0},$ (98)
@@
 By applying this formula to the action of the i.i.d.
-model with the source term $\frac{1}{2}x\mathcal{J}x$ added to the single-site path integral (Eq.
-76), we recover the expression for $\Psi^{x}_{\star}(\bm{\tau})$ from the two-site cavity method (Clark et al.
+model with the source term $\frac{1}{2}x\mathcal{J}x$ added to the single-site path integral [Eq.
+(77)], we recover the expression for $\Psi^{x}_{\star}(\bm{\tau})$ from the two-site cavity method [Clark et al.
 [39]; Eq.
-73).
+(74)].
 For actions of the form $\mathcal{S}[\bm{C};\mathcal{J}]=-\mathcal{J}C_{a}+\text{$\mathcal{J}$-independent terms}$, this formula reduces to $({\delta^{2}}/{\delta\mathcal{J}^{2}})\log Z|_{\mathcal{J}=0}=[\bm{H}^{-1}]_{aa},$ as expected.
-We now apply Eq. 96 to the action of the i.i.d. connectivity model with a source term for correlations among pre-activations. The action is
+We now apply Eq. (98) to the action of the i.i.d. connectivity model with a source term for correlations among preactivations. The action is
 $\mathcal{S}[C^{\phi},\hat{C}^{\phi}]=-\frac{1}{2}C^{\phi}\hat{C}^{\phi}-\log\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\bigg(i\hat{x}T[x]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi-\frac{1}{2}x\mathcal{J}x\bigg).$ (99)
-Computing the necessary quantities and taking the temporal limits of Eq. 34, we obtain, in Fourier space,
+Computing the necessary quantities and taking the temporal limits of Eq. (35), we obtain, in Fourier space,
 $\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\quad\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}g^{2}\left(S_{12}^{x}\right)^{*}\\ -C_{12}^{x\phi}\end{pmatrix}$, (100)
@@
 Substituting these into Eq.
-96 and using the frequency-dependent Hessian from Eq.
-88, we obtain
+(98) and using the frequency-dependent Hessian from Eq.
+(90), we obtain
 $\Psi_{\star}^{x}=C^{x}_{12}+|U|^{2}C^{\phi}_{12}+UC^{x\phi}_{12}+\text{H.c.},\quad\text{where }U=\frac{g^{2}S^{x}_{12}}{1-g^{2}S^{\phi}_{12}},$ (101)
-in agreement with the two-site cavity result of Clark et al. [30], given in Eq. 73.
+in agreement with the two-site cavity result of Clark et al. (Ref. [39]), given in Eq. (74).
 F.4 Other kinds of path integrals
-Different path-integral formalisms have been applied to neural network models depending on the underlying dynamics.
+Different path-integral formalisms have been applied to neural-network models depending on the underlying dynamics.
 The Doi-Peliti formalism [100] derives path integrals from master equations for systems with non-negative integer degrees of freedom (such as spiking neural networks) that evolve according to master equations, using an operator formalism as an intermediate step.
-In contrast, we use the MSRDJ formalism for continuous-time stochastic differential equations, which naturally describes our rate-based network dynamics.
-For a comprehensive discussion of different path-integral approaches in neural network theory, see [60].
+In contrast, we use the MSRJD formalism for continuous-time stochastic differential equations, which naturally describes our rate-based network dynamics.
+For a comprehensive discussion of different path-integral approaches in neural-network theory, see Ref.
+[69].
 Appendix G Temporal limits and inverses
 Here, we demonstrate, for the i.i.d.
-model that the temporal limits in Eq.
-34 can be taken before the inverses.
+model, that the temporal limits in Eq.
+(35) can be taken before the inverses.
 This allows us to work with a frequency-dependent Hessian.
@@
 We would like to evaluate
-$\Psi(\tau_{1},\tau_{2})=U^{-1}(t_{1},t_{2},s_{1},s_{2})K(s_{1},s_{2},s_{3},s_{4})U^{-1}(t_{1}+\tau_{1},t_{2}+\tau_{2},s_{3},s_{4})$ (100)
-in the limits Eq.
-34, where $U(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ and $K(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ (note that $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0$ due to the vanishing of correlation functions involving only the conjugate field).
+$\Psi^{\phi}(\tau_{1},\tau_{2})=U^{-1}(t_{1},t_{2},s_{1},s_{2})K(s_{1},s_{2},s_{3},s_{4})U^{-1}(t_{1}+\tau_{1},t_{2}+\tau_{2},s_{3},s_{4})$ (102)
+in the limits of Eq.
+(35), where $U(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ and $K(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ [note that $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0$ due to the vanishing of correlation functions involving only the conjugate field].
 Restating Eq.
-81, we have
+(82), we have
 $U(t_{1},t_{2},t_{3},t_{4})$ $=-\left[\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})\right]-g^{2}\langle\langle\phi(t_{1})\phi(t_{2}),\hat{x}(t_{3})\hat{x}(t_{4})\rangle\rangle_{\star},$ (103)
@@
 We start by examining the properties of $(t_{1},t_{2},t_{3},t_{4}).
-UU$ vanishes if one of the four time points is far (measured on the timescale of $C^{\phi}_{\star}(\tau)$) from the other three.
+UU$ vanishes if one of the four time points is far [measured on the timescale of $C^{\phi}_{\star}(\tau)$] from the other three.
 For $t_{1}$ or $t_{2}$, this is a consequence of the perturbation associated with $\hat{x}(t_{1})$ or $\hat{x}(t_{2})$ being far away; for $t_{3}$ or $t_{4}$, this is a consequence of $\langle\phi(t_{i})\rangle_{\star}=0$ separating from the expectation.
@@
 However, if $t_{1},t_{2}$ are close and $t_{3},t_{4}$ are close, but the pairs are distant, $U$ still vanishes since the perturbations associated with $\hat{x}(t_{i})$ are both far.
-This leaves three possibilities for a non-vanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
-Separating the three possibilities into two subdomains where all time points are close ($c$) and where the time points belong to either of the two pairwise separate ($s$) possibilities, we write $U=U_{s}+U_{c}$ where $U_{s}$ and $U_{c}$ are defined to vanish on the respective other domain.
+This leaves three possibilities for a nonvanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
+Separating the three possibilities into two subdomains where all time points are close ($c$) and where the time points belong to either of the two pairwise separate ($s$) possibilities, we write $U=U_{s}+U_{c}$, where $U_{s}$ and $U_{c}$ are defined to vanish on the respective other domain.
 On the separated domain, $U$ simplifies to
 $U_{s}(t_{1},t_{2},t_{3},t_{4})=-\left[\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})\right]\\ +g^{2}\left[S^{\phi}_{\star}(t_{1}-t_{3})S^{\phi}_{\star}(t_{2}-t_{4})+S^{\phi}_{\star}(t_{1}-t_{4})S^{\phi}_{\star}(t_{2}-t_{3})\right],$ (105)
-where we used that the two distant contributions are non-overlapping to add them.
+where we used that the two distant contributions are nonoverlapping to add them.
 We assume that we can also decompose the inverse into a contribution for separate pairs and a contribution where all time points are close, $U^{-1}=U_{s}^{-1}+U_{c}^{-1}$, where again $U_{s}^{-1}$ and $U_{c}^{-1}$ vanish on the respective other domain.
@@
 For Eq.
-100, we need $U^{-1}_{s}$ due to the limits Eq.
-34.
-The fact that $U$ reduces to $U_{s}$ in the limits Eq.
-34 justifies taking the limits before inverting $U$.
+(102), we need $U^{-1}_{s}$ due to the limits of Eq.
+(35).
+The fact that $U$ reduces to $U_{s}$ in the limits of Eq.
+(35) justifies taking the limits before inverting $U$.
 Furthermore, these limits imply that, in Eq.
-100, $s_{1}$ and $s_{2}$, as well as $s_{3}$ and $s_{4}$, are far apart.
+(102), $s_{1}$ and $s_{2}$, as well as $s_{3}$ and $s_{4}$, are far apart.
 In this case, $K$ simplifies to
 $K_{s}(s_{1},s_{2},s_{3},s_{4})=C^{\phi}_{\star}(s_{1}-s_{3})C^{\phi}_{\star}(s_{2}-s_{4})+C^{\phi}_{\star}(s_{1}-s_{4})C^{\phi}_{\star}(s_{2}-s_{3}).$ (106)
-This justifies taking the limits Eq. 34 for $K$.
+This justifies taking the limits of Eq. (35) for $K$.
 Appendix H Partially symmetric disorder
@@
 model in which we introduce correlations between reciprocal couplings, characterized by the following statistics:
-$\left\langle J_{ij}J_{k\ell}\right\rangle=g^{2}\delta_{ik}\delta_{j\ell}+g^{2}\rho\delta_{i\ell}\delta_{jk}.$ (105)
+$\left\langle J_{ij}J_{k\ell}\right\rangle=\frac{g^{2}}{N}\delta_{ik}\delta_{j\ell}+\frac{g^{2}\rho}{N}\delta_{i\ell}\delta_{jk}.$ (107)
 Here, $\rho\in[-1,1]$ is a parameter that controls the degree of symmetry in the connectivity.
@@
 We express this partially symmetric connectivity as a linear combination of two matrices, $\bm{X}$ and $\bm{Y}$, that are i.i.d.
-random with mean zero and variance $1/N$:
-$J_{ij}=\sqrt{{g^{2}-|\sigma|}}X_{ij}+\sqrt{\frac{|\sigma|}{2}}(Y_{ij}+\text{sgn}(\sigma)Y_{ji}),$ (106)
+random with mean zero and variance $1/N$,
+$J_{ij}=\sqrt{{g^{2}-|\sigma|}}X_{ij}+\sqrt{\frac{|\sigma|}{2}}[Y_{ij}+\text{sgn}(\sigma)Y_{ji}],$ (108)
 where $\sigma=\rho g^{2}$. In terms of $\bm{X}$ and $\bm{Y}$, the path integral for this partially symmetric model is
-$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg{(}i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{ij}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{ij}Y_{ij}(\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j})\bigg{)}.$ (107)
+$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg(i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{i,j=1}^{N}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{i,j=1}^{N}Y_{ij}[\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j}]\bigg).$ (109)
 We integrate out $\bm{X}$ and $\bm{Y}$. To factorize the action over indices spanning extensive dimensions, we introduce
@@
 $\frac{\delta\mathcal{S}}{\delta\hat{S}^{\phi}(t_{1},t_{2})}$ $=\sigma S^{\phi}(t_{1},t_{2})+i\sigma\left\langle\phi(t_{1})\hat{x}(t_{2})\right\rangle_{W}.$ (118)
-Setting these derivatives to zero yields the saddle-point conditions:
+Setting these derivatives to zero yields the saddle-point conditions,
 $\hat{C}^{\phi}_{\star}(t_{1},t_{2})$ $=0,$ (119)
@@
 The single-site process at the saddle point is described by
-$T[x](t)$ $=\eta^{x}(t)+\sigma\left[S^{\phi}_{\star}\ast\phi\right](t),$ (120)
+$T[x](t)$ $=\eta^{x}(t)+\sigma\left[S^{\phi}_{\star}\circ\phi\right](t),$ (122)
 $\eta^{x}$ $\sim\mathcal{GP}(0,g^{2}C^{\phi}_{\star}),$ (123)
-where $\ast$ denotes convolution.
-The symmetric structure provides a convolutional, nonlinear self-coupling $\sigma[S^{\phi}_{\star}\ast\phi](t)$ in the single-site dynamics.
-The two-point correlation and response functions, $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$, must be determined self-consistently within this single site picture.
+where $\circ$ denotes convolution.
+The symmetric structure provides a convolutional, nonlinear self-coupling $\sigma[S^{\phi}_{\star}\circ\phi](t)$ in the single-site dynamics.
+The two-point correlation and response functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ must be determined self-consistently within this single-site picture.
 H.2 Four-point functions
@@
 (124)
-Inverting this matrix and isolating the upper-left element, we obtain, in agreement with Clark et al. [30],
+Inverting this matrix and isolating the upper-left element, we obtain, in agreement with Clark et al. (Ref. [39]),
 $\Psi^{\phi}_{\star}=\frac{1}{|1-g^{2}S^{\phi}_{12}|^{2}}\frac{1-|\sigma(S^{\phi}_{1})^{*}S^{\phi}_{2}|^{2}}{|1-\sigma(S^{\phi}_{1})^{*}S^{\phi}_{2}|^{2}}C^{\phi}_{12}.$ (125)
@@
 $z_{a}(t)=D_{a}\sum_{i=1}^{N}R_{ia}\phi_{i}(t),$ (127)
-and their conjugates $\hat{z}_{a}(t)$. We use indices $i,j$ for neurons and $a,b$ for latent variables. The introduction of these latent variables makes the action linear in $\bm{L}$ and $\bm{R}$:
+and their conjugates $\hat{z}_{a}(t)$. We use indices $i,j$ for neurons and $a,b$ for latent variables. The introduction of these latent variables makes the action linear in $\bm{L}$ and $\bm{R}$,
 $Z[\bm{L},\bm{R}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\int\mathcal{D}\bm{z}\int\mathcal{D}\hat{\bm{z}}\exp\Big(i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]+i\sum_{a=1}^{M}z_{a}\hat{z}_{a}-i\sum_{i,a}L_{ia}\hat{x}_{i}z_{a}-i\sum_{i,a}D_{a}R_{ia}\phi_{i}\hat{z}_{a}\Big).$ (128)
 We can now easily integrate out $\bm{L}$ and $\bm{R}$.
-To factorize the action over indices spanning extensive dimensions, we introduce the field $C^{\phi}(t_{1},t_{2})$ (Eq.
-17) with conjugate $\hat{Q}(t_{1},t_{2})$ as well as
+To factorize the action over indices spanning extensive dimensions, we introduce the field $C^{\phi}(t_{1},t_{2})$ [Eq.
+(17)] with conjugate $\hat{Q}(t_{1},t_{2})$ as well as
 $Q(t_{1},t_{2})=\frac{1}{N}\sum_{a=1}^{M}z_{a}(t_{1})z_{a}(t_{2})$ (129)
@@
 $\eta^{z}$ $\sim\mathcal{GP}(0,D^{2}C^{\phi}_{\star}),$ (144)
-showing that the neuronal and latent-variable single-site processes have mutually referential statistics  [^8] Note that the variance of $\eta^{z}(t)$ is scaled by $D^{2}$; equivalently, we could factor out $D^{2}$ from the variance and include $D$ as a multiplicative factor in the latent-variable single-site dynamics (Eq.
-140).
-Solving for $Q_{\star}(t_{1},t_{2})$ by combining Eq.
-138, Eq.
-140, and Eq.
-142, we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq.
-10.
+showing that the neuronal and latent-variable single-site processes have mutually referential statistics  [^8] Note that the variance of $\eta^{z}(t)$ is scaled by $D^{2}$; equivalently, we could factor out $D^{2}$ from the variance and include $D$ as a multiplicative factor in the latent-variable single-site dynamics [Eq.
+(142)].
+Solving for $Q_{\star}(t_{1},t_{2})$ by combining Eqs.
+(140), (142), and (144), we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq.
+(10).
 Consolidating these results yields Eq.
-36c.
+(37c).
 I.2 Four-point functions
@@
 Component strengths $D_{a}$ are constant for $a/N\leq\text{PR}^{D}$ and 0 otherwise.
-Small dots: individual simulations; large dots: means over 10 simulations; solid lines: theoretical predictions; dotted lines: small-$\text{PR}^{D}$ expansion, $K^{\phi}(g_{\text{eff}})\mathrm{PR}^{D}$.
+Small dots, individual simulations; large dots: means over ten simulations; solid lines, theoretical predictions; dotted lines, small-$\text{PR}^{D}$ expansion, $K^{\phi}(g_{\text{eff}})\mathrm{PR}^{D}$.
 Inset: ratio of proportionality factor $K^{\phi}(g_{\text{eff}})$ to $\text{PR}^{\phi}_{\text{i.i.d.}}$ versus $g_{\text{eff}}$.
-Bottom: Same analysis for pre-activations.
+Bottom: same analysis for preactivations.
 (b) Top: rank-ordered eigenvalue plots for activation covariance matrix $C^{\phi}_{ij}(0)$.
 Dotted vertical lines indicate $N\,\text{PR}^{D}$.
-Bottom: Same for pre-activations.
+Bottom: same for preactivations.
 All simulations use $N=6500$ neurons.
@@
 network with linear dynamics and a heterogeneous distribution of single-neuron scale factors $D_{a}$; the lower-right block, corresponding to the neurons, is related to the Hessian of an i.i.d.
-network of nonlinear neurons (Eq.
-88).
-They are coupled through the simple off-diagonal blocks, reflecting the interaction between neurons and latent variables.
-This yields the expression for $\Psi^{\phi}_{\star}$ (Eq. 38). To obtain $\Psi^{x}_{\star}$, we add $\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral (Eq. 130) and compute
+network of nonlinear neurons [Eq.
+(90)].
+They are coupled through the simple off diagonal blocks, reflecting the interaction between neurons and latent variables.
+This yields the expression for $\Psi^{\phi}_{\star}$ [Eq. (39)]. To obtain $\Psi^{x}_{\star}$, we add $\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral [Eq. (132)] and compute
 $\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C_{12}^{x},\>\>\>\frac{\delta^{2}S}{\delta\mathcal{J}\delta\bm{C}}=\begin{pmatrix}0\\ 0\\ (S_{12}^{x})^{*}\\ -C_{12}^{x\phi}\\ \end{pmatrix}$. (146)
-Inserting these into Eq. 96 we obtain
+Inserting these into Eq. (98) we obtain
 $\Psi_{\star}^{x}=C^{x}_{12}+\left(1+\frac{1}{\alpha\text{PR}^{D}}\right)|U|^{2}C^{\phi}_{12}+UC^{x\phi}_{12}+\text{H.c.},$ (147)
-where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ (Eq. 73b with $g=g_{\text{eff}}$). Here, we have used the following formula to invert the Hessian:
+where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ [Eq. (74b) with $g=g_{\text{eff}}$]. Here, we have used the following formula to invert the Hessian:
 $\begin{pmatrix}0&a&0&-1\\ a^{*}&b&-1&0\\ 0&-1&0&c\\ -1&0&c^{*}&d\end{pmatrix}^{-1}=\begin{pmatrix}-\frac{b+|c|^{2}d}{|1-ac|^{2}}&-\frac{c^{*}}{1-(ac)^{*}}&-\frac{ad+bc^{*}}{|1-ac|^{2}}&-\frac{1}{1-(ac)^{*}}\\ -\frac{c}{1-ac}&0&-\frac{1}{1-ac}&0\\ -\frac{a^{*}d+bc}{|1-ac|^{2}}&-\frac{1}{1-(ac)^{*}}&-\frac{|a|^{2}d+b}{|1-ac|^{2}}&-\frac{a^{*}}{1-(ac)^{*}}\\ -\frac{1}{1-ac}&0&-\frac{a}{1-ac}&0\end{pmatrix}$.
@@
 $Q(t_{1},t_{2})$ $=\frac{1}{N}\sum_{a=1}^{M}z_{a}(t_{1})z_{a}(t_{2}),$ $\hskip-72.26999pt\text{conjugate: }\hat{C}^{\phi}(t_{1},t_{2}),$ (150)
-where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}(x_{i}(t))$. The resulting action and single-site path integrals for this system are
+where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}\bm{(}x_{i}(t)\bm{)}$. The resulting action and single-site path integrals for this system are
 $\mathcal{S}[C^{\Phi},\hat{C}^{\Phi},Q,\hat{Q}$] $=-\frac{1}{2}C^{\Phi}\hat{Q}-\frac{1}{2}Q\hat{C}^{\phi}-\left\langle\log W_{\bm{\theta}}[Q,\hat{Q};\mathcal{J}]\right\rangle_{\bm{\theta}}-\alpha\left\langle\log W^{z}_{D}[C^{\Phi},\hat{C}^{\Phi}]\right\rangle_{D},$ (151)
@@
 The self-consistency condition Eq.
-155c is identical to that of an i.i.d.
+(157c) is identical to that of an i.i.d.
 network, but with an average over $\bm{\theta}\sim P(\bm{\theta})$, in addition to the usual Gaussian average over $\eta^{x}(t)$, to account for single-neuron heterogeneity.
 J.2 Four-point functions
-To write down the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$:
+To write down the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$,
 $\mathfrak{C}^{\Phi}_{12}$ $=\left\langle\left\langle\Phi_{\bm{\theta}}\Phi_{\bm{\theta}}\right\rangle_{\star}(\omega_{1})\left\langle\Phi_{\bm{\theta}}\Phi_{\bm{\theta}}\right\rangle_{\star}(\omega_{2})\right\rangle_{\bm{\theta}},$ (158a)
@@
 $\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}{C_{{12}}^{\Phi}}&-1&0\\ \hline\cr 0&-1&0&\mathfrak{s}^{\Phi}_{12}\\ -1&0&(\mathfrak{s}^{\Phi}_{12})^{*}&-\mathfrak{C}^{\Phi}_{12}\end{array}\right),$ (159)
-where the new $\mathfrak{F}$raktur variables are defined in Eq. 156. We invert the Hessian using Eq. 146 and identify the upper-left element as $\Phi_{\bm{\theta}_{i}}(\bm{\omega})$:
+where the new Fraktur variables are defined in Eq. (158). We invert the Hessian using Eq. (148) and identify the upper-left element as $\Phi_{\bm{\theta}_{i}}(\bm{\omega})$,
 $\Psi^{\Phi}_{\star}=\frac{\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\Phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}|\alpha r_{2}\mathfrak{s}^{\Phi}_{12}|^{2}}{|1-\alpha r_{2}\mathfrak{s}^{\Phi}_{12}|^{2}}C^{\Phi}_{12}.$ (160)
@@
 $S^{\phi}_{12}$ $=\left\langle\frac{\delta\phi}{\delta I}\right\rangle_{\star}(\omega_{1})\left\langle\frac{\delta\phi}{\delta I}\right\rangle_{\star}(\omega_{2}),$ (163)
-we obtain for the normalized variables $\phi_{i}(t)$:
+we obtain for the normalized variables $\phi_{i}(t)$,
 $\Psi^{\phi}_{\star}=\Bigg[|A|^{2}\left(\frac{\mathfrak{C}^{\Phi}_{12}}{C^{\phi}_{12}}+\frac{1}{\alpha\text{PR}^{D}}\frac{C^{\Phi}_{12}}{C^{\phi}_{12}}\right)+\left(A\frac{\mathfrak{C}^{\Phi\phi}_{12}}{C^{\phi}_{12}}+\text{H.c.}\right)+1\Bigg]C^{\phi}_{12},$ (164)
@@
 J.3 Four-point functions for firing-rate heterogeneity
-Upon specializing the above general results to the case where $\bm{\theta}_{i}$ contains a single gain parameter, $G_{i}$, we obtain the four-point functions for unnormalized (Eq.
-46a) and normalized (Eq.
-46b) variables given in the main text.
+Upon specializing the above general results to the case where $\bm{\theta}_{i}$ contains a single gain parameter $G_{i}$, we obtain the four-point functions for unnormalized [Eq.
+(47a)] and normalized [Eq.
+(47b)] variables given in the main text.
 Note that, like the distribution over component strengths $D_{a}$, two-point functions of activity depend only on the second moment of $G_{i}$, and four-point functions depend only on the second and fourth moments of $G_{i}$.
-For “random readout” gains with the same distribution as the recurrent gains (see main text), the four-point function is
+For “random-readout” gains with the same distribution as the recurrent gains (see main text), the four-point function is
 $\Psi^{\Phi^{\text{readout}}}=\Bigg[\left(\frac{1}{\text{PR}^{G}}-1\right)\left(|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}+|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\right)+1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\Bigg]\frac{C^{\Phi^{\text{readout}}}_{12}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}},$ (166)
@@
 Here, we consider random gains rather than judiciously chosen or learned gains.
-In particular, they are sampled independently of the connectivity.
+In particular, they are sampled independent of the connectivity.
 However, our framework could be extended to model gains that are related to the connectivity and, more ambitiously, could be extended to model learned gain parameters (or, as explored in the Discussion, learned connectivity).
 Appendix K Path-integral calculation of two- and four-point functions for the random-mode model with diagonally structured overlaps
-To generate matrices $\bm{L}$ and $\bm{R}$ with appropriate correlation structure, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$:
+To generate matrices $\bm{L}$ and $\bm{R}$ with appropriate correlation structure, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$,
 ${L}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{1}_{ia}+\sqrt{{|\rho_{a}|}}{Y}_{ia},$ (167)
 ${R}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{2}_{ia}+\text{sgn}(\rho_{a})\sqrt{{|\rho_{a}|}}{Y}_{ia}.$ (168)
-We begin with the path integral containing latent variables, insert this parameterization of $\bm{L}$ and $\bm{R}$, then and integrate out $\bm{X}_{1}$, $\bm{X}_{2}$, and $\bm{Y}$.
+We begin with the path integral containing latent variables, insert this parametrization of $\bm{L}$ and $\bm{R}$, and then integrate out $\bm{X}_{1}$, $\bm{X}_{2}$, and $\bm{Y}$.
 To factorize the action over indices spanning extensive dimensions, we introduce
@@
 $R(t_{1},t_{2})$ $=\frac{1}{N}\sum_{a=1}^{M}D_{a}\rho_{a}\frac{\delta z_{a}(t_{1})}{\delta I_{a}(t_{2})},$ $\hskip-72.26999pt\text{conjugate: }\hat{S}^{\phi}(t_{1},t_{2}).$ (172)
-The resulting path integral, action, and single-site path integrals are given by:
+The resulting path integral, action, and single-site path integrals are given by
 […]
@@
 K.1 Two-point functions
-We first calculate the saddle point conditions by taking the derivatives of the action:
+We first calculate the saddle-point conditions by taking the derivatives of the action,
 $\frac{\delta\mathcal{S}}{\delta C^{\phi}(t_{1},t_{2})}$ $=-\hat{Q}(t_{1},t_{2})+\alpha\left\langle D^{2}\left\langle\hat{z}(t_{1})\hat{z}(t_{2})\right\rangle_{W^{z}_{D,\rho}}\right\rangle_{D,\rho},$ (178)
@@
 $\frac{\delta\mathcal{S}}{\delta\hat{R}(t_{1},t_{2})}$ $=S^{\phi}(t_{1},t_{2})+i\left\langle\phi(t_{1})\hat{x}(t_{2})\right\rangle_{W^{x}}.$ (185)
-The saddle-point conditions yield:
+The saddle-point conditions yield
 $\hat{C}^{\phi}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (186)
@@
 where $\left\langle\cdots\right\rangle_{\star D,\rho}$ and $\left\langle\cdots\right\rangle_{\star}$ are averages within the dynamic processes described by $W^{z}_{D,\rho}[C^{\phi}_{\star},0,S^{\phi}_{\star},\hat{S}^{\phi}_{\star}]$ and $W^{x}[Q_{\star},0,R_{\star},\hat{R}_{\star}$], respectively.
-The yields neuronal and latent-variable single-site processes described by
-$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\ast\phi](t)$ (189)
-$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\ast z](t),$ (190)
+This yields neuronal and latent-variable single-site processes described by
+$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\circ\phi](t)$ (191)
+$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\circ z](t),$ (192)
 where the Gaussian fields have statistics
@@
 $\sigma_{k}$ $=\frac{1}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (200)
-Isolating the upper-left element of the inverse of this matrix yields Eq. 51.
-K.3 Sigmoidal parameterization of overlaps
-To model structured left–right mode overlaps, we used the following sigmoidal form:
-$\rho_{a}=\gamma_{\rho}\left(1-\frac{2}{1+\exp(-\beta_{\rho}(a/N-u_{\text{flip}}))}\right),$ (199)
+Isolating the upper-left element of the inverse of this matrix yields Eq. (52).
+K.3 Sigmoidal parametrization of overlaps
+To model structured left-right mode overlaps, we used the following sigmoidal form:
+$\rho_{a}=\gamma_{\rho}\left(1-\frac{2}{1+\exp[-\beta_{\rho}(a/N-u_{\text{flip}})]}\right),$ (201)
 where $\rho_{a}=0$ at $a/N=u_{\text{flip}}$.
@@
 While the path-integral approach provides a systematic framework for analyzing neural-network dynamics, the cavity method offers an intuitive, complementary perspective.
-This section provides an overview of the cavity analysis for the random-mode model, with the full mathematical details available in Appendix L.
+This section provides an overview of the cavity analysis for the random-mode model.
 While these approaches yield equivalent results, they differ in their methodology: The path-integral approach focuses on fluctuations around the saddle point in a field theory, whereas the cavity approach examines perturbations induced by a subset of held-out variables.
-Here we compute the four-point function using a two-site cavity approach, noting that the two-point function (which, if desired, could be computed using a single-site cavity method) is the same as that of an i.i.d.
+Here, we compute the four-point function using a two-site cavity approach, noting that the two-point function (which, if desired, could be computed using a single-site cavity method) is the same as that of an i.i.d.
 network with appropriate $g_{\text{eff}}$.
 Leveraging Eq.
-125, we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al.
+(127), we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al.
 [39] to both groups.
@@
 As in the path-integral formalism, self-consistent expressions in one picture are defined using averages from the other.
-Unlike the path-integral approach, which derives $\Psi^{\phi}(\bm{\tau})$ using its time-by-time definition (Eq.
-25), the cavity method derives this function using its neuron-by-neuron definition (Eq.
-24).
+Unlike the path-integral approach, which derives $\Psi^{\phi}(\bm{\tau})$ using its time-by-time definition [Eq.
+(25)], the cavity method derives this function using its neuron-by-neuron definition [Eq.
+(24)].
 The cavity method we employ, schematized in Fig.
@@
 Second, it incorporates a bipartite structure: We treat both neurons and latent variables as dynamic objects, creating separate but mutually referential cavity pictures for each group.
-Such bipartite structure, without the two-site structure, has been used in cavity calculations in the Hopfield model [96, 97].
+Such bipartite structure, without the two-site structure, has been used in cavity calculations in the Hopfield model [103, 104, 69].
 Note that these two features are distinct: The two-site aspect refers to the number of neurons removed in each cavity, while the bipartite aspect refers to the types of variables considered (neurons and latent variables).
@@
 as done in the path-integral approach.
-In the two-site cavity calculation, we keep track various intermediate quantities to order $1/\sqrt{N}$, leading to an expression for $\Psi^{\phi}(\bm{\tau})$ accurate to leading order (order one), which we identify as $\Psi^{\phi}_{\star}(\bm{\tau})$.
+In the two-site cavity calculation, we keep track of various intermediate quantities to order $1/\sqrt{N}$, leading to an expression for $\Psi^{\phi}(\bm{\tau})$ accurate to leading order (order one), which we identify as $\Psi^{\phi}_{\star}(\bm{\tau})$.
 To distinguish between different types of variables and their indices, we use the following notation.
@@
 We begin by introducing two neurons, $x_{0}(t)$ and $x_{0^{\prime}}(t)$. The leading-order ($1/\sqrt{N}$) effect on the latent variables is
-$\delta z_{a}(t)=\int^{t}dt^{\prime}\sum_{b}S^{z}_{ab}(t_{1},t_{2})D_{b}\sum_{\mu}R_{\mu b}\phi_{\mu}(t^{\prime}),$ (202)
+$\delta z_{a}(t)=\int^{t}dt^{\prime}\sum_{b=1}^{M}S^{z}_{ab}(t_{1},t_{2})D_{b}\sum_{\mu\in\{0,0^{\prime}\}}R_{\mu b}\phi_{\mu}(t^{\prime}),$ (204)
 where $S^{z}_{ab}(t_{1},t_{2})$ is the response function of the latent variables. The dynamic equations for the cavity neurons are
-$T[x_{\mu}](t)=\eta_{\mu}(t)+\sum_{\nu}\int^{t}dt^{\prime}F_{\mu\nu}(t_{1},t_{2})\phi_{\nu}(t^{\prime}),$ (203)
+$T[x_{\mu}](t)=\eta_{\mu}(t)+\sum_{\nu\in\{0,0^{\prime}\}}\int^{t}dt^{\prime}F_{\mu\nu}(t_{1},t_{2})\phi_{\nu}(t^{\prime}),$ (205)
 where
-$\eta_{\mu}(t)$ $=\sum_{a}L_{\mu a}z_{a}(t),$ (204)
-$F_{\mu\nu}(t_{1},t_{2})$ $=\sqrt{N}\sum_{a,b}L_{\mu a}D_{b}R_{\nu b}S^{z}_{ab}(t_{1},t_{2}).$ (205)
+$\eta_{\mu}(t)$ $=\sum_{a=1}^{M}L_{\mu a}z_{a}(t),$ (206)
+$F_{\mu\nu}(t_{1},t_{2})$ $=\sqrt{N}\sum_{a,b=1}^{M}L_{\mu a}D_{b}R_{\nu b}S^{z}_{ab}(t_{1},t_{2}).$ (207)
 Defining the cavity-field time-average cross-covariance as
-$C^{\eta}_{00^{\prime}}(\omega)=\sum_{a,b}L_{0a}L_{0^{\prime}b}C^{z}_{ab}(\omega),$ (206)
+$C^{\eta}_{00^{\prime}}(\omega)=\sum_{a,b=1}^{M}L_{0a}L_{0^{\prime}b}C^{z}_{ab}(\omega),$ (208)
 we obtain the following expression for the cross-covariance of the cavity units, up to order $1/\sqrt{N}$:
-$C^{\phi}_{00^{\prime}}(\omega)=|S_{\star}^{\phi}(\omega)|^{2}C^{\eta}_{00^{\prime}}(\omega)+\frac{1}{\sqrt{N}}\left[(S_{\star}^{\phi}(\omega)F_{00^{\prime}}(\omega))^{*}+S_{\star}^{\phi}(\omega)F_{0^{\prime}0}(\omega)\right]C^{\phi}_{\star}(\omega).$ (207)
+$C^{\phi}_{00^{\prime}}(\omega)=|S_{\star}^{\phi}(\omega)|^{2}C^{\eta}_{00^{\prime}}(\omega)+\frac{1}{\sqrt{N}}\left[\left(S_{\star}^{\phi}(\omega)F_{00^{\prime}}(\omega)\right)^{*}+S_{\star}^{\phi}(\omega)F_{0^{\prime}0}(\omega)\right]C^{\phi}_{\star}(\omega).$ (209)
 Our goal is to compute the parameter
@@
 where $\left\langle\cdot\right\rangle_{\bm{L},\bm{D},\bm{R}}$ denotes an average over $\bm{L}$, $\bm{D}$, and $\bm{R}$.
-To evaluate this, we need to square and disorder-average Eq.
-207.
+To evaluate this, we need to square and disorder average Eq.
+(209).
 This requires us to consider several two-frequency correlation functions, which we define as
@@
 $\Gamma_{C^{\eta}_{00^{\prime}}C^{\eta}_{00^{\prime}}}(\bm{\omega})$ $=N\left\langle C^{\eta}_{00^{\prime}}(\omega_{1})C^{\eta}_{00^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}}.$ (214)
-These have all been scaled to be order-one.
+These have all been scaled to be order one.
 We can evaluate these $\Gamma_{\cdots}(\bm{\omega})$ functions due to the independence of the couplings and dynamic variables in the expressions defining them, which is a consequence of the cavity construction.
@@
 We now introduce two new latent variables, $z_{\hat{0}}(t)$ and $z_{\hat{0}^{\prime}}(t)$. The leading-order ($1/\sqrt{N}$) effect on the neurons is
-$\delta\phi_{i}(t)=\int^{t}dt^{\prime}\sum_{j=1}^{N}S^{\phi}_{ij}(t_{1},t_{2})\sum_{\alpha}L_{j\alpha}z_{\alpha}(t^{\prime}),$ (215)
+$\delta\phi_{i}(t)=\int^{t}dt^{\prime}\sum_{j=1}^{N}S^{\phi}_{ij}(t_{1},t_{2})\sum_{\alpha\in\{\hat{0},\hat{0}^{\prime}\}}L_{j\alpha}z_{\alpha}(t^{\prime}),$ (217)
 where $S^{\phi}_{ij}(t_{1},t_{2})$ is the response function of the neurons. The dynamic equations for the cavity latent variables are
-$z_{\alpha}(t)=D_{\alpha}\left[\gamma_{\alpha}(t)+\frac{1}{\sqrt{N}}\sum_{\beta}\int^{t}dt^{\prime}G_{\alpha\beta}(t_{1},t_{2})z_{\beta}(t^{\prime})\right],$ (216)
+$z_{\alpha}(t)=D_{\alpha}\left[\gamma_{\alpha}(t)+\frac{1}{\sqrt{N}}\sum_{\beta\in\{\hat{0},\hat{0}^{\prime}\}}\int^{t}dt^{\prime}G_{\alpha\beta}(t_{1},t_{2})z_{\beta}(t^{\prime})\right],$ (218)
 where
@@
 $\psi^{z}(\bm{\omega})=N\left\langle C^{z}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{z}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\bm{L},\bm{D},\bm{R}}$ (223)
-by squaring and disorder-averaging Eq. 220. This requires us to consider several two-frequency correlation functions, which we define as
+by squaring and disorder averaging Eq. (222). This requires us to consider several two-frequency correlation functions, which we define as
 $\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\left\langle G_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (224)
@@
 $S^{z}_{\hat{0}\hat{0}^{\prime}}(\omega)$ $=\frac{D_{\hat{0}}}{\sqrt{N}}G_{\hat{0}\hat{0}^{\prime}}(\omega),$ (231)
-$S^{\phi}_{00^{\prime}}(\omega)$ $=\frac{1}{\sqrt{N}}(S^{\phi}_{\star}(\omega))^{2}F_{00^{\prime}}(\omega).$ (230)
-Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions (Eqs. 209-212 and 222225), we obtain
+$S^{\phi}_{00^{\prime}}(\omega)$ $=\frac{1}{\sqrt{N}}[S^{\phi}_{\star}(\omega)]^{2}F_{00^{\prime}}(\omega).$ (232)
+Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions [Eqs. (211)–(214) and (224)–(227)], we obtain
 $\Gamma_{C^{\eta}_{00^{\prime}}C^{\eta}_{00^{\prime}}}(\bm{\omega})$ $=\alpha r_{4}C_{\star}^{\phi}(\omega_{1})C_{\star}^{\phi}(\omega_{2})+\alpha^{2}\psi^{z}(\bm{\omega}),$ (233)
@@
 $\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=S_{\star}^{\phi}(\omega_{1})S_{\star}^{\phi}(\omega_{2})\left[1+S_{\star}^{\phi}(\omega_{1})S_{\star}^{\phi}(\omega_{2})\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}(\bm{\omega})\right].$ (236)
-We solve Eq. 232 and Eq. 234 simultaneously, yielding
+We solve Eqs. (234) and (236) simultaneously, yielding
 $\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}(\bm{\omega})$ $=\frac{\alpha r_{2}}{1-\alpha r_{2}S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})},$ (237)
@@
 $\psi^{\phi}_{\star}$ $=|S^{\phi}_{12}|^{2}\Gamma_{C^{\eta}_{00^{\prime}}C^{\eta}_{00^{\prime}}}+\left(\Gamma_{F_{00^{\prime}}F_{00^{\prime}}}S^{\phi}_{12}+\text{H.c.}\right)C^{\phi}_{12}.$ (240)
-Now, we substitute Eq. 237 into Eq. 238. Using $\Psi^{\phi}_{\star}=\psi^{\phi}_{\star}+C^{\phi}_{12}$, we obtain
+Now, we substitute Eq. (239) into Eq. (240). Using $\Psi^{\phi}_{\star}=\psi^{\phi}_{\star}+C^{\phi}_{12}$, we obtain
 $\Psi^{\phi}_{\star}=|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}\left[\Psi^{\phi}_{\star}+\frac{1}{\alpha\text{PR}^{D}}\frac{1-|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12}\right]+\frac{1-|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12},$ (241)
-where we used the previously defined quantities $g_{\text{eff}}^{2}=\alpha r_{2}$ and $\text{PR}^{D}=r_{2}^{2}/r_{4}$. Eq. 239 can be solved for $\Psi^{\phi}_{\star}$, yielding
+where we used the previously defined quantities $g_{\text{eff}}^{2}=\alpha r_{2}$ and $\text{PR}^{D}=r_{2}^{2}/r_{4}$. Equation (241) can be solved for $\Psi^{\phi}_{\star}$, yielding
 $\Psi^{\phi}_{\star}=\frac{1+\frac{1}{\alpha\text{PR}^{D}}|g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}{|1-g_{\text{eff}}^{2}S^{\phi}_{12}|^{2}}C^{\phi}_{12}.$ (242)
-This result is identical to Eq. 38 derived using the path-integral approach, demonstrating the consistency between the two methods.
+This result is identical to Eq. (39) derived using the path-integral approach, demonstrating the consistency between the two methods.
 Appendix M Numerical details
@@
 VII), the single-site two-point functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ depend only on $g_{\text{eff}}$ and can be numerically calculated as in i.i.d.
-networks (see, e.g., [10, 98]).
+networks (see, e.g., Refs.
+[10, 105]).
 In summary:
@@
     Since $\eta(t)$ is Gaussian, so is $x(t)$ for this linear form of $T[\cdot]$, allowing us to write $C^{\phi}_{\star}(\tau)$ in terms of Gaussian integrals over $x(t)$ and $x(t+\tau)$, which have marginal variance $C^{\phi}_{\star}(0)$ and covariance $C^{\phi}_{\star}(\tau)$.
-Due to our choice of $\phi(x)$, this expression simplifies to
+Because of our choice of $\phi(x)$, this expression simplifies to
     $C^{\phi}_{\star}(\tau)=\frac{2}{\pi}\tan^{-1}\left[\frac{C^{x}_{\star}(\tau)}{\sqrt{\left(C^{x}_{\star}(0)+\frac{2}{\pi}\right)^{2}-C^{x}_{\star}(\tau)^{2}}}\right].$ (243)
   2. 2.
-    Squaring the single-site picture gives a second-order ODE $\partial^{2}_{\tau}C^{x}_{\star}(\tau)=C^{x}_{\star}(\tau)-g_{\text{eff}}^{2}C^{\phi}_{\star}(\tau)$.
+    Squaring the single-site picture gives a second-order ordinary differential equation (ODE) $\partial^{2}_{\tau}C^{x}_{\star}(\tau)=C^{x}_{\star}(\tau)-g_{\text{eff}}^{2}C^{\phi}_{\star}(\tau)$.
 Since the rhs depends only on $C^{x}_{\star}(\tau)$ and $C^{x}_{\star}(0)$, we can consider it to be the negative derivative $\partial V/\partial C^{x}_{\star}(\tau)$ of a potential $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$ with an explicit dependence on the initial condition $C^{x}_{\star}(0)$.
   3. 3.
-    The rhs is integrated w.r.t. $C^{x}_{\star}(\tau)$ to provide an expression for $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$, which due to our choice of $\phi(x)$ simplifies to
+    The rhs is integrated with respect to $C^{x}_{\star}(\tau)$ to provide an expression for $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$, which due to our choice of $\phi(x)$ simplifies to
     $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))=-\frac{1}{2}C^{x}_{\star}(\tau)^{2}+g_{\text{eff}}^{2}\left[\frac{2}{\pi}\sqrt{\left(C^{x}_{\star}(0)+\frac{2}{\pi}\right)^{2}-C^{x}_{\star}(\tau)^{2}}+C^{x}_{\star}(\tau)C^{\phi}_{\star}(\tau)\right].$ (244)
@@
     Finally, Euler integration of the $C^{x}_{\star}(\tau),\partial_{\tau}C^{x}_{\star}(\tau)]$ dynamics with these initial conditions gives its full time course, along with $C^{\phi}_{\star}(\tau)$ via Eq.
-241.
+(243).
 We integrated the Newtonian ODE from $\tau=0$ to $\tau_{\text{max}}=200$ with step size $d\tau=0.025$.
-When then use an FFT to obtain the frequency-space representation.
+We then use an FFT to obtain the frequency-space representation.
 The frequency-space representation of the response function $S^{\phi}_{\star}(\omega)$ was computed directly as
@@
 with $\left\langle\phi^{\prime}\right\rangle_{\star}=1/\sqrt{1+(\pi/2)C^{x}_{\star}(0)}$.
-Various expressions for $\Psi^{a}_{\star}(\omega_{1},\omega_{2})$ were then computed in frequency space before transforming back to the time domain by a 2D inverse FFT.
+Various expressions for $\Psi^{a}_{\star}(\omega_{1},\omega_{2})$ were then computed in frequency space before transforming back to the time domain by a two-dimensional inverse FFT.
 For structured $\bm{L}$-$\bm{R}$ overlaps, the $\mathcal{O}(1)$ self-coupling in the single-site process makes $x(t)$ non-Gaussian, preventing the use of Gaussian integrals in the DMFT solution.
-We there obtain the solution via standard numerical methods, enforcing self-consistency among the $\eta^{x}(t)$ autocovariance function, $Q_{\star}(\tau)$; the self-coupling kernel, $R_{\star}(\tau)$; the response function, $S^{\phi}_{\star}(\tau)$; and the $\phi(t)$ autocovariance function, $C^{\phi}_{\star}(\tau)$.
+We therefore obtain the solution via standard numerical methods, enforcing self-consistency among the $\eta^{x}(t)$ correlation function $Q_{\star}(\tau)$, the self-coupling kernel $R_{\star}(\tau)$, the response function $S^{\phi}_{\star}(\tau)$, and the $\phi(t)$ correlation function $C^{\phi}_{\star}(\tau)$.
 In summary:
@@
   2. 2.
-    $Q_{0}(\tau)$ and $R_{0}(\tau)$ are computed in Fourier space using the current values of $C^{\phi}_{0}(\tau)$ and $S^{\phi}_{0}(\tau)$ via Eq. 49, with the average taken over $(D,\rho)$.
+    $Q_{0}(\tau)$ and $R_{0}(\tau)$ are computed in Fourier space using the current values of $C^{\phi}_{0}(\tau)$ and $S^{\phi}_{0}(\tau)$ via Eq. (50), with the average taken over $(D,\rho)$.
   3. 3.
@@
     Each of $N_{\text{sample}}$ single-site processes $x_{s}(t)$ is Euler integrated according to Eq.
-47a with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi(x_{s}(t))$.
+(48a) with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi\bm{(}x_{s}(t)\bm{)}$.
   5. 5.
-    These time series $\phi(x_{s}(t))$ and $\eta^{x}_{s}(t)$ are used to compute empirical values $(C^{\phi}_{1}(\tau),S^{\phi}_{1}(\tau))$ for the autocovariance and response.
+    These time series $\phi\bm{(}x_{s}(t)\bm{)}$ and $\eta^{x}_{s}(t)$ are used to compute empirical values $(C^{\phi}_{1}(\tau),S^{\phi}_{1}(\tau))$ for the autocovariance and response.
 The latter is computed by the Furutsu-Novikov theorem: $S^{\phi}(\omega)=C^{\phi\eta}(\omega)/C^{\eta}(\omega)$.
@@
 With these quantities, the various components of Eq.
-51 and the expression for $\Psi^{\phi}_{\star}$ are straightforward to compute.
+(52) and the expression for $\Psi^{\phi}_{\star}$ are straightforward to compute.
 We took $\left\langle\cdot\right\rangle_{D,\rho}$ averages in this case by averaging over every other discrete value of $D_{a}$ and $\rho_{a}$.
 M.2 Simulation details
-Each network was integrated via Runga-Kutta (RK4) with $dt=0.05$.
-The empirical lagged covariance matrix was computed by averaging over $N_{\text{loop}}$ individual covariance estimates, each of which is based on $T=2000$ discrete, evenly spaced time samples $t^{i_{\text{loop}}}_{k}$, with $t^{i_{\text{loop}}}_{k}-t^{i_{\text{loop}}}_{k-1}=1$:
+Each network was integrated via Runge-Kutta with $dt=0.05$.
+The empirical lagged covariance matrix was computed by averaging over $N_{\text{loop}}$ individual covariance estimates, each of which is based on $T=2000$ discrete, evenly spaced time samples $t^{i_{\text{loop}}}_{k}$, with $t^{i_{\text{loop}}}_{k}-t^{i_{\text{loop}}}_{k-1}=1$,
 $\widehat{C^{\phi}_{ij}(\tau)}=\left\langle\frac{1}{T-1}\sum_{k=1}^{T}\phi_{i}(t^{i_{\text{loop}}}_{k})\phi_{j}(t^{i_{\text{loop}}}_{k}-\tau)\right\rangle_{i_{\text{loop}}=1}^{N_{\text{loop}}}.$ (248)